A Selected Bibliography of Publications by, and about, Benoît Mandelbrot

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

05 June 2024 Version 1.62

Title word cross-reference

(2 + 1) [Zha09]. 1/f [Ano99, Bak00, Gar78, Man67a, Man99b, Man02d, VC75, VC78, Vos79]. $14.00 [Man61f]. $14.95 [Kir79, Sto79]. 2 [Man85d, Mar04, NCS04, OdlCA02]. 2 + 1 [Mac08]. 20 [FMS96]. 3 [Bou09b, Car96, CB97, Gin02, HSK89, HD91, HKHP11, Maj98, Mar10, Nor82]. $3.00 [Man60c]. $30.00 [Tay13]. $32.50 [Whe83]. $39.95 [Ano98a, Ano03a]. 3x + 1 [XX07]. $42.95 [Ano99]. $5.00 [GM61, Man60e, Man60e, Man60b]. 5/3 + B [Man76b]. $54.95 [Ano03b]. $6.75 [Man60d]. $8.75 [Man62a]. 80 [FMS96]. [0, 1] [BJM10b]. ^{th} [Wei05]. C [CZ98]. M [Man83c, Man83d]. ϵ [Pic86]. exp(−αζ + Z/ζ − Z) − 1 = 0 [YHHS93]. f(α) [Man89b, Man90b]. ≫ [CB59]. H [Man69]. J [CZ98]. k [AS92, STI99]. L [Bra94, AO97, AO01, OdlCA02]. λ [Man80]. ≪ [CB59]. M [Car99a, CZ98, Man85a, Man85b, Man85d, Old01]. μ [Man83c, Man83d]. n [Cha11]. n^{2} [Man85a]. q [MML01]. Q_{α}(R^{n}) [Yue10]. R [Man85b]. R/S [MW69c, Man72b, MT79, Man02d]. ρ [TW11]. σ [Ate93]. Z [CZ98, Man80, Man83c, Man83d]. z ⟵ zw + c (w = a + ib) [CZ98]. z → z^{−n} + c [Shi93]. z → λ(z + 1/z) [Man84h]. z → λz(1 − z) [Man80, Man85c]. z → z^{2} − μ [Man83c, Man83d].

-Automorphisms [Ate93]. -D [CB97, Car96, HSK89, HD91, Nor82]. -dimensional [OdlCA02, Zha09]. -exponential [MML01]. -gram [Cha11]. -martingales [BJM10b]. -regular [AS92]. -set [Car99a, Man85a, Man85b, Man85d, Old01]. -spectra [Man69]. -system [Bra94].

0-7546-3933-9 [Pic07]. 0-902205-58-7 [Fri04].

1 [Man01d]. 10 [CHR54]. 11th [CMR06]. 14 [BGL67, LBG63]. 150 [GM61]. 1960 [Man61a]. 1974/75 [Man75f]. 1977 [MS78]. 1980 [AGP81]. 1982 [CR83]. 1983. [SMR84]. 1985 [PT86]. 1989 [AF90]. 1992 [KK93]. 1993 [Ano93b]. 1995 [EPV96].

2002 [LvF04a, LvF04b, Man02c]. 2003 [ACM03]. 2005 [CMR06]. 200th [MS78]. 2010 [And11]. 20th [Man02b]. 2nd [Man62a].

3 [Kur90, Voj89]. 3-7643-1771-X [Kur90, Voj89]. 3.1 [Lar90]. 31 [Man75f]. 32.50 [Ano98a]. 33 [Seu12]. 34.00 [Ano99]. 3D [Hug98].

40 [BCM92]. 41 [Man73d]. 44th [IEE03]. 459 [Man62a].

65th [AF90]. 6th [DLM99].

7 [Tay13]. 75 [Bri85].

’89 [RW89].

’92 [KK93]. ’96 [BDLS96]. 978 [Tay13]. 978-0-307-37735-7 [Tay13].

= [M^{+}85].

A. [Man55b, Man61a]. Abb [Sto79, Voj89]. Able [Man10b]. aboard [Ano89b, Dew89]. Abstract [Kel00a, Kel00b]. abstracts [GMA^{+}98, LMP85, MP84]. Abundance [BATL88]. abuse [Old01]. acceleration [SH96]. Accelerator [Chu11, Chu12]. accelerators [Mar10]. According [Man61d]. ACM [ACM03]. acquisition [EW96]. across [Man01a]. activities [PJS99]. Activity [GM64, BBM79]. Adaptation [Man51a, Man51b, Fla98, Man51a, Man51b]. Adaptive [LLP96]. additifs [Man60g, Man62e]. Additional [Man84a]. Additive [yWjCnG07, Man60g, Man62e, XRZ09]. addresses [KLPS06]. Administration [PR84b]. Advanced [BBBH92]. Advances [FpDS99, DLM99]. advancing [DDH10]. Adventurers [Gar88b]. Aeolian [Pic94]. aesthetic [SCNT03]. aesthetics [MRM^{+}05]. Affine [CD93b, LVEB09, NR95, BV11, Bou09a, Bra94, DV93, DHKT95, Man85g, Man86c, Man86d, Man86e, Mar04]. affinity [Ano99, Ano03b, Man99b, Man02d]. after [DK06, KNBZ12]. against [VHA05]. Age [VBM98, EHH05]. aged [Ano10a]. agglomerates [CGC^{+}09]. agglomeration [MML01]. aggregates [BRRK94, EMW92a, EMW92b, MP84, Man91a, ME91, NFT^{+}01, NCS04, Vor02]. Aggregation [MKA02, KVMW95, Vor02]. Airspaces [Hel92]. Alamos [CR83]. Albert [Lee59]. Aldershot [Pic07]. aléatoire [Man59c, Man74a, Man74b]. aléatoires [KM65, Man74a, Man74b, Man75b, Man78b]. algebra [Gin02, Old01]. Algorithm [All90, MH94, BCM91, BCM92, GJV17, Gar88a, GAH93, GTL11, JT96a, LPZ08, PU08, Ran91, SL97, UP07, Woo94]. Algorithms [Lew89, RBM96]. Allen [Man60d]. Allendoerfer [Man60d]. Allometric [WBE99]. along [FPR92]. alphabetic [FM02a]. Alternative [CGM96, Man96a, Man78b]. alternatives [Man95e]. am [PR84b]. America [Ano03a]. American [LvF04a, LvF04b, ALW^{+}88]. among [Est11]. Amount [Man64d]. AMS [DK06]. AMS-IMS-SIAM [DK06]. analog [Wes94]. analogue [SH92]. analyse [BBB^{+}10]. Analysis [BB99, BP17, BBBH92, CL87, Fis95, GT07, JSM06, LL97, Man59d, Man60e, Man60c, Man06c, MS78, TMJ06, VBM98, VEJA89, ADdM^{+}12, BM04c, BBB^{+}10, Bin24, CGC^{+}09, CP98, DR99, DSB96, DdMdO^{+}99, EPV96, Hig11, JT96a, JTEA91, LKH97, Man67d, Man72b, MT79, Man06b, MV02, NIS^{+}03a, NIS^{+}03b, OMK^{+}05, RRB06, Sig96, SCC^{+}07, TMJ99, TKR^{+}97, TdRGD09, VC86, Man61h]. analytic [BB79, BBM79]. analyze [Cas96]. Angeles [Hou95]. Angular [MKA02]. Animated [Dev92, Man01a, SSK93, ZPSJ90]. Animation [Bur02]. animées [Man01a]. Annals [Man60e]. annealing [Sah93]. anniversary [ALW^{+}88, MS78]. Annual [IEE03, LvF04a, LvF04b]. Anomalies [Man03d, Man86e, Man93c]. anomalou [Man90b]. Anthony [Man74e]. antialiased [HD91]. antipersistent [CGM95, Man95b]. Antipodal [MS94]. Anwendung [Man57a]. Anwendungsbeispielen [Kau92]. anxiety [TKR^{+}97]. aortic [PM94]. Apollonian [Bou06]. Apostel [CC58, Lee59]. Apparent [Man63d, Man01d]. appearance [DK06]. appeared [Man61a]. Application [BATL88, Man96b, ZWZL06, BB79, BM93, Izs06a, JM96b, Kau92, Man57a, Man65d, Mas05b, Thi89, VM95, WGKS12]. Applications [BBBH92, La 86, MV68, CHR54, DR99, Esp97, Fer95, GMA^{+}98, Jaf98, LvF04a, LvF04b, Suz91, Rai03]. Applied [GM61, Jak84]. apprenticeship [Man02c]. Approach [Pic86, BK03, CL87, EB73, EPV95, FHD09, LKH97, MH05c, MH09, Mar04, MC02, PH91, SW92, Vla94, VM95, YJ09]. approche [MH05c, MH09]. approx [Man61f]. Approximant [BR78, Bar73, BR74b, Bar76]. approximants [BR74a, Bar75, BB76, EB73].

Approximate [You13, BR74b]. approximation [BA78, KGN^{+}24, Man72a, Man75b]. approximations [AK13]. April [Rai03]. Arbeiten [Man61a]. Arbitraged [Man71c]. architecture [NFT^{+}01, PM94, YJ09]. area [AK13, ES92, Imr06, Man75j]. Aref [Ano93a, Ano93c]. arguments [RPÁM04, RPAM06]. Arising [Dev88, Man49, Man00]. arithmetic [PH90]. Arithmetics [Ano89a]. Arrow [Man61e]. Art [Man89c, RW89, Ent89, Kel00c, Man89a, Man93b, Man01a, Man02a, MM91, Rad96]. arterial [SCC^{+}07]. artery [SCC^{+}07]. arthropod [MLDW85]. Arthur [Lev05]. Artificial [Imr06, DKW94a]. Artist [Gin02]. artistic [Car99a, Loy91, Ye02]. Arts [Man81b, Man07]. Ashgate [Pic07]. ASI [Fis95]. Aspects [Man68, HWM87, Izs06b, LMP85, Man80, MP84, WSM88]. Assessment [NIS^{+}03a, NIS^{+}03b]. Assisted [Vrs86]. Associated [LVEB09, Vrs86, dACSS12]. Association [Ano03a]. astronomy [FF91]. Asymptotics [MDM02]. atmospheric [KK93]. attractions [Ber86]. Attractors [GOY87, Man77c, Man79b, PS88b, CM91, Ghi86, Man77a]. August [KK93, RW89]. Ausstellung [PR84b]. Author [Man60e]. authorship [Aus14]. automata [AvHP^{+}97, CD93b, Sah93]. Automatic [Pic94]. Automaticity [AvHP^{+}97]. automaton [SS91]. Automorphisms [Ate93]. Autumn [Loy91]. aux [DT89, Man56b, Man78b, Ste98]. avalanche [GDC^{+}10]. Averages [GT07]. Away [Sel10]. axial [Man49].

B [AF90, Ano98a, Ano99, Ano03a, Ano03b, Ano05, Bak00, BF12b, Boy83, Bre79, Bur84, Cam96, Can84, Cha84, CC58, CB59, Dys78, Fer83, Goo78, Goo84a, Goo84b, Goo91, Hir89, How78, Ios79, Kac78, Kal03, Kil78, Kir79, Kir83, Kur90, Lan89, Man60d, Man61b, Man62a, Man74e, Man01d, Mir01, Pei10a, Pei10b, Sch05, Seu11, Seu12, Spa84, Sto79, Tay13, Voj89, Whe83, Man60a]. B. [Bri85, KNBZ12, Man62a]. backbone [GAMK81]. Bacteria [Man74c]. balancing [JH97]. Balaton [JV98]. bands [MM08, PRÁM04]. bandwidth [MdF05]. Barcelona [DLM99]. Barnsley [Goo91, Lan89, Lev05, Jon90a]. base [Gof91, Man62e, Man79a]. Based [HLTP01, BB79, CP98, Dav95, Dre01, EW96, GAH93, KPY96, LPZ08, Man57c, Man59c, Man62e, Man79a, PH91, Pum96, SL97, TQ10, TTMB96, WY96, WCH03, WW08, XX07]. basée [Man57c, Man59c]. Basel [Kur90, Voj89]. Bases [VBMS96, Gil86]. Basic [GFB^{+}10, Man86c, Van10]. Basin [GOY87, CGH02]. Batrachion [Pic95]. BBC [Ano10a]. be [Man71c, Wal96]. beat [Ste88]. Beating [Fla91]. Beautiful [MPM02]. Beauty [Ano87, Dew87, Gar91, Har90, PR86, Fla98, LG10b]. become [ALW^{+}88]. becomes [Man92a]. behavior [CS87, GK03, MH04a]. behavioral [Man57a]. behaviors [Zha09]. Behaviour [EJM91, MH08b, Gru05, KGN^{+}24, MH05b]. Behind [Mag04, Man88b]. Beneath [Pop22]. benefit [Zaa10]. Benoît [AF90, Ano98a, Ano99, Ano03a, Ano05, Bak00, Boy83, Bre79, Bur84, Cam96, Can84, Cha84, CHR54, Dev92, Dys78, FR98, Fer83, Goo78, Goo84a, Goo84b, Hir89, How78, Ios79, Kac78, Kal03, Kil78, Kir79, Kir83, LvF04a, LvF04b, Lee59, Lev05, Mir01, Sch05, Tay13, LMNW05, Whe83, And11, Ano95, Ano10b, Ano10c, Ano10a, Bar85, Bar08, BF12a, BF12b, Baw11, Bou11, Cam00, Con10, Dav84, DrKP10, Est11, Far11, Gil04, Gom10, Gru05, Gui87, Jaf11, Kah74, KP76, Kir11, Kur90, LG10a, Man06a, Man12, Pei10a, Pei10b, Per10, SBvB^{+}12, Sel10, Seu11, Seu12, Shl05, Sim61b, Taq13, Tay11, Ten10, Val11, Voj89, Wei05]. Benoít [Hof10]. benzodiazepine [TKR^{+}97]. Berlin [Ano98a]. Berry [Man57b, SH92]. Better [Adl12, BS88, KRP^{+}91]. Between [Man03c, BBB^{+}10, BL93, CZ98, EG94, Man67a, Man84c, MH05a, MH08a, Orl76]. Beyond [Mon01, Man04a, Sch05]. Biaccessible [Zak06]. Bibl [CB59]. Bibliometric [Fai69, Fai05]. Bicentenary [MS78]. Bifurcation [Pic87, Pro97]. Bifurcations [Bur02]. Bilder [M^{+}85, PR84a, PR84b]. binary [Pro97, Sen90, YM94]. binding [TKR^{+}97]. binomial [MJ97]. biological [GDC^{+}10, Man56b, MS78]. biologiques [Man56b]. Biology [SG00, Wil97]. biomorphs [OdlCA02]. Biophysical [PR84b]. Biophysikalische [PR84b]. bird [HRO02]. Birkhäuser [Kur90, Voj89]. Birth [Man74c]. Birth-and-Mutation [Man74c]. Birthday [Wei05, AF90, Shl05, LMNW05]. bisection [Ran91]. bit [Slo94]. Bivariational [BB76, BR76, BR77, RB79]. black [FHD09]. block [GAH93, KM11, KPY96, Liu95a]. block-coding [GAH93]. blood [SCC^{+}07]. Bochner [Man60e]. body [HRO02, MLDW85, Suz91]. Boltzmann [BC80, NCS04, Man61d]. Bombay [VBMS96]. Book [Ano93a, Ano98a, Ano99, Ano03a, Ano03b, Bak00, Bri85, Dev92, Fri04, GM61, Goo84a, Gru05, Ios79, Kir79, Kir83, Kur90, Liu06, Man60e, Man60a, Man60b, Man60c, Man60d, Man61b, Man61e, Man61c, Man61d, Man61f, Man61a, Man62a, Man63a, Pic07, Spa84, Sto79, Tay13, Voj89, Whe83, Ano91]. Books [Tay13]. Bose [Mas05a]. Boston [RW89, Voj89]. bottom [EJM91]. boucle [Man78b]. Boundaries [GOY87, ADdM^{+}12, Bis95, CGH02, DdMdO^{+}99]. Boundary [BR76, BH89, Shi98, AST07, Man85b, Man85d]. Boundary-Value [BR76]. Bounded [GKL03]. Bounding [BT81, Bou85, Bar73]. Bounds [BR76, BR78, BR74b, BR75, BB76, Bar76, BR77, RB79]. box [CP98, FHD09, FpDS99]. box-counting [CP98, FpDS99]. Bradford [Fai69, Fai05]. Braid [dLKHS06]. brain [HKS^{+}05, HTHE05, RPS^{+}06]. Branched [Dha05, Fle97]. Branching [HKE01, YMK94]. breakdown [Lan97]. breaking [VJ06]. Breaks [MB94]. Bremen [AGP81]. bridge [Man67a]. Brief [Mit03, Vos79, WLWL22]. Britain [Man67b, Man91c]. Broken [Man72a]. brown [Gar78]. Brownian [BB99, CGM95, CGM96, MV68, Man75b, Man82b, Man95b, Man96a, Taq13]. brownien [Man75b]. Buchbesprechung [Ano05]. Bud [CZ98]. Bud-sequence [CZ98]. buffalo [Fin89]. Buffon [MS78]. building [PR84b]. Bureau [SMR84]. Burridge [Has07].

C [Dev92, Lev05, Man60d, Roo10, Esp97, Man04a]. C/C [Esp97]. CA [ACM03, LvF04a, LvF04b]. Cache [VMH^{+}83, Thi88, Thi89]. cache-miss [Thi88]. Calculations [Pic86]. Calculator [For04]. Called [Ano87, Dew87]. Cambridge [Fri04, IEE03, Man60a, Man60c]. Can [Ewi95, Man71c]. cancer [TdRGD09]. Canonical [Man74e, Man59c, Man56d]. canoniques [Man59c]. Canopy [MF99]. Cantor [DM90, Man82c, YJ09]. capabilities [Esp97]. Capacity [Man61a]. Captures [BATL88]. Card [Sch90]. cardiovascular [NIS^{+}03a, NIS^{+}03b]. career [Tem96]. Carlo [Pic86]. Carpets [BB99]. carrier [Man74d]. cartoons [Man05b]. Carus [Man60c]. cas [Man75b]. Cascades [LR00, BPW09, BJM10a, Man74d]. Case [CGMST95, Man75b, Tem96, VHA05]. Cat [HKE01]. Catalog [RW89]. catalyst [MP84]. catchments [KFN00]. categories [Man56c]. cauliflowers [RPÁM04]. cavities [DdCVR03]. Cayley [PSvH84]. CBF [HTHE05, OMK^{+}05]. CBV [EHH05]. celebrates [ALW^{+}88]. céleste [Man75e]. celestial [Man75e]. Cell [Ald07]. cellar [Cia88a]. cells [TdRGD09]. cellular [AvHP^{+}97, Sah93]. Center [CR83, BA78]. Central [Per96]. century [Cha90, Man02b]. Ceramic [Sch89]. Cerebral [HLTP01, TKR^{+}97, SCC^{+}07]. Certain [Man63e, Man63f, Man72c, BM99, Gar88a, Man62e, Man64b, Man85e]. certaines [KP76]. certains [Man62e]. Chadwick [Man60a]. chain [LO96]. Chains [Man60b]. Challenge [Man03c]. Chance [Bre79, Dys78, Goo78, Ios79, Kac78, Kil78, Man70c, MA79, Sto79, Kir79, Man64b, Man73a, Man75d, Man77b, Man84f, Man89d, Man98c, Man09a, Man10c, Man75c]. Chandler [MM70]. Change [How78]. changements [Man66d]. changes [JV98, Man66d]. Chaos [Ano93a, Ano93c, Are93, BGR94, BJM^{+}88, CFPS86, Gin02, GOY87, Hir89, Jon90a, M^{+}85, Mir90, OdlCA02, PR84a, PJS92a, PJS92b, PJS92d, PJS93, PJS04a, PJS04b, Pic87, Pop22, Sch05, Swe91, Ano91, CR83, Dev90, EG94, Fla98, Gul12, LK98, Man84h, Man86a, Man04a, PH90, SBvB^{+}12, Ste98, TE92, Abb95, Nat91]. Chaotic [BD86, Dev88, Mac08, PRÁM04, BF95, GM88]. Chapters [BGL67, LBG63]. character [MPP84]. Characterization [Vos86, Vos88, LSL86, NDS85]. characterizations [BB79]. Characterizing [Dys78]. charge [FBVdlN99]. Charles [Pic07]. CHD [Hig11]. Chelsea [Man62a]. chemical [DL87]. Chemie [PR84b]. Chemistry [PR84b, KFN00, Ano93b]. chiffrées [Man67c]. choice [Ye02]. Christos [Win05]. Ciarcia [Cia88a]. Cie [Man61f]. circle [Car99b, FC00, Man72d]. circuit [Cia88a]. Circuits [BM63]. citations [Van90]. clamped [Ska99]. Clark [Man73d]. Clarke [Lev05]. Class [BR78, Sim55, BC80, CGM95, HJ94, Man59d, Man61h, Man65d, Man89b, Man95b, Sim60]. classe [Man65d]. classical [Bau96]. classification [MRM^{+}05]. classify [Cas96]. Classifying [MPM02]. Classroom [Abb95, FL91, Nat91, PJS91, PJS92d, PJS92e, PJS99]. clay [VOL^{+}86]. clef [Man75c]. climatologique [Man65d]. Clinical [HLTP01, Mai90]. cloistered [Bai97]. closeness [AK10]. cloud [DLS93]. Clouds [MSLG00, Man10b, SSK93]. clus [Man65c]. cluster [FBVdlN99, GTL11, UH96, Vor02]. cluster-cluster [Vor02]. Clustering [BM63, DM73, Ham95]. Clusters [HASM96, AAM^{+}03, GM84b, Man77d, Man83a, Man84e, MG84, ME90, MS94, YMK94]. Co [Bri85, Kir79, Kir83, Aus14, Van90]. co-authorship [Aus14]. co-citations [Van90]. Coast [Man67b, Man91c]. coastlines [Kap86, Man75j]. Codebook [Ham95, Kom95, HS93, Sig96]. Codes [Sta96]. Coding [Ano89a, GAH93, Hür93, HS93, Hür95, KM11, LKH97, Liu95a, Man55a, RBM96, Slo94, TTMB96, WY96, WW08, ZZZ^{+}96]. coefficients [CD93a]. cognitive [OMK^{+}05]. coherent [Mac08, Man84b]. collaborative [Bou09b]. collage [DHKT95]. Collapse [TSS85]. Collection [GM61]. College [Man02c]. Colliers [Man78b]. Colloids [Rar89]. Colonies [Ric94]. color [CCC00]. Colours [Cla10a, LG10c, Lev05, LG10b]. Combination [Van10]. Combinatorial [VEJA89]. Comment [GM84a, Man61g, Man68, Man71b, Man82a, Man84b, Man84c, Man01d]. Comments [Man73d]. Communication [CHR54, Man53a, Man54b, Wu99]. communications [Man52a, Man54a, Man54c, Man65c]. compact [Man92a]. Company [Sto79]. comparative [TPVG06]. Comparison [EW96, FSR94]. Competitions [SL11]. complete [Bra94]. completeness [STI99]. Complex [Abb95, Adl12, DK06, Nat91, PJS92e, Sch11, BB76, BJM10a, BJM10b, BSV92, Car96, CD93b, Dew85a, DM92, Eng93, Ent89, Fla98, Gil86, Gof91, Man80, Man83c, Man83d, Man85c, M^{+}85, PR84b, DK06]. Complexity [Gar91, Har90, Man06c, SG00, Man06b]. Components [GF95, Man85e]. comportement [CHR54]. composing [XRZ09]. Comprehensible [SO93]. Compress [BS88]. Compression [Fis92, MFCM96, BE95a, BE95b, CK96, CD93a, Dav96, DV93, EW96, Ham95, HMS96a, HMS96b, JT96a, JM96a, KKL95, Kom95, KPY96, LV96, LLP96, PH91, SR96, Sau96a, SH96, Sau96b, Sig96, UH96, WCH03, Woo94, WM94, Wu99, ZY94]. computable [Her05]. Computation [DAB99, Bin24, BL93, BCM91, BCM92, CCRR91, MdF05, Sau87]. Computational [TdRGD09, Fla98]. compute [GJV17]. Computer [Dew85a, Dew85b, Dew87, Dew89, IEE03, MW69b, Man82a, PH90, Rad96, RW89, VEJA89, Vrs86, CCRR91, Fla98, Gar88a, HCF95, JTEA91, Kok94, M^{+}85, Roj91, Thi89]. computer-generated [JTEA91]. Computers [MPM02, Wai88]. Computing [ACM03, MC02, PU08, Rai03, RAA^{+}08, UP07]. Concentration [MG98, Mir01, Ano98a, Man97b]. concept [Man65c]. Conceptional [VBM98]. concepts [Smi91]. concerning [Man72e]. condition [Sta96]. conditional [Man65c, Man67d, Man75a]. conditions [AST07]. Conference [CMR06, DK06, GMA^{+}98, IEE95, VBMS96, AF90, BDLS96, CR83, Ano89a]. Configuration [GT07]. conjecture [CZ98, Man85a, SH92, Man85d, Man85e]. connected [Wai88]. connection [Orl76]. Connectivity [CCD88]. consequences [DSS90]. conservation [CJK11, Dro10]. constant [Man67c]. constantes [Man67c]. constrained [Izs06a]. construct [Nar07]. Constructing [Cha11]. Construction [Bou09a, DHN85, BM93]. constructions [Man84i]. Constructive [BB79]. contact [BBB^{+}03]. Contacting [MF99]. contaminant [KFN00]. content [Man95e]. contes [Man75c]. Context [DKW94b, RW89]. continental [DSS90]. Continued [Rat93]. Continuous [GT07, Man97a, WGKS12, RM97, Man85c]. Continuous-time [WGKS12]. continuum [Sah93]. Contractivity [Hür93]. Contribution [Man52a, Man53a, Man52a, Man53a]. Contributions [Man63a, Man61b]. control [MC02]. convection [DdCVR03]. Convergence [BJM10a, BJM10b, Ska99]. Convex [DM92]. convolution [SH96]. copper [BB84]. copulas [dACSS12]. cordonnets [Man78b]. cords [Man78b]. Corneal [KNBZ12]. corner [YHHS93]. corrected [PJS93]. Correction [Bar75, Man72c, BB76, Man55b]. Corrections [Man54a, Man54a]. correlation [CCRR91, EG94, HKS^{+}05, MT79, MV02, WWGF05]. Corrélations [Man79a, Man75e, MS94, Man75e, Man79a]. Corrigendum [BCM92, Man75f]. cortex [HWEH05]. cosas [FR98]. cosine [FSR94, ZY94]. cosmic [Ohn90]. cosmic-ray [Ohn90]. cosmographic [Man75a]. Cosmological [LS92]. cosmos [PR84a]. cost [Wu99]. could [Wal96]. Coulton [Sel10]. counting [CP98, FpDS99]. coupled [IKP01]. Cours [Lef82]. course [LK98]. Cousins [Ano87, Dew87]. covariance [Man72b]. covered [Man93a, SAM93]. coverings [Man72d]. CQUATS [Gin02]. cracks [TM86]. Craig [Man60d]. creases [Man88c]. Creates [PJS04b]. creation [Van09]. Critical [GMA80, Man03d, AGMK81, AAM^{+}03, FR92, MS94]. critique [Man59d, Man61h]. crying [Pic95]. CUG305 [Sch90]. cultural [ABC^{+}94]. Culture [Man74c]. Curacao [EPV96]. current [KK93, Man67a, Man91a]. Curves [McK19, PB89, PS88b, AO97, BGR94, BSV92, Dek12, Gar78, Nar07, ODA03]. Cut [Gai06]. Cut-out [Gai06]. cutouts [Man72f]. cybernetics [Man57a]. cybernétique [CHR54]. Cycles [Man66d, Man72b, Man73b]. cycling [Ald07]. Cylindrical [CGMST95, BM02].

D [Abb95, CB97, Dev92, Maj98, Man62a, Mar10, Nat91, PR84a, Bou09b, Car96, Gin02, HSK89, HD91, HKHP11, Mar04, NCS04, Nor82]. D.C [Ano03a]. Damage [Man64d]. dans [CHR54, Man56b, Man59b, Man67f, Man73a, Man79a]. Darkest [Ano90]. Data [VBMS96, Bur81, CD93a, DKW94a, HKS^{+}05, KPY96, TPVG06, Win05, WWGF05, Woo94, vW91]. David [Lev05]. DC [IEE95]. DCT [FSR94]. decay [Bra94]. decipherability [Man65b]. Decipherment [Man60a]. Decision [Man61c]. decoder [KKL95]. decomposable [Man75e, Man75e]. Decomposition [Man74e]. deduction [Man75e, Man75e]. Defined [Kaj83a, Kaj83b, Man89g]. Definition [Man57a, Mil86, Man76c, CHR54]. degeneracy [BM04c]. degree [DM73]. degrees [MF09]. Della [Man60h]. demi [Cha90]. demi-siècle [Cha90]. demons [Man52b, Man52b]. dendritic [BRRK94, NFT^{+}01]. Density [CM91]. d’Entretien [Lef82]. Dependence [Man03c, MW69c, Man70b, Man01b, RVA00]. Dependent [DKW94b, Man75g, GH98, Man75f, Man75h]. d’Épistémologie [CB59]. deposition [Fle97]. Derivation [Man64c]. derived [Gil86, Man72a]. Description [Fai69, Fai05, Man59b, BSV92, KCC89]. descriptions [Mor96]. design [ODA03]. d’espèces [Man56b]. Detecting [WCH03]. Determination [AO01, MS78]. determines [PM94]. determining [DR99, Gar88a]. Deterministic [PJS04b, HSK89]. Deutscher [Man61a]. deux [Man54a, Man54c]. Devaney [Ano93a, Ano93c, Goo91, Lan89, Are93, Hal96, Swe91]. Developing [MVW^{+}03]. Development [Esp97]. Deviations [MKA02]. Diagnosis [OMK^{+}05]. Diagram [Ano93a, Ano93c, Are93, Dev91, Swe91, Dev90]. Dictionary [DG99]. did [Had88]. Diego [ACM03, LvF04a, LvF04b]. Dies [Hof10, Ano10b, Ano10a, Ano10d]. Dietmar [Goo91, Lan89]. Dieu [Ste98]. difference [Bar76, DL92]. Differences [MT67]. different [BP84]. differential [BB79, CT91]. Diffusion [GM83, MKA02, EMW92a, EMW92b, KVMW95, VCM^{+}11]. Diffusion-Limited [MKA02, EMW92a, EMW92b, KVMW95]. diffusions [Man84e]. Digital [DKW94b, BM93, KM11]. digitalized [Imr06]. Dimargaris [ME72]. Dimension [Hel92, Man67b, Man79b, MA79, Pic86, Shi98, VBM98, WBE99, ADdM^{+}12, BM04c, BM97, Cah89, Cas96, CCRR91, DdMdO^{+}99, EG94, FpDS99, Gar88a, Ghi86, Gil86, GBR^{+}09, GFB^{+}10, Imr06, MJ86, Man74d, Man75a, Man75i, Man75d, Man75j, Man76a, Man76b, Man77b, Man77d, Man84d, Man84f, Man85b, Man85d, Man85g, Man86e, Man89b, Man89d, Man90c, Man98c, Man08, Man10c, MLDW85, NM09, ODA03, PU08, Sau87, Thi88, Thi89, TdRGD09, UP07, Van90, Vor02, WWGF05, Man76a, Man08, Sto79, Bre79, Dys78, Goo78, How78, Ios79, Kac78, Kil78, Kir79]. dimensional [AvHP^{+}97, GMAK84, Has07, OdlCA02, VJ06, Zha09]. dimensionality [GMMA83, Man84c, Man84e]. Dimensions [Man91d, MKA02, Man03d, AO01, Bur81, CCGS91, DSB96, Gué01, Mac08, Man86a, Man86c, Man86d, Man90a, Man95a, MF09]. d’information [Man51a, Man52c]. Diophantine [BBM79]. direct [Man67a, VM95]. Directed [MV89]. dirt [Zaa10]. discontinuité [Man73c]. Discontinuity [Ano98a, MG98, Mir01, Man97b]. Discontinuous [MR97, CJK11, RM98]. discours [Man67c]. discourse [Man61i, Man70a]. discover [Had88]. Discovering [For04]. discrete [FSR94, Man78b, Man79a, Man85c, Ska99, ZY94]. discrets [Man78b]. discs [Man82d, Man83e, Man85f]. Discussion [Dev92, Fis92, Man79b, MB89, ZPSJ90, Man70a]. Disequilibrium [FSD89]. Disorder [Man06c, Man06b, TKR^{+}97, VM95]. Disordered [Pen88, HWM87, WSM88, dVdRL93]. dispersion [Man76a, Man77a, Man77c]. Dispersional [VBM98]. display [FP89, Nor82]. displays [Fin89]. dissipation [CDP07, Man72e]. dissolution [DL87]. distance [BL93, BCM91, BCM92]. Distinction [Man81b, Man03c, Man07]. distinguishing [TE92]. distortion [GKL03, WM94]. distributed [JT96a].

Distribution [BATL88, FMS96, Hou95, Man63d, MT67, Sim55, BZ96, MML01, Man56b, Man59d, Man60f, Man61h, Man72e, Man88a, ME90, MLDW85, Orl76, Sim60, TKR^{+}97]. Distributions [Fai69, Fai05, Man62b, Man74c, Man03d, Mit03, DrKP10, Izs06a, Man74a, Man74b, Man09c, Mas05a, Rac03, You13]. distributive [Gin02]. divergence [Man74d]. divisibilité [Man59c]. divisibility [Man59c]. division [Ald07, Per96]. DLA [Man90b, Man92a, Man92b, YMK94]. DM [Ano98a, Ano99]. DNA [Fin89]. do [MB89]. Documentation [Fai69, Fai05]. Does [MB89]. domain [Sau96b, Man85f]. Domain-filling [Man85f]. domains [AST07]. Dopant [EKKR06]. Douady [RPÁM04]. double [AvHP^{+}97]. down [Man99a]. Dr. [Man61g, Sim61a]. Dragon [WGKS12]. Drawing [RAA^{+}08]. drip [JSM06, TMJ99, TMJ06]. driven [GTL11]. Drivers [MH94]. Drum [Fla91, Ste88]. Dual [BR74a]. Due [MB94, Man61h]. Dugas [Man61d]. d’un [Man56a]. d’une [Gui87, Man52c]. d’univers [Man75e, Man79a]. duration [Man52c]. durée [Man52c]. Dürer [Jon90a]. during [HTHE05, JV98]. dusts [BM97]. Dutch [Man61f]. Dvoretzky [Man72d]. Dynamic [Dav90, Man93c, VM95]. Dynamical [Dev88, NYO94, ADdM^{+}12, DdMdO^{+}99, MC02, PR84a, PR84b, Sir97]. Dynamically [SO93]. Dynamics [BD86, BPW09, GOY87, Orb86, Bai97, DK06, EHH05, GV89, dLKHS06, Man84h, Man85a, Man85b, Man85d, Man85e, Man85f, Man85c, M^{+}85, Sch11, WLWL22, XX07]. dynamischer [PR84a, PR84b].

Each [Man84d]. Early [WT10, Man86a]. earth [Man02d, Man75j, WT10]. earthquake [Has07]. Earthquakes [Hou95, MB94]. Easy [CM99]. echoes [VM98]. Eco [Fin89]. ecological [WY92]. Econometrica [Man73d]. Economic [Mir90]. Economics [Man63b, Man63c, Man09c, Man64a]. ectasia [KNBZ12]. ed [Man61c, Man62a]. edge [Hol18]. Edited [Ano03a]. edition [Man60e, Gar91, Har90]. Editorial [CMBL77]. Editors [Man63a]. editorship [Man60d]. eds [Man61b, Man61e, Pic07]. education [FM02b, MF02, Kal03, Ano03a]. Edward [Dev92]. Effect [EKKR06, FBVdlN99]. effective [GJV17, JH97]. effectiveness [MF02]. Effects [Mil57, NFT^{+}01]. Efficient [CCRR91, CD93a, HD91, Sau87, Roj91]. Efficiently [Man71c]. eigendimensional [MGAP85]. eigenfunction [Man82b]. eigenmode [SL97]. eigenvalue [BV11]. eightieth [Shl05]. Einführung [Kau92]. Einstein [GM83, Hol18, Mas05a]. EKG [DR99]. electrochemical [Fle97]. electrodeposits [BB84]. electronic [IKP01]. electrostatic [Man91a]. element [ZWZL06]. Elevation [DKW94b, BM93]. Ellipsoids [Bou85]. elliptic [BGR94, BB79]. embed [Esp97]. embeddings [GKL03]. Embryological [Mai90]. Empirical [Fai69, Fai05, DrKP10, Man62e, Man85a]. empiriques [Man62e]. emptiness [MF09]. Encoded [MPM02]. Encoding [Fis95, PJS92c, COK96, TQ10]. Encounters [Gul12]. encryption [JT96b]. end [FHD09]. end-to-end [FHD09]. énergie [Man56a, Man59c]. Energy [CDP07, BA78, HKHP11, Man56a, Man59c, Man72e, VM95, dSVTM99]. energy-minimizing [HKHP11]. engineer [Man57a]. English [DM73]. enhance [SSK93]. enhanced [HMS96b]. Enhancing [BE95a, HMS96a]. Ensembles [KM65, Man56d, Man59c, Man79a]. entre [BBB^{+}10]. Entropid [Man61a]. entropies [MV02]. entropy [KPP19, MV02, Man61a]. Envelope [MdF05]. environmental [Bur81]. Ephremides [Man74e]. épistémologie [Man67f]. epistemology [Man64b]. equal [Man85d]. equation [AST07, BC80, Liu95b, Man01c, Gui87]. equations [AGP81, BR74b, BB79, CT91, DL92, DT91, DM92, Liu95a, Mac08, RB79]. équilibre [Man56a]. equilibrium [Fle97, Man56a, PM94, dSVTM99]. Equipment [Gar88b]. equivalence [Man84c, Mor96, Mor96]. Equivalences [Kel00b, AO97]. equivalent [AO01, MdF05]. equivariant [CCC00]. Ergodic [Jor12]. Eroded [MKM89]. Erratic [Man70c]. Erratum [EMW92a]. erreurs [Man55b]. Error [BM63, Man71b, Man72c, LKH97, Man65c]. errors [Man55b]. Escape [Bur02, NG03]. Escher [Roo10]. especially [Man89e]. essay [Man93e]. Essays [Man61b, Pic07, AF90, Ano91]. Essex [Rai03]. Estimated [VBM98]. estimates [Set92]. estimating [GBR^{+}09, GFB^{+}10]. Estimation [BATL88, GMV09, Man62c, FpDS99, Izs06a, Man56a, RVA00, TPVG06, WWGF05]. estimator [KK90]. Estonia [KK93]. Études [CB59, Man54a, Man54c]. Euclid [Cam00]. Euclidean [GV89]. European [BDLS96, DLM99]. EuroPVM [BDLS96]. eutrophication [JV98]. Evaluating [Bou09b]. evaluation [Hür95]. Evaporate [Man89f]. event [GTL11]. event-driven [GTL11]. events [Man88b, WGKS12]. everything [Shl05]. Everywhere [Bar88a, Bar93b]. evidence [Man62e]. EvoBIO [Rai03]. EvoCOP [Rai03]. EvoIASP [Rai03]. evolution [ODA03]. Evolutionary [MVW^{+}03, SR96, Rai03]. Evolving [NG95]. EvoMUSART [Rai03]. EvoROB [Rai03]. EvoSTIM [Rai03]. EvoWorkshops [Rai03]. ex [MM89]. exact [Bra94, Liu95a]. Exactly [MEH90]. Examination [Man89f]. examines [Ano91]. Example [Man03d]. examples [Bau96, Kau92, Man67d]. Exceptions [RM98]. excimer [KNBZ12]. excursions [Hol18]. Exhaustivité [Man56a]. Exhaustivity [Man56a]. Exhibition [Fri04, PR84b]. expansion [Man82b]. Expansions [Man96b, JM96b]. Expeditionary [Gar88b]. experience [JeJP58, JeJP58]. experiment [IKP01]. Experimental [dLKHS04, Bin24]. experiments [MW69b]. Explained [Man03d]. explanation [TW11, dVSC96]. Explicit [Man74c, Man83a]. exploration [Bou09b]. explorations [Fla98]. Explorer [Sch90]. Exploring [Cla10b, Dew85b, For04, Cha11]. Exponent [Man63d, DR99, Man76b]. exponential [MML01]. exponents [WC06]. extended [LMP85, MP84, ME91]. extension [JH97]. Extensions [Liu93, Man74b]. extensive [EG94]. External [RPÁM04, RPAM06, RAA^{+}08]. Extraits [CHR54]. extreme [GMV09, WGKS12].

F [ABC^{+}94, Goo91, Lan89, Man70a]. fabric [CP98]. face [M^{+}85]. facets [Man78a]. Facial [TQ10]. Fact [Man89f]. Factors [Kel00b]. faits [Man62e]. Faloutsos [Win05]. Fame [Ano03a]. Families [DAB99]. family [Dha05, Eng93, GAMK81]. Farey [Dev99, DM92]. fashions [Mad86]. Fassberg [PR84b, PR84b]. Fast [COK96, HS93, KH92, KPY96, WWGF05, Man71a, RBM96, SH96, YJ09]. Faster [MFCM96]. Father [Gil04, Ano10d]. faulting [DSS90]. Feat [Ano06]. features [FSD89]. Feinstein [Man55b]. ferromagnetic [BBM79]. festschrift [EPV96]. Fever [Ano98b]. few [Man84g]. fiber [PM94]. Fibonacci [Dev99]. Field [Phi92]. fields [VJ06]. Fifth [ACM03, ALW^{+}88]. figures [Man67c, Mor96]. filling [Man85f]. Film [Cla10a, EKKR06, SSK93]. Filming [LG10c]. Filter [BHI11]. Final [Man61h, Man61j, Sim61b]. Finance [MG98, Man01d, Mir01, Ano98a, DrKP10, Man97b, Man99c, Man99d, Man01a, Man02a, Man03a, Man04b, MH05a, Man05a, Man05b, MH08a, Man09a, Rac03]. Financial [Man10a, DDH10, Man01b, Man01c, Man01d]. Finanzen [Ano05, MH05a, MH08a]. fingering [VOL^{+}86]. Fingers [Rob85, NDS85, San85]. Finitary [MC63]. finite [LO96, Man73d, ZWZL06, Man60b]. Fire [Man64d]. First [Man61e]. Fisher [Goo91, Lan89]. FISTE [FHD09]. Fitness [MVW^{+}03]. fits [Sla07]. fitting [Izs06b]. five [ALW^{+}88, DK06]. fixed [EPV95, Man99c]. fixed-scale [EPV95]. fjords [EJM91]. FL [GMA^{+}98]. Flake [Lev05]. flicker [Vos79]. flights [CS87, TG09]. Floating [SO93]. Flock [Ano87, Dew87]. Flow [Sah93, CDP07, CMP07, GMV09, Man84b, SCC^{+}07]. flows [DdCVR03, Ghi86, MdF05]. Fluctuation [Man89g]. Fluctuations [Man70c, Man10a, Gar78]. Fluid [Bai97, SO93, Ghi86, Man76a, NDS85]. fluides [Man76a]. Fluids [Rob85, Man84b]. fMRI [HKS^{+}05]. FOCS [IEE03]. focusing [MT10]. fonction [DT89]. fonctionnelle [Gui87]. Fonctions [Man75b, Jul18]. Footprints [Fri04]. forced [CDP07]. forecasting [SSK93]. Forecasts [Man66a]. forest [Edg00]. Foreword [Man88b]. Form [Bre79, Dys78, Goo78, How78, Ios79, Kac78, Kil78, Kir79, MA79, Sto79, ZT96, EMW92a, EMW92b, Man75d, Man77b, Man84f, Man89d, Man98c, Man10c, VJ06]. Formal [Smi84, Man54a, Man54c]. Formalism [Fra07, Jaf97a, Jaf97b, Man91d, Man95c, RM95, RM98]. formas [FR98]. Formation [TM86, GTL11]. Forme [Man75d, Man84f, Man89d, Man98c, Man10c]. formelle [CHR54, Man54a, Man54c]. Formes [Man73a]. Forms [Man06c, GH04, Man73a, Man06b]. Formula [Man97a, MR97, RM97]. fossils [Wal96]. Foundations [IEE03]. four [DSB96, Man75c]. Fourier [Man60e]. Fourth [WBE99]. Fracas [Poo90]. FracMAP [CGC^{+}09]. Fractal [Ald07, Ano98b, Ano06, Ano10a, AAM^{+}03, ABLM98, Bal10, Bar84, Bar88b, BEHM89, Bar93a, BB10, BL80, Ber86, Bia89, Bou85, Boy83, BB84, Bri85, Bur84, Bur81, Cah89, Cam96, Cam00, Can84, CR94, Cha84, CS87, Chu11, Chu12, CGMST95, Cla86, Cla10b, Cor90, DL87, DHN85, Dev88, Dev96, Dur78, EKKR06, EHH05, Esp97, Fer83, Fin89, Fis92, Fis95, FP10, Fla91, Fro97, Gib93, Goo84a, Goo84b, Goo91, GDC^{+}10, GOY87, H.88, Hal96, HRO02, Hel92, HKE01, HKS^{+}05, HTHE05, HWEH05, Hig11, Hir94, Hou95, JSM06, Jon94, KMNW99, KFN00, Kop88, KM89b, La 86, LVEB09, Lan89, LL97, LR85, LM85, LSL86, Man75d, Man79b, Man80, Man82a, Man83b, MPP84, Man84f, Man86a, Man88c, MB89, Man89d, Man91a, Man91b, Man98b, MPB^{+}98, Man98c, MF99, MKA02]. Fractal [Man03c, Man06b, Man06c, MH08b, Man10a, Man10c, Mar99, MB94, MFCM96, MC87, MM10, MLDW85, MKM89, NWM^{+}06, NDS85, Orb86, PS88a, Pic86, Poo90, Ric94, Rob85, Roo10, San86, San87a, San87b, Sla07, SO93, Spa84, SC88, Ste00a, Ste10, TMJ99, TMJ06, Tem96, Tsu08, VCM^{+}11, VOL^{+}86, Van90, VBM98, VC86, VMH^{+}83, WBE99, Whe83, Wil97, WT10, Wu95, ZT96, AST07, AG06, AO97, AO01, Ano10d, ADdM^{+}12, AFP84, Bai97, BE95a, BE95b, BDM^{+}88, Bis95, BZ96, BM97, BF95, Bou09a, Bou06, BM93, BRRK94, BBB^{+}03, Cam87, Cas96, COK96, CK96, CGH02, CGC^{+}09, CZ98, CDP07, CMP07, CJK11, CP98, CD93a, Dav96, DSS90, DR99, Dek12, DdCVR03, DV93, DT91, Dro10, DdMdO^{+}99, Dub89, DT89, Edg00, EG94, EPV95, EW96, Fal85, FBVdlN99]. fractal [FR98, Fer95, FSR94, Fle97, FpDS99, FV95, FP89, FSD89, FC00, Gai06, Gar78, GMA80, GAMK81, GMMA83, GAH93, Ghi86, Gil86, GK03, GM83, GM84a, Gof91, GV89, GTL11, Gon98, GBR^{+}09, GFB^{+}10, GH98, Gué01, Ham95, HMS96a, HMS96b, Hug98, Hür93, HS93, Hür95, Imr06, JT96a, JM96a, Jaf98, Jak84, JTEA91, Jor12, KNBZ12, KCC89, Kel00c, KM11, KKL95, Kir83, Kom95, KVŽŽ08, KPY96, Lan97, Lap08, LKH97, LPZ08, LV96, Liu93, Liu95a, Liu95b, LLP96, Loy91, LO96, MSW08, MJ86, Mac08, Mad86, Mad87, Mai90, Man75i, Man75j, Man76b, Man78b, Man78c, Man82e, Man82c, Man83a, Man83c, Man83d, Man84b, Man84d, Man84g, Man84e, MG84, Man84i, Man85b, Man85d, Man85f, Man85g, Man86c, Man86d, Man86e, Man88b, MV89]. fractal [Man90a, ME90, Man90c, ME91, Man92a, MS94, Man95d, Man95e, Man97c, Man00, Man01a, Man02a, MH04a, MH04b, Man04b, MH05c, Man05a, MH05b, Man05b, MH09, Mei92, MC02, Mil88, MM91, NFT^{+}01, NYO94, Nar07, NIS^{+}03a, NIS^{+}03b, NM09, NCS04, Ohn90, OMK^{+}05, OdlCA02, ODA03, PU08, PC06, Pe04, PH91, PM94, Pic95, Pum96, Rar89, RBM96, RPS^{+}06, SSK93, San85, Sau87, SR96, Sau96a, SH96, Sau96b, Sch87, Set92, Sig96, Sir97, Ska99, SH92, Slo94, SCC^{+}07, SS00, SS91, Ste88, Ste00b, Suz91, TQ10, TG09, TM86, Thi88, Thi89, TKR^{+}97, TdRGD09, TTMB96, TE92, TSS85, UP07, VJ06, Van09, Van10, Vla94, VM95, Vor02, Wai88, Wal96, WY92, WY96, WCH03, WC06, WW08, WLWL22, WWGF05, Woo94, WM94, Wu99, XX07, Ye02]. fractal [YHHS93, YJ09, Yue10, Zaa10, ZZZ^{+}96, ZWZL06, dACSS12, dVSC96, dVdRL93, Esp97, EPV96, HWM87, LMP85, LvF04a, LvF04b, Man76a, Man78d, MP84, WSM88]. fractal-based [EW96, GAH93, LPZ08, PH91, Pum96]. Fractal-Dimension [Pic86]. fractal-forced [CDP07]. fractal-imaging [Esp97]. fractal-like [CGC^{+}09, CD93a, Wai88]. fractal-shaped [DT91]. fractal/wavelet [CK96]. fractale [Man76a, Man82c, MH05c, MH09]. Fractales [Man09a, Cha90, DT89]. Fractalist [Tay13, Man09b]. fractality [KRP^{+}91]. fractally [SL97]. Fractals [Abb95, Adl12, Aha86, BD86, Bar88a, Bar93b, BSV02, BP17, Caw89, FdL10, FM02b, Gro92, JPS90, LS92, Man77b, Man77a, Man77c, Man79b, MA79, Man81a, Man84a, Man84e, MGAP85, Man86b, Man89c, Man89a, Man93b, Man97b, MG98, MF03, Man04a, Man09a, Mea89, Nat91, NR95, Oli91, PR86, PJS91, PJS92a, PJS92b, PJS92d, PJS92e, PJS93, PJS99, PT86, Pop22, Sch89, SM89b, SMR84, Smi84, ST89, Tau98, Tay85, Vos88, Wei05, ZPSJ90, AGMK81, ADR01, Ano90, Ano91, Ano10b, BBB^{+}10, Bau96, Bou09b, Bro07, CM91, Car96, Car99a, Car99b, Cha90, CB97, DL92, Dav95, Dav90, Dha05, DHKT95, Dre01, Ent89, Fla98, FM02a, Fra98, GPS92, Gar88a, GAM83, GASM84, GAM84, GH04, Gil04, GBCK93, Gul12, GKL03, HSK89, HD91, HCF95, HJ94, Kap86]. fractals [LK98, LG10b, Ley05, Maj98, Man75d, Man78b, Man78a, Man78d, Man83e, Man84b, Man84f, Man85g, Man89d, Man93a, Man98c, Man01e, Man02d, MF02, Man10c, Mar04, Mar10, Mur93, NG95, NG03, Nor82, Nov04, PH90, PJS04a, Pic94, Sah93, SBvB^{+}12, SS89, Shi93, SCNT03, SAM93, Vos86, LMNW05, Win05, YMK94, ZY94, Zha09, Abb95, AF90, MH05a, MH08a, Man08, Nat91, PT86, SMR84, Sto79, Ano03b, Abb95, Ano98a, Ano03a, Bre79, Dev92, Goo78, How78, Ios79, Kac78, Kal03, Kil78, Kir79, Mir01, Nat91, Sch05]. Fractals. [Dys78]. Fractional [BM09, Man67b, MV68, Taq13, CGM95, CGM96, MW69b, Man71a, Man72a, Man75a, Man76c, Man77d, Man82b, Man95b, Man96a, WLWL22]. Fractions [Rat93]. fracton [Man84c]. fracton/spectral [Man84c]. Fractons [MJ86]. Fracture [CMR06, Man06c, Cah89, MPP84, Man06b, TM86]. Fragmentation [GV89, SS91]. fraktale [Kur90, Man91b, Voj89, Ano05, MH05a, MH08a, Man08]. Frame [Kal03]. France [AF90, BATL88]. Francisco [Bri85, Kir83, Sto79, Whe83]. Franken [Man61a]. Franks [Hir89]. frauds [Wal96]. Fredholm [RB79]. Free [ZT96, Man78b, WCH03]. Freeman [Bri85, Kir79, Kir83, Sto79, Whe83]. French [Ano10e, BBB^{+}10, Cha90, Dev57, DT89, Fat17b, Fat17a, JeJP58, Jul18, KM65, Lef82, Man51a, Man51b, Man52a, Man52b, Man52c, Man53a, Man54a, Man54c, Man55b, Man56b, Man56a, Man57c, Man59c, Man59b, Man59e, Man59f, Man60g, Man61d, Man62e, Man65d, Man66d, Man67c, Man73a, Man73b, Man73c, Man74a, Man74b, Man75b, Man75c, Man75e, Man75d, Man76a, Man78b, Man78d, Man79a, Man82c, Man84f, Man89d, Man98c, Man01a, MH05c, MH09, Man09a, Man10c, Mor96]. frequencies [Man61i]. Fresnel [KVŽŽ08]. Frontiers [M^{+}85, PJS92a, PJS92b, PJS93, PJS04a]. fronts [VCM^{+}11]. Function [AB85, BHI11, BL80, DHN85, Man61a, Man96b, BV11, BZ96, Bra94, JM96b, Lan97, Van09, WY96, XX07, Yue10]. Functional [Rat93, Tsu08, SW92, VM95]. Functions [Bar71, BR78, BEHM89, DAB99, GT07, Jaf97a, Jaf97b, MVW^{+}03, Man60e, Man61k, Sim55, Vrs86, BA78, BM04b, BM04a, Bou09a, CM99, DT89, Fra98, HKHP11, Jaf98, Jul18, Man59d, Man61f, Man61h, Man67d, Man75b, Man88a, Sim60, Man61f].

Funktionalraümen [Man61a]. further [Sim60, Val11]. futility [Sau96a]. Future [Man66a, Win05]. fuzzy [CM91].

G [Man60b, Man61b]. gagner [MH05c, MH09]. Gaithersburg [SMR84]. galactic [Man75e]. galactiques [Man75e]. galaxy [Man77d, MC87]. Galerkin [CJK11]. Game [Jon90a, PJS04b]. Games [Sch90, Gar78, Man53a, Man54b]. Gap [HASM96]. Gaps [MKA02, Man73c, Thi88]. Gary [Lev05]. gas [Mas06]. gaseous [SW92]. Gasket’ [NWM^{+}06]. Gaskets [Jon90a, GASM84]. Gaussian [Man62e, MW69b, Man69, Man71a, Man76c, Man02d, Man09c, Set92, Ano03b]. gaussiens [Man62e]. GBP [Ano98a, Ano99]. Gebiete [Man75f]. gender [EHH05]. General [CCC00, MDM02, Suz91, Woo94, BZ96, Man52c, Man60d, Mas05a, Wai88]. générale [CHR54, Man52c]. généralisations [Man75b]. generalization [MV02, Orl70, Vla94]. generalizations [Man75b]. Generalized [DAB99, KVŽŽ08, Lew87, Mei92, SVVRGA09, AK13, Bou09a, FV95, FR92, Gué01, Orl76, TKR^{+}97, WC06, yWjCnG07, XX07, XRZ09]. generalized-affine [Bou09a]. generate [Ye02]. generated [AvHP^{+}97, JTEA91, Shi93]. Generating [DHKT95, Miy90, VM95, Wai88, Bra94, SS89]. Generation [DSB96, NG03, Nor82, Cas96, CM99, CD93b, CK93, DKW94a, Dre01, KH92, KGN^{+}24, Pic94, VJ06, YHHS93, ZWZL06]. Generative [Mit03]. generator [Man71a, Man01d]. Generic [Fra07]. Genetic [RBM96]. génétique [CB59]. genomic [Fin89]. genres [Man56b]. Gentle [Mad86]. geometer [Man09b]. geometría [FR98, Man97c]. Geometric [GMMA83, LO96, Man78a, GBCK93, Nor82]. Geometrical [MS78, Sig96, MS78]. Geometrie [Kur90, Man91b, Voj89, Man82c, Man01a, Man76a]. geometries [GKL03]. Geometry [ABLM98, Bar84, Bia89, Boy83, Bri85, Bur84, Cam96, Cam00, Can84, Cha84, Fer83, Goo84a, Goo84b, Hal96, Man83b, MB89, Man91b, Man97c, Man10b, Spa84, Ste10, Wei05, WBE99, Whe83, Wil97, Ano10d, AAM^{+}03, Dev96, EPV96, Fal85, Fer95, HRO02, Kap86, KCC89, Kir83, KM89b, LvF04a, LvF04b, Man75i, Man76a, Man78c, Man81a, Man82e, Man82c, Man84g, Man00, Man01a, Man04b, MM10, Mil88, Tem96, LMNW05]. geophysical [LSL86, MW69a]. geophysicist [Man89e]. geophysics [SM89b]. German [Man60e, Man61a, Kau92, Man57a, Man91b, MH05a, MH08a, Man08, PR84a, PR84b]. Germany [BDLS96]. Gets [Wil97]. Ghurye [Man61b]. Gibbs [KM89a, Man59c, Mas05a]. given [ODA03, Dha05]. glasses [MSW08]. Gleick [Hir89]. Glimpses [BF12a]. Global [Man03c, HTHE05]. Globality [Man02d]. Gnedenko [Man62a]. Gödel [Hol18]. Gordon [Lev05]. Göttingen [PR84b, PR84b]. gram [Cha11]. grammars [DHKT95]. Grammatical [ODA03]. grand [Man59c]. grand-canoniques [Man59c]. Graphic [Hug98, Roj91]. graphical [DSB96]. Graphics [GH04, Gin02, Kal03, OdlCA02, Pic87, Ano03a, CB97, FM02b, M^{+}85, Mar10]. Graphing [For04]. Graphs [Pic86, UP07]. Gravitation [AG06]. gravity [dVSC96]. gray [WW08]. gray-level [WW08]. Greek [Est11]. Grenzen [PR84b]. Group [DLM99, CCC00, Vla94]. groups [Man82d, Man83e]. growing [ME90]. grows [Man92a]. Growth [San87a, San87b, Bra94, BB84, BRRK94, EPV95, Fle97, MV89, NDS85, San85, San86]. growth-decay [Bra94]. Guide [Ste00a, Ste00b]. Gustav [Man61f].

H [Abb95, Dev92, Gar91, Har90, Kir79, Kir83, Man60e, Nat91, Sto79, Man02d]. H. [Man59d, Man61h, Man65d]. H.-O [Dev92]. hadron [Ohn90]. half [Cha90]. half-century [Cha90]. Hamiltonian [GM88, NM09]. Handbook [BGL67, LBG63, Rac03]. Hard [Caw89]. hardback [Kir83, Pic07]. Harmonic [BB99, EM92, Man60e, EJM91, EMW92a, EMW92b, ME91, Man92b]. harmonics [Cla86, CM99]. Harmonie [PR84a]. Harmony [PR84a]. Harnessing [BJM^{+}88]. Harold [Man61b]. hasard [Man67f, Man73a, Man75d, Man78b, Man84f, Man89d, Man98c, Man09a, Man10c]. Hasards [Man75c]. Hashing [HCF95]. Hassan [Ano93a, Ano93c]. Hausdorff [BL93, Gar88a, Man76a, Man86e, Shi98]. Hazes [WT10]. Health [Pic07]. Heart [Bia89, VBM98, NIS^{+}03a, NIS^{+}03b, PM94]. heat [Set92]. Heavy [DrKP10, Man03a, Rac03]. Heinz [Goo91, Lan89]. Heinz-Otto [Goo91, Lan89]. held [CR83, CMR06, LvF04a, LvF04b, MS78, SMR84]. Helmholtz [AST07]. hemorrhage [SCC^{+}07]. Henry [Man61b]. Herdan [Man61f, Man61g]. Heterogeneous [LL97]. Heuristic [BF95]. Hexagonal [NWM^{+}06]. HGA [Sch90]. Hidden [BEHM89, Pop22, Man08]. hierarchical [HS93, Man75e, Man79a, WGKS12]. hiérarchisé [Man75e, Man79a]. High [Zaa10, Man74d, SL97]. high-order [SL97]. High-temperature [Zaa10]. Higher [Vrs86]. Hilbert [BB76, McK19]. History [Mit03, Mur93]. Hoeffding [Man61b]. Hogg [Man60d]. Hölder [Man95a]. Hölders [Man95a]. homage [Ano10e]. home [HRO02]. hommage [Ano10e]. homogeneous [Man75i, RPS^{+}06]. homomorphisms [GDH96]. homothetic [Man65d]. homothétiques [Man65d]. Honor [Ano95, Man61b]. honoring [DrKP10]. honour [AF90, CMR06]. honouring [AF90]. Horton [YM94]. Host [ME72]. Hotelling [Man61b]. Hubbard [Gar91, Har90]. Hudson [Gru05, Ano05]. Hugh [Ano90]. hull [DM92]. human [EW96]. humans [EHH05]. Hungary [JV98]. Hurst [DR99, Man65d]. Hutchinson [BCM92, BCM91]. Hybrid [ZZZ^{+}96]. Hydrodynamic [NCS04]. Hydrology [SS00, MW68]. Hype [Hir89]. Hyperbolic [Fai69, Fai05, GF95, HJ94]. hypercubic [GMMA83]. hypotension [HTHE05]. hypotheses [Man67f]. Hypothesis [Fam63, Man72e].

Ian [Lev05]. IBM [Pic94]. ICF11 [CMR06, CMR06]. ICIP’95 [IEE95]. ICTP [PT86]. Idea [Wol16]. ideal [Mas06]. ideas [Kir11, Wol16]. identification [TdRGD09]. IEEE [IEE03]. IFS [CB97, Mar04, Mar10]. IFSs [Hor90b]. ihre [Man57a]. II [BGL67, LBG63, BM04c, BM04a, DL92, GASM84, Jaf97b, Man51b, Man86d, Man01c]. III [Man61a, GAM84, Man85a, Man86e]. Illustration [RHD^{+}06]. im [M^{+}85, PR84b]. Image [BB90, BJM^{+}88, Fis92, Fis95, IEE95, Liu95a, MS78, Par89, ZY94, BE95a, BE95b, COK96, CK96, CZ98, CP98, CK93, Dav96, EW96, GAH93, Ham95, HMS96a, HMS96b, HKHP11, Hor90b, Hür93, JT96a, JM96a, JT96b, KM11, KKL95, Kom95, KPY96, LKH97, LPZ08, LV96, LLP96, PH91, RBM96, SR96, Sau96a, SH96, Sau96b, Sig96, Slo94, TQ10, TTMB96, UH96, WY96, WW08, WM94, Wu99, YHHS93, ZZZ^{+}96]. Imagery [VEJA89, JTEA91, Pum96]. Images [BS88, BHI11, LVEB09, PS88a, PJS92c, Bar88b, BDM^{+}88, Bar93a, BL93, CD93b, FP89, GBR^{+}09, GFB^{+}10, HS93, Imr06, Loy91, Man88b, Man01a, OdlCA02, WCH03, Woo94, Ye02, Goo91, Lan89]. Imaging [Hig11, Esp97]. impairment [OMK^{+}05]. Imperfection [BB90]. Imperial [Man02c]. implementation [FpDS99, GMMA83]. implementations [JT96a]. implications [KFN00, Mai90, Man86e]. Implicit [Dav96, HKHP11]. Impossible [Bro07]. improved [WW08]. IMS [DK06]. Include [Man10b, Man85e]. Income [Man61k, Man62b, Man63d, Man60f]. incomes [Man59f]. incompressible [Ghi86]. increasingly [Man92a]. Incremental [CT91]. increments [Man03a]. Independence [HASM96, Man60c]. independent [BM04b, BM04a, Man03a, NM09]. index [CL87, FM02a]. India [VBMS96]. indices [Man61f]. individual [Man85a]. inequality [Yue10]. inescapable [Man05a]. infinie [Man59c, Man73c]. Infinite [Man95c, YHHS93, BM04b, BM04a, DL92, FC00, Man59c, Man69, Man73c, Man86a, RM95]. Infinite-corner-point [YHHS93]. Infinitely [GAM84]. Infinity [Cla10a, LG10c, LG10b, Lev05]. Inflationary [Lin94]. Influence [BF12b, EHH05]. Inform [BCM92]. Information [Man59b, Man65a, Man66b, Man66c, AMM57, AMM74, Baw11, Man51a, Man52c, Man57c, Man59b, Man61c]. Informational [Man53b]. Informationstheorie [Man61a]. Informazione [Man60h]. Ingenieur [Man57a]. Ingram [Man61b]. inhomogeneous [LSL86]. insistence [Man57b]. inspection [CP98]. instability [NDS85]. instantaneous [Man66d]. instantanés [Man66d]. Institut [PR84b]. Institute [PR84b, BBBH92]. integrable [Mac08, Zha09]. integral [BR74b, RB79, WWGF05]. Integrals [Man60e, KVŽŽ08, Suz91, Man60e]. intelligent [MC02]. Interaction [VMH^{+}83]. interactions [BBB^{+}10]. Interactive [HKHP11, CGC^{+}09, Hor90b, LK98]. interest [Man70b]. interface [DLM99]. intermiss [Thi88]. Intermittent [Man74d, Man76b, Mil57, Man72e]. intern [CB59]. internal [Hoo91]. International [CR83, CMR06, IEE95, KK93, PT86, AF90, VBMS96]. Internet [Tau98, TG09]. Interoccurrence [Has07]. Interpolation [BEHM89, DKW94b, Bou09a, Man85c, Pum96]. Interpretation [Man59b, SVVRGA09, LO96, Man51b]. Interprétations [Man51b].

Interscience [GM61]. Intersection [Bou85]. interstellar [dVSC96]. interval [Per96]. Interview [Dav84]. Interviews [AA85, AA08]. intrinsèque [Man52c]. intrinsic [Man52c, MJ97]. Introduction [Abb95, BM04b, BM04a, Man94, Man95d, Nat91, PJS92d, SM89a, Bou06, Kau92, Man88a, Man60d]. intuitive [Man85f]. invariance [Man67f, MC87]. Invariant [GM88, Kel00b, Man74a, Man74b, Man93c, San85]. invariantes [Man74a, Man74b]. Inverse [MR97, CM91, HCF95, Man83e]. Inversion [Man97a, MR97, RM97, FC00, Ley05, Man83e]. inversions [Man82d]. investigation [BF95, Shi93]. Invitation [SL11]. IP [KLPS06, VHA05]. irregular [Nar07, UP07]. Irregularity [Dys78]. IRS [KK93]. Isaac [Fri04]. ISBN [Fri04, Kur90, Pic07, Tay13, Voj89]. Ising [BBM79]. islands [Man75j, SBvB^{+}12]. iso [Man75i]. iso-surfaces [Man75i]. isometries [Sau96a]. Isothermal [DdCVR03]. issue [DrKP10]. Italy [CMR06, PT86]. Iterated [BHI11, DHN85, BV11, Man74a, Man74b, Man84h, Man85a, Man85b, Man85d, Man85e, Man85f, Man85c, Van09, WY96, Esp97]. Iteration [DAB99, MH94, Rat93, Vrs86, Jul18, Man80, SL97, Van10, WCH03]. iteration-free [WCH03]. iterative [KKL95, KGN^{+}24, Mor96]. iterées [Mor96, Man74a, Man74b]. IV [Man61f, Man85b]. IX [Man85c].

J [CB59, Man60b, Man61e, WC06]. Jaffe [ABC^{+}94]. James [Hir89]. Janua [Man61f]. January [LvF04a, LvF04b]. Java [Hug98]. jeux [Man53a]. John [Gar91, Har90, Hir89, Man60c, Man74e, Man60a, Man60b, Yue10]. Joint [DK06]. Jonathan [Sel10]. Jonckheere [CB59]. Joseph [Man63a, MW68, Man73b]. joue [Ste98]. joue-t-il [Ste98]. JPEG [EW96]. jubilee [LvF04a, LvF04b]. Julia [Ano87, ARC14, CCC00, Dew87, DSB96, DAB99, DMT03, Eng93, Ent89, GH04, Gin02, Kel00b, KGN^{+}24, Man85e, Man85f, OdlCA02, PSvH84, Sau87, Shi98, Vrs86, yWjCnG07]. July [Fis95, GMA^{+}98, PT86, RW89]. June [ACM03, DK06, MS78, PR84b, RW89]. Juni [PR84b]. Jurgens [Dev92, Abb95, Nat91]. justification [Man59c].

Kac [Man60c]. Kapazität [Man61a]. Karlin [Man61e]. Kemeny [Man60b]. Kenner [Ano90, Ano91]. Kenneth [Man61e]. Keplerian [Dav95]. kernel [Set92]. key [Man75c]. kinetic [TM86]. Kinetics [Kop88]. Kings [WGKS12]. Kirkwood [BR74b]. Kit [Esp97]. knew [Man95f]. Knopoff [Has07]. known [DDH10]. Koch [PS88b]. Kolmogorov [MV02]. Kolmogorov-like [MV02]. komplexer [M^{+}85, PR84b]. Kosmos [PR84a]. Krook [BC80]. kurtosis [Man76b]. Kybernetik [Man57a].

L [Ano93a, Ano93c, Ano03a, Ano05, Are93, Boy83, CC58, Goo91, Gru05, Hal96, Kal03, Lan89, Swe91]. Lacunarity [HASM96, BM97, GMMA83, MS94, Man95e, Mas06]. Lagoon [BATL88]. Lake [JV98]. landscapes [Bur81, Hir94, Hug98, Man88c, Mil88, MM91]. langage [AMM57, AMM74, CC58, Lee59]. Language [JPS90, MC63, Cha11, Man56c]. Languages [Man53b, Smi84, Dre01, Fer95, Man54b, Man59b]. Langues [Man59b]. L’anneau [Man01a]. Laplacian [Set92]. Large [LS92, VBMS96, Vor02, YMK94]. Large-Scale [LS92]. Laser [KMNW99]. latent [Man89b, MF09]. lateral [Man95a]. lattice [EMW92a, EMW92b, Fin89, NCS04]. lattices [GMA80, GMMA83, GAM83, GAM84, GM83]. Laurie [Man60b]. Law [FMS96, Man03d, Mit03, SVVRGA09, ARMG03, BP84, Cha11, CL87, Dev57, Egg99, Gai06, MSW08, MRM^{+}05, Man57b, Man60f, Man60g, Man62d, Man09c, Mas05a, Mas05b, Mas06, Mon01, Orl70, Orl76, Per96, TW11, Lef82, MPM02, MVW^{+}03, Man57c]. laws [CJK11, Dro10, Man01d, Win05, Par89]. Layering [Kel00c]. leads [Mas05a]. leaf [VC86]. leaflets [PM94]. Lean [Sau96b]. Learning [KK90]. LeBaron [Man01d]. Lecture [Man77c, CB59, JeJP58]. lectures [CMR06, Man60e]. Leffler [BM97]. Lefschetz [Man63a]. Left [Man90b, MEH90]. left-sided [MEH90]. Lei [Mud02]. L’Emploi [Lef82]. length [EG94, Man81a, Man86d]. lengths [MLDW85]. lents [Man66d, Man73b]. Léo [Lee59, Man65b]. Lesmoir [Lev05]. Lesmoir-Gordon [Lev05]. Lett [BCM92]. letter [Man61g]. Letters [KM89a]. leur [Man59c]. level [WW08]. levels [BP84]. Lévy [BM97, CS87, Man59e, Man59f, Man60f, Man62e, Man95f, PC06, TG09]. l’expérience [CB59]. Library [Fri04, Esp97]. Life [SG00, Wei05, WBE99, Wil97, LMNW05, Bou09b]. ligne [Man51a, Man51b]. Like [DAB99, Vrs86, CGC^{+}09, CD93a, MV02, Wai88]. likelihood [Izs06a]. Limit [LR00, Man71c, Man75f, Man75h, Man75g, Per96, FC00, Man82d, Man83e]. limitation [Cha11]. Limitations [Man91d, DV93]. Limited [MKA02, EMW92a, EMW92b, KVMW95]. limiting [Man55a]. Limits [Roo10, Man67d, PR84b, You13]. line [Bar73, Man51a, Man51b, Man72a, Ran91]. linéaire [Gui87]. Linear [Man60a, Man71b, MKA02, AvHP^{+}97, BR77, BV11, HD91, HCF95, KGN^{+}24, LO96, Man84h, MC02, Van09]. linearity [Man69]. linearly [EM92]. lines [Man91a, Phi92]. l’information [CC58, Lee59]. L’ingénieur [CHR54]. Linguarum [Man61f]. Linguistics [Man61f, Man57a, Man57c, Man59a, Mon01]. Linguistik [Man57a]. Linguistique [Man57c]. linguistiques [CHR54]. Lion [Fri04]. literature [Ano90]. lives [Wol16]. LIW [CCRR91]. load [JH97]. loaded [TSS85]. Local [JM96b, Man96b, Man03c, HTHE05, DL92]. localization [dVdRL93]. locally [Man69]. Logique [AMM57, AMM74, CC58, Lee59]. logistic [IKP01]. Lognormal [Mit03, Man72e]. Loi [Lef82, Dev57, Man57c, Man60g, Man65d]. lois [Man67f]. London [Man61f]. Long [Man67b, Man69, MW69c, MW69a, MT79, Man91c, Man01d, Man09c, RVA00]. long-range [RVA00]. Long-run [Man69, MW69a, MT79]. look [Dew85a]. Looking [Bri88]. loop [Man78b]. loop-free [Man78b]. Lorenz [Dev92].

Lossless [SH96, LV96]. Louisa [Lev05]. Low [Slo94, GMMA83, GKL03]. Low-bit-rate [Slo94]. low-distortion [GKL03]. Lungs [Hel92]. LXII [CHR54]. Lyapunov [WC06].

M [GM61, Goo91, Lan89, Man60e, WC06]. M. [Roo10]. M.Tichomirow [Man61a]. Machina [MM89]. machine [BDLS96, DLM99]. Machol [Man61c]. Macintosh [Esp97, Kat86]. Macmillan [Man60d, Man63a]. macro [Man59a]. macro-linguistics [Man59a]. macroscopic [Man57c]. macroscopique [Man57c]. Madow [Man61b]. Magic [Lar90, Ano88]. magmatic [Cor90]. Mai [PR84b]. Maintenance [Lef82]. Maior [Man61f]. make [Fra98]. makers [Wol16]. makes [Esp97]. mammal [HRO02]. Manage [Adl12]. management [DDH10, FHD09]. Mandelbrot [Abb95, AF90, And11, Ano93a, Ano98a, Ano99, Ano03a, Ano03b, Ano05, Are93, Aus14, BATL88, Bri85, CHR54, CC58, CB59, DK06, EPV96, Gar91, Gom10, Goo91, Gru05, Hal96, Har90, Hir89, Kir79, Kir83, Kur90, Lan89, LvF04a, LvF04b, Lee59, Lef82, Lev05, MVW^{+}03, Mud02, Nat91, Pei10a, Sch90, Sch05, Swe91, Tay13, Voj89, LMNW05, Whe83, ARMG03, AK10, AK13, Ano87, Ano88, Ano89b, Ano93b, Ano93c, Ano95, Ano10b, Ano10c, Ano10a, Ano10d, Ano10e, ARC14, Ate93, AB85, Bak00, Bar84, Bar85, Bar08, BH89, BF12a, BF12b, BPW09, Baw11, BJ19, BL80, Bin24, BP84, Bou11, Boy83, Bre79, Bri88, Bur84, Bur02, Cam96, Cam00, Can84, Cha84, Cha11, CCD88, CCGS91, Chu11, Chu12, CCC00, CHR54, Con10, Dav84, DM90, Dev90]. Mandelbrot [Dev91, Dev96, Dev99, Dev04, Dev57, Dew85b, Dew87, Dew89, Dha05, DSB96, DAB99, DMT03, DrKP10, Dur78, Dys78, Egg99, Est11, ES92, Ewi95, FL91, Fai69, Fai05, Fam63, Far11, FR98, Fer83, For04, FR92, FPR92, GJV17, Gar88b, GF95, Gil04, Gin02, GMV09, Goo78, Goo84a, Goo84b, Gui87, Gün24, Her05, Hof10, Hoo91, Hor90a, How78, Ios79, IKP01, Izs06a, Izs06b, Jaf11, Jon90b, JV98, Kac78, KNBZ12, Kah74, KP76, Kal03, Kat86, Kel00a, Kel00b, KPP19, Kil78, Kir11, KM89a, dLKHS04, dLKHS06, KGN^{+}24, Lan97, Lef82, LG10a, LR00, Mag04, MPM02, Man04a, Man06a, Man12, Mas05a, Mas05b, Mas06, MM08, Mei92, Met94, Mir90, Mir01, Mon01, Orl70, Orl76, OdlCA02, PRÁM04, PR85, PJS92e, Pei10b, Per10, Phi92, PFR94, Pou86, Pro95]. Mandelbrot [Pro97, Rad96, Rat93, Roj91, RPÁM04, RPAM06, RAA^{+}08, Roo10, SBvB^{+}12, SVVRGA09, Sch05, Sch11, Sch86, Sel10, Sen90, Seu11, Seu12, Shi98, Shl05, Sim61a, Sim61b, Smi91, Spa84, Sto79, TW11, Taq13, Tay11, Ten10, Val11, Vrs86, VM98, yWjCnG07, WLWL22, Wei05, LMNW05, WGKS12, Wes94, XRZ09, You13, Zak06, Dev92, Lar90]. Mandelbrot-Like [DAB99, Vrs86]. Mandelbrote [Fri04, Pic07]. Mandelbus [Ano89b, Dew89]. Mann [Man61b]. many [Man81a, Suz91]. many-body [Suz91]. map [Car99b, Man84h, Man85c]. mapping [BF95, Man83c, Man83d]. mappings [CCC00]. Maps [Mil86, Car99b, Eng93, Ham95, IKP01, Man84h, Man85a, Man85b, Man85d, Man85e, Man85f, Man85c]. March [CMR06, Fri04]. marchés [MH05c, MH09]. Margaret [Pic07]. Mark [Man60c]. Market [Man68, Mas05a, Mas05b]. Markets [Man66a, MH08b, MH04a, MH04b, MH05c, MH05b, MH09, MH05a, MH08a, Gru05]. Markov [Man60b, Man60g]. Markovian [Man61i]. markoviens [Man60g]. Märkte [Ano05, MH05a, MH08a]. Martingale [Man66a, Man71c, BCM03, Man03b]. martingales [BJM10b, KP76]. Maryland [SMR84]. mass [BM97]. Massachusetts [Man60c, RW89, IEE03]. Mat [Man61a]. matching [KPY96]. Material [Bal10]. materiality [SBvB^{+}12]. Materials [Sch89, Tsu08, Cla86, HWM87, LMP85, MP84, WSM88]. Math [Tau98]. MATHEMATICA [Bis95, Bau96, Bau05, GH04, Kau92]. Mathematical [AA85, AA08, Ano03a, BGL67, Gar78, GM61, LvF04a, LvF04b, LBG63, Man60c, Man60d, Man61f, MS78, Ste00b, Man52a, Man53a, Man84g, Man00, Man57c, Man61e]. Mathematician [Hof10, Ano10a]. Mathematics [BF12b, Dev04, GM61, Kal03, Man60e, Man02b, Pop22, SL11, Ano03a, ABC^{+}94, Dew85a, FF91, FM02b, Jak84, Man60d, Man01e, Man02a, MF02, Man04b, Sch11, Man61f, Man63a]. mathématique [Man52a, Man53a, Man57c]. mathématiques [Ste98]. matrices [DL92, MGAP85]. matter [MT10]. Maverick [Ten10, Man02c, Tay13]. Max [PR84b]. Max-Planck-Institut [PR84b]. Maximization [Man62b]. Maximum [Izs06a]. Maxwell [DT91, Man52b]. May [CR83, PR84b]. McGuire [Goo91, Lan89]. mean [Man74a, Man74b]. meanings [Man90c]. Measure [LVEB09, Wil97, EJM91, EMW92a, EM92, EMW92b, Gué01, ME91, Man92b, Man93c]. Measure-Valued [LVEB09]. measurement [DDH10, MW69c, Man84h, Vos86]. Measures [Man95c, Man95e, MR97, BCM03, BM04b, BM04c, BM04a, GH98, GM88, Jor12, Man84e, Man86a, Man89b, Man89e, MEH90, Man93c, Man03b, Pey91, RM95, RM98, Set92, dACSS12]. Measuring [Mil88, LSL86]. Mechanical [PM94]. mechanics [Bau96, BBB^{+}03]. Mechanism [BRRK94]. Mechanisms [WGKS12]. medallions [RPAM06, RAA^{+}08]. media [DT91, PC06]. medical [HKHP11]. Medicine [Pic07]. medium [MSW08, dVSC96]. medium-range [MSW08]. Meeting [DLM99, LvF04a, LvF04b]. membranes [GDC^{+}10, Lap08]. Memoir [Tay13, Man09b, Jul18]. Mémoire [Jul18]. memoriam [Jaf11]. memory [Man01d]. Mengen [Man61a]. mesh [Wai88, ZWZL06]. mesh-connected [Wai88]. mesoscopic [FBVdlN99]. message [DLM99, Man51a, Man51b]. metal [MP84]. metallic [MSW08]. metals [MPP84]. metaphilosophy [Mur93]. metaphor [Old01]. metaphors [MM08]. MetaPost [Gün24]. Method [DR99, Miy90, Bis95, BM93, Car99b, CJK11, CP98, Dro10, FpDS99, JT96b, KKL95, KGN^{+}24, KPY96, Nar07, Sch11, WW08, YHHS93, ZZZ^{+}96]. méthodes [Man57c]. methodology [Man72b, VJ06]. Methods [BBBH92, Man61e, Man63b, Man63c, Man09c, BB79, BT81, Car96, Car99a, COK96, DMT03, FSR94, KH92, Man57c, Man64a, SS89, TPVG06]. Metropolitan [Hou95]. Mexico [CR83]. Michael [Ano03a, Kal03, Lev05]. micro [YJ09]. micro-architecture [YJ09]. microgravity [VCM^{+}11]. microparticle [dLKHS06]. micropulses [CGM95, CGM96, Man95b, Man96a]. microscope [Dew85a]. microvascular [RPS^{+}06]. middle [SCC^{+}07]. midgets [PFR94]. MIDI [MPM02]. MIDI-Encoded [MPM02]. Mild [MT10, OMK^{+}05]. mineral [FSD89]. minimizing [HKHP11]. Minimum [Wu99]. mining [Win05]. Minkowski [Man93c, Man95e]. Miranda [Par89]. Mis [MH08b, Gru05, MH04a, MH05b]. misbehaviour [MH04b]. miss [Thi88, Thi89]. Mittag [BM97]. Mittag-Leffler [BM97]. Möbius [MM08]. Model [BATL88, BK02, BM63, LM85, Man68, Aus14, BBM79, BC80, BM93, CB97, GM84b, Has07, Izs06b, JV98, LLP96, Man61h, Man62e, Man73d, Man75e, Man79a, Man83a, MG84, PR85, TM86, WY92]. model-based [Man62e]. modèle [Kah74, Man62e, Man75e, Man79a]. modéles [Man66d]. Modeling [BK03, CMP07, FMS96, Liu06, WGKS12, ZT96, Gro92, TG09]. Modelling [MC02, Bar88b, Bar93a, Man90b]. Models [DKW94b, GM64, Man66a, Man71c, Man82a, MC63, Mit03, Cha11, DLS93, Dav96, Izs06a, KPY96, Man61i, Man66d, Man75j, MV89, Man05b, Sah93]. modes [KMNW99]. modified [WW08]. Modifying [CGH02]. modular [CCC00]. Molecular [AFP84, NWM^{+}06, NFT^{+}01]. molecules [Man85a]. moment [BBM79, Man86a]. Moments [dACSS12, BM04c, Man74d]. money [Jon94]. monitor [FP89]. Monograph [Man60c]. Monotonic [Man60e]. monsters [Man82c]. monstres [Man82c]. Monte [Pic86]. Morf [CC58, Lee59]. morphing [Mar04]. morphisms [Dek12]. Morphologie [PR84b]. morphology [CGC^{+}09, PR84b]. Morphosis [Eng93]. mortality [Hig11]. most [Dew85a]. Motion [BB99, MM70, Taq13, CGM95, CGM96, Man76a, Man95b, Man96a]. Motions [MV68, Man82b, MJ97, PC06]. Motivate [DAB99]. mountain [BM93]. Mountains [Man10b]. Mouton [Man61f]. mouvement [Man76a]. Moving [SO93, Gon98]. moyenne [Man74a, Man74b]. MPI [DLM99]. MRI [HLTP01]. multi [Bou09a, Egg99, KM11, Man75b]. multi-purpose [KM11]. multi-temporal [Man75b]. multi-wavelets [Bou09a]. multi-word [Egg99]. multichannel [VM95]. Multifractal [BK02, BM99, BM02, Fra07, GT07, Jaf97a, Jaf97b, LL97, Liu06, Man89e, Man95c, Man99a, Man03d, MDM02, Pey91, RM95, BK03, BCM03, BM04b, BM04c, BM04a, DLS93, GM88, Jaf98, Man88a, Man89b, MEH90, Man93c, MJ97, Man03a, Man03b, MdF05, MV02, RM98, TPVG06]. Multifractality [ME91, VHA05, Man90c, Man99d, dSVTM99]. Multifractals [BM04a, FMS96, Man86b, Man91d, Man97a, MR97, Man99b, CM99, Man90b, Man90a, Man95a, RM97, STI99, Man01c, Ano99, Bak00]. Multimedia [BSV02]. Multinomial [Man95c, Izs06a, Man89b, RM95]. Multiperiodic [BCM03, Man03b]. multiple [RPAM06, RAA^{+}08]. multiple-spiral [RPAM06, RAA^{+}08]. Multiplications [Man74a, Man74b]. Multiplicative [LR00, Man61k, BM04b, BM04c, BM04a, BM09, BJM10a, Man86a, Man90b, XRZ09]. multiplicité [KM65]. Multiplicity [Man74e, KM65]. Multiplying [CM99]. multipoint [Bar73, EB73]. Multiresolution [JTEA91]. Multiscale [RRB06]. Multivariate [AB85]. Mumbai [VBMS96]. Munich [BDLS96]. Museum [RW89]. Music [MPM02, MVW^{+}03, VC75, VC78, Cam87, Gar78, MRM^{+}05, Sch87]. Mutants [Man74c]. Mutation [Man74c]. Mycoparasite [ME72]. Mystery [Mag04, Met94].

N [Man70a, Man99b]. N.Kolmogoroff [Man61a]. nano [CGC^{+}09]. nano-agglomerates [CGC^{+}09]. Nanoassembly [NWM^{+}06]. Nanoscale [EKKR06]. National [SMR84]. NATO [BBBH92, Fis95]. Natur [Kur90, Man91b, Voj89]. Natura [MM89]. Natural [Man81b, Man07, Cha11, CM99, DdCVR03, GH04, KH92, Man54b, Man56c, Man62d, Man78c]. naturaleza [Man97c, FR98]. Nature [ABLM98, Man83b, Man91b, Man97c, Man98b, MPB^{+}98, Ste10, VMH^{+}83, Vos88, DSS90, Fla98, KRP^{+}91, MSW08, Man76a, Man82e, Man82c, Nov04, WY92, Bri85, Boy83, Bur84, Cam96, Can84, Cha84, Fer83, Goo84a, Goo84b, Spa84, Whe83, Kir83]. Nauk [Man61a]. Nawrotzki [Man61a]. Ndynamics [ADdM^{+}12]. Near [Man63d, KK90]. near-optimal [KK90]. Necklaces [Man78b]. need [Man05a]. Negative [Man90a, Man91d, Man95a, Man03d, Man70a, Man89b, Man90c, MF09]. neighborhoods [Man93a, SAM93]. Neonatal [VBM98]. Net [Ano98b]. Networks [Orb86, LSL86, NYO94, RPS^{+}06]. Neuron [GM64]. neurophysiologic [HWEH05]. Never [Sel10]. newer [Man64b]. News [Ano10a]. Newton [Fri04, Bis95, Car99a, Car99b, Sch11, Smi91, YHHS93]. NGC5506 [MC87]. Nigel [Lev05]. Nirenberg [Yue10]. No [Kir79, Kir83, Man60e, Man60c, Man73d, Man01d, WW08]. no-search [WW08]. Noah [MW68]. Noise [Lew89, VC75, Ano99, Man55a, Man67a, Man71a, Man72a, Man76c, Man99b, Man02d, VC78, Vos79, XRZ09, vW91, Bak00]. Noises [MV68, Man67a, MW69b]. Noisy [GPS92, EHH05]. nombres [Man56b]. Non [BM04c, Man71b, Van09, BR77, Fle97, FSD89, Gin02, Gui87, KKL95, KGN^{+}24, Man60g, Man62e, MC02]. non-additive [Man60g]. Non-degeneracy [BM04c]. non-distributive [Gin02]. non-equilibrium [Fle97]. non-fractal [FSD89]. non-Gaussian [Man62e]. non-iterative [KKL95]. Non-linear [Van09, BR77, KGN^{+}24, MC02]. Non-Normal [Man71b]. noncommutative [Lap08]. noncyclic [MW69c]. nonfractal [CS87]. noninteger [GMMA83]. Nonlinear [GOY87, TE92, AGP81, BB79, CR83, NFT^{+}01]. nonperiodic [Man72b]. Nonsmooth [Man96b, DT91, JM96b]. Nontransient [MKA02]. Noor [ARC14]. Normal [Man71b, RPS^{+}06]. Normalized [Man75g, Man75f, Man75h, Man85b]. Norway [Fis95]. nos [CHR54]. Nostrand [Man60b]. notable [Wol16]. Note [FP89, Hoo91, Man76c, Gué01, Man57b, Man59d, Man61h, Man61j, Sim61b, WLWL22]. notes [Sim60, Val11]. Notion [Man89g, Man52c, Man85b, Man90c]. nouveau [Man79a]. Nouveaux [Man66d]. nouvelles [Man73a, Ste98]. Novel [KKL95]. November [And11, SMR84]. NP [STI99]. NP-completeness [STI99]. nuclear [BT81]. nucleation [Fle97]. nucleus [MM10]. Number [GM61, Gof91, Man60c, Man74c, DM92, Man56b, Man75j]. number-area [Man75j]. numbers [Sen90]. numeration [DT89, DT89]. Numerical [ADdM^{+}12, Bin24, DdMdO^{+}99, TPVG06, AK13, CMP07, Dro10, NYO94, AGP81].

O [Abb95, Dev92, Nat91]. Obituary [And11, Seu12, Ano10c, LG10a, Seu11, Tay11]. object [Dew85a]. Objects [DHN85, Kaj83a, Kaj83b, BF95, DT91, Jak84, Man75d, Man77d, Man78d, Man84f, Man89d, Man98c, Man10c]. objets [Man75d, Man78d, Man84f, Man89d, Man98c, Man10c]. observation [dLKHS04]. Observed [KLPS06]. observers [EW96]. obstacle [SH92]. obtained [Imr06]. occurring [Man54b]. October [AF90, And11, BDLS96, BCM92, Fri04, IEE95, IEE03]. off [EMW92a, EMW92b]. off-off-lattice [EMW92a, EMW92b]. Offer [Adl12]. Old [Man74c, Wal96, Man02d]. Olkin [Man61b]. One [Abb95, Nat91, PJS92d, AvHP^{+}97]. one-dimensional [AvHP^{+}97]. open [Man86a, Sta96, VCM^{+}11]. operational [MW68]. operator [PC06, Ska99]. operator-stable [PC06]. ophthalmology [KNBZ12]. Opinion [Hir89]. Opinions [Man93d, Man93e]. optical [NFT^{+}01]. optics [KMNW99]. Optimal [MH94, KK90]. Orbit [Ano93a, Ano93c, Are93, Dev91, Swe91, ARC14, Car99a, Dev90, Ye02]. orchestra [VM98]. Order [Pic87, Pop22, Vrs86, CR83, MSW08, NFT^{+}01, SL97, CR83]. ordered [SL97]. Ordering [All90, YM94]. Organic [WT10]. Organisms [WBE99]. organization [BP84, FP10]. organizing [Ham95]. orientation [CGC^{+}09]. orientation-specific [CGC^{+}09]. origami [Van10]. original [Man60e, Man61a]. originating [Man04b]. Orlando [GMA^{+}98]. orthogonal [Bou09a]. Oscillation [Jaf98]. oscillatory [Fle97]. osculate [Man82d]. osculated [Man83e]. Other [Man64d, Man67g, Man03d, Bur81, CM91, DM90, KH92, Man78c, Man84i, YMK94]. Otto [Goo91, Lan89]. Outils [Lef82]. outline [Man56c]. overview [DMT03].

P [Man61a, Man63a]. package [ADdM^{+}12, CGC^{+}09]. Padé [BB76, Bar73, BR74a, BR74b, Bar75, Bar76, EB73]. Padé-approximant [BR74b]. Padé-type [BR74a]. pages [Ano98a, Ano99, Ano03a, Ano03b, Kir79, Kir83]. paintings [JSM06, TMJ99, TMJ06]. Panorama [FM02a]. Pantheon [Tay13]. Paper [Man79b, Man59d, Man70a]. paperback [Fri04]. Paperfolding [Dek12]. papers [Man12]. Parabolic [Bur02, BB79, Man93c]. Parallel [BDLS96, BL93, JM96a, KVMW95, Man05b, SC88, UH96, DLM99, DMT03, Pic94, VM95]. Parallelism [GDH96]. parameter [Eng93, GJV17, TW11]. Parameterization [GF95]. Parameters [SVVRGA09, Izs06a, RVA00]. Parametric [OdlCA02]. Parametrized [CK93]. Paretian [DrKP10, Fam63, Man61k, Man62b, Man63d, Man64d]. Pareto [Aus14, Man59e, Man59f, Man60f, Man62d, Mas05a, Orl76]. Paris [MS78]. Parsimonious [BK02]. Part [Abb95, LvF04a, LvF04b, Nat91, PJS92d, PJS92e, BM04a, BM04a, Jaf97a, Jaf97b]. Partial [GMAK84, WY96, CT91]. Partial-dimensional [GMAK84]. partially [DV93]. particle [FBVdlN99]. Particles [SO93]. passing [DLM99]. Path [MF99, Suz91]. Patrick [Man61e]. Pattern [HKE01]. Patterns [Dev88, Miy90, VEJA89, DL87, Fle97, GH04, HTHE05, Nov04, SS91, Van09]. Paul [Man62e, Man95f]. pays [Ano10e]. Peano [Man82c, Man85f, MJ97]. Peitgen [Abb95, Dev92, Goo91, Lan89, Nat91]. Pelling [Pic07]. Pennock [Lev05]. People [AA85, AA08, Man88b, Wol16]. perceived [BM97]. Perceptual [Ska99]. Perceptually [LV96]. Percolation [AGMK81, BJ19, HASM96, CCD88, CCGS91, DM90, GAMK81, GMAK84, GM84b, Man83a, MG84, MS94, Sah93]. perdre [MH05c, MH09]. Performance [JT96a, WM94]. performing [CGC^{+}09]. perimeter [Imr06]. Period [GF95]. periodic [CGH02, CM99, WC06]. periodontal [Lan97]. permanent [Man60g]. permanents [Man60g]. personal [Bou11, Wol16]. Perspectives [Man84a, Wol16]. perturbed [AK10, yWjCnG07, XRZ09]. Pervades [SG00]. Peter [Man73d]. pg [Seu12]. Phase [CCGS91, GAM83, GASM84, GAM84, PR85, SH92]. Phenomena [Man64d, AGMK81, GMA80, KH92, Man78c, Nov04, Sah93, San85]. Phenomenological [Man64c, Man56d]. phenomenon [NDS85]. Philosophical [Mur93]. phrases [Egg99]. Phys [EMW92a]. Physical [Man61d, Man77d, MG84, SO93, NYO94, SMR84, Man51b]. Physics [Ano93b, Bau96, Man02c, AF90, Ano99, BT81, Bau05, FF91, Man78a, Man84e, Man99b, Man01e, Man04b, PT86, Smi91, Suz91, ABC^{+}94]. Physique [Man61d]. physiques [Man51b]. Piaget [CB59]. Pial [HKE01]. Pics [Ano98b]. Pictures [PR84b, PR84a]. Piecewise [BH89]. PIFS [LLP96]. pioneer [Ano10b]. pipelined [JM96a]. Pixel [All90]. Placement [Pic86]. Planar [PB89, Man84h]. Planck [PR84b]. Plane [Man92a, Car96, Ent89]. planefilling [Dek12]. planes [CZ98, GJV17]. Plants [Smi84]. Pleasant [MVW^{+}03]. plenary [CMR06]. Plotting [Bur02, Sch86]. pluri [Man75b]. pluri-temporelles [Man75b]. Point [BR76, ADR01, Vla94, YHHS93]. pointlike [DT91]. Points [For04, Zak06, FR92, Man99c]. Pointwise [RB79]. Poisson [Man75b, Vla94]. poissonienne [Man75b]. Pole [MM70]. Pollard [Man60e]. Pollock [JSM06, TMJ99, TMJ06]. Pólya [JM96b, Man96b, MJ97]. polygons [Van09]. Polymer [NWM^{+}06]. Polymers [Sch89, Dha05]. polynomials [Bin24]. Polypyrrole [EKKR06]. pondérée [Man74a, Man74b]. pools [Sau96b]. popularity [Aus14]. Population [Man74c]. Porosities [BJ19]. porous [PC06]. positive [Man60g]. Possible [Hou95, Man72e, Esp97, GAMK81]. Post [Man61j, VBM98, Sim61a]. Post-Conceptional [VBM98]. potential [BA78, Man49, ME90]. potential-energy [BA78]. potentials [Man91a]. Potts [AAM^{+}03]. pour [Man56a]. powders [MP84]. Power [MSW08, Man03d, Mit03, Pic94, LG10b, Man01d, Man09c, SL97, Win05]. Power-law [MSW08, Man09c]. powers [DM92, Fra98]. Pp [Fri04, Pic07, BCM92, Man60e, Man60b, Man60d, Man61f, Man61a, Tay13, Whe83]. pp. [Man60c, Man63a]. Practical [Gib93, Man03c, PH91, FP89, Izs06b]. Practice [Pic07, DDH10]. praxis [MM08]. pré [Man55b]. pré-correction [Man55b]. preasymptotics [Man95a]. Prechter [Lev05]. precorrection [Man55b]. Prediction [Fai69, Fai05, Thi89, TE92]. predicts [HRO02]. Preface [Man98a]. Preis [Bri85, Kur90]. presented [CMR06]. President [Ano10e]. Press [Man60e, Man02c, Man60a]. pressure [SCC^{+}07]. Preterm [Hel92]. Price [Kir79, Kir83, MT67, Man71c, Whe83, Man61a, Man66d, Man73c, Man03a]. Prices [Man63e, Man63f, Man66a, Man67g, Man68, Man72c, Man62e, Man70b, Man73d, Man01b, Man01c]. primer [MF09]. Princeton [Man60e]. principle [Man75a]. Principles [Man64c, BR74a]. printing [PJS93]. prix [Man62e, Man66d, Man73c]. Prizes [Ano93b, Man02c]. Probabilistic [BBBH92]. probabilités [BBB^{+}10]. probabilities [BR75]. Probability [BP17, Man60c, Man61b, BBB^{+}10, Gué01, MS78, Man62a]. Problem [PB89, Pro95, BBM79, BV11, CM91, HCF95, Man73b, PSvH84, Pro97, VM95]. problème [Man73b]. Problems [BR76, BR77, KK93, Man49, Man86a, GM61]. Procedurally [Kaj83a, Kaj83b]. Procedure [CR94]. Proceedings [BBBH92, IEE95, LvF04a, LvF04b, Man61e, SMR84, VBMS96, AF90, AGP81, BDLS96, CR83, DK06, DLM99, KK93, MS78, PT86, ACM03, IEE03, Rai03]. Process [BCM92, Man74c, Man75g, CCD88, CCGS91, DM90, Lan97, Man69, Man72a, Man73d, MdF05]. Processes [Man74e, BM09, GM84a, Man59e, Man59f, Man60g, Man62e, Man64e, Man65d, Man75f, Man75h, San86, Vla94, Man61c]. Processing [IEE95]. Processus [Man60g, Man59e, Man59f, Man62e, Man65d]. producing [BF95]. products [BM99, BM02, BM04b, BM04a, DL92]. Prof. [Man70a]. Professor [Ano91]. Profiles [AA85, AA08, BZ96]. Profundity [Ano87, Dew87]. Program [Kat86]. programs [Kok94, Thi89]. Progress [Fai69, Fai05, MPM02]. promenades [Man78b]. Properties [LM85, Tsu08, Bar73, CCD88, HWEH05, KVŽŽ08, Mai90, MW69a, MG84, Man85a, NFT^{+}01, NYO94, NCS04, Jaf98]. property [Bra94]. proposed [DSS90]. proteins [LR85, MJ86]. Provided [WT10]. Pseudo [Car96, SSK93]. Pseudo- [Car96]. Pseudo-satellite [SSK93]. psycholinguistics [Man65a, Man66b, Man66c]. Psychology [BGL67, LBG63]. pulse [NIS^{+}03a, NIS^{+}03b]. Pulses [CGMST95, Man03c, BM99, BM02, Man95d]. Pure [GM61]. Purely [Man64c, Man56d]. purpose [KM11, Wai88]. puzzle [Edg00]. PVM [BDLS96, DLM99, JH97, UH96, Wu99]. PVM/MPI [DLM99].

QoS [FHD09]. quad [Gin02]. quadratic [Eng93, Man83c, Man83d, Met94]. quadtree [JM96a]. Quant [Man01d]. Quanta [Man51a]. Quantification [VBM98, NIS^{+}03a, NIS^{+}03b]. quantitative [Mon01, NDS85]. Quantization [HMS96a, Mas05b]. Quantum [FdL10]. Quasi [NR95]. Quasi-Affine [NR95]. quasicrystals [ADR01]. Quasilinear [GAM83]. quatre [Man75c]. que [CHR54]. Quelques [BBB^{+}10, Man74b]. questions [Man84g, Man00]. Queue [MDM02]. queueing [RRB06]. Quinn [ABC^{+}94]. Quotients [Ate93].

R [Goo91, Lan89]. Rabbit [Hel92]. Radial [MKA02]. radiation [KK93]. radiative [DLS93]. radical [Man85b]. Radiosity [Wu95]. Rain [LM85]. raised [Man84g]. Ramified [VEJA89, AST07, GAM84]. Ramsey [Man70a]. Random [GM64, Man61k, Man64d, Man68, Man71c, Man91d, MH94, Vos86, WGKS12, Wu95, BM04b, BM04c, BM04a, Bra94, DM90, KM65, Man59c, Man64e, Man67d, Man72f, Man74a, Man74b, Man78b, Man93a, Per96, SAM93, TE92, Vla94, YM94, BM04b, BM04c, Man74a, Man74b, Man75b, Man74e]. Randomness [PJS04b, MT10]. Range [Man75g, MSW08, Man64e, MW69c, Man75f, Man75h, RVA00]. ranges [HRO02]. Raphson [Bis95]. rapid [Tem96]. rapports [Man73c]. raster [ADR01]. rat [Fin89, HKS^{+}05, HTHE05, HWEH05]. Rate [VBM98, WM94, NIS^{+}03a, NIS^{+}03b, Slo94]. Rate-distortion [WM94]. rates [Man70b]. ratio [Pro97, Thi88, Thi89]. Rational [BR78, Vrs86, Fat17b, Fat17a, Jul18]. rationale [Man85f]. rationnelles [Fat17b, Fat17a, Jul18]. Rauzy [Sir97]. Ray [Bou85, HSK89, Kaj83a, Kaj83b, MC87, Ohn90]. Ray-Fractal [Bou85]. rays [RAA^{+}08]. Reaction [Kop88]. Reactions [LL97]. reactivity [NIS^{+}03a, NIS^{+}03b]. Reading [JeJP58]. Real [RHD^{+}06, RVA00, Bar73, Bar88b, Bar93a, CB97, Eng93, Hir94, Man59b, Mar10, Oli91]. Real-Time [RHD^{+}06, RVA00, CB97, Mar10]. Realistic [BB90, Mar10, Maj98]. réalité [Man73b]. Reality [Pop22, Man73b]. realizations [NYO94]. reasons [ALW^{+}88, MF02]. receptor [TKR^{+}97]. recipe [Gon98, Loy91]. recognition [TQ10]. Recognizing [MPM02]. recomposition [JM96a]. records [ARMG03, MW69a]. Recreations [Dew85a, Dew85b, Dew87, Dew89, Ste00b]. rectifiable [UP07]. Recurrence [MB94, Man84c, NIS^{+}03a, NIS^{+}03b]. recurrent [CK93, Man55a]. Recursive [Ran91, MV89, OdlCA02]. Reducing [Sig96]. reduction [Kom95]. Réelles [Man59b]. Reemerge [Tau98]. refinement [Man72e, Mas06, ZWZL06]. Reform [Pic07]. Refugee [Pop22]. Region [Hou95, BM93]. Regression [Man71b]. regular [AS92, Van09]. Regularity [Man96b, DL92, JM96b]. regulation [FP10]. Related [Man06c, CD93b, GM88, KVŽŽ08, Man61i, Man85c, Man06b, Nov04, Yue10]. relatifs [DT89]. Relation [EG94, GAMK81, GM83, dLKHS04, dLKHS06]. relationship [Imr06, Man73c]. relationships [HRO02]. relative [DT89, Man56b]. Relativity [Had88]. relaxation [LPZ08, VM95]. Relevance [MS78]. relief [Man75j]. Remarques [Gui87]. rend [Ano10e].

Render [Hug98]. Rendering [Man82a, Mil86, MKM89, SC88, Car96, Car99a, HD91, Mar10]. renderings [Gin02]. Rendite [MH05a, MH08a]. René [Man61d]. Renewal [Man72f]. Renormalization [Man99c, Vla94]. renormalized [CMP07]. repartition [Man59f, Man59e, Man59f]. Repellers [PS88b]. Reply [Sim61a, Sim61b, TMJ06]. representation [AO97]. Representations [Par89, Jon90b, Roj91]. Reproducing [Lin94]. requests [BK03]. rescaled [MW69c]. Research [DK06, Egg99, SL11, WC06]. reshaping [CGH02]. resistance [Fro97]. resonators [KMNW99]. resources [Man62d]. Responses [ABC^{+}94]. restoring [WCH03]. Resulting [Man91d]. Results [Jaf97a, CK96]. retinal [Mai90]. Retrospective [Pei10b]. returns [GMV09]. Rev [EMW92a]. revealed [JV98]. revenus [Man59e, Man59f]. Review [Abb95, Ano93a, Ano98a, Ano99, Ano03a, Ano03b, Are93, Bak00, Boy83, Bre79, Bri85, Bur84, Cam96, Can84, Cha84, CHR54, CC58, CB59, Dev92, Dys78, FR98, Fer83, Fri04, Gar91, GM61, Goo78, Goo84a, Goo84b, Goo91, Gru05, Hal96, Har90, How78, Ios79, Kac78, Kal03, Kil78, Kir79, Kir83, Kur90, Lan89, Lar90, Lee59, Lev05, Mag04, Man60e, Man60a, Man60b, Man60c, Man60d, Man61b, Man61e, Man61c, Man61d, Man61f, Man61a, Man62a, Man63a, Man74e, Mir01, Mud02, Nat91, Pic07, Sch05, Spa84, Sto79, Swe91, Tay13, Voj89, Whe83, Vos79]. Reviews [Liu06]. revisited [Man09c]. Revisiting [JSM06, TMJ06]. Revolutionized [Pop22]. Revue [CHR54]. Reward [MH08b, MH04a, MH04b, MH05c, MH05a, MH05b, MH08a, MH09]. RI [Fin89]. Richard [Ano05, Gru05]. Riemann [Lap08, Sla07]. Rigor [Hir89]. ring [AS92, Man01a, Man02a]. Riseman [BR74b]. Risiko [Ano05, MH05a, MH08a]. Risk [Man64d, MG98, MH08b, DDH10, Man97b, MH04a, MH04b, MH05c, MH05a, MH05b, MH08a, MH09, Mir01, Ano98a]. risks [MT10]. risquer [MH05c, MH09]. rivers [Man88c]. Robert [Ano93a, Are93, Hal96, Lev05, Man61c, Swe91, Ano93c, Man60d]. Robust [MT79]. Robustness [MW69c]. rocks [Sah93]. Role [Man62c, FR92, Man95a]. Romans [Est11]. Rood [Lev05]. roots [Bin24]. Rosie [Ano91]. rough [BZ96]. Roughness [Man06c, Man06b]. roundoff [LKH97]. rugged [Cla86]. Ruin [MH08b, MH04a, MH04b, MH05c, MH05a, MH05b, MH08a, MH09, Ano05, MH05a, MH08a]. Rule [Bar71, Man75j, Man85a, Mei92]. rules [Cha11, Rar89]. run [Man69, MW69c, MW69a, MT79]. Rupture [Hou95]. ruptures [GDC^{+}10]. Russian [Man61a].

S [Sto79, Voj89]. S. [GM61, Man60e]. sa [Man56a]. sabbaticals [Win05]. Sake [Man89c, Man89a, Man93b]. Salesman [PB89]. Salle [Man63a]. sampled [TPVG06]. samples [DM73]. Samuel [Man61e]. San [ACM03, Bri85, Kir83, LvF04a, LvF04b, Sto79, Whe83]. sans [Man59b, Man78b]. Sarkozy [Ano10e]. SAT [STI99]. satellite [SSK93]. Saupe [Abb95, Dev92, Goo91, Lan89, Nat91, Abb95, Nat91]. scalars [Man75i]. Scale [LS92, San85, DL92, EPV95, Vor02]. Scale-invariant [San85]. Scalebound [Man81b, Man07]. Scales [Wil97, Jor12, Man81a]. Scaling [LL97, Man81b, MG98, Man01b, Man01c, Man07, MB94, Mir01, NYO94, WBE99, ARMG03, Ano98a, Bis95, Bra94, FPR92, HRO02, MSW08, Man78a, Man83c, Man83d, Man97b]. scanning [WC06]. Scattering [Jak84, GH98, SH92]. schemes [Hür95]. Schönheit [M^{+}85]. School [Gar91]. Schröder [DAB99, Vrs86]. Science [Caw89, Goo91, IEE03, Lan89, Man89c, PS88a, PJS92a, PJS92b, PJS93, Pic07, BDM^{+}88, Man88b, Man89a, Man93b, Man95a, Man01a, PJS04a]. Sciences [CHR54, Man61e, Man81b, Man07, Man64b, Man67f, Man73a, Man01a, Man02a, SMR84]. Scientific [Tay13]. scientist [ALW^{+}88, ALW^{+}88]. Scientists [DG99]. Scott [Pic07, Fri04]. scratched [Shl05]. Script [Cla10a]. scriptum [Man61j, Sim61a]. seacoasts [Man77d]. search [HS93, Lap08, RBM96, WW08]. Seckler [Man62a]. second [PJS93, VBMS96, Bou09b]. sections [Mei92]. Secular [MM70]. security [JT96b]. See [Ewi95]. Seeing [FdL10]. segmentation [HKHP11, KCC89]. Seiten [Bri85, Kur90]. Selecta [Ano98a, Man97b, Man99b, Man02d, Man04a]. Selected [Man04b, Bau96]. Self [CR94, FP10, Jaf97b, LVEB09, Lin94, Man64e, Man65c, Man67b, Man70c, Man75g, Man83e, Man85g, Man86c, Man86d, Man86e, Man92b, Man93a, MF99, MKA02, Roo10, SAM93, YMK94, dVSC96, Ano99, Ano03b, Baw11, Bra94, DV93, DHKT95, EM92, Ham95, HJ94, Kok94, Man65d, Man74d, Man75f, Man75h, Man85a, MEH90, ME91, Man92a, Man99b, Man02d, Ano90, Man67d]. Self-affine [Man85g, Man86c, Man86d, Man86e, Bra94, DV93, DHKT95]. self-affinity [Ano99, Ano03b, Man99b, Man02d]. Self-Contacting [MF99]. Self-gravity [dVSC96]. self-homothetic [Man65d]. Self-inverse [Man83e]. Self-Normalized [Man75g, Man75f, Man75h]. Self-organization [FP10]. self-organizing [Ham95]. Self-regulation [FP10]. Self-Reproducing [Lin94]. Self-Similar [CR94, Jaf97b, Roo10, Man64e, EM92, HJ94, Man74d, MEH90, Man92a, Man67d]. Self-Similarity [LVEB09, Man67b, MKA02, FP10, Man92b, Man93a, SAM93, YMK94, Baw11, Kok94, Man85a, ME91, Ano90]. Sens [Man61d]. September [DLM99, RW89, VBMS96]. sequence [CZ98, Dev99, DM92]. sequences [AS92, AvHP^{+}97, GMAK84, Man85f, MGAP85, Man86a, TE92]. serial [MT79]. Series [Man61f, Bar73, Bar76, CCRR91, DR99, Man60d]. ses [Man73c]. Session [LvF04a, LvF04b]. Set [Ano87, Ano89b, Ano93c, Ate93, BH89, Bri88, Bur02, Chu11, Chu12, Dev91, Dew85b, Ewi95, FL91, For04, Gar91, GF95, Har90, Hoo91, Hor90a, Kel00b, Mag04, Mud02, PJS92e, Roo10, Sch05, Sch86, Shi98, Zak06, Car99a, Dev90, Dev96, Dev99, DK06, Dew87, Dew89, Eng93, Ent89, ES92, FR92, FPR92, FC00, Gün24, Her05, IKP01, Jon90b, Kel00a, Man83c, Man83d, Man84d, Man85a, Man85b, Man85d, Man04a, Mas05a, Met94, Old01, OdlCA02, PRÁM04, PR85, Phi92, PFR94, Rad96, Roj91, RPÁM04, RPAM06, RAA^{+}08, Sch11, Sen90, Sta96, XRZ09, Abb95, Nat91, Ano93a, Are93, Hal96, Swe91].

Set-To [Hor90a]. Sets [CR94, DAB99, Man61a, Shi98, Vrs86, ADR01, AK10, AK13, ARC14, CM91, Cas96, CCC00, DM90, DSB96, DMT03, Fal85, Gai06, GJV17, GH04, Gil86, Gin02, KGN^{+}24, Man72f, Man79a, Man82d, Man83e, Man85e, Man85f, Man86c, Man86d, Man86e, MF09, MM08, OdlCA02, PSvH84, Sau87, WC06, yWjCnG07, WLWL22, KM65]. Seventy [ALW^{+}88]. seventy-fifth [ALW^{+}88]. Seventy-five [ALW^{+}88]. Seyfert [MC87]. Sfr [Kur90, Ano98a, Ano99]. SFr98 [Voj89]. shading [KK90]. shape [KK90, LPZ08, Man75j]. shaped [DT91]. Shapes [Man81b, Man07, PJS04b, VC86]. sheets [Man84i]. Shield [WT10]. Shift [Ate93, TW11]. shifts [SL97]. Short [Man09c]. Shortest [MF99]. Shorthand [H.88]. Show [RW89]. shower [Ohn90]. SIAM [DK06]. sided [MEH90, Man90b]. siècle [Cha90]. Siegel [Man85f]. Sierpinski [BB99, GASM84, Maj98, NWM^{+}06]. SIGGRAPH [GMA^{+}98, RW89]. sigma [Man82d, Man83e]. sigma-discs [Man83e]. sign [FBVdlN99]. signal [DR99]. signals [HWEH05]. Significant [EKKR06]. Signs [SG00]. Silence [Mil57]. silicate [FSD89]. Similar [CR94, Jaf97b, Roo10, EM92, HJ94, Man64e, Man65c, Man74d, MEH90, Man92a, Man67d]. Similarity [HCF95, LVEB09, Man67b, Man70c, MKA02, Baw11, FP10, Kok94, Man85a, ME91, Man92b, Man93a, SAM93, YMK94, Ano90]. Simon [Man59d, Man61h]. Simple [Man54b, Man03d, PJS92c, YJ09, BSV92, JH97, Man01d]. simplifying [Sta96]. simulated [Sah93]. Simulation [Vos88, CGC^{+}09, CMP07, DLS93, MC02, NCS04, SW92, YJ09]. simulations [NYO94, Lan97]. since [Man99c]. Single [GM64]. singularités [Man76a]. singularities [Man76a, Man84b]. singularity [TPVG06]. Sixth [PT86]. Size [EKKR06, Hou95, HRO02, Sig96]. Skew [Sim55, Man59d, Man61h, Sim60]. Skewed [FMS96, HJ94]. slow [Man66d, Man73b]. slurries [VOL^{+}86]. small [EMW92a, EMW92b]. Smooth [BH89, CD93a, Man85e]. Snell [Man60b]. Snowbird [DK06]. soap [Man77d]. Social [Man61e]. sociales [Man67f]. Society [LvF04a, LvF04b, Man02b]. soft [MC02]. Software [VMH^{+}83, JH97]. Software-Cache [VMH^{+}83]. soi [Man65d]. Solid [Lew89, CGC^{+}09, Gro92]. solids [Dav95]. Solomon [Man63a]. Solution [Liu95b, Pro95, Wu95, AGP81, CM91, HCF95, SL97]. Solutions [BR76, BR74b, BC80, Gui87, Liu95a, Mac08, RB79]. Solvable [GAMK81]. solving [CT91, YHHS93]. somatosensory [HWEH05]. Some [DSS90, Izs06b, Man49, Man67a, Man67g, Man68, MW69a, Man89f, Man00, MF02, Mil57, Sim60, Zha09, AST07, Hoo91, Man74b, Man84i, Sir97, Wol16, BBB^{+}10]. Soviet [Man63a]. Space [GT07, AG06, BB76, LKH97, Woo94]. space-time [AG06]. Spacefilling [PB89]. spaces [Jaf98, Man61a]. spacetimes [Lap08]. Spain [DLM99]. Spanish [Man97c]. spatial [EG94]. speaks [Win05]. Special [DrKP10, LvF04a, LvF04b]. species [Man56b]. specific [CGC^{+}09]. specification [Bra94]. SPECT [OMK^{+}05]. spectra [Man69, Man95a, TPVG06]. spectral [Man67d, Man76b, Man84c, MC87].

Spectrum [ME72, dSVTM99, Man67a]. spéculatifs [Man62e]. Speculative [Man63e, Man63f, Man67g, Man72c, Man62e, Man73d]. Speech [VC75, Man67c]. speed [BE95a]. Speeding [BE95b]. Sphere [Ley05, Man75e, Man75e]. spheres [MSLG00]. Spike [GM64, FPR92]. spiral [RPAM06, RAA^{+}08]. spirals [KVŽŽ08]. Splines [ST89]. Sporadic [Man67d, Man67e]. Spot [vW91]. Springer [Ano98a, Ano99, Ano03b]. Springer-Verlag [Ano99]. square [Sau96a]. Squig [Man84e, Man84i]. Stable [CGMST95, Fam63, Man61k, Man63d, Man60g, Man62e, PC06]. stables [Man62e]. Standards [SMR84]. standstill [Bal10]. Stanford [Man61e]. star [Man01c]. State [LKH97]. State-space [LKH97]. statement [Gin02]. stationarity [Man65c, Man76c]. stationary [Man60g]. stationnaires [Man60g]. Statistical [Hür95, Man53b, Man59a, Man63b, Man63c, Man64c, Man70b, Man70c, Man72b, Man09c, JTEA91, Man56c, Man56d, Man64a, MW69c, Man78a, Suz91, Man57c, Man60c, Man67b]. Statistics [Man61b, Man62d, BM97, Has07, Vla94, Man60d]. statistique [Man57c]. step [CM91]. Stereology [MS78]. Stewart [Lev05]. stick [Rar89]. Stieltjes [Man60e, Bar73, Bar76, BR78]. still [HS93]. sting [SS00]. Stochastic [BBBH92, Lew87, Man75j, Man82a, Man01d, Vla94, Man65d, Man73d, Man60g]. stochastics [Man59e, Man59f]. stochastiques [Man59e, Man59f, Man60g, Man65d]. Stock [MT67, Man68, Mas05a, Mas05b]. Stone [Miy90]. stories [Man75c]. Story [Pop22]. Strahler [YM94]. Strange [GOY87]. Stratege [Man57a, CHR54]. strategic [PJS99]. stratégie [Man52c]. strategist [Man57a]. strategy [Man52c, Man54b]. stream [KFN00, WWGF05]. Street [Man99a]. stress [Man57b, Man75i]. strings [Lap08]. Strongly [Man75g, Man75f, Man75h]. Structural [DV93, JV98]. Structure [EKKR06, LS92, Man53b, Man54a, Man54c, TTMB96, BBM79, DM90, DKW94a, HKS^{+}05, KLPS06, Mad87, Man56c, Man83c, Man83d, Man84b, WC06, YMK94, dVSC96, CHR54]. Structure-based [TTMB96]. Structures [RHD^{+}06, FBVdlN99, GM84a, GM88, Hoo91, Man84b, MS78]. structuring [KH92]. studies [CR83, Man54a, Man54c]. Study [BBBH92, Man60e, Vrs86, AO97, Kap86, TPVG06]. subarachnoid [SCC^{+}07]. Subdivision [Lew87]. subordinated [Man73d]. subsets [Sir97]. substitutions [DT89, Fat17b, Fat17a]. such [Cam87, Sch87]. Sudhist [Man61b]. Sufficiency [Man62c]. Sum [Bar71]. Summary [Liu06]. Summer [DK06]. Sums [CGMST95, Man03c, Man95d]. super [WGKS12]. super-extreme [WGKS12]. Supercomputer [DLS93, Cia88a, Cia88b]. superconductivity [Zaa10]. Suppes [Man61e]. Supplement [Man60e]. support [dACSS12]. Surface [Man06c, BZ96, BRRK94, KK90, KH92, Man86d, Man06b, Shl05]. Surfaces [LL97, SC88, Wu95, AFP84, BBB^{+}03, Dub89, LR85, Man75i, MP84, MPP84, Wai88]. surgery [KNBZ12]. Survey [Man99d]. surveys [Ano90]. symmetries [LO96]. Symmetry [Pen88, VJ06, CCC00]. symphony [VM98]. Symposium [ACM03, Ano95, IEE03, KK93, MS78, PT86, SMR84, Man61e]. syndrome [Man73b, Man73c]. Synthesis [BB90, BJM^{+}88, Lew89, Man06c, MKM89, ABC^{+}94, Hor90b, Man06b, vW91]. System [Esp97, Bra94, CP98, JH97, Mac08, Man56a, WY96, Zha09, Pic94]. system-based [WY96]. Systeme [PR84b, Man56a, M^{+}85, PR84a]. Systémes [DT89]. systemic [MM08]. Systems [Abb95, Adl12, BHI11, DHN85, Dev88, Nat91, PJS92e, Pen88, AO97, AO01, ADdM^{+}12, BA78, BV11, Cor90, CK93, DdMdO^{+}99, Fla98, Gof91, GV89, GM88, Man56c, Man65c, Man93c, MC02, NM09, OdlCA02, PR84a, PR84b, Van09, VM95, dVdRL93, Ano89a, DT89, Esp97, HWM87, WSM88]. Szilard [Man65b].

t [CHR54, Ste98]. Tab [Voj89]. tables [Man61f]. Tac [Ste00a, Ste00b]. Taches [Lef82]. Tail [MDM02, SS00]. tailed [Rac03]. tails [DrKP10, Man03a, Man09c, Man01b]. Takagi [Dub89]. Tallinn [KK93]. tampered [WCH03]. Tan [Mud02]. tant [CHR54]. target [EW96]. Tasks [Lef82]. taxonomy [Man56c]. TCP [GK03, VHA05]. TCP/IP [VHA05]. teaching [Gof91]. Techniques [Kaj83a, Kaj83b, ZT96, ADR01, CD93b, Kel00c, WC06, Woo94]. Telephone [BM63]. Temperature [Man89g, Man56a, Man70a, Zaa10]. temporal [Man75b]. Temporally [MB94]. temporelles [Man75b]. Tenenbaum [Man60e]. Teoria [Man60h]. Terms [Man71b, Bar75, BB76, CGH02]. Terrain [Mil86, DKW94a, KH92, LPZ08, Pum96]. Terrains [MKM89]. ters [Man65c]. test [Man93a, MF09, SAM93]. testing [Izs06b]. Tests [DM73]. Text [Man61a, BP84, Man54a, Man54c]. Textbook [Man61f]. textes [CHR54, Man54a, Man54c]. texts [Man60d]. Texture [KCC89, Man79a, MS94, vW91]. textures [FSD89]. th [LMNW05]. Thau [BATL88]. theéories [CHR54]. Their [Tsu08, AO97, BB79, FM02a, Man59c, MGAP85, Man86a, Man86e, Man91a, Man95a, MS78, Sau87]. Theme [Mud02]. Theorem [Man75g, Man55b, Mas05a, Per96]. théorème [Man55b]. Theorems [LR00, Man75f, Man75h]. Theoretical [Bau96, ABC^{+}94, Bau05, Smi91, ABC^{+}94]. Theorie [Man61d, PR84a, PR84b, Man55b, Man57c, M^{+}85, AMM57, AMM74, CC58, Lee59, Man52a, Man53a]. theories [Man57a, Suz91]. Theory [ACM03, CR94, FV95, Man53b, Man61d, Man62a, Man74e, Mir90, AG06, Cas96, DDH10, EB73, KNBZ12, dLKHS06, Liu93, Man49, Man52a, Man53a, Man56c, Man56d, Man57c, Man61i, Man65a, Man66b, Man66c, Mas05a, MC02, PR84a, PR84b, Suz91, ZWZL06, Man55b, Man60c]. there [Cam87, Sch87]. thermo [Man56c]. thermo-statistical [Man56c]. Thermodynamic [Man91d]. Thermodynamics [Man62c, Man64c, Man56d, Man57c, dSVTM99]. thermodynamique [Man57c]. thing [Cam87, Sch87]. Thinking [Nov04, SBvB^{+}12]. Third [NFT^{+}01, BDLS96]. Third-order [NFT^{+}01]. Thirty [ACM03]. Thirty-Fifth [ACM03]. Thomas [Man74e]. those [MT10]. thought [Hol18]. Three [Jon90b, CCGS91, VJ06]. three-dimensional [VJ06]. Tic [Ste00a, Ste00b]. Tic-Tac-Toe [Ste00a, Ste00b]. tiles [Dek12]. Tilings [Roo10]. Time [GH98, GT07, RHD^{+}06, AG06, CB97, CCRR91, DR99, Has07, MJ97, Mar10, MC87, NIS^{+}03a, NIS^{+}03b, NM09, NG03, RVA00, WGKS12]. Time-dependent [GH98]. time-independent [NM09]. Tjon [BC80]. Toe [Ste00a, Ste00b]. Token [Man61f]. tolerance [You13]. tool [Gof91, GBR^{+}09, GFB^{+}10, Kau92]. Tools [Lef82, Man05a]. Topics [Man01e, Man85c, Man04b]. topological [AK10]. topology [Gof91]. total [Man56a]. totale [Man56a]. Tour [Ano89b, Dew89]. tourbillons [Man75c]. Tracing [Kaj83a, Kaj83b, HSK89]. track [ARMG03]. Tracts [GM61]. Traffic [BK02, Liu06, MDM02, KLPS06, TG09, VHA05]. trajectory [Far11]. transfer [DLS93, MGAP85, VM95]. Transform [BHI11, FSR94, WW08, ZY94]. transformation [EPV95]. Transformations [BB10, NR95, PJS92c]. Transforms [LVEB09, FV95, Hür93, WM94]. transistor [Fro97]. transit [NIS^{+}03a, NIS^{+}03b]. Transition [Ano93c, Dev90, BR75, BRRK94, Ano93a, Are93, Swe91]. transitions [CCGS91, GAM83, GASM84, GAM84, PR85]. Translated [Man60e, Man61a, Man62a]. transmission [Man51a, Man51b, Man55b]. Transparent [AST07]. Transport [Pen88, GM84a, KFN00]. trap [Car99a]. traps [Ye02]. Travelling [PB89]. travers [Man01a]. Traversed [GBCK93]. treasury [FF91]. Tree [Dre01, MF99, Dev99, EM92]. Tree-based [Dre01]. Trees [VEJA89, CMP07, Gon98, Man78c, Pro97, TSS85, YM94]. trémas [Man79a]. Trend [Sel10]. Tribute [Wei05, Bou11, LMNW05]. Trieste [PT86]. trinomial [DM92]. trisection [YJ09]. Trondheim [Fis95]. Tsallis [MV02]. tumorous [RPS^{+}06]. tunable [Vor02]. Turbocharging [Pou86]. turbomachines [Man49]. Turbulence [Man67e, Kah74, Man72e, Man74d, Man75i, Man75c, Man76a, Man76b, Man77a, Man77c, Man77d, SW92]. turbulent [CMP07]. Turin [CMR06]. tutorial [Gon98, Maj98, Roj91]. twenty [DK06, VBMS96]. twenty-five [DK06]. twenty-second [VBMS96]. Two [Abb95, BR76, BR78, Car99a, DL92, Dav95, Fra98, Man63d, Man90c, MKA02, Nat91, PJS92e, SS89, Bar76, BA78, Has07, Man54a, Man54c]. two-center [BA78]. two-dimensional [Has07]. Two-Point [BR76]. Two-scale [DL92]. Two-Variable [BR78]. type [BR74a, KGN^{+}24, Man61f]. types [Man56b].

u.a. [PR84b]. U.S [Kir79, SMR84]. UK [Rai03]. Ulam [GM61]. Ultrafractal [Sin07]. Ultraviolet [WT10]. Unavoidable [Man89g]. Unbiased [Man66a]. unconventional [Jon90b]. understanding [KRP^{+}91]. unicité [Man59c]. unicity [Man84e]. Uniform [BJM10b, MJ97]. unique [Man65b, Man84d]. unit [Per96]. units [BP84]. Universal [SCNT03]. Universality [NM09]. Universe [Cla10b, Lin94, Mar99, Man75a, Man75e, Man79a, Mad87]. University [Fri04, Har90, Man60e, Man60a]. unknowable [DDH10]. unknown [DDH10, Man61a]. unknowns [CT91]. unsmooth [Far11]. Unstable [MB94, KMNW99]. Upon [Man89f]. upscaling [PC06]. urban [MML01]. USA [ACM03, CR83, IEE95]. Use [Lef82, FP89, GMMA83]. used [dLKHS06]. Useful [Man81b, Man07, Sch11]. useless [Sch11]. User [DLM99, CGC^{+}09]. user-interactive [CGC^{+}09]. Users [MC63]. uses [FM02a]. Using [BHI11, Cas96, FMS96, VBM98, ADR01, Bou09a, GMV09, Imr06, JM96a, KM11, Lan97, Liu95a, MC02, NCS04, OMK^{+}05, Par89, Pum96, SSK93, UH96, Vor02, Woo94, ZT96, ZY94]. Uspekhi [Man61a]. Utah [DK06]. utilized [WC06].

V [Man60d, Man62a, Man77c, Man85d]. Valid [Jaf97a]. Validity [Fra07, Man71c]. Valuations [Fer95]. Value [BR76]. Valued [LVEB09]. values [GMV09, Man89b]. valve [PM94]. Vamping [Sin07]. Variability [EMW92a, EMW92b, SCC^{+}07, VBM98, MC87]. Variable [BR78, BEHM89]. Variables [Man59e, Man59f]. variance [Man73d, Man73c]. variances [Man69]. Variation [Man61k, Man63e, Man63f, Man67g, Man72c, Man66d]. Variational [BR75, BR74a, EB73]. Variations [Mud02]. varying [HRO02]. Vascular [RHD^{+}06]. Vasculature [HKE01]. Vector [GT07, HMS96a]. vegetation [MLDW85]. velocity [SCC^{+}07]. Vence [AF90]. Ventricles [HLTP01]. verborgene [Man08]. Verhaltenstheorien [Man57a]. verification [Man67f]. Verlag [Ano99, Kur90, Man61a, Voj89]. version [GFB^{+}10]. versus [Man09c]. verticillata [ME72]. Verw [Man75f]. Verwaltungsgebäude [PR84b]. Very [Man70c, VBMS96, YMK94]. vessels [Mai90]. VGA [FP89]. VI [Man85e]. via [GDH96]. View [MH08b, MH04a, MH04b, MH05b]. VII [Man85f]. viii [Man60b, Man84h]. vindication [Kir11]. virtual [BDLS96, DLM99]. viscosity [KGN^{+}24]. Viscous [Rob85, San85, Ghi86, NDS85, VOL^{+}86]. Vision [Dev04, HCF95]. Visual [GBR^{+}09, Man81b, Man07, SW92, GFB^{+}10]. visualisation [DMT03]. Visualization [Pic94, Maj98, vW91]. Visualizing [ADR01, Gün24, XX07]. voids [Gai06]. vol [CHR54, Seu12]. volatility [Man01d]. Volume [Ano98a, Man97b, Man99b, Man02d, Man04a, DrKP10, BGL67, LBG63]. volumetric [NG03]. Volumetry [HLTP01]. vom [PR84b]. Voss [Goo91, Lan89]. VQ [HMS96b]. VQ-enhanced [HMS96b]. vs [MT10].

W [Kir79, Kir83, Man61a, Sto79]. Wahrscheinlichkeitstheorie [Man75f]. Walk [GM64, Man71c, Man99a, WGKS12, Bra94, Man68, Man93a, SAM93]. walked [Hol18]. Walks [Man64d, Man78b]. Wall [Miy90, Man99a]. warning [Mad86]. Warped [PFR94]. Washington [Ano03a, IEE95]. Wassily [Man61b]. watermarking [KM11]. Wave [dVdRL93]. Waveform [Pic86]. waveforms [PU08]. Wavelet [Man96b, CK96, CD93a, JM96b]. wavelets [Bou09a, FSR94]. waves [Bai97]. Way [Adl12, BS88, TE92]. Weakly [Man75g, Man75f, Man75h]. weather [SSK93]. Web [FP10, BK03]. Webster [Pic07]. Weeks [VBM98]. Weierstrass [AB85, BL80, Lan97, WGKS12]. weighted [Man74a, Man74b, Set92]. Well [Man89g]. Well-Defined [Man89g]. Werkzeug [Kau92]. which [Man61a]. White [Gar78, Man67a]. Who [Pop22]. whose [BH89, Fra98]. Wild [Ano99, MT10, Man99b]. Wiley [Man60c]. Will [Lev05]. William [Man61b]. Willis [Man56b, Man56c]. Win [Ano93b]. Windows [Esp97]. Wissenschaften [Man61a]. without [Man59b, Man88c]. witness [Bal10]. Wobble [MM70]. Wolf [Ano93b, Man02c]. word [CHR54, DM73, Egg99, Man61i]. wordbook [FM02a]. wordbook-index [FM02a]. words [Sta96]. Work [Fri04, Mea89]. works [Man86a]. workstation [UH96]. world [ARMG03, Bai97, Bar88b, Bar93a, FF91, Hir94, Oli91]. written [DM73]. Wu [BC80]. WWW [BK02].

x [Man60d, Pe04, Kur90, MC87, Voj89]. X-ray [MC87]. Xeometría [Dur78]. xv [Pic07].

Yale [KM89a]. years [DK06]. Yekutieli [Pro95, Pro97]. York [Ano99, Ano03b, GM61, Man60e, Man60b, Man61f, Man62a, Man60d]. Yule [Man56b].

Z [Man75f]. zahlreiche [Voj89]. Zahlten [Dev92]. zeros [Lap08, Sla07]. Zewail [Ano93b]. Zipf [Fai69, Fai05, Man57c, ARMG03, Aus14, BP84, Cha11, CL87, Dev57, Egg99, Gai06, GMV09, Izs06a, Izs06b, JV98, KPP19, dLKHS04, dLKHS06, Lef82, MPM02, MVW^{+}03, MRM^{+}05, Man57c, Mas05a, Mas05b, Mas06, Mon01, Orl70, Orl76, Per96, SVVRGA09, TW11, You13]. Zipf-Mandelbrot [MPM02]. Zooming [Pum96]. zooms [Dew85a]. zur [Man61a]. zwischen [Ano05, MH05a, MH08a].

References

Albers:1985:MPP

[AA85] Donald J. Albers and Gerald L. Alexanderson, editors. Mathematical People: Profiles and Interviews. Birkhäuser Verlag, Basel, Switzerland, 1985. ISBN 0-8176-3191-7, 3-7643-3191-7. xvi + 2 + 372 pp. LCCN QA28 .M37 1985.

Albers:2008:MPP

[AA08] Donald J. Albers and Gerald L. Alexanderson, editors. Mathematical people: profiles and interviews. A. K. Peters, Ltd., Wellesley, MA, USA, second edition, 2008. ISBN 1-56881-340-6. xxvi + 386 pp. LCCN QA28 .M37 2008.

Asikainen:2003:FGC

[AAM^{+}03] J. Asikainen, A. Aharony, B. B. Mandelbrot, E. Rausch, and J.-P. Hovi. Fractal geometry of critical Potts clusters. European Physical Journal B: Condensed Matter and Complex Systems, 34(4): 479–487, ???? 2003. CODEN EPJBFY. ISSN 1434-6028 (print),

1434-6036 (electronic). URL http://www.springerlink.com/
content/x3m42hyyg2rt88qd/.

Ausloos:1985:MWM

[AB85] M. Ausloos and D. H. Berman. A multivariate Weierstrass– Mandelbrot function. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 400(1819): 331–350, August 8, 1985. CODEN PRLAAZ. ISSN 0080-4630.

URL http://www.jstor.org/stable/2397968.

Abbott:1995:RBC

[Abb95] Steve Abbott. Review: Fractals for the Classroom. Part One: Introduction to Fractals and Chaos by H. O. Peitgen, H. Jürgens, D. Saupe; Fractals for the Classroom. Part Two: Complex Systems and Mandelbrot Set by H. O. Peitgen, H. Jürgens, D. Saupe. Mathematical Gazette, 79(484):200, March 1995. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

http://www.jstor.org/stable/3620080.

Atiyah:1994:RBM

[ABC^{+}94] Michael Atiyah, Armand Borel, G. J. Chaitin, Daniel Friedan, James Glimm, et al. Responses to Theoretical mathematics: toward a cultural synthesis of mathematics and theoretical physics, by A. Jaffe and F. Quinn. Bulletin of the American Mathematical Society, 30(??):178–207, ???? 1994. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Avnir:1998:GNF

[ABLM98] David Avnir, Ofer Biham, Daniel Lidar, and Ofer Malcai. Is the geometry of nature fractal? Science, 279(5347):39–40, January 2, 1998. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/279/
5347/39.full.

ACM:2003:PTF

[ACM03] ACM, editor. Proceedings of the Thirty-Fifth ACM Symposium on Theory of Computing, San Diego, CA, USA, June 9–11, 2003. ACM Press, New York, NY 10036, USA, 2003. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Avellar:2012:NPN

[ADdM^{+}12] J. Avellar, L. G. S. Duarte, L. A. C. P. da Mota, N. de Melo, and J. E. F. Skea. The Ndynamics package — numerical analysis of dynamical systems and the fractal dimension of boundaries. Computer Physics Communications, 183(9):2019–2020, September 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465512001294.

Adler:2012:FOB

[Adl12] Paulo Adler. Fractals offer a better way to manage complex systems. Research & Development, November 12, 2012. CODEN REDEEA. ISSN 0746-9179. URL http://www.

rdmag.com/articles/2012/11/fractals-offer-better-way-
manage-complex-systems.

Allis:2001:VPS

[ADR01] Ned W. Allis, Jeffrey P. Dumont, and Clifford A. Reiter. Visualizing point sets, fractals, and quasicrystals using raster techniques. Computers and Graphics, 25(3):519–527, June 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.elsevier.nl/gej-ng/10/13/20/57/
32/40/abstract.html;   http://www.elsevier.nl/gej-ng/
10/13/20/57/32/40/article.pdf.

Aharony:1990:FPE

[AF90] Amnon Aharony and Jens Feder, editors. Fractals in physics: essays in honour of Benoît B. Mandelbrot: proceedings of the international conference honouring Benoît B. Mandelbrot on his 65th birthday, Vence, France, 1–4 October, 1989. North-Holland Publishing Co., Amsterdam, The Netherlands, 1990. ISBN 0-444-88646-X (U.S.). LCCN QC20.7.G44 F73 1990.

Avnir:1984:MFS

[AFP84] David Avnir, Dina Farin, and Peter Pfeifer. Molecular fractal surfaces. Nature, 308(5956):261–263, March 15, 1984. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v308/
n5956/pdf/308261a0.pdf.

Agop:2006:GTF

[AG06] M. Agop and I. Gottlieb. Gravitation theory in a fractal spacetime. Journal of Mathematical Physics, 47(5):053503, May 2006.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i5/p053503_s1.

Aharony:1981:PCP

[AGMK81] A. Aharony, Y. Gefen, B. Mandelbrot, and S. Kirkpatrick. Percolation, critical phenomena and fractals. Disordered Systems and Localization, 11(3):56–58, ???? 1981. CODEN ????

ISSN ???? URL http://www.springerlink.com/content/
yq56048um078t851/.

Allgower:1981:NSN

[AGP81] E. L. (Eugene L.) Allgower, Klaus Glashoff, and Heinz-Otto Peitgen, editors. Numerical solution of nonlinear equations: proceedings, Bremen, 1980, volume 878 of Lecture notes in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. ISBN 0-387-10871-8. LCCN QA3 .L471 no.878; QA3.L471.

Aharony:1986:F

[Aha86] Amnon Aharony. Fractals. Science, 232(4752):891–892, May 16, 1986. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/232/
4752/891.1.full.pdf.

Andreadis:2010:TCPa

[AK10] Ioannis Andreadis and Theodoros E. Karakasidis. On a topological closeness of perturbed Mandelbrot sets. Applied Mathematics and Computation, 215(10):3674–3683, January 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300309009801.

Andreadis:2013:NAA

[AK13] Ioannis Andreadis and Theodoros E. Karakasidis. On numerical approximations of the area of the generalized Mandelbrot sets. Applied Mathematics and Computation, 219(23):10974–10982, August 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300313004876.

Alderton:2007:CDF

[Ald07] Gemma K. Alderton. Cell division: Fractal cycling. Nature Reviews Molecular Cell Biology, 8(11):850–851, November 2007.

CODEN NRMCBP. ISSN 1471-0072 (print), 1471-0080 (elec-

tronic). URL http://www.nature.com/nrm/journal/v8/n11/
full/nrm2284.html.

Allen:1990:POA

[All90] Norton T. Allen. A pixel ordering algorithm. Dr. Dobb’s Journal of Software Tools, 15(6):56–61, 116–117, June 1990. CODEN DDJOEB. ISSN 1044-789X.

Arnheim:1988:SFR

[ALW^{+}88] Rudolf Arnheim, Thomas E. Lovejoy, David Gordon Wilson, Freeman Dyson, Jane Goodall, Ian Shelton, Kenneth H. Olsen, Irene C. Peden, Richard W. Hamming, Thomas Eisner, Preston Cloud, Matt Cartmill, Samuel C. Florman, Jeremy Bernstein, George A. Miller, Robert M. May, G. Evelyn Hutchinson, Jerome Bruner, Priscilla C. Grew, William Bevan, Elisabeth S. Vrba, Myrdene Anderson, Kevin Padian, Harry Shipman, Victor F. Weisskopf, Walter A. Hill, Patricia D. Moehlman, Melvin Kranzberg, Malak Kotb, Raymond Kurzweil, Marcia McNutt, Masakazu Konishi, Miriam Rothschild, Edward Teller, Alison Jolly, H. Jane Brockmann, Keith Stewart Thomson, Peter J. Denning, Benoît B. Mandelbrot, Abraham Pais, Paul MacCready, Kip S. Thorne, Ruth Sager, Gerald J. Wasserburg, Neal E. Miller, Rita Levi-Montalcini, Stephen Jay Gould, Edwin H. Land, Michel Boudart, Anne Kernan, Douglas R. Hofstadter, Rosalyn S. Yalow, Bruce H. Tiffney, Mimi Koehl, Walter E. Massey, David P. Billington, John A. W. Kirsch, Abner Shimony, J. Donald Fernie, Brian J. Skinner, Lynn Margulis, Sheldon Lee Glashow, Michael LaBarbera, J. Tuzo Wilson, E. R. Ward Neale, Rudolf Peierls, Roald Hoffmann, Mary L. Good, Donald R. Griffin, Vaclav Smil, Michael S. Turner, Sarah Ann Woodin, Luis Alvarez, George A. Bartholomew, and George B. Schaller. Seventy-five reasons to become a scientist: American Scientist celebrates its seventyfifth anniversary. American Scientist, 76(5):450–463, September 1988. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786

(electronic). URL http://www.jstor.org/stable/27855384;
 http://www.jstor.org/stable/pdfplus/27855384.pdf.

Apostel:1957:LLT

[AMM57] Leo Apostel, Benoît B. Mandelbrot, and Albert Morf. Logique, langage et théorie de l’information. Presses Universitaires de France, Paris, France, 1957. 207 pp. LCCN BC199.E7 A6.

Apostel:1974:LLT

[AMM74] Leo Apostel, Benoît B. Mandelbrot, and Albert Morf. Logique, langage et théorie de l’information. Kraus reprint, Nendeln, Liechtenstein, 1974. vi + 207 pp. LCCN ????

Andrew:2011:OBM

[And11] Alex M. Andrew. Obituary: Benoît Mandelbrot: 20 November 1924–14 October 2010. Kybernetes, 40(1–2):??, ???? 2011. CODEN KBNTA3. ISSN 0368-492X (print), 1758-7883 (elec-

tronic). URL http://en.wikipedia.org/wiki/ernst_von_
glasersfeld;   http://www.boston.com/news/world/europe/
articles/2010/11/12/intellectual_ernst_von_glasersf;
http://www.emeraldinsight.com/journals.htm?issn=0368-
492X&volume=40&issue=1/2&articleid=1921825;   http://
www.guardian.co.uk/science/2010/oct/17/benoit-mandelbrot-
obituary;   http://www.math.yale.edu/mandelbrot/;   http:
//www.telegraph.co.uk/news/obituaries/science-obituaries/
8069558/benoit-mandelbrot.html.

Anonymous:1987:BPM

[Ano87] Anonymous. Beauty and profundity: The Mandelbrot set and a flock of its cousins called Julia. Scientific American, 257(5):118– ??, November 1987. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1988:MM

[Ano88] Anonymous. Mandelbrot magic. Left Coast Software, Cupertino, CA, USA, version 3.1 edition, 1988. LCCN QA614.86 .M36 1989. One computer disk.

Anonymous:1989:CAC

[Ano89a] Anonymous. Conference on Arithmetics and Coding Systems. Theoretical Computer Science, 65(2):??, June 28, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:1989:TMS

[Ano89b] Anonymous. A tour of the Mandelbrot set aboard the Mandelbus. Scientific American, 260(2):88–??, February 1989. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1990:DSS

[Ano90] Anonymous. In darkest Self-Similarity: Hugh Kenner surveys the literature on fractals. BYTE Magazine, 15(6):382–??, June 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:RPK

[Ano91] Anonymous. Up from Rosie: Professor Kenner examines a new book of essays about fractals and chaos. BYTE Magazine, 16(5): 386–??, May 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:BRB

[Ano93a] Anonymous. Book review: Transition to Chaos: The Orbit Diagram and the Mandelbrot Set, by Robert L. Devaney and Hassan Aref. SIAM Review, 35(2):308–??, June 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1993:MZW

[Ano93b] Anonymous. Mandelbrot and Zewail win 1993 Wolf Prizes in Physics, Chemistry. Physics Today, 46(6):109–??, June 1993. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://link.aip.org/link/phtoad/v46/i6/p109/s1.

Anonymous:1993:TCO

[Ano93c] Anonymous. Transition to chaos: The orbit diagram and the Mandelbrot set, with Robert L. Devaney and Hassan Aref. SIAM Review, 35(2):308–??, June 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:SHBb

[Ano95] Anonymous, editor. Symposium in Honor of Benoît Mandelbrot. World Scientific Publishing, Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1995. ISSN 0218-348X. i–xii + 629–914 pp. LCCN ???? Held in Curaçao, February 2–4, 1995, Fractals 3 (1995), no. 4.

Anonymous:1998:FSF

[Ano98a] Anonymous. Book review: Fractals and scaling in finance: Discontinuity, concentration, risk: Selecta Volume E. By Benoît B. Mandelbrot. Springer, Berlin. (1997). 551 pages. $39.95, DM 84.00, öS 613.20, sFr 76.50, GBP 32.50. Computers and Mathematics with Applications, 35(5):135, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0898122198907453.

Anonymous:1998:NPF

[Ano98b] Anonymous. Net pics: Fractal fever. Science, 280(5363):491, April 24, 1998. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.sciencemag.org/content/
280/5363/491.3.full.

Anonymous:1999:BRMq

[Ano99] Anonymous. Book review: Multifractals and 1/f noise: Wild self-affinity in physics (1963–1976): By Benoît B. Mandelbrot. Springer-Verlag, New York. (1999). 442 pages. $42.95, DM 89.00, öS 650.00, sFr 81.00, GBP 34.00. Computers and Mathematics with Applications, 38(11–12):283, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0898122199911951.

Anonymous:2003:BRFb

[Ano03a] Anonymous. Book review: Fractals, graphics. & mathematics education: Edited by Michael L. Fame & Benoît B. Mandelbrot. The Mathematical Association of America, Washington, D.C. (2002). 206 pages. $39.95. Computers and Mathematics with Applications, 45(10–11):1776, May/June 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0898122103801360.

Anonymous:2003:BRGc

[Ano03b] Anonymous. Book review: Gaussian self-affinity and fractals: By B. B. Mandelbrot. Springer, New York. (2002). 654 pages. $54.95. Computers and Mathematics with Applications, 46(2–3):505, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0898122103900472.

Anonymous:2005:BBF

[Ano05] Anonymous. Buchbesprechung: Fraktale und Finanzen. Märkte zwischen Risiko und Ruin. Von Benoît B. Mandelbrot und Richard L. Hudson. RiskNews, 2(4):56–57, August 2005. CODEN ???? ISSN 1612-8931 (print), 1616-0045 (electronic). URL http:

//onlinelibrary.wiley.com/doi/10.1002/risk.200590086/
abstract.

Anonymous:2006:FF

[Ano06] Anonymous. Fractal feat. Science, 312(5781):1712, June 23, 2006. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/312/
5781/1712.12.full.pdf.

Anonymous:2010:BNF

[Ano10a] Anonymous. BBC news: ‘fractal’ mathematician Benoît Mandelbrot dies aged 85. BBC Online, ??(??):??, October 17, 2010.

CODEN ???? ISSN ???? URL http://www.bbc.co.uk/news/
world-europe-11560101.

Anonymous:2010:BMF

[Ano10b] Anonymous. Benoît Mandelbrot, fractals pioneer, dies. United Press International, ??(??):??, October 16, 2010. CODEN ????

ISSN ???? URL http://www.upi.com/Science_News/2010/10/
16/Benoit-Mandelbrot-fractals-pioneer-dies/UPI-11551287266964/

.

Anonymous:2010:BMO

[Ano10c] Anonymous. Benoît Mandelbrot’s obituary. The Economist, ?? (??):??, October 21, 2010. CODEN EONOEH. ISSN 0013-0613 (print), 1476-8860 (electronic). URL http://www.economist.

com/node/17305197.

Anonymous:2010:MFF

[Ano10d] Anonymous. Mandelbrot, father of fractal geometry, dies. The Gazette, ??(??):??, October 16, 2010. CODEN ???? ISSN

???? URL http://www.upi.com/Science_News/2010/10/16/
Benoit-Mandelbrot-fractals-pioneer-dies/UPI-11551287266964/

.

Anonymous:2010:SRH

[Ano10e] Anonymous. Sarkozy rend hommage à Mandelbrot. (French) [[President] Sarkozy pays homage to Mandelbrot]. Le Figaro, ??(??):??, October 17, 2010. CODEN ???? ISSN

????  URL http://www.lefigaro.fr/flash-actu/2010/
10/16/97001-20101016FILWWW00611-sarkozy-rend-hommage-
a-mandelbrot.php.

Alfonseca:1997:SRF

[AO97] M. Alfonseca and A. Ortega. A study of the representation of fractal curves by L systems and their equivalences. IBM Journal of Research and Development, 41(6):727–736, November 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.almaden.ibm.com/journal/rd/416/
alfonseca.html.

Alfonseca:2001:DFD

[AO01] M. Alfonseca and A. Ortega. Determination of fractal dimensions from equivalent L systems. IBM Journal of Research and Development, 45(6):797–805, ???? 2001. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://

www.research.ibm.com/journal/rd/456/alfonseca.html;
http://www.research.ibm.com/journal/rd/456/alfonseca.
pdf.

Ashish:2014:JSM

[ARC14] Ashish, Mamta Rani, and Renu Chugh. Julia sets and Mandelbrot sets in Noor orbit. Applied Mathematics and Computation, 228(??):615–631, February 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300313012460.

Aref:1993:RBC

[Are93] Hassan Aref. Review: Transition to Chaos: The Orbit Diagram and the Mandelbrot Set, with Robert L. Devaney. SIAM Review, 35(2):308–309, June 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://link.aip.org/link/siread/v35/i2/p308/s1;
http://www.jstor.org/stable/2133149.

Alvarez-Ramirez:2003:ZMS

[ARMG03] Jose Alvarez-Ramirez, Monica Meraz, and Gustavo Gallegos. Zipf–Mandelbrot scaling law for world track records. Physica A, 328(3–4):545–560, October 15, 2003. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037843710300579X.

Allouche:1992:RRS

[AS92] J.-P. Allouche and J. Shallit. The ring of k-regular sequences. Theoretical Computer Science, 98(2):163–197, May 18, 1992. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Achdou:2007:TBC

[AST07] Yves Achdou, Christophe Sabot, and Nicoletta Tchou. Transparent boundary conditions for the Helmholtz equation in some ramified domains with a fractal boundary. Journal of Computational Physics, 220(2):712–739, January 10, 2007. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999106002567.

Atela:1993:MSA

[Ate93] Pau Atela. The Mandelbrot set and σ-automorphisms of quotients of the shift. Transactions of the American Mathematical Society, 335(2):683–703, February 1993. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic). URL http://

www.jstor.org/stable/2154400.

Ausloos:2014:ZMP

[Aus14]  Marcel Ausloos.  Zipf–Mandelbrot–Pareto model for co-

authorship popularity. Scientometrics, 101(3):1565–1586, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/s11192-014-1302-y.

Allouche:1997:ADS

[AvHP^{+}97] J.-P. Allouche, F. von Haeseler, H.-O. Peitgen, A. Petersen, and G. Skordev. Automaticity of double sequences generated by onedimensional linear cellular automata. Theoretical Computer Science, 188(1–2):195–209, November 30, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:/

/www.elsevier.com/cgi-bin/cas/tree/store/tcs/cas_sub/
browse/browse.cgi?year=1997&volume=188&issue=1-2&aid=
2498.

Barnsley:1978:APE

[BA78] Michael F. Barnsley and Jacques G. Aguilar. On the approximation of potential-energy functions for two-center systems. International Journal of Quantum Chemistry, 13(5):641–677, May 1978. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Baines:1997:FDF

[Bai97] Peter G. Baines. Fluid dynamics: a fractal world of cloistered waves. Nature, 388(6642):518–519, August 7, 1997. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v388/
n6642/full/388518a0.html.

Bak:2000:BRB

[Bak00] Per Bak. Book review: Multifractals and 1/f noise, by Benoît B. Mandelbrot. Complexity, 5(3):46–47, January/February 2000. CODEN COMPFS. ISSN ????

Ball:2010:MWF

[Bal10] Philip Ball. Material witness: Fractal standstill. Nature Materials, 9(12):964, November 23, 2010. CODEN NMAACR. ISSN 1476-1122 (print), 1476-4660 (electronic). URL http://www.

nature.com/nmat/journal/v9/n12/full/nmat2910.html.

Barnsley:1971:SRF

[Bar71] Michael Barnsley. Sum rule functions. Journal of Mathematical Physics, 12(6):957–964, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v12/i6/p957_s1.

Barnsley:1973:BPM

[Bar73] Michael Barnsley. The bounding properties of the multipoint Padé approximant to a series of Stieltjes on the real line. Journal of Mathematical Physics, 14(3):299–313, March 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i3/
p299_s1.

Barnsley:1975:CTP

[Bar75] M. F. Barnsley. Correction terms for Padé approximants. Journal of Mathematical Physics, 16(4):918–928, April 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i4/
p918_s1.

Barnsley:1976:PAB

[Bar76] M. F. Barnsley. Padé approximant bounds for the difference of two series of Stieltjes. Journal of Mathematical Physics, 17(4): 559–565, April 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i4/p559_s1.

Barcellos:1984:FGM

[Bar84] Anthony Barcellos. The fractal geometry of Mandelbrot. College Mathematics Journal, 15(2):98–114, March 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/00494925.1984.11972758.

Barcellos:1985:BM

[Bar85] Anthony Barcellos. Benoît Mandelbrot. In Albers and Alexanderson [AA85], page ?? ISBN 0-8176-3191-7, 3-7643-3191-7. LCCN QA28 .M37 1985.

Barnsley:1988:FE

[Bar88a] Michael Barnsley. Fractals Everywhere. Academic Press, New York, NY, USA, 1988. ISBN 0-12-079062-9. xii + 394 pp. LCCN QA614.86 .B37 1988.

Barnsley:1988:FMR

[Bar88b] Michael F. Barnsley. Fractal modelling of real world images. In Peitgen and Saupe [PS88a], chapter 5, pages 219–242. ISBN 0-387-96608-0 (New York), 1-4612-8349-3, 3-540-96608-0 (Berlin). LCCN QA614.86 .S35 1988.

Barnsley:1993:FMR

[Bar93a] Michael F. Barnsley. Fractal modelling of real world images. In Fractals Everywhere [Bar93b], pages 219–239. ISBN 0-12-079061-

0. LCCN ????

Barnsley:1993:FE

[Bar93b] Michael F. Barnsley. Fractals Everywhere. Academic Press, Boston, MA, USA, second edition, 1993. ISBN 0-12-079061-0. xiv + 531 pp. LCCN ????

Barcellos:2008:BM

[Bar08] Anthony Barcellos. Benoît Mandelbrot. In Albers and Alexanderson [AA08], pages 213–234. ISBN 1-56881-340-6. LCCN QA28 .M37 2008.

Bach:1988:AMM

[BATL88] P. Bach, M. Amanieu, Lam Hoai T., and G. Lasserre. Application of mandelbrot’s model of abundance distribution to the estimation of captures in the thau lagoon France. Journal du Conseil International pour l’Exploration de la Mer, 44(3):235–246, ???? 1988. ISSN 0020-6466.

Baumann:1996:MTP

[Bau96] Gerd Baumann. Mathematica in Theoretical Physics: selected examples from classical mechanics to fractals. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94424-9. xi + 384 pp. LCCN QC20.7.E4B3813 1996. US$59.00. Revised, translated, expanded and updated edition of [Kau92].

Baumann:2005:MTP

[Bau05] Gerd Baumann. Mathematica in theoretical physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2005. ISBN 0-387-01674-0. xvi + 544 pp. LCCN QC20.7.E4 B3813 2005.

Bawden:2011:BMS

[Baw11] David Bawden. Benoît Mandelbrot and the self-similarity of information. Journal of Documentation, 67(2):??, ???? 2011. CO-DEN JDOCAS. ISSN 0022-0418 (print), 1758-7379 (electronic).

URL http://www.emeraldinsight.com/journals.htm?issn=
0022-0418&volume=67&issue=2&articleid=1911717.

Barnsley:1976:BBC

[BB76] M. F. Barnsley and George A. Baker, Jr. Bivariational bounds in a complex Hilbert space, and correction terms for padé approximants. Journal of Mathematical Physics, 17(6):1019–1027, June 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v17/i6/p1019_s1.

Barnsley:1979:CMB

[BB79] M. Barnsley and D. Bessis. Constructive methods based on analytic characterizations and their application to nonlinear elliptic and parabolic differential equations. Journal of Mathematical Physics, 20(6):1135–1145, June 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v20/i6/p1135_s1.

Brady:1984:FGC

[BB84] R. M. Brady and R. C. Ball. Fractal growth of copper electrodeposits. Nature, 309(5965):225–229, May 17, 1984. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v309/
n5965/pdf/309225a0.pdf.

Badler:1990:IRI

[BB90] Norman I. Badler and Welton Becket. Imperfection for realistic image synthesis. Journal of Visualization and Computer Animation, 1(1):26–32, August 1990. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Barlow:1999:BMH

[BB99] Martin T. Barlow and Richard F. Bass. Brownian motion and harmonic analysis on Sierpinski carpets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(4):673– 744, August 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barnsley:2010:FT

[BB10] Michael Barnsley and Louisa Barnsley. Fractal transformations. In Lesmoir-Gordon [LG10b], chapter 4, pages 58–73. ISBN 1-84996-485-8. LCCN QA614.86 .C65 2010. URL http://www.

springerlink.com/content/978-1-84996-485-2. With contri-

butions by Ian Stewart, Sir Arthur C. Clarke, Benoît B. Mandelbrot, Michael and Louisa Barnsley, Will Rood, Gary Flake, and David Pennock.

Buzio:2003:CMF

[BBB^{+}03] Renato Buzio, Corrado Boragno, Fabio Biscarini, Francesco Buatier De Mongeot, and Ugo Valbusa. The contact mechanics of fractal surfaces. Nature Materials, 2(4):233–236, March 9, 2003. CODEN NMAACR. ISSN 1476-1122 (print), 1476-4660 (elec-

tronic). URL http://www.nature.com/nmat/journal/v2/n4/
full/nmat855.html.

Barral:2010:QIE

[BBB^{+}10] J. Barral, J. Berestycki, J. Bertoin, et al. Quelques interactions entre analyse, probabilités et fractals. (French) [Some interactions between analysis, probability, and fractals]. Société Mathématique de France, Paris, France, 2010. x + 243 pp. LCCN ???? Introduction by Stéphane Jaffard.

Byrnes:1992:PSM

[BBBH92] J. S. Byrnes, Jennifer L. Byrnes, Karl Berry, and Kathryn A. Hargreaves, editors. Probabilistic and Stochastic Methods in Analysis, with Applications: Proceedings of the NATO Advanced Study Institute, volume 372 of NATO ASI series C: Mathematical and physical sciences. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-1804-8. LCCN ????

Barnsley:1979:DMP

[BBM79] M. Barnsley, D. Bessis, and P. Moussa. The Diophantine moment problem and the analytic structure in the activity of the ferromagnetic Ising model. Journal of Mathematical Physics, 20(4): 535–546, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i4/p535_s1.

Barnsley:1980:CSK

[BC80] M. Barnsley and H. Cornille. On a class of solutions of the Krook– Tjon–Wu model of the Boltzmann equation. Journal of Mathematical Physics, 21(5):1176–1193, May 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v21/i5/p1176_s1.

Brandt:1991:ACH

[BCM91] Jonathan Brandt, Carlos Cabrelli, and Ursula Molter. An algorithm for the computation of the Hutchinson distance. Information Processing Letters, 40(2):113–117, October 25, 1991. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigendum [BCM92].

Brandt:1992:CAC

[BCM92] Jonathan Brandt, Carlos Cabrelli, and Ursula Molter. Corrigendum: “An algorithm for the computation of the Hutchinson distance” [Inform. Process. Lett. 40(2), 25 October 1991, pp. 113–117]. Information Processing Letters, 41(6):347, April 17, 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [BCM91].

Barral:2003:MMM

[BCM03] J. Barral, M.-O. Coppens, and B. B. Mandelbrot. Multiperiodic multifractal martingale measures. Journal de Mathématiques Pures et Appliquées, 82(12):1555–1589, ???? 2003. CODEN JM-PAAM. ISSN 0021-7824 (print), 1776-3371 (electronic).

Barnsley:1986:CDF

[BD86] Michael F. Barnsley and Stephen G. Demko, editors. Chaotic Dynamics and Fractals, volume 2 of Notes and reports in mathematics in science and engineering. Academic Press, New York, NY, USA, 1 edition, 1986. ISBN 0-12-079060-2. xi + 292 pp. LCCN QA843 .C471 1986.

Bode:1996:PVM

[BDLS96] Arndt Bode, Jack Dongarra, T. Ludwig, and V. Sunderam, editors. Parallel virtual machine, EuroPVM ’96: third European PVM conference, Munich, Germany, October 7–9, 1996: proceedings, volume 1156 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN ???? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Barnsley:1988:SFI

[BDM^{+}88] Michael F. Barnsley, Robert L. Devaney, Benoît B. Mandelbrot, Heinz-Otto Peitgen, Dietmar Saupe, and Richard F. Voss. The science of fractal images. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-96608-0. xiv + 312 pp. LCCN ???? With contributions by Yuval Fisher and Michael McGuire. Based on notes for the course Fractals — introduction, basics and perspectives, given as part of the SIGGRAPH ’87, Anaheim, California course program.

Bani-Eqbal:1995:ESF

[BE95a] Behnam Bani-Eqbal. Enhancing the speed of fractal image compression. Optical Engineering, 34(6):1705–1710, June 1995. CODEN OPEGAR. ISSN 0091-3286 (print), 1560-2303

(electronic). URL ftp://links.uwaterloo.ca/pub/Fractals/
Papers/bani95.ps.gz.

Bani-Eqbal:1995:SFI

[BE95b] Behnam Bani-Eqbal. Speeding up fractal image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2418(??):67–74, February 1995. CODEN PSISDG. ISSN ????

Barnsley:1989:HVF

[BEHM89] M. F. Barnsley, J. Elton, D. Hardin, and P. Massopust. Hidden variable fractal interpolation functions. SIAM Journal on Mathematical Analysis, 20(5):1218–1242, September 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Berry:1986:FA

[Ber86] Michael Berry. Fractal attractions. Nature, 323(6089):590, October 16, 1986. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v323/n6089/pdf/323590a0.pdf.

Bohman:1995:HIC

[BF95] J. Bohman and C.-E. Fröberg. Heuristic investigation of chaotic mapping producing fractal objects. BIT Numerical Mathematics, 35(4):609–615, December 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://

www.mai.liu.se/BIT/contents/bit35.html;   http://www.
springerlink.com/openurl.asp?genre=article&issn=0006-
3835&volume=35&issue=4&spage=609.

Barnsley:2012:GBM

[BF12a] Michael F. Barnsley and Michael Frame. Glimpses of Benoît Mandelbrot (1924–2010). Notices of the American Mathematical Society, 59(8):1056–1063, September 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/201208/rtx120801056p.pdf.

Barnsley:2012:IBB

[BF12b] Michael F. Barnsley and Michael Frame. The influence of Benoît B. Mandelbrot on mathematics. Notices of the American Mathematical Society, 59(9):1208–1221, October 2012. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL http://www.ams.org/notices/201209/rtx120901208p.

pdf. Other contributors are Roger Howe, Ian Stewart, David Mumford, Hillel Furstenberg, Kenneth Falconer, Bruce J. West, Marc-Olivier Coppens, Nathan Cohen, Stéphane Jaffard, and Sir Michael Berry.

Bush:1967:HMP

[BGL67] Robert R. Bush, Eugene Galanter, and R. Duncan Luce, editors. Handbook of Mathematical Psychology. Volume II, Chapters 9– 14. Wiley, New York, NY, USA, second edition, 1967. vii + 606 pp. LCCN ????

Balkin:1994:CEC

[BGR94] Sandy D. Balkin, Elizabeth L. Golebiewski, and Clifford A. Reiter. Chaos and elliptic curves. Computers and Graphics, 18 (1):113–117, January–February 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Barnsley:1989:MSW

[BH89] M. F. Barnsley and D. P. Hardin. A Mandelbrot set whose boundary is piecewise smooth. Transactions of the American Mathematical Society, 315(2):641–659, October 1989. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic). URL http://

www.jstor.org/stable/2001299.

Barnsley:2011:HTF

[BHI11] Michael F. Barnsley, Brendan Harding, and Konstantin Igudesman. How to transform and filter images using iterated function systems. SIAM Journal on Imaging Sciences, 4(4):1001–1028, ???? 2011. CODEN SJISBI. ISSN 1936-4954. URL http://

epubs.siam.org/siims/resource/1/sjisbi/v4/i4/p1001_s1.

Bialy:1989:FGH

[Bia89] Harvey Bialy. The fractal geometry of the heart. Nature Biotechnology, 7(12):1219, December 1989. CODEN NABIF9. ISSN 1087-0156 (print), 1546-1696 (electronic). URL http://www.

nature.com/nbt/journal/v7/n12/full/nbt1289-1219.html.

Bini:2024:NCR

[Bin24] Dario A. Bini. Numerical computation of the roots of Mandelbrot polynomials: an experimental analysis. Electronic Transactions on Numerical Analysis (ETNA), 61:1– 27, 2024. CODEN ???? ISSN 1068-9613 (print), 1097-

4067 (electronic). URL https://etna.math.kent.edu/vol.61.
2024/pp1-27.dir/pp1-27.pdf;   https://etna.math.kent.
edu/volumes/2021-2030/vol61/abstract.php?vol=61&pages=
1-27.

Bisoi:1995:NRM

[Bis95] A. K. Bisoi. Newton Raphson method, scaling at fractal boundaries and MATHEMATICA. Mathematical and Computer Modelling, 21(10):91–102, ???? 1995. CODEN MCMOEG. ISSN 0895-7177 (print), 1872-9479 (electronic).

Berlinkov:2019:PMP

[BJ19] Artemi Berlinkov and Esa Järvenpää. Porosities of Mandelbrot percolation. Journal of Theoretical Probability, 32(2):608–632, June 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).

Barnsley:1988:HCI

[BJM^{+}88] Michael F. Barnsley, Arnaud Jacquin, François Malassenet, Laurie Reuter, and Alan D. Sloan. Harnessing chaos for image synthesis. Computer Graphics, 22(4):131–140, August 1988. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569

(electronic). URL http://www.acm.org:80/pubs/citations/
proceedings/graph/54852/p131-barnsley/.

Barral:2010:CCM

[BJM10a] Julien Barral, Xiong Jin, and Benoît Mandelbrot. Convergence of complex multiplicative cascades. Annals of applied probability, 20(4):1219–1252, ???? 2010. CODEN ???? ISSN 1050-5164.

Barral:2010:UCC

[BJM10b] Julien Barral, Xiong Jin, and Benoît Mandelbrot. Uniform convergence for complex [0, 1]-martingales. Annals of applied probability, 20(4):1205–1218, ???? 2010. CODEN ???? ISSN 1050-5164.

Balamash:2002:PMM

[BK02] Abdullah Balamash and Marwan Krunz. A parsimonious multifractal model for WWW traffic. Lecture Notes in Computer Science, 2376(??):1–??, ???? 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.de/link/service/series/0558/bibs/2376/
23760001.htm;   http://link.springer.de/link/service/
series/0558/papers/2376/23760001.pdf.

Balamash:2003:MWR

[BK03] Abdullah Balamash and Marwan Krunz. Modeling Web requests: a multifractal approach. Computer Networks (Amsterdam, Netherlands: 1999), 43(2):211–226, October 7, 2003. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Berry:1980:WMF

[BL80] M. V. Berry and Z. V. Lewis. On the Weierstrass–Mandelbrot fractal function. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 370(1743):459–484, April 24, 1980. CODEN PRLAAZ. ISSN 0080-4630. URL http:

//www.jstor.org/stable/2397195.

Bossomaier:1993:PCH

[BL93] Terry Bossomaier and Adrian Loeff. Parallel computation of the Hausdorff distance between images. Parallel Computing, 19(10): 1129–1140, October 1993. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Berger:1963:NME

[BM63] J. M. Berger and Benoît Mandelbrot. A new model for error clustering in telephone circuits. IBM Journal of Research and Development, 7(3):224–236, ???? 1963. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5392305.

Brivio:1993:FMD

[BM93] P. A. Brivio and D. Marini. A fractal method for digital elevation model construction and its application to a mountain region. Computer Graphics Forum, 12(5):297–309, December 1993. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Blumenfeld:1997:LDM

[BM97] Raphael Blumenfeld and Benoît B. Mandelbrot. Lévy dusts, Mittag-Leffler statistics, mass fractal lacunarity, and perceived dimension. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 56(1, part A):112–118, ???? 1997. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-

2376 (electronic). URL http://link.aps.org/abstract/PRE/
v56/e112.

Barral:1999:MPC

[BM99] Julien Barral and Benoît B. Mandelbrot. Multifractal products of certain pulses. Technical report, UM2—Département des sciences mathématiques, Montpellier, France, 1999. 33 pp.

Barral:2002:MPC

[BM02] Julien Barral and Benoît B. Mandelbrot. Multifractal products of cylindrical pulses. Probability theory and related fields, 124(3): 409–430, ???? 2002. CODEN PTRFEU. ISSN 0178-8051 (print),

1432-2064 (electronic). URL http://www.springerlink.com/
content/ggbdvw9n6bh590up/.

Barral:2004:PIM

[BM04a] J. Barral and Benoît B. Mandelbrot. Part II: Multifractals: Introduction to infinite products of independent random functions (random multiplicative multifractal measures, part I). In Lapidus and van Frankenhuijsen [LvF04b], page ?? ISBN 0-8218-3638-2 (part 2), 0-8218-3292-1 (set). LCCN QA325 .F73 2004. Multifractals, probability and statistical mechanics, applications.

Barral:2004:IIP

[BM04b] Julien Barral and Benoît B. Mandelbrot. Introduction to infinite products of independent random functions (Random multiplicative multifractal measures. I). In Fractal geometry and applications: a jubilee of Benoît Mandelbrot, Part 2, volume 72 of Proc. Sympos. Pure Math., pages 3–16. American Mathematical Society, Providence, RI, USA, 2004.

Barral:2004:NDM

[BM04c] Julien Barral and Benoît B. Mandelbrot. Non-degeneracy, moments, dimension, and multifractal analysis for random multiplicative measures (Random multiplicative multifractal measures. II). In Lapidus and van Frankenhuijsen [LvF04b], pages 17–52. ISBN 0-8218-3638-2 (part 2), 0-8218-3292-1 (set). LCCN QA325 .F73 2004. Multifractals, probability and statistical mechanics, applications.

Barral:2009:FMP

[BM09] Julien Barral and Benoît Mandelbrot. Fractional multiplicative processes. Annales de l’Institut Henri Poincaré, Probabilités et Statistiques, 45(4):1116–1129, ???? 2009. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic).

Bouville:1985:BER

[Bou85] Christian Bouville. Bounding ellipsoids for ray-fractal intersection. Computer Graphics, 19(3):45–52, July 1985. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Bourke:2006:IAF

[Bou06] Paul Bourke. An introduction to the Apollonian fractal. Computers and Graphics, 30(1):134–136, February 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0097849305002189.

Bouboulis:2009:COM

[Bou09a] P. Bouboulis. Construction of orthogonal multi-wavelets using generalized-affine fractal interpolation functions. IMA Journal of Applied Mathematics, 74(6):904–933, December 2009. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634

(electronic). URL http://imamat.oxfordjournals.org/cgi/
content/abstract/74/6/904;   http://imamat.oxfordjournals.
org/cgi/reprint/74/6/904.

Bourke:2009:ESL

[Bou09b] Paul Bourke. Evaluating Second Life for the collaborative exploration of 3D fractals. Computers and Graphics, 33(1):113– 117, February 2009. CODEN COGRD2. ISSN 0097-8493 (print),

1873-7684 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0097849308000964.

Bouchaud:2011:BMP

[Bou11] Jean-Philippe Bouchaud. Benoît Mandelbrot: a personal tribute. Quantitative Finance, 11(2):161, ???? 2011. CODEN ???? ISSN 1469-7688 (print), 1469-7696 (electronic).

Boyd:1983:RBF

[Boy83] James N. Boyd. Review: The Fractal Geometry of Nature, (L) by Benoît B. Mandelbrot. The Mathematics Teacher, 76(4):288, April 1983. CODEN ???? ISSN 0025-5769. URL http://www.

jstor.org/stable/27963479.

Boroda:1984:ZML

[BP84] M. G. Boroda and A. A. Polikarpov. The Zipf–Mandelbrot law and units of different levels of text organization. Tartu Riikl. Ül. Toimetised, ??(689):35–60, ???? 1984. CODEN ???? ISSN ????

Bishop:2017:FPA

[BP17] Christopher J. Bishop and Y. (Yuval) Peres. Fractals in Probability and Analysis, volume 162 of Cambridge studies in advanced mathematics. Cambridge University Press, Cambridge, UK, 2017. ISBN 1-107-13411-0 (hardcover), 1-316-86692-0, 1-316-46023-1 (e-book). ix + 402 pp. LCCN QA614.86 .B57 2017.

Barral:2009:DMC

[BPW09] Julien Barral, Jacques Peyrière, and Zhi-Ying Wen. Dynamics of Mandelbrot cascades. Probability theory and related fields, 144 (3–4):615–631, July 2009. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s00440-008-0156-8.

Barnsley:1974:DVP

[BR74a] M. F. Barnsley and P. D. Robinson. Dual variational principles and Padé-type approximants. Journal of the Institute of Mathematics and its Applications, 14(??):229–249, ???? 1974. CODEN JMTAA8. ISSN 0020-2932.

Barnsley:1974:PAB

[BR74b] M. F. Barnsley and P. D. Robinson. Padé-approximant bounds and approximate solutions for Kirkwood–Riseman integral equations. Journal of the Institute of Mathematics and its Applications, 14(??):251–265, ???? 1974. CODEN JMTAA8. ISSN 0020-2932.

Barnsley:1975:VBT

[BR75] Michael F. Barnsley and Peter D. Robinson. Variational bounds on transition probabilities. International Journal of Quantum Chemistry, 9(3):479–487, May 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Barnsley:1976:BBS

[BR76] M. F. Barnsley and P. D. Robinson. Bivariational bounds on solutions of two-point boundary-value problems. Journal of Mathematical Analysis and Applications, 56(??):172–184, ???? 1976. CODEN JMANAK. ISSN 0022-247x (print), 1096-0813 (electronic).

Barnsley:1977:BBN

[BR77] M. F. Barnsley and P. D. Robinson. Bivariational bounds for non-linear problems. Journal of the Institute of Mathematics and its Applications, 20(4):485–504, ???? 1977. CODEN JMTAA8. ISSN 0020-2932.

Barnsley:1978:RAB

[BR78] M. F. Barnsley and P. D. Robinson. Rational approximant bounds for a class of two-variable Stieltjes functions. SIAM Journal on Mathematical Analysis, 9(2):272–290, April 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bradley:1994:CSS

[Bra94] James Bradley. A complete L-system specification for generating an exact self-affine growth-decay function with a random walk scaling property. The Computer Journal, 37(1):58–66, ???? 1994.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/
Volume_37/Issue_01/Vol37_01.body.html#AbstractBradley.

Bresson:1979:RBF

[Bre79] Denis Bresson. Review: Fractals: Form, Chance and Dimension, by Benoît B. Mandelbrot. Leonardo (Oxford, England), 12(3): 248–249, ???? 1979. CODEN LEONDP. ISSN 0024-094X (print),

1530-9282 (electronic). URL http://www.jstor.org/stable/
1574228.

Brickmann:1985:BRB

[Bri85] J. Brickmann. Book review: B. Mandelbrot: The Fractal Geometry of Nature, Freeman and Co., San Francisco 1982. 460 Seiten, Preis: £22,75. Berichte der Bunsen-Gesellschaft fuer Physikalische Chemie, 1963–1990, 89(2):209, February 1985. CODEN

BBPCAX. ISSN ???? URL http://onlinelibrary.wiley.
com/doi/10.1002/bbpc.19850890223/abstract.

Bridger:1988:LMS

[Bri88] Mark Bridger. Looking at the Mandelbrot set. College Mathematics Journal, 19(4):353–363, September 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic).  URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973139.

Browne:2007:IF

[Bro07] Cameron Browne. Impossible fractals. Computers and Graphics, 31(4):659–667, August 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0097849307000799.

Brune:1994:MTF

[BRRK94] Harald Brune, Christoph Romainczyk, Holger Röder, and Klaus Kern. Mechanism of the transition from fractal to dendritic growth of surface aggregates. Nature, 369(6480):469–471, June 9, 1994. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v369/n6480/pdf/369469a0.pdf.

Barnsley:1988:BWC

[BS88] Michael F. Barnsley and Alan D. Sloan. A better way to compress images. BYTE Magazine, 13(??):215–223, January 1988. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bouras:1992:SDC

[BSV92] A. Bouras, B. Shariat, and D. Vandorpe. A simple description of complex curves. Computer Graphics Forum, 11(3):C425–C433, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barnsley:2002:FM

[BSV02] Michael F. Barnsley, Dietmar Saupe, and Edward R. Vrscay, editors. Fractals in Multimedia, volume 132 of The IMA Volumes in Mathematics and its Application. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 1-4419-3037-X. ISSN 0940-6573. LCCN QA273.A1-274.9; QA274-274.9.

Barnsley:1981:BMN

[BT81] M. Barnsley and G. Turchetti. Bounding methods in nuclear physics. Computer Physics Communications, 22(2–3):325– 343, April 1981. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465581900680.

Burrough:1981:FDL

[Bur81] P. A. Burrough. Fractal dimensions of landscapes and other environmental data. Nature, 294(5838):240–242, November 19, 1981. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v294/
n5838/pdf/294240a0.pdf.

Burn:1984:RBF

[Bur84] R. P. Burn. Review: The Fractal Geometry of Nature, by Benoît B. Mandelbrot. Mathematical Gazette, 68(443):71–72, March 1984. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL http://www.jstor.org/stable/3615422.

Burns:2002:PEA

[Bur02] Anne M. Burns. Plotting the escape: An animation of parabolic bifurcations in the Mandelbrot set. Mathematics Magazine, 75 (2):104–116, April 2002. CODEN MAMGA8. ISSN 0025-570X.

URL http://www.jstor.org/stable/3219143.

Barnsley:2011:EPL

[BV11] Michael Barnsley and Andrew Vince. The eigenvalue problem for linear and affine iterated function systems. Linear Algebra and Its Applications, 435(12):3124–3138, December 15, 2011. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Blackmore:1996:GFD

[BZ96] Denis Blackmore and Jack G. Zhou. A general fractal distribution function for rough surface profiles. SIAM Journal on Applied Mathematics, 56(6):1694–1719, December 1996. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL

http://epubs.siam.org/sam-bin/dbq/article/28312.

Cahn:1989:FDF

[Cah89] Robert W. Cahn. Fractal dimension and fracture. Nature, 338 (6212):201–202, March 16, 1989. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v338/n6212/pdf/338201a0.pdf.

Campbell:1987:TST

[Cam87] Philip Campbell. Is there such a thing as fractal music? Nature, 325(6107):766, February 26, 1987. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v325/n6107/pdf/325766a0.pdf.

Camp:1996:RBF

[Cam96] Dane R. Camp. Review: The Fractal Geometry of Nature, by Benoît B. Mandelbrot. The Mathematics Teacher, 89(3):256, March 1996. CODEN ???? ISSN 0025-5769. URL http://

www.jstor.org/stable/27969730.

Camp:2000:BME

[Cam00] Dane R. Camp. Benoît Mandelbrot: The Euclid of fractal geometry. The Mathematics Teacher, 93(8):708–712, November 2000. CODEN ???? ISSN 0025-5769. URL http://www.jstor.org/

stable/27971555.

Cannon:1984:RBF

[Can84] J. W. Cannon. Review: The Fractal Geometry of Nature, by Benoît B. Mandelbrot. American Mathematical Monthly, 91(9): 594–598, November 1984. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2323761.

Carlson:1996:PDR

[Car96] Paul W. Carlson. Pseudo-3-D rendering methods for fractals in the complex plane. Computers and Graphics, 20(5):751–758, September–October 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http://www.elsevier.

com/cgi-bin/cas/tree/store/cag/cas_sub/browse/browse.
cgi?year=1996&volume=20&issue=5&aid=9600048.

Carlson:1999:TAO

[Car99a] Paul W. Carlson. Two artistic orbit trap rendering methods for Newton M-set fractals. Computers and Graphics, 23(6): 925–931, December 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http://www.elsevier.nl/

gej-ng/10/13/20/24/34/47/abstract.html;   http://www.
elsevier.nl/gej-ng/10/13/20/24/35/47/article.pdf.

Cartwright:1999:NMF

[Car99b] Julyan H. E. Cartwright. Newton maps: fractals from Newton’s method for the circle map. Computers and Graphics, 23 (4):607–612, August 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http://www.elsevier.nl/

gej-ng/10/13/20/24/34/40/abstract.html;   http://www.
elsevier.nl/gej-ng/10/13/20/24/34/40/article.pdf.

Casey:1996:UDT

[Cas96] Stephen D. Casey. Using dimension theory to analyze and classify the generation of fractal sets. Computers and Graphics, 20(5):731–749, September–October 1996. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.elsevier.com/cgi-bin/cas/tree/
store/cag/cas_sub/browse/browse.cgi?year=1996&volume=
20&issue=5&aid=9600047.

Cawley:1989:HSF

[Caw89] Robert Cawley. The hard science of fractals. Science, 245 (4920):873, August 25, 1989. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/245/4920/873.1.full.pdf.

Colnort-Bodet:1959:RBL

[CB59] Suzanne Colnort-Bodet. Review: La lecture de l’expérience, ≪Bibl. sc. intern.≫, Études d’Épistémologie génétique by (A.) Jonckheere, (B.) Mandelbrot, (J.) Piaget. Les Études philosophiques, ??(1):85, January 1959. CODEN ???? ISSN ????

URL http://www.jstor.org/stable/20842901.

Chen:1997:DIF

[CB97] Yan Qiu Chen and Guoan Bi. 3-D IFS fractals as realtime graphics model. Computers and Graphics, 21(3):367– 370, May–June 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http://www.elsevier.

com/cgi-bin/cas/tree/store/cag/cas_sub/browse/browse.
cgi?year=1997&volume=21&issue=3&aid=9700014.

Cohen:1957:RBL

[CC58] Marcel Cohen and David Cohen. Review: Logique, langage et théorie de l’information, by L. Apostel, B. Mandelbrot, A. Morf. L’Année sociologique (1940/1948-), Troisième Série, 9(??):508– 509, ???? 1957–1958. CODEN ???? ISSN ???? URL http://

www.jstor.org/stable/27885595.

Chung:2000:GMS

[CCC00] K. W. Chung, H. S. Y. Chan, and N. Chen. General Mandelbrot sets and Julia sets with color symmetry from equivariant mappings of the modular group. Computers and Graphics, 24(6): 911–918, December 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http://www.elsevier.nl/

gej-ng/10/13/20/47/34/34/abstract.html;   http://www.
elsevier.nl/gej-ng/10/13/20/47/34/34/article.pdf.

Chayes:1988:CPM

[CCD88] J. T. Chayes, L. Chayes, and R. Durrett. Connectivity properties of Mandelbrot’s percolation process. Probability theory and related fields, 77(3):307–324, March 1988. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://

link.springer.com/article/10.1007/BF00319291.

Chayes:1991:PTM

[CCGS91] J. T. Chayes, L. Chayes, E. Grannan, and G. Swindle. Phase transitions in Mandelbrot’s percolation process in three dimensions. Probability theory and related fields, 90(3):291–300, ???? 1991. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01193747.

Corana:1991:ECC

[CCRR91] A. Corana, A. Casaleggio, C. Rolando, and S. Ridella. Efficient computation of the correlation dimension from a time series on a LIW computer. Parallel Computing, 17(6–7):809–820, September 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Culik:1993:ECW

[CD93a] K. Culik and S. Dube. Efficient compression of wavelet coefficients for smooth and fractal-like data. Lecture Notes in Computer Science, 665(??):343–??, ???? 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Culik:1993:AAR

[CD93b] Karel Culik II and Simant Dube. Affine automata and related techniques for generation of complex images. Theoretical Computer Science, 116(2):373–398, August 16, 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.com/cgi-bin/cas/tree/
store/tcs/cas_sub/browse/browse.cgi?year=1993&volume=
116&issue=2&aid=1241.

Cheskidov:2007:EDF

[CDP07] Alexey Cheskidov, Charles R. Doering, and Nikola P. Petrov. Energy dissipation in fractal-forced flow. Journal of Mathematical Physics, 48(6):065208, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v48/i6/p065208_s1.

Crutchfield:1986:C

[CFPS86] James P. Crutchfield, J. Doyne Farmer, Norman H. Packard, and Robert S. Shaw. Chaos. Scientific American, 255(6):46–?? (Intl. ed. 38–49), December 1986. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Chakrabarty:2009:FUI

[CGC^{+}09] Rajan K. Chakrabarty, Mark A. Garro, Shammah Chancellor, Christopher Herald, and Hans Moosmüller. FracMAP: a user-interactive package for performing simulation and orientation-specific morphology analysis of fractal-like solid nanoagglomerates. Computer Physics Communications, 180(8):1376– 1381, August 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509000411.

Chacon:2002:MFB

[CGH02] R. Chacón and A. Martínez García-Hoz. Modifying fractal basin boundaries by reshaping periodic terms. Journal of Mathematical Physics, 43(7):3586–3591, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cioczek-Georges:1995:CMA

[CGM95] R. Cioczek-Georges and B. B. Mandelbrot. A class of micropulses and antipersistent fractional Brownian motion. Stochastic Processes and Their Applications, 60(1):1–18, ???? 1995. CODEN STOPB7. ISSN 0304-4149 (print), 1879-209X (electronic).

Cioczek-Georges:1996:AMF

[CGM96] R. Cioczek-Georges and B. B. Mandelbrot. Alternative micropulses and fractional Brownian motion. Stochastic Processes and Their Applications, 64(2):143–152, ???? 1996. CODEN STOPB7. ISSN 0304-4149 (print), 1879-209X (electronic).

Cioczek-Georges:1995:SFS

[CGMST95] Renata Cioczek-Georges, Benoît B. Mandelbrot, Gennady Samorodnitsky, and Murad S. Taqqu. Stable fractal sums of pulses: The cylindrical case. Bernoulli: official journal of the Bernoulli Society for Mathematical Statistics and Probability, 1 (3):201–216, September 1995. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://www.jstor.org/

stable/3318477.

Chakerian:1984:RBF

[Cha84] Don Chakerian. Review: The Fractal Geometry of Nature, by Benoît B. Mandelbrot. College Mathematics Journal, 15(2): 175–177, March 1984. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.jstor.org/stable/
2686529.

Chabert:1990:DSF

[Cha90] Jean-Luc Chabert. Un demi-siècle de fractales: 1870–1920. (French) [A half-century of fractals: 1870–1920]. Historia Mathematica, 17(4):339–365, November 1990. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

URL http://www.sciencedirect.com/science/article/pii/
031508609090026A.

Chang:2011:CGR

[Cha11] Harry M. Chang. Constructing n-gram rules for natural language models through exploring the limitation of the Zipf– Mandelbrot law. Computing: Archiv für Informatik und Numerik, 91(3):241–264, March 2011. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0010-
485X&volume=91&issue=3&spage=241.

Cohen:1953:RBF

[CHR54] Marcel Cohen, André-G. Haudricourt, and Maxime Rodinson. Review: Structure formelle des textes et Communication, dans word, vol 10 by Benoît Mandelbrot; L’ingénieur en tant que stratége: theéories du comportement. Une définition de la cybernétique; applications linguistiques. Extraits de la Revue générale des Sciences, t. LXII, nos 9–10, 1955 by Benoît Mandelbrot. L’Année sociologique (1940/1948-), Troisième Série, 7 (??):501–502, ???? 1953–1954. CODEN ???? ISSN ???? URL

http://www.jstor.org/stable/27885319.

Chu:2011:MSF

[Chu11] Pong P. Chu. Mandelbrot set fractal accelerator. In Embedded SoPC Design with Nios II Processor and VHDL Examples, chapter 21, pages 637–670. Wiley, New York, NY, USA, 2011. ISBN 1-118-00888-X (print), 1-118-14653-0 (online). LCCN TK7895.E42 C48 2011.

Chu:2012:MSF

[Chu12] Pong P. Chu. Mandelbrot set fractal accelerator. In Embedded SoPC Design with Nios II Processor and Verilog Examples, pages 681–714. Wiley, New York, NY, USA, 2012. ISBN 1-118-01103-1 (print), 1-118-30972-3 (e-book). URL http://onlinelibrary.

wiley.com/doi/10.1002/9781118309728.ch22/summary.

Ciarcia:1988:CCCc

[Cia88a] S. Ciarcia. Ciarcia’s circuit cellar. 3 A supercomputer. BYTE Magazine, 13(13):327–339, December 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ciarcia:1988:Sb

[Cia88b] S. Ciarcia. A supercomputer. 2. BYTE Magazine, 13(12):399– 406, November 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cifani:2011:DGM

[CJK11] Simone Cifani, Espen R. Jakobsen, and Kenneth H. Karlsen. The discontinuous Galerkin method for fractal conservation laws. IMA Journal of Numerical Analysis, 31(3):1090–1122, July 2011. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (elec-

tronic). URL http://imajna.oxfordjournals.org/content/
31/3/1090.full.pdf+html.

Culik:1993:PRS

[CK93] Karel Culik II and Jarkko Kari. Parametrized recurrent systems for image generation. Information Processing Letters, 48(6):267– 274, December 20, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Caso:1996:NRF

[CK96] Gregory Caso and C.-C. J. Kuo. New results for fractal/wavelet image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2727(??):536–547, ???? 1996. CODEN PSISDG. ISSN ????

Chen:1987:AZL

[CL87] Ye-Sho Chen and Ferdinand F. Leimkuhler. Analysis of Zipf’s law: An index approach. Information Processing and Management, 23 (??):171–182, ???? 1987. CODEN IPMADK. ISSN 0306-4573 (print), 1873-5371 (electronic).

Clark:1986:FHR

[Cla86] Nigel N. Clark. Fractal harmonics and rugged materials. Nature, 319(6055):625, February 20, 1986. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v319/n6055/pdf/319625a0.pdf.

Clarke:2010:CIF

[Cla10a] Arthur C. Clarke. The colours of infinity — the film script. In Lesmoir-Gordon [LG10b], chapter 9, pages 144–172. ISBN 1-84996-485-8. LCCN QA614.86 .C65 2010. URL http://www.

springerlink.com/content/978-1-84996-485-2. With contri-

butions by Ian Stewart, Sir Arthur C. Clarke, Benoît B. Mandelbrot, Michael and Louisa Barnsley, Will Rood, Gary Flake, and David Pennock.

Clarke:2010:EFU

[Cla10b] Arthur C. Clarke. Exploring the fractal universe. In Lesmoir-Gordon [LG10b], chapter 2, pages 24–37. ISBN 1-84996-485-8.

LCCN QA614.86 .C65 2010. URL http://www.springerlink.
com/content/978-1-84996-485-2. With contributions by Ian

Stewart, Sir Arthur C. Clarke, Benoît B. Mandelbrot, Michael and Louisa Barnsley, Will Rood, Gary Flake, and David Pennock.

Cabrelli:1991:DFA

[CM91] Carlos Cabrelli and Ursula Molter. Density of fuzzy attractors: a step towards the solution of the inverse problem for fractals and other sets. In Byrnes et al. [BBBH92], pages 163–173. ISBN 0-7923-1804-8. LCCN ????

Coppens:1999:ENG

[CM99] Marc-Olivier Coppens and Benoît B. Mandelbrot. Easy and natural generation of multifractals: Multiplying harmonics of periodic functions. In Michel Dekking, editor, Fractals: Theory and applications in engineering, pages 113–122. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999.

Cairns:1977:E

[CMBL77] Stewart S. Cairns, Benoît B. Mandelbrot, I. Barsotti, and Andrew Lenard. Editorial. The Mathematical Intelligencer, 1(3):125–126, September 1977. CODEN MAINDC. ISSN 0343-6993 (print),

1866-7414 (electronic). URL http://www.springerlink.com/
content/n1205964h8854601/.

Chester:2007:MTF

[CMP07] Stuart Chester, Charles Meneveau, and Marc B. Parlange. Modeling turbulent flow over fractal trees with renormalized numerical simulation. Journal of Computational Physics, 225(1):427– 448, July 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999106005997.

Carpinteri:2006:IHP

[CMR06] Alberto Carpinteri, Yiu-Wing Mai, and Robert O. Ritchie, editors. ICF11 2005: honour and plenary lectures presented at the 11th International Conference on Fracture (ICF11), held in Turin, Italy, on March 20–25, 2005, volume 131(1–2). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 1-4020-4626-X (cloth), 1-4020-5423-8 (e-book). LCCN TA409 .I44 2005eb.

Caso:1996:FMF

[COK96] Gregory Caso, Pere Obrador, and C.-C. J. Kuo. Fast methods for fractal image encoding. Proceedings of the SPIE — The International Society for Optical Engineering, 2501(??):583–594, ???? 1996. CODEN PSISDG. ISSN ????

Cont:2010:MB

[Con10] Rama Cont. Mandelbrot, Benoît. In Encyclopedia of Quantitative Finance, page ?? Wiley, New York, NY, USA, 2010. ISBN 0-470-06160-X, 0-470-05756-4. LCCN HG106 .E53 2010. Four volumes.

Cortini:1990:FMS

[Cor90] Massimo Cortini. Fractal magmatic systems. Nature, 346(6281): 226, July 19, 1990. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v346/n6281/pdf/346226b0.pdf.

Conci:1998:FIA

[CP98] Aura Conci and Claudia Belmiro Proença. A fractal image analysis system for fabric inspection based on a box-counting method. Computer Networks and ISDN Systems, 30(20–21):1887–1895, November 12, 1998. CODEN CNISE9. ISSN 0169-7552 (print),

1879-2324 (electronic). URL http://www.elsevier.com/cas/
tree/store/comnet/sub/1998/30/20-21/2019.pdf.

Campbell:1983:OCP

[CR83] David R. Campbell and Harvey Rose, editors. Order in chaos: proceedings of the International conference on order in chaos, held at the Center for nonlinear studies, Los Alamos, New Mexico 87545, USA, 24–28 May 1982. North-Holland Publishing Co., Amsterdam, The Netherlands, 1983. ISBN 0-444-86727-9. LCCN QA843 .I57 1982.

Casey:1994:SSF

[CR94] Stephen D. Casey and Nicholas F. Reingold. Self-similar fractal sets: Theory and procedure. IEEE Computer Graphics and Applications, 14(3):73–82, May 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Cheng:1987:FNB

[CS87] Zheming Cheng and Robert Savit. Fractal and nonfractal behavior in Lévy flights. Journal of Mathematical Physics, 28(3):592– 597, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i3/p592_s1.

Chen:1991:IUS

[CT91] Min Chen and Roger Temam. Incremental unknowns for solving partial differential equations. Numerische Mathematik, 59 (3):255–271, June 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chen:1998:BSC

[CZ98] Ning Chen and Weiyong Zhu. Bud-sequence conjecture on M fractal image and M-J conjecture between C and Z planes from z ⟵ zw + c (w = a + ib). Computers and Graphics, 22(4): 537–546, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http://www.elsevier.

com/cas/tree/store/cag/sub/1998/22/4/585.pdf.

Drakopoulos:1999:GCS

[DAB99] V. Drakopoulos, N. Argyropoulos, and A. Böhm. Generalized computation of Schröder iteration functions to motivate families of Julia and Mandelbrot-like sets. SIAM Journal on Numerical Analysis, 36(2):417–435, April 1999. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (elec-

tronic). URL http://epubs.siam.org/sam-bin/dbq/article/
31736;   http://www.jstor.org/stable/2587202.

deAmo:2012:MAM

[dACSS12] Enrique de Amo, Manuel Díaz Carrillo, Juan Fernández Sánchez, and Antonio Salmerón. Moments and associated measures of copulas with fractal support. Applied Mathematics and Computation, 218(17):8634–8644, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300312001476.

Davis:1984:IBM

[Dav84] Monte Davis. Interview with Benoît Mandelbrot. Omni (New York), 6(5):64–??, February 1984. ISSN 0149-8711. URL http:

//archive.org/details/omni-magazine-1984-02;   http:/
/omnimagindex.wordpress.com/author-indices/authors-d/.

Davidoff:1990:DF

[Dav90] Frank Davidoff. Dynamic fractals. Computers and Graphics, 14 (1):135–136, ???? 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

David:1995:TFB

[Dav95] Hollister David. Two fractals based on Keplerian solids. Computers and Graphics, 19(6):885–888, November–December 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.elsevier.com/cgi-bin/cas/tree/
store/cag/cas_sub/browse/browse.cgi?year=1995&volume=
19&issue=6&aid=9500075.

Davis:1996:IIM

[Dav96] Geoffrey Davis. Implicit image models in image fractal compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2569(??):88–97, ???? 1996. CODEN PSISDG. ISBN 0-8194-2213-4. ISSN ????

Doescher:2003:INC

[DdCVR03] Erwin Doescher, Haroldo F. de Campos Velho, and Fernando M. Ramos. Isothermal and natural convection flows in fractal cavities. Applied Numerical Mathematics: Transactions of IMACS, 47(3–4):407–419, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Diebold:2010:KUU

[DDH10] Francis X. Diebold, Neil A. Doherty, and Richard Herring, editors. The known, the unknown, and the unknowable in financial risk management: measurement and theory advancing practice. Princeton University Press, Princeton, NJ, USA, 2010. ISBN 0-691-12883-9 (hardcover). viii + 380 pp. LCCN HD61 .K598 2010.

Duarte:1999:NAD

[DdMdO^{+}99] L. G. S. Duarte, L. A. C. P. da Mota, H. P. de Oliveira, R. O. Ramos, and J. E. F. Skea. Numerical analysis of dynamical systems and the fractal dimension of boundaries. Computer Physics Communications, 119(2–3):256–271, June 2, 1999. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465599002040.

Dekking:2012:PMP

[Dek12] Michel Dekking. Paperfolding morphisms, planefilling curves, and fractal tiles. Theoretical Computer Science, 414(1):20–37, January 13, 2012. CODEN TCSCDI. ISSN 0304-3975 (print),

1879-2294 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304397511008085.

Devooght:1957:LZM

[Dev57] J. Devooght. Sur la loi de Zipf–Mandelbrot. (French) [on the Zipf–Mandelbrot law]. Acad. Roy. Belg. Bull. Cl. Sci. (5), 43 (??):244–251, ???? 1957. CODEN ???? ISSN 0001-4141.

Devaney:1988:FPA

[Dev88] R. L. Devaney. Fractal patterns arising in chaotic dynamical systems. In Peitgen and Saupe [PS88a], pages 137–167. ISBN 0-387-96608-0 (New York), 1-4612-8349-3, 3-540-96608-0 (Berlin). LCCN QA614.86 .S35 1988.

Devaney:1990:TCO

[Dev90] Robert L. Devaney. Transition to chaos: the orbit diagram and the Mandelbrot set. Science television, New York, NY, USA, 1990. ISBN 1-878310-08-9. LCCN QA447 .T7 1990. One VHS videocassette (65 min.).

Devaney:1991:ODM

[Dev91] Robert L. Devaney. The orbit diagram and the Mandelbrot set. College Mathematics Journal, 22(1):23–38, January 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973355.

Devaney:1992:BRB

[Dev92] Robert L. Devaney. Book review: Fractals: An Animated Discussion with Edward Lorenz and Benoît Mandelbrot (H.-O. Peitgen, H. Jurgens, D. Saupe, and C. Zahlten). SIAM Review, 34(2):333– 335, June 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-

7200 (electronic). URL http://link.aip.org/link/siread/
v34/i2/p333/s1;   http://www.jstor.org/stable/2132866.

Devaney:1996:FGM

[Dev96] Robert L. Devaney. Fractal geometry of the Mandelbrot set. Key Curriculum Press, Berkeley, CA, USA, 1996. ISBN 1-55953-224-6. LCCN QA614.86 .F732 1996. One videocassette (70 min.).

Devaney:1999:MSF

[Dev99] Robert L. Devaney. The Mandelbrot set, the Farey tree, and the Fibonacci sequence. American Mathematical Monthly, 106 (4):289–302, April 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2589552.

Devaney:2004:MVM

[Dev04] Robert L. Devaney. Mandelbrot’s vision for mathematics. Proceedings of Symposia in Pure Mathematics, 72(1):??, ???? 2004.

CODEN ???? ISSN ???? URL http://www.math.yale.edu/
mandelbrot/web_pdfs/jubileeletters.pdf.

Dewdney:1985:CRC

[Dew85a] A. K. Dewdney. Computer recreations: a computer microscope zooms in for a look at the most complex object in mathematics. Scientific American, 253(2):17–21, 24, August 1985. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). Exploring the Mandelbrot set.

Dewdney:1985:CRE

[Dew85b] A. K. Dewdney. Computer recreations: Exploring the Mandelbrot set. Computer Graphics, 20(4):16–??, ???? 1985. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Dewdney:1987:CRBb

[Dew87] A. K. Dewdney. Computer recreations: Beauty and profundity: the Mandelbrot set and a flock of its cousins called Julia. Scientific American, 257(5):140–145, November 1987. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Dewdney:1989:CRT

[Dew89] A. K. Dewdney. Computer recreations: a tour of the Mandelbrot set aboard the Mandelbus. Scientific American, 260(2):108–111, February 1989. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Daintith:1999:DS

[DG99] John Daintith and Derek Gjertsen, editors. A Dictionary of Scientists. Oxford paperback reference. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-585-11047-6 (e-book), 0-19-280086-8 (paperback). 586 pp. LCCN Q141 .D52 1994. URL http://www.loc.gov/

catdir/enhancements/fy0637/99488304-d.html;   http://
www.loc.gov/catdir/enhancements/fy0637/99488304-t.html;
 http://www.netLibrary.com/urlapi.asp?action=summary&
v=1&bookid=12306.

Dhar:2005:BPG

[Dha05] Deepak Dhar. Branched polymers on the Given–Mandelbrot family of fractals. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 71(3 (part 1)):031801, ???? 2005. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic).

Drewes:1995:GSA

[DHKT95] F. Drewes, A. Habel, H.-J. Kreowski, and S. Taubenberger. Generating self-affine fractals by collage grammars. Theoretical Computer Science, 145(1–2):159–187, July 10, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.com/cgi-bin/cas/tree/
store/tcs/cas_sub/browse/browse.cgi?year=1995&volume=
145&issue=1-2&aid=1759.

Demko:1985:CFO

[DHN85] Stephen Demko, Laurie Hodges, and Bruce Naylor. Construction of fractal objects with iterated function systems. Computer Graphics, 19(3):271–278, July 1985. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Devaney:2006:CDT

[DK06] Robert L. Devaney and Linda Keen, editors. Complex dynamics: twenty-five years after the appearance of the Mandelbrot set: proceedings of an AMS-IMS-SIAM Joint Summer Research Conference on Complex Dynamics — Twenty-five Years after the Appearance of the Mandelbrot Set, June 13–17, 2004, Snowbird, Utah, volume 396 of Contemporary mathematics. American Mathematical Society, Providence, RI, USA, 2006. ISBN 0-8218-3625-0. ISSN 0271-4132 (print), 1098-3627 (electronic). LCCN QA614.86.A47 2004.

Dixon:1994:DSA

[DKW94a] A. R. Dixon, G. H. Kirby, and D. P. M. Wills. A data structure for artificial terrain generation. Computer Graphics Forum, 13(1): 37–47, March 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dixon:1994:TCD

[DKW94b] A. R. Dixon, G. H. Kirby, and D. P. M. Wills. Towards context dependent interpolation of digital elevation models. Computer Graphics Forum, 13(3):C/23–C/32, ???? 1994. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Daccord:1987:FPC

[DL87] Gérard Daccord and Roland Lenormand. Fractal patterns from chemical dissolution. Nature, 325(6099):41–43, January 1, 1987. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v325/
n6099/pdf/325041a0.pdf.

Daubechies:1992:TSD

[DL92] Ingrid Daubechies and Jeffrey C. Lagarias. Two-scale difference equations. II. Local regularity, infinite products of matrices and fractals. SIAM Journal on Mathematical Analysis, 23(4):1031– 1079, July 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kristiansen:2004:EOZ

[dLKHS04] Kai de Lange Kristiansen, Geir Helgesen, and Arne T. Skjeltorp. Experimental observation of Zipf–Mandelbrot relation. Physica A, 335(3-4):413–420, ???? 2004. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

Kristiansen:2006:BTZ

[dLKHS06] K. de Lange Kristiansen, G. Helgesen, and A. T. Skjeltorp. Braid theory and Zipf–Mandelbrot relation used in microparticle dynamics. European Physical Journal B: Condensed Matter and Complex Systems, 51(3):363–371, June 2006. CODEN EPJBFY. ISSN 1434-6028 (print), 1434-6036 (electronic).

Dongarra:1999:RAP

[DLM99] J. J. Dongarra, E. Luque, and Tomas Margalef, editors. Recent advances in parallel virtual machine and message passing interface: 6th European PVM/MPI Users’ Group Meeting, Barcelona, Spain, September 26–29, 1999: proceedings, volume 1697 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN ???? ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Davis:1993:SSR

[DLS93] A. B. Davis, S. Lovejoy, and D. Schertzer. Supercomputer simulation of radiative transfer inside multifractal cloud models. In Keevallik and Karner [KK93], pages 112–115. ISBN 0-937194-28-X. LCCN QC912.3.I57 1992.

Damerau:1973:TDW

[DM73] F. J. Damerau and Benoît Mandelbrot. Tests of the degree of word clustering in samples of written English. Linguistics, 102 (??):58–75, ???? 1973. CODEN ???? ISSN ????

Dekking:1990:SMP

[DM90] F. M. Dekking and R. W. J. Meester. On the structure of Mandelbrot’s percolation process and other random Cantor sets. Journal of Statistical Physics, 58(5–6):1109–1126, March 1990. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01026566.

Dubuc:1992:CHP

[DM92] Serge Dubuc and Abdul Malik. Convex hull of powers of a complex number, trinomial equations and the Farey sequence. Numerical Algorithms, 2(1):1–32, February 1992. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Drakopoulos:2003:OPV

[DMT03] V. Drakopoulos, N. Mimikou, and T. Theoharis. An overview of parallel visualisation methods for Mandelbrot and Julia sets. Computers and Graphics, 27(4):635–646, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

DePetrillo:1999:MDH

[DR99] P. B. DePetrillo and U. Ruttimann. Method for determining the Hurst exponent of a fractal time series and its applications to EKG signal analysis. Clinical Pharmacology and Therapeutics, 65(2):158, February 1999. CODEN CLPTAT. ISSN

???? URL http://www.nature.com/clpt/journal/v65/n2/
pdf/clpt1999171a.pdf.

Drewes:2001:TBG

[Dre01] Frank Drewes. Tree-based generation of languages of fractals. Theoretical Computer Science, 262(1–2):377–414, July 6, 2001. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://www.elsevier.nl/gej-ng/10/41/
16/204/21/41/abstract.html;   http://www.elsevier.nl/
gej-ng/10/41/16/204/21/41/article.pdf.

Dufour:2010:SIH

[DrKP10] Jean-Marie Dufour, Jeong ryeol Kim, and Franz C. Palm. Special issue: Heavy tails and Paretian distributions in empirical finance: a volume honoring Benoît Mandelbrot, volume 17.2010,2 of Journal of empirical finance. Elsevier, Amsterdam, The Netherlands, 2010. 177–282 pp. LCCN ????

Droniou:2010:NMF

[Dro10] Jérôme Droniou. A numerical method for fractal conservation laws. Mathematics of Computation, 79(269):95–124, January 2010. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-

6842 (electronic). URL http://www.ams.org/journals/mcom/
2010-79-269/S0025-5718-09-02293-5/home.html;   http:/
/www.ams.org/journals/mcom/2010-79-269/S0025-5718-09-
02293-5/S0025-5718-09-02293-5.pdf.

Dixon:1996:GGA

[DSB96] Stephen L. Dixon, Kevin L. Steele, and Robert P. Burton. Generation and graphical analysis of Mandelbrot and Julia sets in more than four dimensions. Computers and Graphics, 20(3): 451–456, May–June 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http://www.elsevier.

com/cgi-bin/cas/tree/store/cag/cas_sub/browse/browse.
cgi?year=1996&volume=20&issue=3&aid=9600015.

Davy:1990:SCP

[DSS90] Ph. Davy, A. Sornette, and D. Sornette. Some consequences of a proposed fractal nature of continental faulting. Nature, 348 (6296):56–58, November 1, 1990. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v348/n6296/pdf/348056a0.pdf.

daSilva:1999:SEM

[dSVTM99] Luciano R. da Silva, Raúl O. Vallejos, Constantino Tsallis, and Renio S. Mendes. Spectrum energy multifractality and equilibrium thermodynamics. Computer Physics Communications, 121–122(??):737, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465506701400.

Dumont:1989:SNF

[DT89] J.-M. Dumont and Alain Thomas. Systémes de numération et fonction fractales relatifs aux substitutions. (French) [Systems of numeration and fractal functions relative to substitutions]. Theoretical Computer Science, 65(2):153–169, June 28, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dorren:1991:MEN

[DT91] H. J. S. Dorren and A. Tip. Maxwell’s equations for nonsmooth media: fractal-shaped and pointlike objects. Journal of Mathematical Physics, 32(11):3060–3070, November 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i11/p3060_s1.

Dubuc:1989:TFS

[Dub89] Benoît Dubuc. On Takagi fractal surfaces. Canadian Mathematical Bulletin = Bulletin canadien de mathématiques, 32(??): 377–384, ???? 1989. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Duran:1978:XFM

[Dur78] Carlos Durán. A xeometría fractal de Mandelbrot. Grial, ??(62): 465–468, ???? 1978. CODEN ???? ISSN ???? URL http://

www.jstor.org/stable/29749665.

Domaszewicz:1993:SLS

[DV93] Jaroslaw Domaszewicz and Vinay A. Vaishampayan. Structural limitations of self-affine and partially self-affine fractal compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2094(??):1498–1507, ???? 1993. CODEN PSISDG. ISSN ????

deVries:1993:WLD

[dVdRL93] Pedro de Vries, Hans de Raedt, and Ad Lagendijk. Wave localization in disordered and fractal systems. Computer Physics Communications, 75(3):298–310, May 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
001046559390046F.

deVega:1996:SGE

[dVSC96] H. J. de Vega, N. Sánchez, and F. Combes. Self-gravity as an explanation of the fractal structure of the interstellar medium. Nature, 383(6595):56–58, September 5, 1996. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v383/
n6595/pdf/383056a0.pdf.

Dyson:1978:RCI

[Dys78] Freeman Dyson. Review: Characterizing irregularity: Fractals. Form, Chance, and Dimension, by Benoît B. Mandelbrot. Science, 200(4342):677–678, May 12, 1978. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

jstor.org/stable/1746976.

Epstein:1973:VAT

[EB73] Saul T. Epstein and Michael F. Barnsley. A variational approach to the theory of multipoint Padé approximants. Journal of Mathematical Physics, 14(3):314–325, March 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v14/i3/p314_s1.

Edgar:2000:FFP

[Edg00] G. A. Edgar. The forest fractal puzzle. Computers and Graphics, 24(1):133–141, February 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http://www.

elsevier.nl/gej-ng/10/13/20/47/27/38/abstract.html;
http://www.elsevier.nl/gej-ng/10/13/20/47/27/38/article.
pdf.

Egolf:1994:RBF

[EG94] David A. Egolf and Henry S. Greenside. Relation between fractal dimension and spatial correlation length for extensive chaos. Nature, 369(6476):129–131, May 12, 1994. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v369/n6476/pdf/369129a0.pdf.

Egghe:1999:RLZ

[Egg99] L. Egghe. Research: On the law of Zipf–Mandelbrot for multiword phrases. Journal of the American Society for Information Science, 50(3):233–241, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Eke:2005:FNC

[EHH05] Andras Eke, Péter Hermán, and Márton Hajnal. Fractal and noisy CBV dynamics in humans: influence of age and gender. Journal of Cerebral Blood Flow and Metabolism, 26(7):891–898, November 16, 2005. CODEN JCBMDN. ISSN 0271-678X (print),

1559-7016 (electronic). URL http://www.nature.com/jcbfm/
journal/v26/n7/full/9600243a.html.

Evertsz:1991:BHM

[EJM91] Carl J. G. Evertsz, Peter W. Jones, and Benoît B. Mandelbrot. Behaviour of the harmonic measure at the bottom of fjords. Journal of Physics A (Mathematical and General), 24(8):1889–1901, ???? 1991. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-

6447 (electronic). URL http://stacks.iop.org/0305-4470/
24/1889.

Eftekhari:2006:SED

[EKKR06] Ali Eftekhari, Mahmood Kazemzad, and Mansoor Keyanpour-Rad. Significant effect of dopant size on nanoscale fractal structure of polypyrrole film. Polymer Journal (Tokyo, Japan), 38(8):781–785, June 30, 2006. CODEN POLJB8. ISSN

???? URL http://www.nature.com/pj/journal/v38/n8/abs/
pj2006105a.html.

Evertsz:1992:HMA

[EM92] Carl J. G. Evertsz and Benoît B. Mandelbrot. Harmonic measure around a linearly self-similar tree. Journal of Physics A (Mathematical and General), 25(7):1781–1797, ???? 1992. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic). URL

http://stacks.iop.org/0305-4470/25/1781.

Evertsz:1992:EVF

[EMW92a] Carl J. G. Evertsz, Benoît B. Mandelbrot, and Lionel Woog. Erratum: Variability of the form and of the harmonic measure for small off-off-lattice diffusion-limited aggregates [phys. rev. A 4 5, 5798 (1992)]. Physical Review A (Atomic, Molecular, and Optical Physics), 45:8985–??, 1992. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL http://link.

aps.org/abstract/PRA/v45/e8985.

Evertsz:1992:VFH

[EMW92b] Carl J. G. Evertsz, Benoît B. Mandelbrot, and Lionel Woog. Variability of the form and of the harmonic measure for small off-off-lattice diffusion-limited aggregates. Physical Review A (Atomic, Molecular, and Optical Physics), 45:5798–??, 1992. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-

4519. URL http://link.aps.org/abstract/PRA/v45/e5798.

Engel:1993:MJS

[Eng93] Alejandro B. Engel. Morphosis of the Julia set of the real parameter family of complex quadratic maps. Computers and Graphics, 17(3):315–319, May–June 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Entwistle:1989:JSA

[Ent89] Ian D. Entwistle. Julia set art and fractals in the complex plane. Computers and Graphics, 13(3):389–392, ???? 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Erzan:1995:FST

[EPV95] A. Erzan, L. Pietronero, and A. Vespignani. The fixed-scale transformation approach to fractal growth. Reviews of Modern Physics, 67(3):545–604, July 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.67.545;   http:
//rmp.aps.org/abstract/RMP/v67/i3/p545_1.

Evertsz:1996:FGA

[EPV96] C. J. G. (Carl J. G.) Evertsz, Heinz-Otto Peitgen, and Richard Frederick Voss, editors. Fractal geometry and analysis: the Mandelbrot festschrift, Curaçao 1995. World Scientific Publishing, Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1996. ISBN 981-02-2434-6. ISSN 0218-348X. LCCN QA614.86 .F677 1996.

Ewing:1992:AMS

[ES92] John H. Ewing and Glenn Schober. The area of the Mandelbrot set. Numerische Mathematik, 61(1):59–72, February 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Esposito:1997:IIS

[Esp97] Dino Esposito. Inside Iterated Systems’ Fractal Development Kit — the Fractal Development Kit from Iterated Systems is a library that makes it possible for you to embed fractal-imaging capabilities into C/C++ Windows and Macintosh applications. Dr. Dobb’s Journal of Software Tools, 22(1):74–??, January 1997. CODEN DDJOEB. ISSN 1044-789X.

EstradaGallego:2011:BMG

[Est11] Fernando Estrada Gallego. Benoît Mandelbrot (1924–2010): a Greek among Romans. History of Economic Ideas, 19(1):9–14, 2011. CODEN ???? ISSN 1122-8792 (print), 1724-2169 (electronic).

Ewing:1996:CJF

[EW96] Gary J. Ewing and Christopher J. Woodruff. Comparison of JPEG and fractal-based image compression on target acquisition by human observers. Optical Engineering, 35(1):284–288, January 1996. CODEN OPEGAR. ISSN 0091-3286 (print), 1560-2303 (electronic).

Ewing:1995:CWS

[Ewi95] John Ewing. Can we see the Mandelbrot set? College Mathematics Journal, 26(2):90–99, March 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/07468342.1995.11973676.

Fairthorne:1969:PDE

[Fai69] Robert A. Fairthorne. Progress in documentation: Empirical hyperbolic distributions (Bradford–Zipf–Mandelbrot) for bibliometric description and prediction. Journal of Documentation, 25(4): 313–343, December 1969. CODEN JDOCAS. ISSN 0022-0418 (print), 1758-7379 (electronic). Reprinted in [Fai05].

Fairthorne:2005:PDE

[Fai05] Robert A. Fairthorne. Progress in documentation: Empirical hyperbolic distributions (Bradford–Zipf–Mandelbrot) for bibliometric description and prediction. Journal of Documentation, 61(2):

171–193, ???? 2005. CODEN JDOCAS. ISSN 0022-0418 (print), 1758-7379 (electronic). Reprint of [Fai69].

Falconer:1985:GFS

[Fal85] K. J. Falconer. The geometry of fractal sets, volume 85 of Cambridge tracts in mathematics. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-25694-1. xiv + 162 pp. LCCN QA248 .F274 1985. URL http://www.loc.

gov/catdir/description/cam031/84012091.html;   http:/
/www.loc.gov/catdir/samples/cam031/84012091.html;   http:
//www.loc.gov/catdir/toc/cam031/84012091.html.

Fama:1963:MSP

[Fam63] Eugene F. Fama. Mandelbrot and the stable Paretian hypothesis. The Journal of Business at the University of Chicago, 36(4): 420–429, October 1963. CODEN ???? ISSN 0021-9398 (print),

1537-5374 (electronic). URL http://www.jstor.org/stable/
2350971.

Farmer:2011:UTB

[Far11] J. Doyne Farmer. The unsmooth trajectory of Benoît Mandelbrot. Quantitative Finance, 11(2):157–158, ???? 2011. CODEN ???? ISSN 1469-7688 (print), 1469-7696 (electronic).

Fatou:1917:SRb

[Fat17a] Pierre Fatou. Sur les substitutions rationnelles. (French) [On rational substitutions]. Comptes rendus de l’Académie des sciences, Paris, 165(??):992–995, ???? 1917. CODEN ???? ISSN ????

Fatou:1917:SRFa

[Fat17b] Pierre Fatou. Sur les substitutions rationnelles. (French) [On rational substitutions]. Comptes rendus de l’Académie des sciences, Paris, 164(??):806–808, ???? 1917. CODEN ???? ISSN ????

Fernandez-Barbero:1999:EPS

[FBVdlN99] A. Fernández-Barbero, B. Vincent, and F. J. de las Nieves. Effect of the particle sign of charge on mesoscopic cluster fractal structures. Computer Physics Communications, 121–122 (??):635, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506700406.

Frame:2000:ICI

[FC00] Michael Frame and Tatiana Cogevina. An infinite circle inversion limit set fractal. Computers and Graphics, 24(5):797– 804, October 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http://www.elsevier.nl/

gej-ng/10/13/20/47/33/38/abstract.html;   http://www.
elsevier.nl/gej-ng/10/13/20/47/33/38/article.pdf.

Fiete:2010:SQF

[FdL10] Gregory A. Fiete and Alex de Lozanne. Seeing quantum fractals. Science, 327(5966):652–653, February 5, 2010. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/327/5966/652.full.pdf.

Ferson:1983:RBF

[Fer83] Scott Ferson. Review: The Fractal Geometry of Nature, by Benoît B. Mandelbrot. Quarterly review of biology, 58(3):412– 413, September 1983. CODEN QRBIAK. ISSN 0033-5770 (print),

1539-7718 (electronic). URL http://www.jstor.org/stable/
2828658.

Fernau:1995:VLA

[Fer95] Henning Fernau. Valuations of languages, with applications to fractal geometry. Theoretical Computer Science, 137(2):177– 217, January 23, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.elsevier.

com/cgi-bin/cas/tree/store/tcs/cas_sub/browse/browse.
cgi?year=1995&volume=137&issue=2&aid=1690.

Ferris:1991:WTP

[FF91] Timothy Ferris and Clifton Fadiman, editors. The world treasury of physics, astronomy, and mathematics. Little, Brown and Co., Boston, MA, USA, 1991. ISBN 0-316-28129-8. xv + 859 pp. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Feng:2009:FBB

[FHD09] Benjamin Zhong Ming Feng, Changcheng Huang, and Michael Devetsikiotis. FISTE: a black box approach for end-to-end QoS management. ACM Transactions on Modeling and Computer Simulation, 19(4):16:1–16:??, October 2009. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Finkel:1989:FDG

[Fin89] David L. Finkel. Fractal displays of genomic DNA. I. Eco RI fractal lattice of buffalo rat. International Journal of Quantum Chemistry, 36(5):575–586, November 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fisher:1992:DFI

[Fis92] Yuval Fisher. A discussion of fractal image compression. In Heinz-Otto Peitgen, Hartmut Jürgens, and Dietmar Saupe, editors, Chaos and Fractals: New Frontiers of Science, app. Appendix A, pages 903–919. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97903-4.

Fisher:1995:NAF

[Fis95] Yuval Fisher, editor. NATO ASI on Fractal Image Encoding and Analysis: July 8–17, 1995, Trondheim, Norway. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-63196-8. LCCN TA1637 .N38 1998.

FRANTZ:1991:MSC

[FL91] Marny Frantz and Sylvia Lazarnick. The Mandelbrot set in the classroom. The Mathematics Teacher, 84(3):173–177, ???? 1991. CODEN ???? ISSN 0025-5769. URL http://www.jstor.org/

stable/27967086.

Flam:1991:BFD

[Fla91] Faye Flam. Beating a fractal drum. Science, 254(5038): 1593, December 13, 1991. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/254/5038/1593.1.full.pdf.

Flake:1998:CBN

[Fla98] Gary William Flake. The computational beauty of nature: computer explorations of fractals, chaos, complex systems, and adaptation. MIT Press, Cambridge, MA, USA, 1998. ISBN 0-262-06200-3 (hardcover). xviii + 493 pp. LCCN QA76.6 .F557 1998.

Fleury:1997:BFP

[Fle97] Vincent Fleury. Branched fractal patterns in non-equilibrium electrochemical deposition from oscillatory nucleation and growth. Nature, 390(6656):145–148, November 13, 1997. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v390/n6656/
full/390145a0.html.

Frame:2002:PFT

[FM02a] Michael Frame and Benoît Mandelbrot. Panorama of fractals and their uses: an alphabetic wordbook-index. In Fractals, graphics, and mathematics education [FM02b], page ?? ISBN 0-88385-169-5. LCCN QA614.86 .F68424 2002. URL http://www.loc.gov/

catdir/description/cam022/2001097388.html;   http://www.
loc.gov/catdir/toc/cam027/2001097388.html. Foreword by

Lynn Arthur Steen.

Frame:2002:FGM

[FM02b] Michael Frame and Benoît B. Mandelbrot. Fractals, graphics, and mathematics education. MAA notes. The Mathematical Association of America (MAA), Washington, DC, 2002. ISBN 0-88385-169-5. xiii + 206 pp. LCCN QA614.86 .F68424 2002. URL http:/

/www.loc.gov/catdir/description/cam022/2001097388.html;
 http://www.loc.gov/catdir/toc/cam027/2001097388.html.

Foreword by Lynn Arthur Steen.

Faloutsos:1996:MSD

[FMS96] Christos Faloutsos, Yossi Matias, and Abraham Silberschatz. Modeling skewed distribution using multifractals and the ‘80– 20’ law. In Vijayaraman et al. [VBMS96], pages 307–317. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996. URL http://www.

vldb.org/dblp/db/conf/vldb/FaloutsosMS96.html.

Ford:2004:DEM

[For04] Roger Ford. Discovering and exploring Mandelbrot set points with a graphing calculator. The Mathematics Teacher, 98(1): 38–46, ???? 2004. CODEN ???? ISSN 0025-5769. URL http://

www.jstor.org/stable/27971613.

Fortescue:1989:NPU

[FP89] Peter D. Fortescue and Geoffrey S. Puterbaugh. Note on the practical use of the VGA monitor for the display of fractal images. Computers and Graphics, 13(4):559–560, ???? 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Flake:2010:SOS

[FP10] Gary William Flake and David M. Pennock. Self-organization, self-regulation, and self-similarity on the fractal web. In Lesmoir-Gordon [LG10b], chapter 6, pages 88–119. ISBN 1-84996-485-8.

LCCN QA614.86 .C65 2010. URL http://www.springerlink.
com/content/978-1-84996-485-2. With contributions by Ian

Stewart, Sir Arthur C. Clarke, Benoît B. Mandelbrot, Michael and Louisa Barnsley, Will Rood, Gary Flake, and David Pennock.

Foroutan-pour:1999:AIB

[FpDS99] K. Foroutan-pour, P. Dutilleul, and D. L. Smith. Advances in the implementation of the box-counting method of fractal dimension estimation. Applied Mathematics and Computation, 105(2–3):195–210, November 12, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/store/amc/sub/1999/105/2-3/
6317.pdf;   http://www.elsevier.com/cgi-bin/cas/tree/
store/amc/cas_sub/browse/browse.cgi?year=1999&volume=
105&issue=2-3&aid=6317.

Frame:1992:NSA

[FPR92] Michael Frame, A. G. Davis Philip, and Adam Robucci. A new scaling along the spike of the Mandelbrot set. Computers and Graphics, 16(2):223–234, Summer 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Frame:1992:GMS

[FR92] Michael Frame and James Robertson. A generalized Mandelbrot set and the role of critical points. Computers and Graphics, 16(1): 35–40, ???? 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fernandez-Ranada:1998:RFC

[FR98] Antonio Fernández-Rañada. Review: Las formas de las cosas: La geometría fractal de la naturaleza by Benoît Mandelbrot. Revista de libros, ??(24):21, December 1998. CODEN ???? ISSN ????

URL http://www.jstor.org/stable/30228845.

Frantz:1998:TFW

[Fra98] Marc Frantz. Two functions whose powers make fractals. American Mathematical Monthly, 105(7):609–617, August/September 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fraysse:2007:GVM

[Fra07] A. Fraysse. Generic validity of the multifractal formalism. SIAM Journal on Mathematical Analysis, 39(2):593–607, ???? 2007. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Friesen:2004:BRS

[Fri04] John Friesen. Book review: Scott Mandelbrote, Footprints of the Lion: Isaac Newton at Work. Exhibition at Cambridge University Library 9 October 2001–23 March 2002. Cambridge: Cambridge University Library, 2001. Pp. 142. ISBN 0-902205-58-7. £7.50 (paperback). British Journal for the History of Science, 37(3): 349–350, September 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.org/

stable/4028130;   http://www.jstor.org/stable/4028432.

Fromhold:1997:FRT

[Fro97] Mark Fromhold. Fractal resistance in a transistor. Nature, 386 (6621):123–125, March 13, 1997. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v386/n6621/pdf/386123a0.pdf.

Fowler:1989:DSM

[FSD89] Anthony D. Fowler, H. Eugene Stanley, and Gérard Daccord. Disequilibrium silicate mineral textures: fractal and non-fractal features. Nature, 341(6238):134–138, September 14, 1989. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v341/
n6238/pdf/341134a0.pdf.

Fisher:1994:CFM

[FSR94] Y. Fisher, T. P. Shen, and D. Rogovin. A comparison of fractal methods with discrete cosine transform (DCT) and wavelets. Proceedings of the SPIE — The International Society for Optical Engineering, 2304-16(??):132–143, July 28–29, 1994. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic). URL

http://inls.ucsd.edu/y/Fractals/spie.ps.

Forte:1995:TGF

[FV95] Bruno Forte and Edward Vrscay. Theory of generalized fractal transforms. In Fisher [Fis95], pages xiii + 368. ISBN 3-540-63196-8. LCCN TA1637 .N38 1998.

Gharavi-Alkhansari:1993:FBI

[GAH93] Mohammad Gharavi-Alkhansari and Thomas S. Huang. A fractal-based image block-coding algorithm. In Anonymous, editor, PCS’93: 1993 Picture Coding Symposium proceedings, 17–19 March, 1993, Swiss Federal Institute of Technology, Lausanne, Switzerland, page 1.7. Swiss Federal Institute of Technology, Lausanne, Switzerland, March 1993. LCCN TA1632.P611p 1993.

Gaite:2006:CSZ

[Gai06] José Gaite. Cut-out sets and the Zipf law for fractal voids. Physica D, Nonlinear phenomena, 223(2):248–255, ???? 2006. CODEN PDNPDT. ISSN 0167-2789 (print), 1872-8022 (electronic).

Gefen:1983:PTF

[GAM83] Yuval Gefen, Amnon Aharony, and Benoît B. Mandelbrot. Phase transitions on fractals. I. Quasilinear lattices. Journal of Physics A (Mathematical and General), 16(6):1267–1278, ???? 1983. CO-DEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic).

URL http://stacks.iop.org/0305-4470/16/1267.

Gefen:1984:PTFc

[GAM84] Yeval Gefen, Amnon Aharony, and Benoît B. Mandelbrot. Phase transitions on fractals. III. Infinitely ramified lattices. Journal of Physics A (Mathematical and General), 17(6):1277–1289, ???? 1984. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-

6447 (electronic). URL http://stacks.iop.org/0305-4470/
17/1277.

Gefen:1981:SFF

[GAMK81] Yuval Gefen, Amnon Aharony, Benoît B. Mandelbrot, and Scott Kirkpatrick. Solvable fractal family, and its possible relation to the backbone at percolation. Physical Review Letters, 47(25): 1771–1774, ???? 1981. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.

aps.org/abstract/PRL/v47/e1771.

Gardner:1978:MGW

[Gar78] M. Gardner. Mathematical games: White and brown music, fractal curves and 1/f fluctuations. Scientific American, 238(4):16– 32, April 1978. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Garnett:1988:CAD

[Gar88a] Lucy Garnett. A computer algorithm for determining the Hausdorff dimension of certain fractals. Mathematics of Computation, 51(183):291–300, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Garrison:1988:EEM

[Gar88b] Peter Garrison. Expeditionary equipment for Mandelbrot adventurers. C Users Journal, 6(6):75–??, June 1988. CODEN ???? ISSN 0898-9788.

Garrison:1991:RBB

[Gar91] Lionel Garrison. Review: The Beauty and the Complexity of the Mandelbrot Set: School Edition, by John H. Hubbard. The Mathematics Teacher, 84(4):334, ???? 1991. CODEN ???? ISSN

0025-5769. URL http://www.jstor.org/stable/27967167.

Gefen:1984:PTFb

[GASM84] Yuval Gefen, Amnon Aharony, Yonathan Shapir, and Benoît B. Mandelbrot. Phase transitions on fractals. II. Sierpiński gaskets. Journal of Physics A (Mathematical and General), 17(2):435–444, ???? 1984. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447

(electronic). URL http://stacks.iop.org/0305-4470/17/435.

Gujar:1993:TGF

[GBCK93] Uday G. Gujar, Virendra C. Bhavsar, Stephen Y. M. Choi, and Prem K. Kalra. Traversed geometric fractals. IEEE Computer Graphics and Applications, 13(5):61–67, September 1993. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Grossu:2009:VTE

[GBR^{+}09] I. V. Grossu, C. Besliu, M. V. Rusu, Al. Jipa, C. C. Bordeianu, and D. Felea. Visual tool for estimating the fractal dimension of images. Computer Physics Communications, 180(10):1999– 2001, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001635.

Gozen:2010:FAR

[GDC^{+}10] Irep Gözen, Paul Dommersnes, Ilja Czolkos, Aldo Jesorka, Tatsiana Lobovkina, Owe Orwar, et al. Fractal avalanche ruptures in biological membranes. Nature Materials, 9(11):908–912, October 10, 2010. CODEN NMAACR. ISSN 1476-1122 (print), 1476-4660

(electronic). URL http://www.nature.com/nmat/journal/v9/
n11/full/nmat2854.html.

Grant-Duff:1996:PH

[GDH96] Z. N. Grant-Duff and P. G. Harrison. Parallelism via homomorphisms. Parallel Processing Letters, 6(2):279–295, June 1996. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Giarrusso:1995:PPH

[GF95] Dante Giarrusso and Yuval Fisher. A parameterization of the period 3 hyperbolic components of the Mandelbrot set. Proceedings of the American Mathematical Society, 123(12):3731–3737, December 1995. CODEN PAMYAR. ISSN 0002-9939 (print),

1088-6826 (electronic). URL http://www.jstor.org/stable/
2161901.

Grossu:2010:NVV

[GFB^{+}10] I. V. Grossu, D. Felea, C. Besliu, Al. Jipa, C. C. Bordeianu, E. Stan, and T. Esanu. A new version of Visual [Basic] tool for estimating the fractal dimension of images. Computer Physics Communications, 181(4):831–832, April 2010. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S001046550900397X.

Guerin:1998:TDS

[GH98] Charles-Antoine Guerin and Matthias Holschneider. Timedependent scattering on fractal measures. Journal of Mathematical Physics, 39(8):4165–4194, August 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Getz:2004:GMF

[GH04] Chonat Getz and Janet Helmstedt. Graphics with Mathematica: fractals, Julia sets, patterns and natural forms. Elsevier, Amsterdam, The Netherlands, 2004. ISBN 0-444-51760-X (hardback), 0-444-51774-X (CD-ROM). vii + 322 pp. LCCN QA76.95 .G48 2004. Includes CD-ROM.

Ghidaglia:1986:FDA

[Ghi86] J.-M. Ghidaglia. On the fractal dimension of attractors for viscous incompressible fluid flows. SIAM Journal on Mathematical Analysis, 17(5):1139–1157, September 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gibbs:1993:PF

[Gib93] W. Wayt Gibbs. Practical fractal. Scientific American, 269 (1):107–108, July 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v269/n1/pdf/scientificamerican0793-
107.pdf.

Gilbert:1986:FDS

[Gil86] William J. Gilbert. The fractal dimension of sets derived from complex bases. Canadian Mathematical Bulletin = Bulletin canadien de mathématiques, 29(??):495–500, ???? 1986. CODEN CM-BUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Giles:2004:BMF

[Gil04] Jim Giles. Benoît Mandelbrot: Father of fractals. Nature, 432(7015):266–267, November 17, 2004. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

http://www.nature.com/nature/journal/v432/n7015/full/
432266a.html.

Gintz:2002:CGA

[Gin02] Terry W. Gintz. Chaos and graphics: Artist’s statement CQUATS—a non-distributive quad algebra for 3D renderings of Mandelbrot and Julia sets. Computers and Graphics, 26(2):367– 370, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print),

1873-7684 (electronic). URL http://www.elsevier.com/gej-
ng/10/13/20/68/41/42/abstract.html.

Garijo:2017:EAC

[GJV17] Antonio Garijo, Xavier Jarque, and Jordi Villadelprat. An effective algorithm to compute Mandelbrot sets in parameter planes. Numerical Algorithms, 76(2):555–571, October 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gilbert:2003:FBT

[GK03] Anna Gilbert and Howard Karloff. On the fractal behavior of TCP. In ACM [ACM03], pages 297–306. ISBN 1-58113-674-9. LCCN QA75.5 .A22 2003. ACM order number 508030.

Gupta:2003:BGF

[GKL03] A. Gupta, R. Krauthgamer, and J. R. Lee. Bounded geometries, fractals, and low-distortion embeddings. In IEEE [IEE03], pages 534–543. CODEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http://ieeexplore.ieee.

org/iel5/8767/27770/01238226.pdf. IEEE Computer Society

Order Number PR02040.

Gelernter:1961:BRM

[GM61] H. L. Gelernter and Benoît Mandelbrot. Book review: S. M. Ulam, A Collection of Mathematical Problems, (Tracts in Pure and Applied Mathematics, Number 8) $5.00, 1960 Interscience, New York 150. Information and Control, 4(1):81–82, March 1961. CODEN IFCNA4. ISSN 0019-9958 (print), 1878-2981

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0019995861800405.

Gerstein:1964:RWM

[GM64] George L. Gerstein and Benoît Mandelbrot. Random walk models for the spike activity of a single neuron. Biophysical Journal, 4(1 (part 1)):41–68, January 1964. CODEN BIOJAU. ISSN 0006-3495 (print), 1542-0086 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0006349564867680.

Given:1983:DFL

[GM83] James A. Given and Benoît B. Mandelbrot. Diffusion on fractal lattices and the fractal Einstein relation. Journal of Physics A (Mathematical and General), 16(15):L565–L569, ???? 1983. CO-DEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic).

URL http://stacks.iop.org/0305-4470/16/L565.

Given:1984:CTP

[GM84a] James A. Given and Benoît B. Mandelbrot. Comment on transport processes on fractal structures. Journal of Physics A (Mathematical and General), 17(9):1937–1939, ???? 1984. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic). URL

http://stacks.iop.org/0305-4470/17/1937.

Given:1984:NMP

[GM84b] James A. Given and Benoît B. Mandelbrot. A new model of percolation clusters. Journal of Statistical Physics, 36(5–6):545, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01012920;   http://link.
springer.com/article/10.1007/BF01012920;   http://www.
springerlink.com/content/l745uv6w7n3m8835/.

Gutzwiller:1988:IMM

[GM88] Martin C. Gutzwiller and Benoît B. Mandelbrot. Invariant multifractal measures in chaotic Hamiltonian systems, and related structures. Physical Review Letters, 60(8):673–676, ???? 1988. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/abstract/PRL/
v60/e673.

Gefen:1980:CPF

[GMA80] Yuval Gefen, Benoît B. Mandelbrot, and Amnon Aharony. Critical phenomena on fractal lattices. Physical Review Letters, 45 (11):855–858, ???? 1980. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.

aps.org/abstract/PRL/v45/e855.

Grisson:1998:CAA

[GMA^{+}98] Scott Grisson, Janet McAndless, Omar Ahmad, Christopher Stapleton, Adele Newton, Celia Pearce, Ryan Ulyate, and Rick Parent, editors. Conference abstracts and applications: SIGGRAPH 98, July 14–21, 1998, Orlando, FL, Computer Graphics. ACM Press, New York, NY 10036, USA, 1998. ISBN 1-58113-046-5, 1-58113-048-1 (CD-ROM). LCCN T385 .S54 1998b. URL http:

//info.acm.org/pubs/contents/proceedings/graph/. ACM

order number for CD-ROM: 435984. ACM order number: 435983.

Gefen:1984:PDS

[GMAK84] Yuval Gefen, B. B. Mandelbrot, A. Aharony, and A. Kapitulnik. Partial-dimensional sequences and percolation. Journal of Statistical Physics, 36(5–6):827–830, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01012942;   http://www.springerlink.com/content/j815gu7668r33v43/

.

Gefen:1983:GIH

[GMMA83] Yuval Gefen, Yigal Meir, Benoît B. Mandelbrot, and Amnon Aharony. Geometric implementation of hypercubic lattices with noninteger dimensionality, by use of low lacunarity fractal lattices. Physical Review Letters, 50(3):145–148, ???? 1983. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic),

1092-0145. URL http://link.aps.org/abstract/PRL/v50/
e145.

Gisin:2009:EEV

[GMV09] Vladimir Gisin, Andrey Markov, and Igor Vinukov. Estimation of extreme values of returns using the Zipf–Mandelbrot flow. International Journal of Pure and Applied Mathematics, 50(2):245– 250, ???? 2009. CODEN ???? ISSN 1311-8080 (print), 1314-3395 (electronic).

Goffinet:1991:NSC

[Gof91] Daniel Goffinet. Number systems with a complex base: a fractal tool for teaching topology. American Mathematical Monthly, 98 (3):249–255, March 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gomory:2010:BM

[Gom10] Ralph Gomory. Benoît Mandelbrot (1924–2010). Nature, 468 (7322):378, November 17, 2010. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v468/n7322/full/468378a.html.

GonzalezRodriguez:1998:TRM

[Gon98] Jose Angel Gonzalez Rodriguez. A tutorial and recipe for moving fractal trees. Computers and Graphics, 22(2–3):301–305, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-

7684 (electronic). URL http://www.elsevier.com/cas/tree/
store/cag/sub/1998/22/2-3/548.pdf.

Good:1978:RBF

[Goo78] I. J. Good. Review: Fractals: Form, Chance, and Dimension, by Benoît B. Mandelbrot. Journal of the American Statistical Association, 73(362):438, June 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2286682.

Good:1984:BRB

[Goo84a] I. J. Good. Book review: The Fractal Geometry of Nature, by Benoît B. Mandelbrot. SIAM Review, 26(1):131–132, January 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-

7200 (electronic). URL http://link.aip.org/link/siread/
v26/i1/p131/s1;   http://www.jstor.org/stable/2029696.

Good:1984:RBF

[Goo84b] I. J. Good. Review: The Fractal Geometry of Nature, by Benoît B. Mandelbrot. SIAM Review, 26(1):131–132, January 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://www.jstor.org/stable/2029696.

Good:1991:RBS

[Goo91] I. J. Good. Review: The Science of Fractal Images. by Heinz-Otto Peitgen, Dietmar Saupe, M. F. Barnsley, R. L. Devaney, B. B. Mandelbrot, R. F. Voss, Y. Fisher, M. McGuire. SIAM Review, 33(3):496–499, September 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://www.jstor.

org/stable/2031463.

Grebogi:1987:CSA

[GOY87] Celso Grebogi, Edward Ott, and James A. Yorke. Chaos, strange attractors, and fractal basin boundaries in nonlinear dynamics. Science, 238(4827):632–638, October 30, 1987. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/238/4827/632.full.pdf.

Garcia-Pelayo:1992:NF

[GPS92] Ricardo García-Pelayo and William C. Schieve. Noisy fractals. Journal of Mathematical Physics, 33(2):570–577, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v33/i2/p570_s1.

Groller:1992:FSM

[Gro92] E. Groller. Fractals and solid modeling. Computer Graphics Forum, 11(3):C415–C424, C484–C485, ???? 1992. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Grunwald:2005:BRB

[Gru05] Eric Grunwald. Book review: Benoît Mandelbrot and Richard L. Hudson, The (mis)behaviour of markets. The Mathematical Intelligencer, 27(3):77–79, September 2005. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic). URL http://

www.springerlink.com/content/32551765mu871utp/.

Gurevich:2007:MAT

[GT07] B. M. Gurevich and A. A. Tempelman. Multifractal analysis of time averages for continuous vector functions on configuration space. Theory of Probability and its Applications, 51(1):78–91, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gonzalez:2011:EDA

[GTL11] S. González, A. R. Thornton, and S. Luding. An eventdriven algorithm for fractal cluster formation. Computer Physics Communications, 182(9):1842–1845, September 2011. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S001046551000456X.

Guerin:2001:NGF

[Gué01] Charles-Antoine Guérin. A note on the generalized fractal dimensions of a probability measure. Journal of Mathematical Physics, 42(12):5871–5875, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guivarch:1987:RSE

[Gui87] Yves Guivarc’h. Remarques sur les solutions d’une équation fonctionnelle non linéaire de Benoît Mandelbrot. Comptes rendus de l’Académie des sciences. Série I, Mathématique, 305(4):139–141, ???? 1987. CODEN CASMEI. ISSN 0249-6291.

Gulick:2012:ECF

[Gul12] Denny Gulick. Encounters with chaos and fractals. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2012. ISBN 1-58488-517-3 (hardcover). xvi + 371 pp. LCCN Q172.5.C45 G85 2012.

Gunther:2024:VMS

[Gün24] Max Günther. Visualizing the Mandelbrot set with MetaPost. TUGboat, 45(1):113–114, 2024. CODEN ???? ISSN 0896-3207.

URL https://tug.org/TUGboat/tb45-1/tb139guenther-mandelbrot.
pdf.

Gomes:1989:FDF

[GV89] M. A. F. Gomes and G. L. Vasconcelos. Fragmentation dynamics in fractal and Euclidean systems. Computer Physics Communications, 54(2–3):257–261, June/July 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465589900891.

H:1988:FS

[H.88] J. H. Fractal shorthand. Scientific American, 258(2): 28, February 1988. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v258/n2/pdf/scientificamerican0288-
28a.pdf.

Hadamard:1988:HDD

[Had88] Jacques Hadamard. How I did not discover Relativity. The Mathematical Intelligencer, 10(2):65–67, June 1988. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic). URL

http://www.springerlink.com/content/l2q468845183710u/.

Translated by I. H. Rose. Edited by Benoît B. Mandelbrot.

Hallee:1996:RBF

[Hal96] Alan P. Hallee. Review: The Fractal Geometry of the Mandelbrot Set by Robert L. Devaney. The Mathematics Teacher, 89(6):518, ???? 1996. CODEN ???? ISSN 0025-5769. URL http://www.

jstor.org/stable/27969886.

Hamzaoui:1995:CCS

[Ham95] Raouf Hamzaoui. Codebook clustering by self-organizing maps for fractal image compression. In Fisher [Fis95], pages xiii + 368. ISBN 3-540-63196-8. LCCN TA1637 .N38 1998. URL ftp://

ftp.informatik.uni-freiburg.de/papers/fractal/Hamz95.
ps.gz.

Hart:1990:RBB

[Har90] Eric W. Hart. Review: The Beauty and the Complexity of the Mandelbrot Set: University Edition, by John H. Hubbard. The Mathematics Teacher, 83(8):686, ???? 1990. CODEN ???? ISSN

0025-5769. URL http://www.jstor.org/stable/27966920.

Hasumi:2007:ITS

[Has07] Tomohiro Hasumi. Interoccurrence time statistics in the twodimensional Burridge–Knopoff earthquake model. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 76(??):026117, August 29, 2007. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRevE.76.026117.

Hovi:1996:GIL

[HASM96] J.-P. Hovi, Amnon Aharony, Dietrich Stauffer, and Benoît B. Mandelbrot. Gap independence and lacunarity in percolation clusters. Physical Review Letters, 77:877–??, 1996. CODEN PRL-TAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

URL http://link.aps.org/abstract/PRL/v77/e877.

Hart:1995:SHC

[HCF95] John C. Hart, Wayne O. Cochran, and Patrick J. Flynn. Similarity hashing: a computer vision solution to the inverse problem of linear fractals. In Fisher [Fis95], pages xiii + 368. ISBN 3-540-63196-8. LCCN TA1637 .N38 1998.

Hart:1991:EAR

[HD91] John C. Hart and Thomas A. DeFanti. Efficient antialiased rendering of 3-D linear fractals. Computer Graphics, 25(4):91– 100, July 1991. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic). URL http://www.acm.org:80/

pubs/citations/proceedings/graph/122718/p91-hart/.

Heldt:1992:FDA

[Hel92] Gregory Heldt. Fractal dimension of airspaces of preterm rabbit lungs. Pediatric Research, 32(5):634, November 1992. CODEN ???? ISSN 0031-3998 (print), 1530-0447 (electronic). URL http:

//www.nature.com/pr/journal/v32/n5/pdf/pr1992452a.pdf.

Hertling:2005:MSC

[Her05] Peter Hertling. Is the Mandelbrot set computable? Mathematical Logic Quarterly: MLQ, 51(1):5–18, January 2005. CODEN MLQUEF. ISSN 0942-5616 (print), 1521-3870 (electronic).

Higgs:2011:IFA

[Hig11] Rowan Higgs. Imaging: Fractal analysis for CHD mortality. Nature Reviews Cardiology, 8(2):62, February 2011. CO-DEN NRCAE6. ISSN 1759-5002 (print), 1759-5010 (elec-

tronic). URL http://www.nature.com/nrcardio/journal/v8/
n2/full/nrcardio.2010.212.html.

Hirsch:1989:OCR

[Hir89] Morris W. Hirsch. Opinion: Chaos, rigor, and hype: James Gleick, Benoît B. Mandelbrot, and John Franks. The Mathematical Intelligencer, 11(3):6–13, September 1989. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic). URL http://

www.springerlink.com/content/v02l6108766n06h1/.

Hirst:1994:FLR

[Hir94] Bill Hirst. Fractal landscapes: from the real world. Cornerhouse, Manchester, UK, 1994. ISBN 0-948797-24-X (cased), 0-948797-23-1 (paperback). ???? pp. LCCN TR647.H57 A4 1994. £24.95. Photographs by Bill Hirst with an introduction by Benoît Mandelbrot.

Huillet:1994:CSS

[HJ94] Thierry Huillet and Bernard Jeannet. On a class of “skewed” selfsimilar and hyperbolic fractals. Journal of Mathematical Physics, 35(12):6511–6524, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v35/i12/p6511_s1.

Herman:2001:FBP

[HKE01] Peter Hermán, László Kocsis, and Andras Eke. Fractal branching Pattern in the pial vasculature in the cat. Journal of Cerebral Blood Flow and Metabolism, 21(6):741–753, June 2001. CODEN JCBMDN. ISSN 0271-678X (print), 1559-7016 (elec-

tronic). URL http://www.nature.com/jcbfm/journal/v21/
n6/full/9591109a.html.

Heckel:2011:IMI

[HKHP11] Frank Heckel, Olaf Konrad, Horst Karl Hahn, and Heinz-Otto Peitgen. Interactive 3D medical image segmentation with energyminimizing implicit functions. Computers and Graphics, 35(2): 275–287, April 2011. CODEN COGRD2. ISSN 0097-8493 (print),

1873-7684 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0097849310001925.

Herman:2005:FCS

[HKS^{+}05] Peter Herman, Ikuhiro Kida, Basavaraju G. Sanganahalli, Fahmeed Hyder, and Andras Eke. Fractal correlation structure in fMRI data of rat brain. Journal of Cerebral Blood Flow and Metabolism, 25(1s):S379, July 15, 2005. CODEN JCBMDN. ISSN 0271-678X (print), 1559-7016 (electronic).

URL http://www.nature.com/jcbfm/journal/v25/n1s/full/
9591524.0379a.html.

Hahn:2001:CMB

[HLTP01] Horst K. Hahn, Markus G. Lentschig, Burckhard Terwey, and Heinz-Otto Peitgen. Clinical MRI based volumetry: The cerebral ventricles. Lecture Notes in Computer Science, 2208(??):1291–??, ???? 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2208/22081291.htm;   http://
link.springer-ny.com/link/service/series/0558/papers/
2208/22081291.pdf.

Hamzaoui:1996:EFI

[HMS96a] Raouf Hamzaoui, Martin Müller, and Dietmar Saupe. Enhancing fractal image compression with vector quantization. In IEEE Digital Signal Processing Workshop. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA,

September 1996. URL ftp://ftp.informatik.uni-freiburg.
de/papers/fractal/HaMuSa96b.ps.gz. Preprint.

Hamzaoui:1996:VEF

[HMS96b] Raouf Hamzaoui, Martin Müller, and Dietmar Saupe. VQenhanced fractal image compression. In Proceedings ICIP-96 (IEEE International Conference on Image Processing), volume I, pages 153–156. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, September

1996. URL ftp://ftp.informatik.uni-freiburg.de/papers/
fractal/HaMuSa96a.ps.gz.

Hoffman:2010:BMM

[Hof10] Jascha Hoffman. Benoít Mandelbrot, mathematician, dies at 85. New York Times, ??(??):??, October 16, 2010. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://

www.nytimes.com/2010/10/17/us/17mandelbrot.html.

Holt:2018:WEW

[Hol18] Jim Holt. When Einstein walked with Gödel: excursions to the edge of thought. Farrar, Strauss, and Giroux, New York, NY, USA, 2018. ISBN 0-374-14670-5 (hardcover), 0-374-71784-2 (ebook). xi + 368 pp. LCCN PS3608.O4943595 A6 2018.

Hooper:1991:NSI

[Hoo91] Kenneth J. Hooper. Note on some internal structures of the Mandelbrot set. Computers and Graphics, 15(2):295–297, ???? 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Horgan:1990:MS

[Hor90a] John Horgan. Mandelbrot set-to. Scientific American, 262 (4):30–34, April 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v262/n4/pdf/scientificamerican0490-
30b.pdf.

Horn:1990:III

[Hor90b] Alastair N. Horn. IFSs and interactive image synthesis. Computer Graphics Forum, 9(2):127–137, June 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hough:1995:EAM

[Hou95] S. E. Hough. Earthquakes in the Los Angeles metropolitan region: A possible fractal distribution of rupture size. Science, 267(5195): 211–213, January 13, 1995. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/267/5195/211.full.pdf.

Howland:1978:RBF

[How78] Howard C. Howland. Review: Fractals: Form, Change, and Dimension, by Benoît B. Mandelbrot. Quarterly review of biology, 53(2):216, June 1978. CODEN QRBIAK. ISSN 0033-5770 (print), 1539-7718 (electronic). URL http://www.jstor.org/

stable/2826364.

Haskell:2002:FGP

[HRO02] John P. Haskell, Mark E. Ritchie, and Han Olff. Fractal geometry predicts varying body size scaling relationships for mammal and bird home ranges. Nature, 418(6897):527–530, August 1, 2002. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v418/n6897/full/nature00840.html.

Hurtgen:1993:FHC

[HS93] Bernd Hürtgen and Christoph Stiller. Fast hierarchical codebook search for fractal coding of still images. Proceedings of the SPIE

— The International Society for Optical Engineering, 1977(??): 397–408, April 1993. CODEN PSISDG. ISBN 0-8194-1223-6. ISSN ????

Hart:1989:RTD

[HSK89] John C. Hart, Daniel J. Sandin, and Louis H. Kauffman. Ray tracing deterministic 3-D fractals. Computer Graphics, 23(3): 289–296, July 1989. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic). URL http://www.acm.org:

80/pubs/citations/proceedings/graph/74333/p289-hart/.

Herman:2005:FPL

[HTHE05] Peter Herman, Hubert Trübel, Fahmeed Hyder, and Andras Eke. Fractal patterns of local and global CBF in rat brain during hypotension. Journal of Cerebral Blood Flow and Metabolism, 25 (1s):S195, July 15, 2005. CODEN JCBMDN. ISSN 0271-678X (print), 1559-7016 (electronic). URL http://www.nature.com/

jcbfm/journal/v25/n1s/full/9591524.0195a.html.

Hughes:1998:GJR

[Hug98] Merlin Hughes. 3D graphic Java: Render fractal landscapes. JavaWorld: IDG’s magazine for the Java community, 3(8):??, August 1998. CODEN ???? ISSN 1091-8906. URL http://www.

javaworld.com/javaworld/jw-08-1998/jw-08-step.htm.

Hurtgen:1993:CFT

[Hür93] Bernd Hürtgen. Contractivity of fractal transforms for image coding. Electronics Letters, 29(20):1749–1750, September 1993. CO-DEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

Hurtgen:1995:SEF

[Hür95] Bernd Hürtgen. Statistical evaluation of fractal coding schemes. In Proceedings ICIP-95 (IEEE International Conference on Image Processing), volume III, pages 280–283. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, October 1995. LCCN TK8315.I222 1995.

Herman:2005:FPN

[HWEH05] Peter Herman, Shaun A. Wahab, Andras Eke, and Fahmeed Hyder. Fractal properties of neurophysiologic signals in rat somatosensory cortex. Journal of Cerebral Blood Flow and Metabolism, 25(1s):S186, July 15, 2005. CODEN JCBMDN. ISSN 0271-678X (print), 1559-7016 (electronic).

URL http://www.nature.com/jcbfm/journal/v25/n1s/full/
9591524.0186a.html.

Hurd:1987:FAM

[HWM87] Alan J. Hurd, David A. Weitz, and Benoît B. Mandelbrot, editors. Fractal aspects of materials: disordered systems, volume EA–13 of Materials Research Society extended abstracts. Materials Research Society, Pittsburgh, PA, USA, 1987. LCCN QA447 .F72 1987.

IEEE:1995:IPI

[IEE95] IEEE, editor. ICIP’95: Proceedings, International Conference on Image Processing: October 23–26, 1995, Washington, DC, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-3122-2 (casebound), 0-8186-7310-9 (softbound), 0-7803-3123-0 (microfiche), 0-7803-2749-7 (CD-ROM). LCCN TK8315.I222 1995. Three volumes. IEEE catalog number 95CB35819.

IEEE:2003:PAI

[IEE03] IEEE, editor. Proceedings: 44th Annual IEEE Symposium on Foundations of Computer Science, FOCS 2003, 11–14 October 2003, Cambridge, Massachusetts. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. CODEN ASFPDV. ISBN 0-7695-2040-5. ISSN 0272-5428. LCCN QA76 .S979 2003. URL http:

//ieeexplore.ieee.org/iel5/8767/27770/01238173.pdf;
http://ieeexplore.ieee.org/xpl/RecentCon.jsp?punumber=

8767&conhome=1000292. IEEE Computer Society Order Number PR02040.

Isaeva:2001:MSC

[IKP01] Olga B. Isaeva, Sergey P. Kuznetsov, and Vladimir I. Ponomarenko. Mandelbrot set in coupled logistic maps and in an electronic experiment. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 64(5 (part 2)):055201, November 2001. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.aps.

org/abstract/PRE/v64/e055201.

Imre:2006:AFD

[Imr06] Attila R. Imre. Artificial fractal dimension obtained by using perimeter–area relationship on digitalized images. Applied Mathematics and Computation, 173(1):443–449, February 1, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300305003644.

Iosifescu:1979:BRBc

[Ios79] Marius Iosifescu. Book review: Fractals: Form, Chance and Dimension, by Benoît B. Mandelbrot. International Statistical Review = Revue Internationale de Statistique, 47(3):299, December 1979. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/1402650.

Izsak:2006:MLE

[Izs06a] F. Izsák. Maximum likelihood estimation for constrained parameters of multinomial distributions — application to Zipf– Mandelbrot models. Computational Statistics & Data Analysis, 51(3):1575–1583, ???? 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Izsak:2006:SPA

[Izs06b] János Izsák. Some practical aspects of fitting and testing the Zipf– Mandelbrot model. Scientometrics, 67(1):107–120, April 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11192-006-0052-x.

Jaffard:1997:MFFa

[Jaf97a] S. Jaffard. Multifractal formalism for functions. Part I: Results valid for all functions. SIAM Journal on Mathematical Analysis, 28(4):944–970, July 1997. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://epubs.siam.org/

sam-bin/dbq/article/28299.

Jaffard:1997:MFFb

[Jaf97b] S. Jaffard. Multifractal formalism for functions. Part II: Selfsimilar functions. SIAM Journal on Mathematical Analysis, 28 (4):971–998, July 1997. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://epubs.siam.org/

sam-bin/dbq/article/28300.

Jaffard:1998:OSP

[Jaf98] Stéphane Jaffard. Oscillation spaces: Properties and applications to fractal and multifractal functions. Journal of Mathematical Physics, 39(8):4129–4141, August 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jaffard:2011:BMM

[Jaf11] Stéphane Jaffard. Benoît Mandelbrot—in memoriam (1924– 2010). Gazette des Mathématiciens, 127(??):77–78, ???? 2011. CODEN ???? ISSN 0224-8999.

Jakeman:1984:AMS

[Jak84] Eric Jakeman. Applied mathematics: Scattering by fractal objects. Nature, 307(5947):110, January 12, 1984. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v307/
n5947/pdf/307110a0.pdf.

Jonckheere:1958:LE

[JeJP58] A. Jonckheere and B. Mandelbrot et J. Piaget. La lecture de l’expérience. (French) [Reading of experience], volume 5 of Etudes d’épistémologie génétique; Bibliothèque scientifique internationale. Presses universitaires de France, Paris, France, 1958. 150 pp. LCCN BF311.

Jackson:1997:SYE

[JH97] David J. Jackson and Chris W. Humphres. A simple yet effective load balancing extension to the PVM software system. Parallel Computing, 22(12):1647–1660, February 21, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336

(electronic). URL http://www.elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/browse/browse.cgi?year=1997&
volume=22&issue=12&aid=1112.

Jackson:1996:PPF

[JM96a] David Jeff Jackson and Wagdy Mahmoud. Parallel pipelined fractal image compression using quadtree recomposition. The Computer Journal, 39(1):1–13, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www.oup.co.uk/jnls/list/comjnl/hdb/Volume_39/Issue_
01/390001.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/Volume_39/Issue_01/Vol39_01.body.html#AbstractJackson.

Jaffard:1996:LRN

[JM96b] Stéphane Jaffard and Benoît B. Mandelbrot. Local regularity of nonsmooth wavelet expansions and application to the Pólya function. Advances in Mathematics, 120(2):265–282, ???? 1996. CO-DEN ADMTA4. ISSN 0001-8708 (print), 1090-2082 (electronic).

Jones:1990:DGB

[Jon90a] Huw Jones. Dürer, gaskets and Barnsley’s chaos game. Computer Graphics Forum, 9(4):327–332, December 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jones:1990:TUR

[Jon90b] John Dewey Jones. Three unconventional representations of the Mandelbrot set. Computers and Graphics, 14(1):127–129, ???? 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jones:1994:FM

[Jon94] David Jones. Fractal money. Nature, 370(6490):510, August 18, 1994. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v370/n6490/pdf/370510a0.pdf.

Jorgensen:2012:ESF

[Jor12] Palle E. T. Jorgensen. Ergodic scales in fractal measures. Mathematics of Computation, 81(278):941–955, April 2012. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic).  URL http://www.ams.org/journals/mcom/2012-
81-278/S0025-5718-2011-02517-2;   http://www.ams.org/
journals/mcom/2012-81-278/S0025-5718-2011-02517-2/S0025-
5718-2011-02517-2.pdf.

Jurgens:1990:LF

[JPS90] Hartmut Jürgens, Heinz-Otto Peitgen, and Dietmar Saupe. The language of fractals. Scientific American, 263(2):60–?? (Intl. ed. 40–??), August 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Jones-Smith:2006:FAR

[JSM06] Katherine Jones-Smith and Harsh Mathur. Fractal analysis: Revisiting Pollock’s drip paintings. Nature, 444(7119):E9–E10, November 29, 2006. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v444/n7119/full/nature05398.html.

Jackson:1996:PAD

[JT96a] David Jeff Jackson and Greg Scott Tinney. Performance analysis of distributed implementations of a fractal image compression algorithm. Concurrency: Practice and Experience, 8(5):357–386, June 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-

9128 (electronic). URL http://www3.interscience.wiley.
com/cgi-bin/abstract?ID=23284.

Jan:1996:SIE

[JT96b] Jinn-Ke Jan and Yuh-Min Tseng. On the security of image encryption method. Information Processing Letters, 60(5):261–265, December 8, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jones:1991:MSA

[JTEA91] J. G. Jones, R. W. Thomas, P. G. Earwicker, and S. Addison. Multiresolution statistical analysis of computer-generated fractal imagery. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 53(4):349–363, July 1991. CODEN CGMPE5. ISSN 1049-9652 (print), 1557-7643 (electronic).

Julia:1918:MIF

[Jul18] Gaston Julia. Mémoire sur l’iteration des fonctions rationnelles. (French) [Memoir on the iteration of rational functions]. Journal de Mathématiques Pures et Appliquées, 8(??):47–245, ???? 1918. CODEN JMPAAM. ISSN 0021-7824 (print), 1776-3371 (electronic).

Juhos:1998:SCD

[JV98] Szilveszter Juhos and Lajos Voros. Structural changes during eutrophication of Lake Balaton, Hungary, as revealed by the Zipf–Mandelbrot model. Hydrobiologia, 369–370(0):237–242, ???? 1998. CODEN HYDRB8. ISSN 0018-8158 (print), 1573-5117 (electronic).

Kac:1978:RBF

[Kac78] Mark Kac. Review: Fractals: Form, Chance, and Dimension, by Benoît B. Mandelbrot. American Scientist, 66(2):250, March 1978. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786

(electronic). URL http://www.jstor.org/stable/27848608.

Kahane:1974:MTB

[Kah74] Jean-Pierre Kahane. Sur le modèle de turbulence de Benoît Mandelbrot. C. R. Acad. Sci. Paris Sér. A, 278(??):621–623, ???? 1974. CODEN ???? ISSN ????

Kajiya:1983:NTRa

[Kaj83a] James T. Kajiya. New techniques for ray tracing procedurally defined objects. Computer Graphics, 17(3):91–102, July 1983. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Kajiya:1983:NTRb

[Kaj83b] James T. Kajiya. New techniques for ray tracing procedurally defined objects. ACM Transactions on Graphics, 2(3):161–181, July 1983. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). Also appeared in SIGGRAPH ’83 Proceedings, and in Tutorial: Computer Graphics: Image Synthesis, Computer Society Press, Washington, 1988, pp. 168–188.

Kallam:2003:RBG

[Kal03] Linda Kallam. Review: Fractals, Graphics, and Mathematics Education by Michael L. Frame, Benoît B. Mandelbrot. The Mathematics Teacher, 96(5):379, ???? 2003. CODEN ???? ISSN

0025-5769. URL http://www.jstor.org/stable/20871345.

Kappraff:1986:GCS

[Kap86] Jay Kappraff. The geometry of coastlines: a study in fractals. Computers and Mathematics with Applications, 12(3–4):655–671, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0898122186904177.

Katz:1986:MPM

[Kat86] Howard Katz. A Mandelbrot program for the Macintosh. Dr. Dobb’s Journal of Software Tools, 11(11):42–45, November 1986. CODEN DDJOEB. ISSN 0888-3076.

Kaufmann:1992:MWE

[Kau92] Stephan Kaufmann. Mathematica als Werkzeug: Eine Einführung mit Anwendungsbeispielen. (German) [Mathematica as a tool: an introduction with application examples]. Birkhäuser Verlag, Basel, Switzerland, 1992. ISBN 3-7643-2832-0. 396 pp. LCCN ???? See [Bau96].

Keller:1989:TDS

[KCC89] James M. Keller, Susan Chen, and Richard M. Crownover. Texture description and segmentation through fractal geometry.

Computer Vision, Graphics, and Image Processing, 45(2):150– 166, February 1989. CODEN CVGPDB. ISSN 0734-189x (print), 1557-895x (electronic).

Keller:2000:AMS

[Kel00a] Karsten Keller. The abstract Mandelbrot set. Lecture Notes in Mathematics, 1732:73–139, 2000. CODEN LNMAA2. ISBN 3-540-67434-9 (print), 3-540-45589-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0104002/.

Keller:2000:IFJ

[Kel00b] Karsten Keller. Invariant Factors, Julia Equivalences and the (Abstract) Mandelbrot Set, volume 1732 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNMAA2. ISBN 3-540-67434-9 (print), 3-540-45589-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 206 pp. LCCN QA3 .L28 no. 1732. URL http:/

/link.springer.com/book/10.1007/BFb0103999;   http://
www.springerlink.com/content/978-3-540-45589-9.

Kelley:2000:LTF

[Kel00c] Alice Kelley. Layering techniques in fractal art. Computers and Graphics, 24(4):611–616, August 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

URL http://www.elsevier.nl/gej-ng/10/13/20/47/27/37/
abstract.html;   http://www.elsevier.nl/gej-ng/10/13/
20/47/32/37/article.pdf.

Kirchner:2000:FSC

[KFN00] James W. Kirchner, Xiahong Feng, and Colin Neal. Fractal stream chemistry and its implications for contaminant transport in catchments. Nature, 403(6769):524–527, February 3, 2000. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v403/
n6769/full/403524a0.html.

Kumari:2024:VAT

[KGN^{+}24] Sudesh Kumari, Krzysztof Gdawiec, Ashish Nandal, Naresh Kumar, and Renu Chugh. On the viscosity approximation type iterative method and its non-linear behaviour in the generation of Mandelbrot and Julia sets. Numerical Algorithms, 96(1):211–236, May 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01644-4.

Koh:1992:FGS

[KH92] Eng-Kiat Koh and D. D. Hearn. Fast generation and surface structuring methods for terrain and other natural phenomena. Computer Graphics Forum, 11(3):C169–C180, C472–C473, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kilmister:1978:RBF

[Kil78] C. W. Kilmister. Review: Fractals: Form, Chance and Dimension, by Benoît B. Mandelbrot. Mathematical Gazette, 62(420): 130–132, June 1978. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL http://www.jstor.org/stable/
3617679.

Kirkby:1979:BRB

[Kir79] M. J. Kirkby. Book review: Fractals: Form, chance, and dimension, Benoît B. Mandelbrot, W. H. Freeman and Co., 1977. No. of pages: 365. Price: U.S. $14.95. Earth Surface Processes, 4(1): 98, January/March 1979. CODEN ESPRDT. ISSN ????

Kirkby:1983:BRB

[Kir83] M. J. Kirkby. Book review: The fractal geometry of nature. Benoît B. Mandelbrot. W. H. Freeman and Co., San Francisco, 1982. No. of pages: 460. Price: £22.75 (hardback). Earth Surface Processes and Landforms, 8(4):406, July/August 1983. CODEN ESPLDB. ISSN ????

Kirman:2011:BMV

[Kir11] Alan Kirman. Benoît Mandelbrot and the vindication of his ideas. Quantitative Finance, 11(2):159–160, ???? 2011. CODEN ???? ISSN 1469-7688 (print), 1469-7696 (electronic).

Knill:1990:LNO

[KK90] David C. Knill and Daniel Kersten. Learning a near-optimal estimator for surface shape from shading. Computer Vision, Graphics, and Image Processing, 50(1):75–100, April 1990. CODEN CVGPDB. ISSN 0734-189x (print), 1557-895x (electronic).

Keevallik:1993:ICP

[KK93] Sirje Keevallik and Olavi Karner, editors. IRS ’92: current problems in atmospheric radiation: proceedings of the International Radiation Symposium, Tallinn, Estonia, 3–8 August 1992. A. Deepak Pub., Hampton, VA, USA, 1993. ISBN 0-937194-28-X. LCCN QC912.3.I57 1992.

Kim:1995:NFI

[KKL95] Chang-Su Kim, Rin-Chul Kim, and Sang-Uk Lee. Novel fractal image compression method with non-iterative decoder. In IEEE [IEE95], pages 268–271. ISBN 0-7803-3122-2 (casebound), 0-8186-7310-9 (softbound), 0-7803-3123-0 (microfiche), 0-7803-2749-7 (CD-ROM). LCCN TK8315.I222 1995. Three volumes. IEEE catalog number 95CB35819.

Kohler:2006:OSA

[KLPS06] Eddie Kohler, Jinyang Li, Vern Paxson, and Scott Shenker. Observed structure of addresses in IP traffic. IEEE/ACM Transactions on Networking, 14(6):1207–1218, December 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kahane:1965:EMA

[KM65] Jean-Pierre Kahane and Benoît Mandelbrot. Ensembles de multiplicité aléatoires. (French) [Sets of random multiplicity]. Comptes rendus de l’Académie des sciences, Paris, 261(??):3931–3933, ???? 1965. CODEN ???? ISSN ????

Kittel:1989:LGM

[KM89a] Charles Kittel and Benoît B. Mandelbrot. Letters: Gibbs and Mandelbrot at Yale. Physics Today, 42(3):154, 156, March 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v42/
i3/p154/s2.

Krantz:1989:FG

[KM89b] Steven G. Krantz and Benoît B. Mandelbrot. Fractal geometry. The Mathematical Intelligencer, 11(4):12–16, December 1989. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414

(electronic). URL http://www.springerlink.com/content/
n751525375q64411/.

Kiani:2011:MPD

[KM11] Soheila Kiani and Mohsen Ebrahimi Moghaddam. A multipurpose digital image watermarking using fractal block coding. The Journal of Systems and Software, 84(9):1550–1562, September 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0164121211000707.

Karman:1999:LOF

[KMNW99] G. P. Karman, G. S. McDonald, G. H. C. New, and J. P. Woerdman. Laser optics: Fractal modes in unstable resonators. Nature, 402(6758):138, November 11, 1999. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v402/
n6758/full/402138a0.html.

Kachanov:2012:CEA

[KNBZ12] Andre Kachanov, Serge Nikulin, Sveta Bauer, and Boris Zimin. Corneal ectasia after excimer surgery and the Mandelbrot B. fractal theory in ophthalmology. Acta Ophthalmologica, 90(s249):??, September 2012. CODEN ???? ISSN 1755-375X (print), 1755-3768 (electronic).

Kokol:1994:SSC

[Kok94] P. Kokol. The self-similarity and computer programs. ACM SIGPLAN Notices, 29(1):9–12, January 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kominek:1995:CRF

[Kom95] John Kominek. Codebook reduction in fractal image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2669(??):33–41, ???? 1995. CODEN PSISDG. ISBN 0-8194-2043-3. ISSN ????

Kopelman:1988:FRK

[Kop88] Raoul Kopelman. Fractal reaction kinetics. Science, 241(4873): 1620–1626, September 23, 1988. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/241/4873/1620.full.pdf.

Kahane:1976:CMB

[KP76] J.-P. Kahane and J. Peyrière. Sur certaines martingales de Benoît Mandelbrot. Advances in Mathematics, 22(2):131–145, ???? 1976. CODEN ADMTA4. ISSN 0001-8708 (print), 1090-2082 (electronic).

Khan:2019:ZME

[KPP19] Muhammad Adil Khan, Dilda Pecarić, and Josip Pecarić. On Zipf–Mandelbrot entropy. Journal of Computational and Applied Mathematics, 346(??):192–204, January 15, 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0377042718304217.

Kuroda:1996:FBM

[KPY96] Hideo Kuroda, Dan C. Popescu, and Hong Yan. Fast block matching method for image data compression based on fractal models. Proceedings of the SPIE — The International Society for Optical Engineering, 2501(??):1257–1266, ???? 1996. CODEN PSISDG. ISSN ????

Klein:1991:TBU

[KRP^{+}91] Michael Klein, Otto E. Rössler, Jürgen Parisi, Joachim Peinke, Gerold Baier, Claus Kahlert, and John L. Hudson. Toward a better understanding of fractality in nature. Computers and Graphics, 15(4):583–596, ???? 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kurths:1990:BRB

[Kur90] J. Kurths. Book review: Benoît B. Mandelbrot: Die fraktale Geometrie der Natur. Birkhäuser Verlag, Basel, 1987, 491 Seiten. Preis: Sfr 98,-. ISBN 3-7643-1771-X. Astronomische Nachrichten, 311(1):88, ???? 1990. CODEN ASNAAN. ISSN 0004-6337 (print), 1521-3994 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/asna.2113110129/abstract.

Kaufman:1995:PDL

[KVMW95] Henry Kaufman, Alessandro Vespignani, Benoît B. Mandelbrot, and Lionel Woog. Parallel diffusion-limited aggregation. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 52:5602–??, 1995. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.

aps.org/abstract/PRE/v52/e5602.

Korkut:2008:GFI

[KVŽŽ08] Luka Korkut, Domagoj Vlah, Darko Žubrinić, and Vesna Županović. Generalized Fresnel integrals and fractal properties of related spirals. Applied Mathematics and Computation, 206(1):236–244, December 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0096300308006723.

LaBrecque:1986:FA

[La 86] M. La Brecque. Fractal applications. Mosaic, 17(4):34–48, ???? 1986. CODEN ???? ISSN ????

Landsberg:1989:RBS

[Lan89] P. T. Landsberg. Review: The Science of Fractal Images. by Heinz-Otto Peitgen, Dietmar Saupe, M. F. Barnsley, R. L. Devaney, B. B. Mandelbrot, R. F. Voss, Y. Fisher, M. McGuire. Leonardo (Oxford, England), 22(3/4):455–456, ???? 1989. CO-DEN LEONDP. ISSN 0024-094X (print), 1530-9282 (electronic).

URL http://www.jstor.org/stable/1575447.

Landini:1997:PBF

[Lan97] G. Landini. Is periodontal breakdown a fractal process? Simulations using the Weierstrass–Mandelbrot function. Journal of Periodontal Research, 32(3):300–307, April 1997. CODEN JP-DRAY. ISSN 0022-3484 (print), 1600-0765 (electronic).

Lapidus:2008:SRZ

[Lap08] Michel L. (Michel Laurent) Lapidus. In search of the Riemann zeros: strings, fractal membranes and noncommutative spacetimes. American Mathematical Society, Providence, RI, USA, 2008. ISBN 0-8218-4222-6. xxix + 558 pp. LCCN QA333 .L37 2008.

Larison:1990:RBM

[Lar90] Ted Larison. Review: Mandelbrot Magic 3.1. The Mathematics Teacher, 83(2):157, ???? 1990. CODEN ???? ISSN 0025-5769.

URL http://www.jstor.org/stable/27966565.

Luce:1963:HMP

[LBG63] R. Duncan Luce, Robert R. Bush, and Eugene Galanter, editors. Handbook of Mathematical Psychology. Volume II, Chapters 9– 14. Wiley, New York, NY, USA, 1963. LCCN ????

Lees:1959:RBL

[Lee59] R. B. Lees. Review: Logique, langage et théorie de l’information by Léo Apostel, Benoît Mandelbrot, Albert Morf. Language: journal of the Linguistic Society of America, 35(2):271–303, June 1959. CODEN ???? ISSN 0097-8507. URL http://www.jstor.

org/stable/410536.

Lefort:1982:LOA

[Lef82] B. Lefort. L’emploi des outils au cours de taches d’entretien et la loi de Zipf–Mandelbrot. (French) [the use of tools in maintenance tasks and the Zipf–Mandelbrot Law]. Le Travail Humain, 45 (2):307–316, ???? 1982. CODEN TRHUAH. ISSN 0041-1868 (print), 2104-3663 (electronic). URL http://www.jstor.org/

stable/40657805.

Leversha:2005:RBC

[Lev05] Gerry Leversha. Review: The Colours of Infinity by Ian Stewart, Arthur C. Clarke, Benoît Mandelbrot, Michael Barnsley, Louisa Barnsley, Will Rood, Gary Flake, David Pennock, Robert Prechter, Nigel Lesmoir-Gordon. Mathematical Gazette, 89(514): 170–171, March 2005. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL http://www.jstor.org/

stable/3620696.

Lewis:1987:GSS

[Lew87] J. P. Lewis. Generalized stochastic subdivision. ACM Transactions on Graphics, 6(3):167–190, July 1987. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0730-0301/35069.html.

Lewis:1989:ASN

[Lew89] J. P. Lewis. Algorithms for solid noise synthesis. Computer Graphics, 23(3):263–270, July 1989. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

URL http://www.acm.org:80/pubs/citations/proceedings/
graph/74333/p263-lewis/.

Leys:2005:SIF

[Ley05] Jos Leys. Sphere inversion fractals. Computers and Graphics, 29(3):463–466, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lesmoir-Gordon:2010:BMO

[LG10a] Nigel Lesmoir-Gordon. Benoît Mandelbrot obituary. The Guardian (London), ??(??):??, October 17, 2010. CODEN ????

ISSN ???? URL http://www.guardian.co.uk/science/2010/
oct/17/benoit-mandelbrot-obituary.

Lesmoir-Gordon:2010:CIB

[LG10b] Nigel Lesmoir-Gordon, editor. The colours of infinity: the beauty and power of fractals. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 1-84996-485-8. 207 pp. LCCN QA614.86 .C65 2010. URL http://extras. springer.com. With contributions by Ian Stewart, Sir Arthur C. Clarke, Benoît B. Mandelbrot, Michael and Louisa Barnsley, Will Rood, Gary Flake, and David Pennock.

Lesmoir-Gordon:2010:FCI

[LG10c] Nigel Lesmoir-Gordon. Filming the colours of infinity. In The colours of infinity: the beauty and power of fractals [LG10b], chapter 8, pages 136–143. ISBN 1-84996-485-8. LCCN QA614.86

.C65 2010. URL http://www.springerlink.com/content/978-

1-84996-485-2. With contributions by Ian Stewart, Sir Arthur C. Clarke, Benoît B. Mandelbrot, Michael and Louisa Barnsley, Will Rood, Gary Flake, and David Pennock.

Linde:1994:SRI

[Lin94] Andrei Linde. The self-reproducing inflationary universe. Scientific American, 271(5):48–?? (Intl. ed. 32–39), November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Liu:1993:EFT

[Liu93] Ying Liu. Extensions of fractal theory. Proceedings of the SPIE — The International Society for Optical Engineering, 1966(??): 255–268, ???? 1993. CODEN PSISDG. ISSN ????

Liu:1995:IBC

[Liu95a] Ying Liu. Image block coding using exact solutions of fractal equations. Proceedings of the SPIE — The International Society for Optical Engineering, 2488(??):135–149, ???? 1995. CODEN PSISDG. ISSN ????

Liu:1995:SFE

[Liu95b] Ying Liu. Solution of fractal equation. Proceedings of the SPIE — The International Society for Optical Engineering, 2418(??): 48–59, ???? 1995. CODEN PSISDG. ISSN ????

Liu:2006:BRS

[Liu06] Nelson Liu. Book reviews: a summary of multifractal traffic modeling. IEEE Distributed Systems Online, 7(3):??, March 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683

(electronic). URL http://csdl.computer.org/comp/mags/ds/
2006/03/o3008.pdf.

Laxer:1998:ICF

[LK98] Cary Laxer and Aaron Klebanoff. An interactive course on fractals and chaos. In Grisson et al. [GMA^{+}98], pages 44–45. ISBN 1-58113-046-5, 1-58113-048-1 (CD-ROM). LCCN T385 .S54 1998b.

URL http://www.acm.org:80/pubs/citations/proceedings/

graph/280953/p44-laxer/. ACM order number for CD-ROM: 435984. ACM order number: 435983.

Lee:1997:SSA

[LKH97] Choong Ho Lee, M. Kawamata, and T. Higuchi. State-space approach to roundoff error analysis of fractal image coding. In Proceedings of 1997 IEEE International Symposium on Circuits and Systems: ISCAS ’97, 9–12 June 1997, volume 2, pages 1341– 1344. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. CODEN ???? ISSN ????

Lee:1997:HRF

[LL97] Shyi-Long Lee and Chung-Kung Lee. Heterogeneous reactions over fractal surfaces: a multifractal scaling analysis. International Journal of Quantum Chemistry, 64(3):337–350, ???? 1997. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract?ID=42691;   http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=42691&PLACEBO=IE.pdf. Special Is-

sue: Second Triennial Congress of the International Society for Theoretical Chemical Physics. Issue Edited by Seán P. McGlynn, Kresimir Rupnik, Janos Ladik.

Loew:1996:APM

[LLP96] Murray H. Loew, Dunling Li, and Raymond L. Pickholtz. Adaptive PIFS model in fractal image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2707 (??):284–293, ???? 1996. CODEN PSISDG. ISSN ????

Lovejoy:1985:FPR

[LM85] S. Lovejoy and B. B. Mandelbrot. Fractal properties of rain, and a fractal model. Tellus, 37A(??):209–232, ???? 1985. CODEN ???? ISSN ????

Weibel:2005:MFGb

[LMNW05] Gabriele A. Losa, Danilo Merlini, Theo F. Nonnenmacher, and Ewald R. Weibel, editors. Mandelbrot’s fractals and the geometry of life: A tribute to Benoît Mandelbrot on his 80(th) birthday, volume IV of Fractals in Biology and Medicine. Birkhäuser Verlag, Basel, Switzerland, 2005. ISBN 3-7643-7172-2 (hardcover). LCCN ????

Laibowitz:1985:FAM

[LMP85] Robert B. Laibowitz, Benoît B. Mandelbrot, and Dann E. Passoja, editors. Fractal aspects of materials: extended abstracts, volume EA–6 of Materials Research Society extended abstracts. Materials Research Society, Pittsburgh, PA, USA, 1985. LCCN QA447 .F72 1985.

Lulek:1996:GIF

[LO96] T. Lulek and R. Olchawa. Geometric interpretation of fractal symmetries of a finite linear chain. Journal of Mathematical Physics, 37(5):2472–2483, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Loyless:1991:ARA

[Loy91] James E. Loyless. Autumn — A recipe for artistic fractal images. Computers and Graphics, 15(1):87–88, ???? 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Liao:2008:FBR

[LPZ08] Iman Yi Liao, Maria Petrou, and Rongchun Zhao. A fractal-based relaxation algorithm for shape from terrain image. Computer Vision and Image Understanding: CVIU, 109(3):227–243, March 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Lewis:1985:FSP

[LR85] M. Lewis and D. C. Rees. Fractal surfaces of proteins. Science, 230(4730):1163–1165, December 6, 1985. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/230/4730/1163.full.pdf.

Liu:2000:LTM

[LR00] Quansheng Liu and Alain Rouault. Limit theorems for Mandelbrot’s multiplicative cascades. Annals of applied probability, 10 (1):218–239, February 2000. CODEN ???? ISSN 1050-5164. URL

http://www.jstor.org/stable/2667193.

Luo:1992:FCL

[LS92] Xiaochun Luo and David N. Schramm. Fractals and cosmological large-scale structure. Science, 256(5056):513–515, April 24, 1992. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/256/
5056/513.full.pdf.

Lovejoy:1986:FCI

[LSL86] S. Lovejoy, D. Schertzer, and P. Ladoy. Fractal characterization of inhomogeneous geophysical measuring networks. Nature, 319 (6048):43–44, January 2, 1986. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v319/n6048/pdf/319043a0.pdf.

Lin:1996:PLF

[LV96] Huawu Lin and Anastasios N. Venetsanopoulos. Perceptually lossless fractal image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2727(??): 1394–1399, ???? 1996. CODEN PSISDG. ISSN ????

LaTorre:2009:MVI

[LVEB09] D. La Torre, E. R. Vrscay, M. Ebrahimi, and M. F. Barnsley. Measure-valued images, associated fractal transforms, and the affine self-similarity of images. SIAM Journal on Imaging Sciences, 2(2):470–507, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Lapidus:2004:FGAa

[LvF04a] Michel L. Lapidus and Machiel van Frankenhuijsen, editors. Fractal geometry and applications: a jubilee of Benoît Mandelbrot.

Part 1. Proceedings of a Special Session of the Annual Meeting of the American Mathematical Society held in San Diego, CA, January 2002, volume 72 of Proceedings of Symposia in Pure Mathematics. American Mathematical Society, Providence, RI, USA, 2004. ISBN 0-8218-3637-4 (part 1), 0-8218-3292-1 (set). LCCN QA325 .F73 2004. Analysis, number theory, and dynamical systems.

Lapidus:2004:FGAb

[LvF04b] Michel L. Lapidus and Machiel van Frankenhuijsen, editors. Fractal geometry and applications: a jubilee of Benoît Mandelbrot. Part 2. Proceedings of a Special Session of the Annual Meeting of the American Mathematical Society held in San Diego, CA, January 2002, volume 72 of Proceedings of Symposia in Pure Mathematics. American Mathematical Society, Providence, RI, USA, 2004. ISBN 0-8218-3638-2 (part 2), 0-8218-3292-1 (set). LCCN QA325 .F73 2004. Multifractals, probability and statistical mechanics, applications.

Mandelbrot:1985:SIC

[M^{+}85] Benoît B. Mandelbrot et al. Schönheit im Chaos: Bilder aus der theorie komplexer Systeme = Frontiers of Chaos: computer graphics face complex dynamics. MAP ART, Forschungsgruppe Komplexe Dynamik, Universität Bremen, Bremen, West Germany, 1985. ISBN 3-920699-65-3. 108 pp. LCCN QA614.86 .S342 1985.

Mandelbrot:1979:FFC

[MA79] B. B. Mandelbrot and Michael Aizenman. Fractals: Form, chance, and dimension. Physics Today, 32(5):65–??, May 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v32/
i5/p65/s1.

Maccari:2008:CFC

[Mac08] Attilio Maccari. Chaotic, fractal, and coherent solutions for a new integrable system of equations in 2 + 1 dimensions. Journal of Mathematical Physics, 49(2):022702, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i2/
p022702_s1.

Maddox:1986:GWF

[Mad86] John Maddox. Gentle warning on fractal fashions. Nature, 322 (6077):303, July 24, 1986. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v322/n6077/pdf/322303a0.pdf.

Maddox:1987:UFS

[Mad87] John Maddox. The Universe as a fractal structure. Nature, 329 (6136):195, September 17, 1987. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v329/n6136/pdf/329195a0.pdf.

Magid:2004:RMB

[Mag04] Aaron Magid. Review: The mystery behind the Mandelbrot set. Math Horizons, 12(2):27, November 2004. CODEN ???? ISSN

???? URL http://www.jstor.org/stable/25678516.

Mainster:1990:FPR

[Mai90] Martin A. Mainster. The fractal properties of retinal vessels: Embryological and clinical implications. Eye, 4(1):235–241, January 1990. CODEN EYEEEC. ISSN 0950-222X (print), 1476-5454

(electronic). URL http://www.nature.com/eye/journal/v4/
n1/pdf/eye199033a.pdf.

Majewski:1998:TRV

[Maj98] Mirek Majewski. A tutorial on the realistic visualization of 3D Sierpiński fractals. Computers and Graphics, 22(1):129–142, February 25, 1998. CODEN COGRD2. ISSN 0097-8493 (print),

1873-7684 (electronic). URL http://www.elsevier.com/cas/
tree/store/cag/sub/1998/22/1/519.pdf.

Mandelbrot:1949:SPP

[Man49] Benoît Mandelbrot. Some potential problems arising in the theory of axial turbomachines. Masters of Science in Aeronautics, California Institute of Technology, Pasadena, CA, USA, 1949. 42 pp.

Mandelbrot:1951:ADMa

[Man51a] Benoît Mandelbrot. Adaptation du message à la ligne de transmission. I. Quanta d’information. (French) [Adaptation of the message to the transmission line. I. Quanta of information].

Comptes rendus de l’Académie des sciences, Paris, 232(??):1638– 1640, ???? 1951. CODEN ???? ISSN ????

Mandelbrot:1951:ADMb

[Man51b] Benoît Mandelbrot. Adaptation du message à la ligne de transmission. II. Interprétations physiques. (French) [Adaptation of the message to the transmission line. II. Physical interpretation]. Comptes rendus de l’Académie des sciences, Paris, 232(??):2003– 2005, ???? 1951. CODEN ???? ISSN ????

Mandelbrot:1952:CTM

[Man52a] Benoît Mandelbrot. Contribution à la théorie mathématique des communications. (French) [Contribution to the mathematical theory of communications]. Docteur d’État ès Sciences Mathématiques, Faculté des Sciences de l’Université de Paris,

Paris, France, 1952. URL http://genealogy.math.ndsu.
nodak.edu/id.php?id=60791.

Mandelbrot:1952:DMF

[Man52b] Benoît Mandelbrot. Les démons de Maxwell. (French) [Maxwell’s demons]. Comptes rendus de l’Académie des sciences, Paris, 234 (??):1842–1844, ???? 1952. CODEN ???? ISSN ????

Mandelbrot:1952:NGI

[Man52c] Benoît Mandelbrot. Sur la notion générale d’information et la durée intrinsèque d’une stratégie. (French) [On the general notion of information and the intrinsic duration of a strategy]. Comptes rendus de l’Académie des sciences, Paris, 234(??):1345– 1347, ???? 1952. CODEN ???? ISSN ????

Mandelbrot:1953:CTM

[Man53a] Benoît Mandelbrot. Contribution à la théorie mathématique des jeux de communication. (French) [Contribution to the mathematical theory of communication games]. Publications de l’Institut de statistique de l’Université de Paris, 2(1–2):124, ???? 1953. CODEN ???? ISSN ????

Mandelbrot:1953:ITS

[Man53b] Benoît B. Mandelbrot. An informational theory of the statistical structure of languages. In Willis Jackson, editor, Communication theory: papers read at a Symposium on “Applications of Communication Theory” held at the Institution of Electrical Engineers, London, September 22nd–26th 1952, pages 486–502. Butterworths, London, UK, 1953. LCCN Q350 .S92

1952. URL http://lccn.loc.gov/53004215;   http://www.
nytimes.com/2010/10/17/us/17mandelbrot.html.

Mandelbrot:1954:CSF

[Man54a] Benoît Mandelbrot. Corrections: Structure formelle des textes et communications [deux études]. (French) [Corrections: Formal structure of text and communications [two studies]]. Word, 11 (??):424–??, ???? 1954. CODEN ???? ISSN ???? See [Man54c].

Mandelbrot:1954:SGS

[Man54b] Benoît Mandelbrot. Simple games of strategy occurring in communication through natural languages. Transaction of the IRE Professional Group on Information Theory PGIT, 3(3):124–137, ???? 1954. CODEN ???? ISSN 2168-2690.

Mandelbrot:1954:SFT

[Man54c] Benoît Mandelbrot. Structure formelle des textes et communications [deux études]. (French) [Formal structure of text and communications [two studies]]. Word, 10(??):1–27, ???? 1954. CO-DEN ???? ISSN ???? See corrections [Man54a].

Mandelbrot:1955:RNL

[Man55a] Benoît Mandelbrot. On recurrent noise limiting coding. In Proceedings of the symposium on information networks. New York, April, 1954, pages 205–221. Polytechnic Institute of Brooklyn, Brooklyn, NY, USA, 1955.

Mandelbrot:1955:TPC

[Man55b] Benoît Mandelbrot. Théorie de la pré-correction des erreurs de transmission (théorème de A. Feinstein). (French) [Theory of the precorrection of transmission errors (theorem of A. Feinstein)]. Annales des Télécommunications, 10(??):122–134, ???? 1955. CODEN ANTEAU. ISSN 0003-4347 (print), 1958-9395

(electronic). URL http://www.springerlink.com/content/
x45p11282n5j9327/fulltext.pdf.

Mandelbrot:1956:EET

[Man56a] Benoît Mandelbrot. Exhaustivité de l’énergie totale d’un système en équilibre, pour l’estimation de sa temperature. (French) [exhaustivity of the total energy of a system in equilibrium for the estimation of its temperature]. Comptes rendus de l’Académie des sciences, Paris, 243(??):1835–1838, ???? 1956. CODEN ???? ISSN ????

Mandelbrot:1956:DWY

[Man56b] Benoît Mandelbrot. La distribution de Willis–Yule, relative aux nombres d’espèces dans les genres biologiques. (French) [The Willis–Yule distribution, relative to the number of species of biological types]. Comptes rendus de l’Académie des sciences, Paris, 242(??):2223–2226, ???? 1956. CODEN ???? ISSN ????

Mandelbrot:1956:LTO

[Man56c] Benoît Mandelbrot. On the language of taxonomy: an outline of a thermo-statistical theory of systems of categories, with Willis (natural) structure. In C. Cherry, editor, Information Theory: the Third London Symposium, pages 135–145. Academic Press, New York, NY, USA, 1956. LCCN ????

Mandelbrot:1956:PPT

[Man56d] Benoît Mandelbrot. A purely phenomenological theory of statistical thermodynamics: I. Canonical ensembles. IRE Transactions on Information Theory, 2(3):190–203, ???? 1956. CODEN IRI-TAY. ISSN 0096-1000.

Mandelbrot:1957:ISV

[Man57a] Benoît Mandelbrot. Der Ingenieur als Stratege: Verhaltenstheorien. (Eine Definition der Kybernetik und ihre Anwendung in der Linguistik). (German) [The engineer as a strategist: behavioral theories. (A definition of cybernetics and its application in linguistics)]. Nachrichtentechnische Fachberichte, 3(??):32–39, ???? 1957. CODEN ???? ISSN ????

Mandelbrot:1957:NLB

[Man57b] Benoît Mandelbrot. A note on a law of Berry and on insistence stress. Information and Control, 1(1):76–81, September 1957. CODEN IFCNA4. ISSN 0019-9958 (print), 1878-2981

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0019995857900943.

Mandelbrot:1957:LSM

[Man57c] Benoît B. Mandelbrot. Linguistique statistique macroscopique. Théorie mathématique de la loi de Zipf (basée sur les méthodes de la théorie de l’information et de la thermodynamique). (French) [Statistical macroscopic linguistics. Mathematical theory of Zipf’s Law (based on methods from the theory of information and from thermodynamics)]. Institut Henri Poincaré, Séminaire de calcul des probabilités, Paris, France, 1957. 58 pp. LCCN ????

Mandelbrot:1959:SML

[Man59a] B. Mandelbrot. Statistical macro-linguistics. Il Nuovo Cimento (10), 13(2 (supplemento)):518–520, ???? 1959. CODEN NU-CIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL

http://www.springerlink.com/content/xt67n3181h87336t/.

Mandelbrot:1959:ISI

[Man59b] Benoît Mandelbrot. Information sans interpretation dans la description des langues Réelles. (French) [Information without interpretation in the description of real languages]. Synthese, 11(2):160–166, June 1959. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/10.1007/BF00485577;   http:
//www.jstor.org/stable/20114289;   http://www.springerlink.
com/content/r10332532461x0r7/.

Mandelbrot:1959:EGC

[Man59c] Benoît Mandelbrot. Les ensembles grand-canoniques de Gibbs; justification de leur unicité, basée sur la divisibilité infinie de leur énergie aléatoire. (French) [Gibbs grand canonical ensembles, based on the infinite divisibility of their random energy]. Comptes rendus de l’Académie des sciences, Paris, 249(??):1464– 1466, ???? 1959. CODEN ???? ISSN ????

Mandelbrot:1959:NCS

[Man59d] Benoît Mandelbrot. A note on a class of skew distribution functions: Analysis and critique of a paper by H. A. Simon. Information and Control, 2(1):90–99, April 1959. CO-DEN IFCNA4. ISSN 0019-9958 (print), 1878-2981 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0019995859900981. See [Sim55].

Mandelbrot:1959:VPSa

[Man59e] Benoît Mandelbrot. Variables et processus stochastiques de Pareto–Lévy, et la répartition des revenus. (French) [Pareto– Lévy variables and stochastics processes]. Comptes rendus de l’Académie des sciences, Paris, 249(??):613–615, ???? 1959. CO-DEN ???? ISSN ????

Mandelbrot:1959:VPSb

[Man59f] Benoît Mandelbrot. Variables et processus stochastiques de Pareto–Lévy, et la répartition des revenus. (French) [Pareto–Lévy variables and stochastics processes, and the repartition of incomes]. Comptes rendus de l’Académie des sciences, Paris, 249 (??):2153–2155, ???? 1959. CODEN ???? ISSN ????

Mandelbrot:1960:BRJa

[Man60a] Benoît Mandelbrot. Book review: John Chadwick, The Decipherment of Linear B (1958) Cambridge University Press. Information and Control, 3(1):95–96, March 1960. CODEN IFCNA4. ISSN 0019-9958 (print), 1878-2981 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0019995860903478.

Mandelbrot:1960:BRJb

[Man60b] Benoît Mandelbrot. Book review: John G. Kemeny and J. Laurie Snell, Finite Markov Chains (1960) van Nostrand, New York, viii + 210 pp., $5.00. Information and Control, 3(2):205–206, June 1960. CODEN IFCNA4. ISSN 0019-9958 (print), 1878-2981

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0019995860908093.

Mandelbrot:1960:BRM

[Man60c] Benoît Mandelbrot. Book review: Mark Kac, Statistical Independence in Probability, Analysis and Number Theory $3.00, Carus Mathematical Monograph No. 12 (1959), John Wiley, Cambridge, Massachusetts 93 pp. Information and Control, 3(1): 98–100, March 1960. CODEN IFCNA4. ISSN 0019-9958 (print),

1878-2981 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S00199958609076.

Mandelbrot:1960:BRR

[Man60d] Benoît Mandelbrot. Book review: Robert V. Hogg and Allen B. Craig, Introduction to Mathematical Statistics (from a series of mathematics texts under the general editorship of C. B. Allendoerfer), x + 245 pp., $6.75 (1959) Macmillan, New York. Information and Control, 3(2):207–208, June 1960. CO-DEN IFCNA4. ISSN 0019-9958 (print), 1878-2981 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0019995860908226.

Mandelbrot:1960:BRB

[Man60e] Benoît Mandelbrot. Book review: S. Bochner, Lectures on Fourier Integrals, (with an Author’s Supplement on Monotonic Functions, Stieltjes Integrals and Harmonic Analysis). (Translated from the original 1932 German edition by M. Tenenbaum and H. Pollard) $5.00. Annals of Mathematics Study No. 421959, Princeton University Press New York, 333 pp., $5.00. Information and Control, 3(1):94–95, March 1960. CO-DEN IFCNA4. ISSN 0019-9958 (print), 1878-2981 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0019995860903326.

Mandelbrot:1960:PLL

[Man60f] Benoît Mandelbrot. The Pareto–Lévy law and the distribution of income. International Economic Review, 1(2):79–106, May 1960. CODEN INERAE. ISSN 0020-6598 (print), 1468-2354 (elec-

tronic). URL http://www.ingenta.com/journals/browse/
bpl/iere;   http://www.jstor.org/stable/2525289.

Mandelbrot:1960:PSL

[Man60g] Benoît Mandelbrot. Processus stochastiques à loi stable positive, permanents, markoviens, stationnaires (non additifs). (French) [Stochastic processes with positive stable law, permanent, Markov, stationary (non-additive)]. Comptes rendus de l’Académie des sciences, Paris, 250(??):451–453, ???? 1960. CO-DEN ???? ISSN ????

Mandelbrot:1960:TDI

[Man60h] Benoît Mandelbrot. Teoria della informazione. Information and Control, 3(4):384–386, December 1960. CODEN IFCNA4. ISSN 0019-9958 (print), 1878-2981 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0019995860910457.

Mandelbrot:1961:BRK

[Man61a] Benoît Mandelbrot. Book review: A. N.Kolmogoroff and W. M.Tichomirow, Entropy and Capacity of Sets in Function Spaces. Translated by P. Franken and K. Nawrotzki from the Russian original which appeared in 1960. 80 pp., price unknown. Text in German: Arbeiten zur Informationstheorie, III Entropid und Kapazität von Mengen in Funktionalraümen. Uspekhi Mat. Nauk 14, 1959 Deutscher Verlag der Wissenschaften 386, 1960. Information and Control, 4(2–3):227–228, September 1961. CO-DEN IFCNA4. ISSN 0019-9958 (print), 1878-2981 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0019995861800120.

Mandelbrot:1961:BRBa

[Man61b] Benoît Mandelbrot. Book review: Contributions to Probability and Statistics: Essays in Honor of Harold Hotelling (Ingram Olkin, Sudhist G. Ghurye, Wassily Hoeffding, William G. Madow, and Henry B. Mann, eds.). SIAM Review, 3(1):80, January 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200

(electronic). URL http://link.aip.org/link/siread/v3/i1/
p80/s1;   http://www.jstor.org/stable/2027260.

Mandelbrot:1961:BRBc

[Man61c] Benoît Mandelbrot. Book review: Information and Decision Processes (Robert E. Machol, ed.). SIAM Review, 3(2):182–183, April 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200

(electronic). URL http://link.aip.org/link/siread/v3/i2/
p182/s1;   http://www.jstor.org/stable/2027402.

Mandelbrot:1961:BRBd

[Man61d] Benoît Mandelbrot. Book review: La Theorie Physique au Sens de Boltzmann (René Dugas). (French) [The Physical Theory According to Boltzmann (René Dugas)]. SIAM Review, 3(2):183– 184, April 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-

7200 (electronic). URL http://link.aip.org/link/siread/
v3/i2/p183/s1;   http://www.jstor.org/stable/2027403.

Mandelbrot:1961:BRBb

[Man61e] Benoît Mandelbrot. Book review: Mathematical Methods in the Social Sciences, 1959, Proceedings of the First Stanford Symposium (Kenneth J. Arrow, Samuel Karlin, and Patrick Suppes, eds.). SIAM Review, 3(1):80–81, January 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://link.aip.org/link/siread/v3/i1/p80/
s2;   http://www.jstor.org/stable/2027261.

Mandelbrot:1961:BRG

[Man61f] Benoît Mandelbrot. Book review: Gustav Herdan, Type Token Mathematics, (Janua Linguarum, Series Maior, IV) (+ 110 pp. of indices and tables of functions). 54 Dutch GL, approx. $14.00. A Textbook of Mathematical Linguistics 1960 Mouton & Cie, New York and London, iv + 110 pp. of indices and tables of functions. Information and Control, 4(1):83–85, March 1961. CO-DEN IFCNA4. ISSN 0019-9958 (print), 1878-2981 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0019995861800417.

Mandelbrot:1961:CDH

[Man61g] Benoît Mandelbrot. Comment on Dr. Herdan’s letter. Information and Control, 4(2–3):241, September 1961. CO-DEN IFCNA4. ISSN 0019-9958 (print), 1878-2981 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S001999586180017X.

Mandelbrot:1961:FNC

[Man61h] Benoît Mandelbrot. Final note on a class of skew distribution functions: Analysis and critique of a model due to H. A. Simon. Information and Control, 4(2–3):198–216, September 1961. CO-DEN IFCNA4. ISSN 0019-9958 (print), 1878-2981 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0019995861800089.

Mandelbrot:1961:TWF

[Man61i] Benoît Mandelbrot. On the theory of word frequencies and on related Markovian models of discourse. In R. Jakobson, editor, Structures of language and its mathematical aspects, pages 120– 219. American Mathematical Society, Providence, RI, USA, 1961. LCCN ????

Mandelbrot:1961:PSF

[Man61j] Benoît Mandelbrot. Post scriptum to “Final note”. Information and Control, 4(2–3):300–304, September 1961. CO-DEN IFCNA4. ISSN 0019-9958 (print), 1878-2981 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0019995861800259.

Mandelbrot:1961:SPR

[Man61k] Benoît Mandelbrot. Stable Paretian random functions and the multiplicative variation of income. Econometrica, 29(4):517–543, October 1961. CODEN ECMTA7. ISSN 0012-9682. URL http:

//www.jstor.org/stable/1911802.

Mandelbrot:1962:BRB

[Man62a] Benoît Mandelbrot. Book review: B. V. Gnedenko, The Theory of Probability, translated by B. D. Seckler, $8.75, 2nd ed.

1962, Chelsea, New York 459. Information and Control, 5(2): 173, June 1962. CODEN IFCNA4. ISSN 0019-9958 (print),

1878-2981 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0019995862903340.

Mandelbrot:1962:PDI

[Man62b] Benoît Mandelbrot. Paretian distributions and income maximization. The Quarterly Journal of Economics, 76(1):57–85, February 1962. CODEN QJECAT. ISSN 0033-5533 (print), 1531-4650

(electronic). URL http://www.jstor.org/stable/1891131.

Mandelbrot:1962:RSE

[Man62c] Benoît Mandelbrot. The role of sufficiency and of estimation in thermodynamics. Annals of Mathematical Statistics, 33(3): 1021–1038, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://www.jstor.

org/stable/2237878.

Mandelbrot:1962:SNR

[Man62d] Benoît Mandelbrot. Statistics of natural resources and the law of Pareto. Research Note NC-146, IBM, Yorktown Heights, NY, USA, June 29, 1962.

Mandelbrot:1962:CPS

[Man62e] Benoît Mandelbrot. Sur certains prix spéculatifs: faits empiriques et modèle basé sur les processus stables additifs non gaussiens de Paul Lévy. (French) [On certain speculative prices: empirical evidence and model-based additive non-Gaussian stable processes of Paul Lévy]. Comptes rendus de l’Académie des sciences, Paris, 254(??):3968–3970, ???? 1962. CODEN ???? ISSN ????

Mandelbrot:1963:BRJ

[Man63a] Benoît Mandelbrot. Book review: Joseph P. La Salle, Solomon Lefschetz, Editors, Recent Soviet Contributions to Mathematics (1962) Macmillan, 324 pp. Information and Control, 6(3):310– 311, September 1963. CODEN IFCNA4. ISSN 0019-9958 (print),

1878-2981 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0019995863904182.

Mandelbrot:1963:NMSa

[Man63b] Benoît Mandelbrot. New methods in statistical economics. Journal of Political Economy, 71(5):421–440, October 1963. CODEN

???? ISSN ???? URL http://www.jstor.org/stable/1829014.

Reprinted in [Man63c].

Mandelbrot:1963:NMSb

[Man63c] Benoît Mandelbrot. New methods in statistical economics. Bulletin of the International Statistical Institute, Ottawa Session, 40 (2):669–720, ???? 1963. CODEN ???? ISSN ???? Reprint of [Man63b].

Mandelbrot:1963:SPI

[Man63d] Benoît Mandelbrot. The stable Paretian income distribution when the apparent exponent is near two. International Economic Review, 4(1):111–115, January 1963. CODEN INERAE. ISSN 0020-6598 (print), 1468-2354 (electronic). URL http://

www.jstor.org/stable/2525463.

Mandelbrot:1963:VCSa

[Man63e] Benoît Mandelbrot. The variation of certain speculative prices. The Journal of Business at the University of Chicago, 36(4):394– 419, October 1963. CODEN ???? ISSN 0021-9398 (print),

1537-5374 (electronic). URL http://www.jstor.org/stable/

2350970. Reprinted in [Man63f].

Mandelbrot:1963:VCSb

[Man63f] Benoît Mandelbrot. The variation of certain speculative prices. In P. H. Cootner, editor, The random character of stock market prices, pages 297–337. MIT Press, Cambridge, MA, USA, 1963. LCCN ???? Reprint of [Man63e].

Mandelbrot:1964:NMS

[Man64a] B. Mandelbrot. New methods in statistical economics. Bull. Inst. Int. Stat., 40(??):699–721, ???? 1964. CODEN ???? ISSN ????

Mandelbrot:1964:ECC

[Man64b] Benoît Mandelbrot. The epistemology of chance in certain newer sciences. Read at The Jerusalem International Congress on Logic, Methodology and the Philosophy of Science, 1964.

Mandelbrot:1964:DST

[Man64c] Benoît Mandelbrot. On the derivation of statistical thermodynamics from purely phenomenological principles. Journal of Mathematical Physics, 5(2):164–171, February 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v5/i2/
p164_s1;   http://link.aip.org/link/jmapaq/v5/i2/p164/
s1.

Mandelbrot:1964:RWF

[Man64d] Benoît Mandelbrot. Random walks, fire damage amount and other Paretian risk phenomena. Operations Research, 12(4):582– 585, August 1964. CODEN OPREAI. ISSN 0030-364X (print),

1526-5463 (electronic). URL http://www.jstor.org/stable/
167704.

Mandelbrot:1964:SSR

[Man64e] Benoît Mandelbrot. Self-similar random processes and the range. Research Report RC-1163, IBM, Yorktown Heights, NY, USA, April 13, 1964.

Mandelbrot:1965:ITP

[Man65a] Benoît Mandelbrot. Information theory and psycholinguistics. In B. B. Wolman and E. N. Nagel, editors, Scientific Psychology: Principles and Approaches, pages 550–562. Basic Books, New York, NY, USA, 1965. LCCN ????

Mandelbrot:1965:LSU

[Man65b] Benoît Mandelbrot. Leo Szilard and unique decipherability. IEEE Transactions on Information Theory, IT-11(3):455–456, ???? 1965. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Mandelbrot:1965:SSE

[Man65c] Benoît Mandelbrot. Self similar error clus ters in communications systems and the concept of conditional stationarity. IEEE Transactions on Communications Technology, 13(1):71–90, ???? 1965. CODEN ???? ISSN ????

Mandelbrot:1965:CPS

[Man65d] Benoît Mandelbrot. Une classe de processus stochastiques homothétiques à soi; application à la loi climatologique H. E. Hurst. (French) [A class of stochastic self-homothetic processes]. Comptes rendus de l’Académie des sciences, Paris, 260(??):3274– 3277, ???? 1965. CODEN ???? ISSN ????

Mandelbrot:1966:FFP

[Man66a] Benoît Mandelbrot. Forecasts of future prices, unbiased markets, and “martingale” models. The Journal of Business at the University of Chicago, 39(1):242–255, January 1966. CODEN ???? ISSN 0021-9398 (print), 1537-5374 (electronic). URL http://

www.jstor.org/stable/2351745.

Mandelbrot:1966:ITPa

[Man66b] Benoît Mandelbrot. Information theory and psycholinguistics. In R. C. Oldfield and J. C. Marshall, editors, Language, Selected Readings, page ?? Penguin, New York, NY, USA, 1966. LCCN ????

Mandelbrot:1966:ITPb

[Man66c] Benoît Mandelbrot. Information theory and psycholinguistics. In P. Lazarsfeld and N. Henry, editors, Readings in Mathematical Social Science, page ?? Science Research Associates, Chicago, IL, USA, 1966. LCCN ???? Paperback edition issued in 1968.

Mandelbrot:1966:NMV

[Man66d] Benoît Mandelbrot. Nouveaux modéles de la variation des prix (cycles lents et changements instantanés). (French) [New models of price changes (slow cycles and instantaneous changes)]. Cahiers du Séminaire d’Économétrie, ??(9):53–66, ???? 1966.

CODEN ???? ISSN ???? URL http://www.jstor.org/stable/
20075411.

Mandelbrot:1967:SNS

[Man67a] B. Mandelbrot. Some noises with 1/f spectrum, a bridge between direct current and white noise. IEEE Transactions on Information Theory, 13(2):289–298, April 1967. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Mandelbrot:1967:HLC

[Man67b] Benoît Mandelbrot. How long is the coast of Britain? Statistical self-similarity and fractional dimension. Science, 156(3775):636– 638, May 5, 1967. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
1721427;   http://www.sciencemag.org/content/156/3775/
636.full.pdf.

Mandelbrot:1967:CCD

[Man67c] Benoît Mandelbrot. Les constantes chiffrées du discours. (French) [the constant figures of speech]. In J. Martinet, editor, Encyclopédie de la Pléiade: Linguistique, pages 46–56. Gallimard, Paris, France, 1967. LCCN ????

Mandelbrot:1967:SRF

[Man67d] Benoît Mandelbrot. Sporadic random functions and conditional spectral analysis: Self-similar examples and limits. In Proceedings of the Fifth Berkeley Symposium on Mathematical Statistics and Probability (Berkeley, California, 1965/66), Vol. III: Physical Sciences, pages 155–179. University of California Press, Berkeley, CA, USA, 1967.

Mandelbrot:1967:ST

[Man67e] Benoît Mandelbrot. Sporadic turbulence. Physics of Fluids, 10(??):5302–5303, ???? 1967. CODEN PHFLE6. ISSN

1070-6631. URL http://link.aip.org/link/pfldas/v10/i9/

pS302/s2. Proceedings: Boundary Layers and Turbulence (Kyoto International Symposium, 1966).

Mandelbrot:1967:EDH

[Man67f] Benoît Mandelbrot. Sur l’épistémologie du hasard dans les sciences sociales: invariance des lois et verification des hypotheses. In J. Piaget, editor, Encyclopédie de la Pléiade: Logique et Connaissance Scientifique, pages 1097–1113. Gallimard, Paris, France, 1967.

Mandelbrot:1967:VSO

[Man67g] Benoît Mandelbrot. The variation of some other speculative prices. The Journal of Business at the University of Chicago, 40(4):393–413, October 1967. CODEN ???? ISSN 0021-9398 (print), 1537-5374 (electronic). URL http://www.jstor.org/

stable/2351623.

Mandelbrot:1968:SAR

[Man68] Benoît Mandelbrot. Some aspects of the random walk model of stock market prices: Comment. International Economic Review, 9(2):258–259, June 1968. CODEN INERAE. ISSN 0020-6598 (print), 1468-2354 (electronic). URL http://www.jstor.org/

stable/2525479.

Mandelbrot:1969:LRL

[Man69] Benoît Mandelbrot. Long-run linearity, locally Gaussian process, H-spectra and infinite variances. International Economic Review, 10(1):82–111, February 1969. CODEN INERAE. ISSN 0020-6598 (print), 1468-2354 (electronic). URL http://www.jstor.org/

stable/2525574.

Mandelbrot:1970:NTD

[Man70a] Benoît Mandelbrot. On negative temperature for discourse. Discussion of a paper by Prof. N. F. Ramsey. In E. B. Stuart et al., editors, Critical Review of Thermodynamics, pages 230– 232. Mono Book, Baltimore, MD, USA, 1970. LCCN ????

Mandelbrot:1970:SDP

[Man70b] Benoît Mandelbrot. Statistical dependence in prices and interest rates. In ????, editor, Papers of the Second World Congress of the Econometric Society, Cambridge, England (8–14 Sept. 1970), page ?? Econometric Society, Cambridge, UK, 1970. LCCN ???? Proceedings published in Econometrica 39(4) 1–266 (1970).

Mandelbrot:1970:SSS

[Man70c] Benoît Mandelbrot. Statistical self similarity and very erratic chance fluctuations. Trumbull Lectures, Yale University., 1970.

Mandelbrot:1971:FFG

[Man71a] Benoît Mandelbrot. A fast fractional Gaussian noise generator. Water Resources Research, 7(??):543–553, ???? 1971. CODEN WRERAQ. ISSN 0043-1397 (print), 1944-7973 (electronic).

Mandelbrot:1971:LRN

[Man71b] Benoît Mandelbrot. Linear regression with non-normal error terms: a comment. Review of Economics and Statistics, 53(2): 205–206, May 1971. CODEN RECSA9. ISSN 0034-6535 (print),

1530-9142 (electronic). URL http://www.jstor.org/stable/
1925721.

Mandelbrot:1971:WCP

[Man71c] Benoît B. Mandelbrot. When can price be arbitraged efficiently? A limit to the validity of the random walk and martingale models. Review of Economics and Statistics, 53(3):225–236, August 1971. CODEN RECSA9. ISSN 0034-6535 (print), 1530-9142

(electronic). URL http://www.jstor.org/stable/1937966.

Mandelbrot:1972:BLP

[Man72a] Benoît Mandelbrot. Broken line process derived as an approximation to fractional noise. Water Resources Research, 8(??): 1354–1356, ???? 1972. CODEN WRERAQ. ISSN 0043-1397 (print), 1944-7973 (electronic).

Mandelbrot:1972:SMN

[Man72b] Benoît Mandelbrot. Statistical methodology for nonperiodic cycles: from the covariance to the R/S analysis. Annals of Economic and Social Measurement, 1(??):259–290, ???? 1972. CO-DEN ???? ISSN ????

Mandelbrot:1972:CEV

[Man72c] Benoît B. Mandelbrot. Correction of an error in “the variation of certain speculative prices” (1963). The Journal of Business at the University of Chicago, 45(4):542–543, October 1972. CODEN ???? ISSN 0021-9398 (print), 1537-5374 (electronic). URL http:

//www.jstor.org/stable/2351576.

Mandelbrot:1972:DCC

[Man72d] Benoît B. Mandelbrot. On Dvoretzky coverings for the circle. Zeitschrift für Wahrscheinlichkeitstheorie und verwandte Gebiete, 22(2):158–160, ???? 1972. CODEN ZWVGAA. ISSN 0044-

3719. URL http://link.springer.com/article/10.1007/
BF00532734.

Mandelbrot:1972:PRL

[Man72e] Benoît B. Mandelbrot. Possible refinement of the lognormal hypothesis concerning the distribution of energy dissipation in intermittent turbulence. In Murray Rosenblatt and Chester Murray Van Atta, editors, Statistical Models and Turbulence. Proceedings of a symposium held at the University of California, San Diego (La Jolla), July 15–21, 1971, volume 12 of Lecture Notes in Physics, pages 333–351. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. ISBN 0-387-05716-1 (New York), 3-387-05716-4 (Berlin). ISSN 0075-8450 (print), 1616-6361 (electronic). LCCN QC1 L47 no. 12. URL http://

www.springerlink.com/content/y0q6173605g21v8j/.

Mandelbrot:1972:RSR

[Man72f] Benoît B. Mandelbrot. Renewal sets and random cutouts. Probability theory and related fields, 22(2):145–157, ???? 1972. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://www.springerlink.com/content/w3423987566u0872/.

Mandelbrot:1973:FND

[Man73a] Benoît Mandelbrot. Formes nouvelles du hasard dans les sciences. (French) [New forms of chance in the sciences]. Économie Appliquée, 26(??):307–319, ???? 1973. CODEN ???? ISSN ????

Mandelbrot:1973:PRC

[Man73b] Benoît Mandelbrot. Le problème de la réalité des cycles lents, et le syndrome de Joseph. (French) [The problem of the reality of slow cycles, and the Joseph syndrome]. Économie Appliquée, 26 (??):349–365, ???? 1973. CODEN ???? ISSN ????

Mandelbrot:1973:SVI

[Man73c] Benoît Mandelbrot. Le syndrome de la variance infinie, et ses rapports avec la discontinuité des prix. (French) [The syndrome of infinite variance, and its relationship with price gaps]. Économie Appliquée, 26(??):321–348, ???? 1973. CODEN ???? ISSN ????

Mandelbrot:1973:CSS

[Man73d] Benoît B. Mandelbrot. Comments on: “A subordinated stochastic process model with finite variance for speculative prices” (Econometrica 41 (1973), no. 1, 135–155), by Peter K. Clark. Econometrica, 41(1):157–159, January 1973. CODEN ECMTA7. ISSN

0012-9682. URL http://www.jstor.org/stable/1913890.

Mandelbrot:1974:MAIa

[Man74a] Benoît Mandelbrot. Multiplications aléatoires itérées et distributions invariantes par moyenne pondérée aléatoire. (French) [Random iterated multiplications and invariant distributions by random weighted mean]. C. R. Acad. Sci. Paris Sér. A, 278(??): 289–292, ???? 1974. CODEN ???? ISSN ????

Mandelbrot:1974:MAIb

[Man74b] Benoît Mandelbrot. Multiplications aléatoires itérées et distributions invariantes par moyenne pondérée aléatoire: quelques extensions. (French) [Random iterated multiplications and invariant distributions by random weighted mean: some extensions]. C. R. Acad. Sci. Paris Sér. A, 278(??):355–358, ???? 1974. CODEN ???? ISSN ????

Mandelbrot:1974:PBM

[Man74c] Benoît Mandelbrot. A population birth-and-mutation process, I: Explicit distributions for the number of mutants in an old culture of bacteria. Journal of Applied Probability, 11(3):437–444, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212688.

Mandelbrot:1974:ITS

[Man74d] Benoît B. Mandelbrot. Intermittent turbulence in self-similar cascades: divergence of high moments and dimension of the carrier. Journal of Fluid Mechanics, 62(??):331–358, ???? 1974. CODEN JFLSA7. ISSN 0022-1120 (print), 1469-7645 (electronic).

Mandelbrot:1974:RBP

[Man74e] Benoît B. Mandelbrot. Review: Random Processes: Multiplicity Theory and Canonical Decomposition by Anthony Ephremides, John B. Thomas. American Scientist, 62(6):745–746, November 1974. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786

(electronic). URL http://www.jstor.org/stable/27845245.

Mandelbrot:1975:CCP

[Man75a] Benoît Mandelbrot. The conditional cosmographic principle and the fractional dimension of the universe. In Les objets fractals: Forme, hasard et dimension. (French) [Fractal objects: form, chance, and dimension] [Man75d], page ?? ISBN 2-08-210647-0, 2-08-211155-5. LCCN QA447 .M36 1984. Nouvelle Bibliothèque Scientifique.

Mandelbrot:1975:FAP

[Man75b] Benoît Mandelbrot. Fonctions aléatoires pluri-temporelles: approximation poissonienne du cas brownien et généralisations. (French) [Random multi-temporal functions: Poisson approximation of the Brownian case and generalizations]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences. Séries A et B, 280(??):A1075–A1078, ???? 1975. CODEN CHASAP. ISSN 0151-0509.

Mandelbrot:1975:IQC

[Man75c] Benoît Mandelbrot. Hasards et tourbillons: quatre contes à clef. (French) [Chance and turbulence: four stories in key]. Annales des Mines, ??(??):61–66, November 1975. CODEN ???? ISSN ????

Mandelbrot:1975:OFF

[Man75d] Benoît Mandelbrot. Les objets fractals: Forme, hasard et dimension. (French) [Fractal objects: form, chance, and dimension]. Flammarion Editeur, Paris, France, 1975. ISBN 2-08-210647-0, 2-08-211155-5. 190 + ii pp. LCCN QA447 .M36 1984. Nouvelle Bibliothèque Scientifique.

Mandelbrot:1975:MDU

[Man75e] Benoît Mandelbrot. Sur un modèle décomposable d’univers hiérarchisé: déduction des corrélations galactiques sur la sphère céleste. (French) [A decomposable model of a hierarchical universe: deduction of galactic correlations on the celestial sphere]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences. Séries A et B, 280(22):Aiii, A1551–A1554, ???? 1975. CODEN CHASAP. ISSN 0151-0509.

Mandelbrot:1975:CLT

[Man75f] Benoît B. Mandelbrot. Corrigendum: “Limit theorems on the self-normalized range for weakly and strongly dependent processes” (Z. Wahrscheinlichkeitstheorie Verw. Gebiete 31 (1974/ 75), 271–285). Zeitschrift für Wahrscheinlichkeitstheorie und verwandte Gebiete, 33(3):220, ???? 1975. CODEN ZWV-

GAA. ISSN 0044-3719. URL http://link.springer.com/
accesspage/article/10.1007/BF00534968. See [Man75h].

Mandelbrot:1975:LTSb

[Man75g] Benoît B. Mandelbrot. Limit theorem on the self-normalized range for weakly and strongly dependent process. Zeitschrift für Wahrscheinlichkeitstheorie und verwandte Gebiete, 33(3): 220, ???? 1975. CODEN ZWVGAA. ISSN 0044-3719.

URL http://link.springer.com/accesspage/article/10.
1007/BF00534968.

Mandelbrot:1975:LTSa

[Man75h] Benoît B. Mandelbrot. Limit theorems on the self-normalized range for weakly and strongly dependent processes. Zeitschrift für Wahrscheinlichkeitstheorie und verwandte Gebiete, 31(4):271– 285, ???? 1975. CODEN ZWVGAA. ISSN 0044-3719. URL

http://link.springer.com/article/10.1007/BF00532867.

Mandelbrot:1975:GHT

[Man75i] Benoît B. Mandelbrot. On the geometry of homogeneous turbulence, with stress on the fractal dimension of the iso-surfaces of scalars. Journal of Fluid Mechanics, 72(??):401–416, ???? 1975. CODEN JFLSA7. ISSN 0022-1120 (print), 1469-7645 (electronic).

Mandelbrot:1975:SME

[Man75j] Benoît B. Mandelbrot. Stochastic models for the Earth’s relief, the shape and the fractal dimension of the coastlines, and the number-area rule for islands. Proceedings of the National Academy of Sciences of the United States of America, 72(10): 3825–3828, October 1975. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.jstor.org/

stable/65184.

Mandelbrot:1976:GFT

[Man76a] Benoît Mandelbrot. Géométrie fractale de la turbulence. Dimension de Hausdorff, dispersion et nature des singularités du mouvement des fluides. (French) [Fractal geometry of turbulence. Hausdorff dimension, dispersion, and nature of the singularities in fluid motion]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences. Séries A et B, 282(2):Aii, A119–A120, ???? 1976. CODEN CHASAP. ISSN 0151-0509.

Mandelbrot:1976:ITF

[Man76b] Benoît Mandelbrot. Intermittent turbulence and fractal dimension: kurtosis and the spectral exponent 5/3 + B. In Turbulence and Navier–Stokes equations (Proceedings of a Conference, Université Paris-Sud, Orsay, 1975), volume 34(5–6) of Lecture Notes in Mathematics, pages 121–145. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., ???? 1976.

CODEN ???? ISSN ???? URL http://www.springerlink.com/
content/x8051gr03w584n68/.

Mandelbrot:1976:NDS

[Man76c] Benoît Mandelbrot. Note on the definition and the stationarity of fractional Gaussian noise. Journal of Hydrology, 30 (4):407–409, August 1976. CODEN JHYDA7. ISSN ????

URL http://www.sciencedirect.com/science/article/pii/
0022169476901220.

Mandelbrot:1977:FTA

[Man77a] Benoît B. Mandelbrot. Fractals and turbulence: attractors and dispersion. In Turbulence Seminar (University of California, Berkeley, California, 1976/1977), volume 615 of Lecture Notes in Mathematics, pages 83–93. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977.

Mandelbrot:1977:FFC

[Man77b] Benoît B. Mandelbrot. Fractals: form, chance, and dimension. W. H. Freeman and Company, New York, NY, USA, revised edition, 1977. ISBN 0-7167-0473-0, 0-7167-0474-9 (paperback). xvi + 365 pp. LCCN QA447 .M3613 1977a. Translated from the French.

Mandelbrot:1977:LVF

[Man77c] Benoît B. Mandelbrot. Lecture V fractals and turbulence: Attractors and dispersion. Turbulence Seminar, 131(1–2):83–93, ???? 1977. CODEN ???? ISSN ???? URL http://www.

springerlink.com/content/w7lk031338u81534/.

Mandelbrot:1977:POF

[Man77d] Benoît B. Mandelbrot. Physical objects with fractional dimension: seacoasts, galaxy clusters, turbulence and soap. Bulletin of the Institute of Mathematics and its Applications, 13(7-8):189– 196, ???? 1977. CODEN IMTABW. ISSN 0950-5628.

Mandelbrot:1978:GFS

[Man78a] B. B. Mandelbrot. Geometric facets of statistical physics: scaling and fractals. In D. Cabib et al., editors, Statistical physics— “STATPHYS 13” (Proc. Thirteenth IUPAP Conf. Statist. Phys., Technion—Israel Inst. Tech., Haifa, 1977), Part 1, volume 2 of Ann. Israel Phys. Soc., pages 225–233. Adam Hilger Ltd., Bristol, UK, 1978.

Mandelbrot:1978:CAA

[Man78b] Benoît Mandelbrot. Colliers aléatoires et une alternative aux promenades au hasard sans boucle: les cordonnets discrets et fractals. (French) [Necklaces and an alternative to loop-free random walks: discrete fractal cords]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences. Séries A et B, 286(20): A933–A936, ???? 1978. CODEN CHASAP. ISSN 0151-0509.

Mandelbrot:1978:FGT

[Man78c] Benoît Mandelbrot. The fractal geometry of trees and other natural phenomena. In Miles and Serra [MS78], pages 235–249. ISBN 3-540-08856-3 (Berlin), 0-387-08856-3 (New York). LCCN ????

Mandelbrot:1978:OFF

[Man78d] Benoît Mandelbrot. Les objets fractals. (French) [Fractal objects]. La Recherche, 9(??):1–13, ???? 1978. CODEN RCCHBV. ISSN 0029-5671 (print), 1625-9955 (electronic).

Mandelbrot:1979:CTN

[Man79a] Benoît Mandelbrot. Corrélations et texture dans un nouveau modèle d’univers hiérarchisé, basé sur les ensembles trémas. (French) [Correlations and texture in a new model of a hierarchical universe, based on [discrete?] sets]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences. Séries A et B, 288(1):A81–A83, ???? 1979. CODEN CHASAP. ISSN 0151-0509.

Mandelbrot:1979:DPF

[Man79b] Benoît B. Mandelbrot. Discussion paper: Fractals, attractors, and the fractal dimension. Annals of the New York Academy of Sciences, 316(1):463–464, February 1979. CODEN ANYAA9. ISSN 0077-8923 (print), 1749-6632 (electronic). Bifurcation theory and applications in scientific disciplines (Papers, Conference, New York, 1977).

Mandelbrot:1980:FAI

[Man80] Benoît B. Mandelbrot. Fractal aspects of the iteration of z → λz(1 − z) for complex λ and z. Annals of the New York Academy of Sciences, 357(1):249–259, December 1980. CODEN ANYAA9. ISSN 0077-8923 (print), 1749-6632 (electronic). International conference on nonlinear dynamics.

Mandelbrot:1981:FGM

[Man81a] Benoît Mandelbrot. Fractals and geometry with many scales of length. In Encyclopedia Britannica 1981 Yearbook of Science and the Future, pages 168–181. Encyclopedia Britannica, Chicago, IL, USA, 1981. ISBN ???? LCCN ????

Mandelbrot:1981:SSS

[Man81b] Benoît B. Mandelbrot. Scalebound or scaling shapes: a useful distinction in the visual arts and in the natural sciences. Leonardo (Oxford, England), 14(1):45–47, ???? 1981. CODEN LEONDP. ISSN 0024-094X (print), 1530-9282 (electronic). URL http://

www.jstor.org/stable/1574481.

Mandelbrot:1982:CCR

[Man82a] B. B. Mandelbrot. Comment on computer rendering of fractal stochastic models. Communications of the Association for Computing Machinery, 25(8):581–584, August 1982. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mandelbrot:1982:EEF

[Man82b] B. B. Mandelbrot. On an eigenfunction expansion and on fractional Brownian motions. Lettere Al Nuovo Cimento (1971–1985), 33(17):549–550, ???? 1982. CODEN LNUCAE. ISSN 0024-1318, 0375-930x, 1827-613x. URL http://www.

springerlink.com/content/d537931242363784/.

Mandelbrot:1982:MCP

[Man82c] Benoît Mandelbrot. Des monstres de Cantor et de Peano à la géométrie fractale de la nature. (French) [Cantor and Peano monsters in the fractal geometry of nature]. In ????, volume 29 of Collect. Points Sér. Sci., pages 226–251. Seuil, Paris, France, 1982.

Mandelbrot:1982:DSD

[Man82d] Benoît Mandelbrot. On discs and sigma discs, that osculate the limit sets of groups of inversions. The Mathematical Intelligencer, 4(??):??, ???? 1982. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Mandelbrot:1982:FGN

[Man82e] Benoît B. Mandelbrot. The fractal geometry of nature. W. H. Freeman and Company, New York, NY, USA, 1982. ISBN 0-7167-1186-9. v + 460 pp. LCCN QA447 .M25 1982; QA447 .M357 1982. Schriftenreihe für den Referenten. [Series for the Referee].

Mandelbrot:1983:EFM

[Man83a] Benoît B. Mandelbrot. An explicit fractal model of percolation clusters. In Percolation structures and processes, volume 5 of Ann. Israel Phys. Soc., pages 59–80. Adam Hilger Ltd., Bristol, UK, 1983.

Mandelbrot:1983:FGN

[Man83b] Benoît B. Mandelbrot. The Fractal Geometry of Nature. W. H. Freeman and Company, New York, NY, USA, 1983. ISBN 0-7167-1186-9. 468 + 1 pp. LCCN QA447 .M271 1983. Updated and augment edition of [Man83b], but with same ISBN.

Mandelbrot:1983:QMCa

[Man83c] Benoît B. Mandelbrot. On the quadratic mapping z → z^{2} − μ for complex μ and z: The fractal structure of its M set, and scaling. In Campbell and Rose [CR83], pages 224–239. ISBN 0-444-86727-9. LCCN QA843 .I57 1982.

Mandelbrot:1983:QMCb

[Man83d] Benoît B. Mandelbrot. On the quadratic mapping z → z^{2} − μ for complex μ and z: The fractal structure of its M set, and scaling. Physica D, Nonlinear phenomena, 7(1–3):224–239, May 1983. CODEN PDNPDT. ISSN 0167-2789 (print), 1872-8022

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0167278983901288. Order in chaos (Los Alamos,

NM, 1982).

Mandelbrot:1983:SIF

[Man83e] Benoît B. Mandelbrot. Self-inverse fractals osculated by sigmadiscs and the limit sets of inversion groups. The Mathematical Intelligencer, 5(2):9–17, June 1983. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic). URL http://www.

springerlink.com/content/b4782275k6415231/.

Mandelbrot:1984:APF

[Man84a] Benoît B. Mandelbrot. Additional perspectives on fractals. College Mathematics Journal, 15(2):115–119, March 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972759.

Mandelbrot:1984:CCS

[Man84b] Benoît B. Mandelbrot. Comment on coherent structures in fluids, fractals, and the fractal structure of flow singularities. In Turbulence and chaotic phenomena in fluids (Kyoto, 1983), pages 207– 208. North-Holland Publishing Co., Amsterdam, The Netherlands, 1984.

Mandelbrot:1984:CEB

[Man84c] Benoît B. Mandelbrot. Comment on the equivalence between fracton/spectral dimensionality, and the dimensionality of recurrence. Journal of Statistical Physics, 36(5– 6):541–543, September 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01012919;   http:
//www.springerlink.com/content/k046263428n04707/.

Mandelbrot:1984:EFS

[Man84d] Benoît B. Mandelbrot. Each fractal set has a unique fractal dimension. In Turbulence and chaotic phenomena in fluids (Kyoto, 1983), pages 203–206. North-Holland Publishing Co., Amsterdam, The Netherlands, 1984.

Mandelbrot:1984:FPS

[Man84e] Benoît B. Mandelbrot. Fractals in physics: Squig clusters, diffusions, fractal measures, and the unicity of fractal dimensionality. Journal of Statistical Physics, 34(5–6):895–930, March 1984. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01009448;   http://www.springerlink.com/content/xq15g34856h83h81/

.

Mandelbrot:1984:OFF

[Man84f] Benoît B. Mandelbrot. Les objets fractals: Forme, hasard et dimension. (French) [Fractal objects: form, chance, and dimension]. Nouvelle bibliothèque scientifique. Flammarion, Paris, France, second edition, 1984. 203 pp. LCCN QA447 .M36 1984.

Mandelbrot:1984:FGF

[Man84g] Benoît B. Mandelbrot. On fractal geometry, and a few of the mathematical questions it has raised. In Proceedings of the International Congress of Mathematicians, Vol. 1, 2 (Warsaw, 1983), pages 1661–1675. PWN, Warsaw, Poland, 1984.

Mandelbrot:1984:DIM

[Man84h] Benoît B. Mandelbrot. On the dynamics of iterated maps. VIII. The map z → λ(z + 1/z), from linear to planar chaos, and the measurement of chaos. In Chaos and statistical methods (Kyoto, 1983), volume 24 of Springer Ser. Synergetics, pages 32–41. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984.

Mandelbrot:1984:SSS

[Man84i] Benoît B. Mandelbrot. Squig sheets and some other squig fractal constructions. Journal of Statistical Physics, 36(5–6):519– 539, ???? 1984. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://www.springerlink.com/
content/k20w5q31555x03nj/.

Mandelbrot:1985:DIMa

[Man85a] Benoît B. Mandelbrot. On the dynamics of iterated maps. III. The individual molecules of the M-set, self-similarity properties, the empirical n^{2} rule, and the n^{2} conjecture. In Chaos, fractals, and dynamics (Guelph, Ont., 1981/1983), volume 98 of Lecture Notes in Pure and Applied Mathematics, pages 213–224. Dekker, New York, NY, USA, 1985.

Mandelbrot:1985:DIMb

[Man85b] Benoît B. Mandelbrot. On the dynamics of iterated maps. IV. The notion of “normalized radical” R of the M-set, and the fractal dimension of the boundary of R. In Chaos, fractals, and dynamics (Guelph, Ont., 1981/1983), volume 98 of Lecture Notes in Pure and Applied Mathematics, pages 225–234. Dekker, New York, NY, USA, 1985.

Mandelbrot:1985:DIMf

[Man85c] Benoît B. Mandelbrot. On the dynamics of iterated maps. IX. Continuous interpolation of the complex discrete map z → λz(1−z), and related topics. Physica Scripta, T9(??):59–63, ???? 1985. CODEN PHSTBO. ISSN 0031-8949 (print), 1402-4896 (electronic). Physics of chaos and related problems: Proceedings of the 59th Nobel Symposium 11–16 June 1984, Gräftåvallen, Sweden.

Mandelbrot:1985:DIMc

[Man85d] Benoît B. Mandelbrot. On the dynamics of iterated maps. V. Conjecture that the boundary of the M-set has a fractal dimension equal to 2. In Chaos, fractals, and dynamics (Guelph, Ont., 1981/1983), volume 98 of Lecture Notes in Pure and Applied Mathematics, pages 235–238. Dekker, New York, NY, USA, 1985.

Mandelbrot:1985:DIMd

[Man85e] Benoît B. Mandelbrot. On the dynamics of iterated maps. VI. Conjecture that certain Julia sets include smooth components. In Chaos, fractals, and dynamics (Guelph, Ont., 1981/1983), volume 98 of Lecture Notes in Pure and Applied Mathematics, pages 239–242. Dekker, New York, NY, USA, 1985.

Mandelbrot:1985:DIMe

[Man85f] Benoît B. Mandelbrot. On the dynamics of iterated maps. VII. Domain-filling (“Peano”) sequences of fractal Julia sets, and an intuitive rationale for the Siegel discs. In Chaos, fractals, and dynamics (Guelph, Ont., 1981/1983), volume 98 of Lecture Notes in Pure and Applied Mathematics, pages 243–253. Dekker, New York, NY, USA, 1985.

Mandelbrot:1985:SAF

[Man85g] Benoît B. Mandelbrot. Self-affine fractals and fractal dimension. Physica Scripta, 32(4):257–260, ???? 1985. CODEN PHSTBO. ISSN 0031-8949 (print), 1402-4896 (electronic).

Mandelbrot:1986:FMT

[Man86a] B. B. Mandelbrot. Fractal measures (their infinite moment sequences and dimensions) and multiplicative chaos: early works and open problems. In Dimensions and entropies in chaotic systems (Pecos River Ranch, NM, 1985), volume 32 of Springer Ser. Synergetics, pages 19–27. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986.

Mandelbrot:1986:MF

[Man86b] Benoît B. Mandelbrot. Multifractals and fractals. Physics Today, 39(9):11–??, September 1986. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/phtoad/v39/i9/p11/s1.

Mandelbrot:1986:SAFa

[Man86c] Benoît B. Mandelbrot. Self-affine fractal sets. I. The basic fractal dimensions. In Fractals in physics (Trieste, 1985), pages 3– 15. North-Holland Publishing Co., Amsterdam, The Netherlands, 1986.

Mandelbrot:1986:SAFb

[Man86d] Benoît B. Mandelbrot. Self-affine fractal sets. II. Length and surface dimensions. In Fractals in physics (Trieste, 1985), pages 17–20. North-Holland Publishing Co., Amsterdam, The Netherlands, 1986.

Mandelbrot:1986:SAFc

[Man86e] Benoît B. Mandelbrot. Self-affine fractal sets. III. Hausdorff dimension anomalies and their implications. In Fractals in physics (Trieste, 1985), pages 21–28. North-Holland Publishing Co., Amsterdam, The Netherlands, 1986.

Mandelbrot:1988:IMD

[Man88a] B. B. Mandelbrot. An introduction to multifractal distribution functions. In Random fluctuations and pattern growth (Cargèse, 1988), volume 157 of NATO Adv. Sci. Inst. Ser. E Appl. Sci., pages 279–291. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1988.

Mandelbrot:1988:FPE

[Man88b] Benoît B. Mandelbrot. Foreword: people and events behind the “science of fractal images”. In Peitgen and Saupe [PS88a], page ?? ISBN 0-387-96608-0 (New York), 1-4612-8349-3, 3-540-96608-0 (Berlin). LCCN QA614.86 .S35 1988.

Mandelbrot:1988:FLC

[Man88c] Benoît B. Mandelbrot. Fractal landscapes without creases and with rivers. In Peitgen and Saupe [PS88a], page ?? ISBN 0-387-96608-0 (New York), 1-4612-8349-3, 3-540-96608-0 (Berlin). LCCN QA614.86 .S35 1988.

Mandelbrot:1989:FASb

[Man89a] B. B. Mandelbrot. Fractals and an art for the sake of science. In Resch and Williams [RW89], pages 21–24. CODEN ???? ISBN 0-08-037921-4. ISSN 0024-094X (print), 1530-9282 (electronic).

LCCN N 7433.8 C66 1989. URL http://www.acm.org:80/pubs/
citations/proceedings/graph/73877/p21-mandelbrot/. Also

published as 1989 Supplement to Leonardo: Journal of the International Society for the Arts, Sciences and Technology.

Mandelbrot:1989:CMM

[Man89b] Benoît B. Mandelbrot. A class of multinomial multifractal measures with negative (latent) values for the “dimension” f(α). In Fractals’ physical origin and properties (Erice, 1988), volume 45 of Ettore Majorana Internat. Sci. Ser. Phys. Sci., pages 3–29. Plenum Press, New York, NY, USA; London, UK, 1989.

Mandelbrot:1989:FASa

[Man89c] Benoît B. Mandelbrot. Fractals and an art for the sake of science. Leonardo (Oxford, England), 2(Supplemental Issue):21–24, ???? 1989. CODEN LEONDP. ISSN 0024-094X (print), 1530-9282

(electronic). URL http://www.jstor.org/stable/1557938.

Mandelbrot:1989:OFF

[Man89d] Benoît B. Mandelbrot. Les objets fractals: Forme, hasard et dimension. (French) [Fractal objects: form, chance, and dimension]. Nouvelle bibliothèque scientifique. Flammarion, Paris, France, third edition, 1989. ISBN 2-08-211188-1. 268 pp. LCCN QA614.86 .M28 1989. 99.00F.

Mandelbrot:1989:MME

[Man89e] Benoît B. Mandelbrot. Multifractal measures, especially for the geophysicist. Pure and Applied Geophysics, 131(1–2):5– 42, ???? 1989. CODEN PAGYAV. ISSN 0033-4553 (print),

1420-9136 (electronic). URL http://www.springerlink.com/
content/x445xm5507251257/.

Mandelbrot:1989:SFE

[Man89f] Benoît B. Mandelbrot. Some ‘facts’ that evaporate upon examination. The Mathematical Intelligencer, 11(4):17–19, December 1989. CODEN MAINDC. ISSN 0343-6993 (print),

1866-7414 (electronic). URL http://www.springerlink.com/
content/08301p9317702218/.

Mandelbrot:1989:TFW

[Man89g] Benoît B. Mandelbrot. Temperature fluctuation: A well-defined and unavoidable notion. Physics Today, 42(1):71–??, January 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v42/
i1/p71/s1.

Mandelbrot:1990:NFD

[Man90a] Benoît B. Mandelbrot. Negative fractal dimensions and multifractals. Physica A, 163(1):306–315, February 1, 1990. CO-DEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

037843719090339T. Statistical physics (Rio de Janeiro, 1989).

Mandelbrot:1990:NAM

[Man90b] Benoît B. Mandelbrot. New ‘anomalous’ multiplicative multifractals: Left sided f(α) and the modelling of DLA. Physica A, 168(1):95–111, September 1, 1990. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037843719090361U.

Mandelbrot:1990:TMM

[Man90c] Benoît B. Mandelbrot. Two meanings of multifractality, and the notion of negative fractal dimension. In Chaos/Khaos (Woods Hole, MA, 1989), pages 79–90. American Institute of Physics, Woodbury, NY, USA, 1990.

Mandelbrot:1991:FAC

[Man91a] Benoît Mandelbrot. Fractal aggregates, and the current lines of their electrostatic potentials. Physica A, 177(1–3):589–592, September 15, 1991. CODEN PHYADX. ISSN 0378-4371 (print),

1873-2119 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037843719190205Q.

Mandelbrot:1991:FGN

[Man91b] Benoît B. Mandelbrot. Die fraktale Geometrie der Natur. (German) [The Fractal Geometry of Nature]. Birkhäuser Verlag, Basel, Switzerland, 1991. ISBN 3-7643-2646-8. 491 pp. LCCN ???? Translated from the English by Reinhilt Zähle and Ulrich Zähle, Edited and with a foreword by U. Zähle.

Mandelbrot:1991:HLC

[Man91c] Benoît B. Mandelbrot. How long is the coast of Britain? In Ferris and Fadiman [FF91], pages 447–455. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Mandelbrot:1991:RMN

[Man91d] Benoît B. Mandelbrot. Random multifractals: Negative dimensions and the resulting limitations of the thermodynamic formalism. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 434(1890):79–88, July 8, 1991. CODEN PRLAAZ. ISSN 0962-8444. URL http://www.jstor. org/stable/51986. Turbulence and stochastic processes: Kolmogorov’s ideas 50 years on.

Mandelbrot:1992:PDS

[Man92a] Benoît Mandelbrot. Plane DLA is not self-similar; is it a fractal that becomes increasingly compact as it grows? Physica A, 191(1–4):95–107, December 15, 1992. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037843719290511N.

Mandelbrot:1992:SSH

[Man92b] Benoît Mandelbrot. Self-similarity of harmonic measure on DLA. Physica A, 185(1–4):77–86, June 15, 1992. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378437192904402.

Mandelbrot:1993:SSC

[Man93a] Benoît Mandelbrot. Self-similarity and covered neighborhoods of fractals: A random walk test. Physica A, 196(1):1–5, May 15, 1993. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037843719390076G.

Mandelbrot:1993:FAS

[Man93b] Benoît B. Mandelbrot. Fractals and an art for the sake of science. In The visual mind, Leonardo Book Ser., pages 11–14. MIT Press, Cambridge, MA, USA, 1993.

Mandelbrot:1993:MMM

[Man93c] Benoît B. Mandelbrot. The Minkowski measure and multifractal anomalies in invariant measures of parabolic dynamic systems. In Chaos in Australia (Sydney, 1990), pages 83–94. World Scientific Publishing, Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1993.

Mandelbrot:1993:O

[Man93d] Benoît B. Mandelbrot. Opinions. Fractals, 1(1):117–123, ???? 1993. CODEN FRACEG. ISSN 0218-348X.

Mandelbrot:1993:OE

[Man93e] Benoît B. Mandelbrot. Opinions (essay). Symmetry Cult. Sci., 4 (3):319–328, ???? 1993. CODEN ???? ISSN 0865-4824.

Mandelbrot:1994:I

[Man94] Benoît Mandelbrot. Introduction. In Fractal landscapes: from the real world [Hir94], page ?? ISBN 0-948797-24-X (cased), 0-948797-23-1 (paperback). LCCN TR647.H57 A4 1994. £24.95. Photographs by Bill Hirst with an introduction by Benoît Mandelbrot.

Mandelbrot:1995:NDH

[Man95a] B. B. Mandelbrot. Negative dimensions and Hölders, multifractals and their Hölder spectra, and the role of lateral preasymptotics in science. Journal of Fourier Analysis and Applications, Special Issue(??):409–432, ???? 1995. CODEN ???? ISSN 1069-5869. Proceedings of the Conference in Honor of Jean-Pierre Kahane (Orsay, 1993).

Mandelbrot:1995:CMA

[Man95b] Benoît Mandelbrot. A class of micropulses and antipersistent fractional Brownian motion. Stochastic Processes and Their Applications, 60(1):1–18, November 1995. CODEN STOPB7. ISSN 0304-4149 (print), 1879-209X (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0304414995000461.

Mandelbrot:1995:MFI

[Man95c] Benoît Mandelbrot. Multifractal formalism for infinite multinomial measures. Advances in Applied Mathematics, 16(2):132–150, June 1995. CODEN ???? ISSN 0196-8858 (print), 1090-2074

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S019688588571007X.

Mandelbrot:1995:IFS

[Man95d] Benoît B. Mandelbrot. Introduction to fractal sums of pulses. In Lévy flights and related topics in physics (Nice, 1994), volume 450 of Lecture Notes in Physics, pages 110–123. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISSN 0075-8450 (print), 1616-6361 (electronic). URL

http://www.springerlink.com/content/y52438vl6j470635/.

Mandelbrot:1995:MFL

[Man95e] Benoît B. Mandelbrot. Measures of fractal lacunarity: Minkowski content and alternatives. In Christoph Bandt et al., editors, Fractal geometry and stochastics. Proceedings of a conference on fractal geometry and stochastics held at Finsterbergen, Germany, June 12–18, 1994, volume 37 of Progress in Probability, pages 15–42. Birkhäuser Verlag, Basel, Switzerland, 1995.

Mandelbrot:1995:PLK

[Man95f] Benoît B. Mandelbrot. The Paul Lévy I knew. In Lévy flights and related topics in physics (Nice, 1994), volume 450 of Lecture Notes in Physics, pages ix–xii. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISSN 0075-8450 (print), 1616-6361 (electronic).

Mandelbrot:1996:AMF

[Man96a] Benoît Mandelbrot. Alternative micropulses and fractional Brownian motion. Stochastic Processes and Their Applications, 64 (2):143–152, November 29, 1996. CODEN STOPB7. ISSN 0304-4149 (print), 1879-209X (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304414996000890.

Mandelbrot:1996:LRN

[Man96b] Benoît Mandelbrot. Local regularity of nonsmooth wavelet expansions and application to the Pólya function. Advances in Mathematics, 120(2):265–282, June 25, 1996. CODEN ADMTA4. ISSN 0001-8708 (print), 1090-2082 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0001870896900390.

Mandelbrot:1997:IFC

[Man97a] Benoît Mandelbrot. Inversion formula for continuous multifractals. Advances in Applied Mathematics, 19(3):332–354, October 1997. CODEN ???? ISSN 0196-8858 (print), 1090-2074

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196885897905509.

Mandelbrot:1997:FSF

[Man97b] Benoît B. Mandelbrot. Fractals and scaling in finance. Discontinuity, concentration, risk. Selecta Volume E. Selected Works of Benoît B. Mandelbrot. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-98363-5. x + 551 pp. LCCN HG176.5 .M36 1997. With a foreword by R. E. Gomory, and contributions by P.H. Cootner and others.

Mandelbrot:1997:GFN

[Man97c] Benoît B. Mandelbrot. La geometría fractal de la naturaleza. (Spanish) [The Fractal Geometry of Nature], volume 49 of Metatemas. Tusquets Editores S.A., Barcelona, Spain, 1997. ISBN 84-8310-549-7. 662 pp. LCCN ???? Translated from the English by Josep Llosa.

Mandelbrot:1998:P

[Man98a] Benoît Mandelbrot. Preface. In Dieu joue-t-il aux dés ?: les nouvelles mathématiques du chaos [Ste98], page ?? ISBN 2-08-081411-7 (hardcover), 2-08-081302-1 (paperback). LCCN ???? Preface by Benoît Mandelbrot. Translation by Marianne Robert from the English edition. Remarks for the second edition by Marcel Filoche.

Mandelbrot:1998:NF

[Man98b] Benoît B. Mandelbrot. Is nature fractal? Science, 279(5352): 783, February 6, 1998. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/279/5352/783.3.full.

Mandelbrot:1998:OFF

[Man98c] Benoît B. Mandelbrot. Les objets fractals: Forme, hasard et dimension. (French) [Fractal objects: form, chance, and dimension]. Flammarion, Paris, France, fourth edition, 1998. ISBN 2-08-081301-3. 212 pp. LCCN QA447 .M36.

Mandelbrot:1999:MWW

[Man99a] Benoît B. Mandelbrot. A multifractal walk down Wall Street. Scientific American, 280(2):70–73, February 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/
v280/n2/pdf/scientificamerican0299-70.pdf.

Mandelbrot:1999:MNW

[Man99b] Benoît B. Mandelbrot. Multifractals and 1/f noise. Wild selfaffinity in physics (1963–1976). Selecta Volume N. Selected Works of Benoît B. Mandelbrot. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98539-5. viii + 442 pp. LCCN QA614.86 .M27 1999. With contributions by J. M. Berger, J.-P. Kahane and J. Peyrière.

Mandelbrot:1999:RFP

[Man99c] Benoît B. Mandelbrot. Renormalization and fixed points in finance, since 1962. Physica A, 263(1–4):477–487, February 1, 1999. CODEN PHYADX. ISSN 0378-4371 (print),

1873-2119 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378437198005202. STATPHYS 20

(Paris, 1998).

Mandelbrot:1999:SMF

[Man99d] Benoît B. Mandelbrot. Survey of multifractality in finance, volume 1238 of Cowles Foundation discussion paper. The Foundation, New Haven, CT, USA, 1999. ???? pp. LCCN ????

Mandelbrot:2000:SMQ

[Man00] Benoît B. Mandelbrot. Some mathematical questions arising in fractal geometry. In Jean-Paul Pier, editor, Development of mathematics 1950–2000, pages 795–811. Birkhäuser Verlag, Basel, Switzerland, 2000.

Mandelbrot:2001:LFA

[Man01a] Benoît Mandelbrot. L’anneau fractal de l’art à l’art à travers la géométrie, la finance et les sciences [images animées]. (French) [The fractal ring of art to art across geometry, finance, and science [animated images]]. Service du Film de Recherche Scientifique, Vanves, France, 2001. LCCN ???? One VHS cassette (72 min).

Mandelbrot:2001:SFPa

[Man01b] Benoît B. Mandelbrot. Scaling in financial prices. I. Tails and dependence. Quantitative Finance, 1(1):113–123, ???? 2001. CO-DEN QFUIAV. ISSN 1469-7688 (print), 1469-7696 (electronic).

Mandelbrot:2001:SFPb

[Man01c] Benoît B. Mandelbrot. Scaling in financial prices. II. Multifractals and the star equation. Quantitative Finance, 1(1):124–130, ???? 2001. CODEN QFUIAV. ISSN 1469-7688 (print), 1469-7696 (electronic).

Mandelbrot:2001:SVP

[Man01d] Benoît B. Mandelbrot. Stochastic volatility, power laws and long memory. comment on: “Stochastic volatility as a simple generator of apparent financial power laws and long memory” [Quant. Finance 1 (2001), no. 6, 621–631; 1 870 018] by B. LeBaron. Quantitative Finance, 1(6):558–559, ???? 2001. CODEN QFUIAV. ISSN 1469-7688 (print), 1469-7696 (electronic).

Mandelbrot:2001:TFM

[Man01e] Benoît B. Mandelbrot. Topics on fractals in mathematics and physics. In Challenges for the 21st century (Singapore, 2000), pages 461–478. World Scientific Publishing, Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2001.

Mandelbrot:2002:FRA

[Man02a] Benoît Mandelbrot. The fractal ring from art to art through mathematics, finance, and the sciences. In Fractals, graphics, and mathematics education [FM02b], page ?? ISBN 0-88385-169-5. LCCN QA614.86 .F68424 2002. URL http://www.loc.gov/

catdir/description/cam022/2001097388.html;   http://www.
loc.gov/catdir/toc/cam027/2001097388.html. Foreword by

Lynn Arthur Steen.

Mandelbrot:2002:MSC

[Man02b] Benoît Mandelbrot. Mathematics and society in the 20th century. In Fractals, graphics, and mathematics education [FM02b], page ?? ISBN 0-88385-169-5. LCCN

QA614.86 .F68424 2002. URL http://www.loc.gov/catdir/
description/cam022/2001097388.html;   http://www.loc.
gov/catdir/toc/cam027/2001097388.html. Foreword by Lynn

Arthur Steen.

Mandelbrot:2002:MAW

[Man02c] Benoît Mandelbrot. A maverick’s apprenticeship. The Wolf Prizes for Physics, Imperial College Press, 2002. Web docu-

ment, 2002. URL http://users.math.yale.edu/~bbm3/web_
pdfs/mavericksApprenticeship.pdf.

Mandelbrot:2002:GSA

[Man02d] Benoît B. Mandelbrot. Gaussian self-affinity and fractals: Globality, the earth, 1/f noise, and R/S. Selecta (old or new) Volume H. Selected Works of Benoît B. Mandelbrot. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-98993-5. x + 654 pp. LCCN QA614.86 .M24 2001. With contributions by F. J. Damerau, M. Frame, K. Mc-Camy, J. W. Van Ness, J. R. Wallis, and others.

Mandelbrot:2003:HTF

[Man03a] Benoît Mandelbrot. Heavy tails in finance for independent or multifractal price increments. In Rachev [Rac03], pages 1–34. ISBN

0-444-50896-1. LCCN ???? URL http://www.sciencedirect.
com/science/article/pii/B9780444508966500030.

Mandelbrot:2003:MMM

[Man03b] Benoît Mandelbrot. Multiperiodic multifractal martingale measures. Journal de Mathématiques Pures et Appliquées, 82 (12):1555–1589, December 2003. CODEN JMPAAM. ISSN 0021-7824 (print), 1776-3371 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021782403000357.

Mandelbrot:2003:FSP

[Man03c] Benoît B. Mandelbrot. Fractal sums of pulses and a practical challenge to the distinction between local and global dependence. Processes with Long-Range Correlations, 34(5–6):118–135, ???? 2003. CODEN ???? ISSN ???? URL http://www.

springerlink.com/content/q5377044322486p7/.

Mandelbrot:2003:MPL

[Man03d] Benoît B. Mandelbrot. Multifractal power law distributions: Negative and critical dimensions and other ‘anomalies,’ explained by a simple example. Journal of Statistical Physics, 110(3–6):739–774, March 2003. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http:/

/link.springer.com/article/10.1023/A%3A1022159802564;
 http://www.springerlink.com/content/j20k474mu4063439/

. Special issue in honor of Michael E. Fisher’s 70th birthday (Piscataway, NJ, 2001).

Mandelbrot:2004:FCM

[Man04a] Benoît B. Mandelbrot. Fractals and chaos: the Mandelbrot set and beyond. Selecta Volume C. Selected Works of Benoît B. Mandelbrot. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 0-387-20158-0. xii + 308 pp. LCCN QA614.86 .M23 2004. With a foreword by P. W. Jones and texts co-authored by C. J. G. Evertsz and M. C. Gutzwiller.

Mandelbrot:2004:STM

[Man04b] Benoît B. Mandelbrot. Selected topics in mathematics, physics, and finance originating in fractal geometry. In Novak [Nov04], pages 1–33. ISBN 981-238-822-2. LCCN QC20.7.F73 T45 2004.

Mandelbrot:2005:INF

[Man05a] Benoît B. Mandelbrot. The inescapable need for fractal tools in finance. Annals of Finance, 1(2):193–195, ???? 2005. ISSN 1614-2446 (print), 1614-2454 (electronic). URL http://www.

springerlink.com/content/7ghf995yeq7k5mtq/.

Mandelbrot:2005:PCF

[Man05b] Benoît B. Mandelbrot. Parallel cartoons of fractal models of finance. Annals of Finance, 1(2):179–192, ???? 2005. CO-

DEN ???? ISSN ???? URL http://www.springerlink.com/
content/j5fcgq4kmntgtkum/.

Mandelbrot:2006:BM

[Man06a] Benoît Mandelbrot. Benoît Mandelbrot. New Scientist, 192 (2578):72, November 18, 2006. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0262407906611457.

Mandelbrot:2006:FASa

[Man06b] Benoît Mandelbrot. Fractal analysis and synthesis of fracture surface roughness and related forms of complexity and disorder. In Carpinteri et al. [CMR06], pages 13–17. ISBN 1-4020-4626-X (cloth), 1-4020-5423-8 (e-book). LCCN TA409 .I44 2005eb. URL

http://www.springerlink.com/content/up4521397545ruw7/.

Mandelbrot:2006:FASb

[Man06c] Benoît B. Mandelbrot. Fractal analysis and synthesis of fracture surface roughness and related forms of complexity and disorder. International Journal of Fracture, 138(1–4):13–17, ???? 2006. CODEN IJFRAP. ISSN 0376-9429 (print), 1573-2673

(electronic). URL http://www.springerlink.com/content/
p017550250307256/.

Mandelbrot:2007:SSS

[Man07] Benoît B. Mandelbrot. Scalebound or scaling shapes: a useful distinction in the visual arts and in the natural sciences. Leonardo (Oxford, England), 40(5):494–496, ???? 2007. CODEN LEONDP. ISSN 0024-094X (print), 1530-9282 (electronic).

Mandelbrot:2008:FVD

[Man08] Benoît Mandelbrot. Fraktale: die verborgene Dimension. (German) [Fractals: the hidden dimension]. Komplett-Media, Grünwald, Germany, 2008. ISBN 3-8312-9945-5. LCCN ???? 1 video DVD (54 min.). Produced by Michael Schwarz.

Mandelbrot:2009:FHF

[Man09a] Benoît B. Mandelbrot. Fractales, hasard et finance: 1959–1997. (French) [Fractals, chance, and finance: 1959–1997]. Flammarion, Paris, France, 2009. 246 pp. LCCN ????

Mandelbrot:2009:FMG

[Man09b] Benoît B. Mandelbrot. The fractalist: memoir of a geometer. Pantheon Books, New York, NY, USA, 2009. ISBN 0-307-37735-0. xviii + 327 pp. LCCN QA29.M34 A3 2009. With an afterword by Michael Frame.

Mandelbrot:2009:NMS

[Man09c] Benoît B. Mandelbrot. “New Methods of Statistical Economics,” revisited: Short versus long tails and Gaussian versus power-law distributions. Complexity, 14(3):55–65, January 2009. CODEN COMPFS. ISSN ????

Mandelbrot:2010:FFF

[Man10a] Benoît Mandelbrot. Fractal financial fluctuations. In Lesmoir-Gordon [LG10b], chapter 7, pages 120–134. CODEN ???? ISBN 1-84996-485-8. ISSN ???? LCCN QA614.86 .C65 2010. URL

http://www.springerlink.com/content/h8w32018354n747q/.

With contributions by Ian Stewart, Sir Arthur C. Clarke, Benoît B. Mandelbrot, Michael and Louisa Barnsley, Will Rood, Gary Flake, and David Pennock.

Mandelbrot:2010:GAI

[Man10b] Benoît Mandelbrot. A geometry able to include mountains and clouds. In Lesmoir-Gordon [LG10b], chapter 3, pages 38–57. ISBN 1-84996-485-8. LCCN QA614.86 .C65 2010. URL http:

//www.springerlink.com/content/h535l123uh2t6643/. With

contributions by Ian Stewart, Sir Arthur C. Clarke, Benoît B. Mandelbrot, Michael and Louisa Barnsley, Will Rood, Gary Flake, and David Pennock.

Mandelbrot:2010:OFF

[Man10c] Benoît B. Mandelbrot. Les objets fractals: Forme, hasard et dimension. (French) [Fractal objects: form, chance, and dimension]. Flammarion, Paris, France, 2010. ISBN ???? 212 pp. LCCN ????

Mandelbrot:2012:BMP

[Man12] Benoît B. Mandelbrot. Benoît Mandelbrot papers, 1932–2010, 2012. Circa 380 linear feet.

Martinez:1999:UF

[Mar99] Vicent J. Martínez. Is the universe fractal? Science, 284(5413): 445–446, April 16, 1999. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/284/5413/445.full.

Martyn:2004:NAM

[Mar04] Tomek Martyn. A new approach to morphing 2D affine IFS fractals. Computers and Graphics, 28(2):249–272, April 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Martyn:2010:RRI

[Mar10] Tomasz Martyn. Realistic rendering 3D IFS fractals in real-time with graphics accelerators. Computers and Graphics, 34(2):167– 175, April 2010. CODEN COGRD2. ISSN 0097-8493 (print),

1873-7684 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0097849309001150.

Maslov:2005:GTS

[Mas05a] V. P. Maslov. On a general theorem of set theory that leads to the Gibbs, Bose–Einstein, and Pareto distributions and to the Zipf– Mandelbrot law for the stock market. Matematicheskie Zametki, 78(6):870–877, ???? 2005. CODEN ???? ISSN 0025-567X.

Maslov:2005:ZML

[Mas05b] V. P. Maslov. The Zipf–Mandelbrot law: quantization and an application to the stock market. Russian Journal of Mathematical Physics, 12(4):483–488, ???? 2005. CODEN RJMPEL. ISSN 1061-9208 (print), 1555-6638 (electronic).

Maslov:2006:RZM

[Mas06] V. P. Maslov. A refinement of the Zipf–Mandelbrot law and lacunarity in an ideal gas. Teoreticheskaya i Matematicheskaya Fizika, 147(3):511–512, ???? 2006. CODEN TMFZAL. ISSN 0564-6162.

Mandelbrot:1989:FGW

[MB89] B. B. Mandelbrot and A. Blumen. Fractal geometry: What is it, and what does it do? [and discussion]. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 423(1864):3–16, May 8, 1989. CODEN PRLAAZ. ISSN 0080-

4630. URL http://www.jstor.org/stable/2398503. Fractals

in the natural sciences.

McCloskey:1994:TUR

[MB94] John McCloskey and C. J. Bean. Temporally unstable recurrence of earthquakes due to breaks in fractal scaling. Science, 266(5184):410–412, October 21, 1994. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/266/5184/410.full.pdf.

Miller:1963:FML

[MC63] George A. Miller and Noam Chomsky. Finitary models of language users. In Luce et al. [LBG63], pages 419–491. LCCN ????

McHardy:1987:FXR

[MC87] Ian McHardy and Bozena Czerny. Fractal X-ray time variability and spectral invariance of the Seyfert galaxy NGC5506. Nature, 325(6106):696–698, February 19, 1987. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v325/n6106/pdf/325696a0.pdf.

Melin:2002:MSC

[MC02] Patricia Melin and Oscar Castillo. Modelling, simulation and control of non-linear dynamical systems: an intelligent approach using soft computing and fractal theory, volume 2 of Numerical insights, 1028-5350. Taylor and Francis, London, UK and Boca Raton, FL, USA, 2002. ISBN 0-415-27236-X. xi + 249 pp. LCCN QA427 .M45 2002.

McKenna:2019:HC

[McK19] Doug McKenna. Hilbert curves. Web site., June 2019. URL

http://www.mathemaesthetics.com/HilbertCurves.html.

Melo:2005:EPC

[MdF05] Cesar A. V. Melo and Nelson L. S. da Fonseca. Envelope process and computation of the equivalent bandwidth of multifractal flows. Computer Networks (Amsterdam, Netherlands: 1999), 48(3):351–375, June 21, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Molnar:2002:QTA

[MDM02] Sándor Molnár, Trang Dinh Dang, and István Maricza. On the queue tail asymptotics for general multifractal traffic. Lecture Notes in Computer Science, 2345(??):105–??, ???? 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2345/23450105.htm;   http://
link.springer-ny.com/link/service/series/0558/papers/
2345/23450105.pdf.

Mandelbrot:1972:HSM

[ME72] A. K. Mandelbrot and K. Erb. Host spectrum of the mycoparasite dimargaris verticillata. Mycologia, 64(5):1124–1129, October 1972. CODEN ???? ISSN ???? URL http://www.jstor.org/

stable/3758077.

Mandelbrot:1990:PDA

[ME90] Benoît B. Mandelbrot and Carl J. G. Evertsz. The potential distribution around growing fractal clusters. Nature, 348(6297): 143–145, November 8, 1990. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v348/n6297/pdf/348143a0.pdf.

Mandelbrot:1991:MHM

[ME91] Benoît B. Mandelbrot and Carl J. G. Evertsz. Multifractality of the harmonic measure on fractal aggregates, and extended selfsimilarity. Physica A, 177(1–3):386–393, September 15, 1991. CO-DEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

037843719190177E. Current problems in statistical mechanics (Washington, DC, 1991).

Meakin:1989:WF

[Mea89] Paul Meakin. Work with fractals. Science, 245(4925):1515– 1516, September 29, 1989. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/245/4925/1515.2.full.pdf.

Mandelbrot:1990:ESS

[MEH90] Benoît B. Mandelbrot, Carl J. G. Evertsz, and Yoshinori Hayakawa. Exactly self-similar left-sided multifractal measures. Physical Review A (Atomic, Molecular, and Optical Physics), 42:4528–??, 1990. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL http://link.aps.org/

abstract/PRA/v42/e4528.

Meisel:1992:GMR

[Mei92] L. V. Meisel. Generalized Mandelbrot rule for fractal sections. Physical Review A (Atomic, Molecular, and Optical Physics), 45:654–??, 1992. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL http://link.aps.org/

abstract/PRA/v45/e654.

Metzler:1994:MQM

[Met94] Wolfgang Metzler. The “mystery” of the quadratic Mandelbrot set. American Journal of Physics, 62:813–??, 1994. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://link.aip.org/link/ajpias/v62/i9/p813/s1.

Mandelbrot:1999:CSP

[MF99] Benoît B. Mandelbrot and Michael Frame. The canopy and shortest path in a self-contacting fractal tree. The Mathematical Intelligencer, 21(2):18–27, June 1999. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic). URL http://

www.springerlink.com/content/136523377672365p/.

Mandelbrot:2002:SRE

[MF02] Benoît Mandelbrot and Michael Frame. Some reasons for the effectiveness of fractals in mathematics education. In Fractals, graphics, and mathematics education [FM02b], page ?? ISBN 0-88385-169-5. LCCN QA614.86 .F68424 2002. URL http://www.

loc.gov/catdir/description/cam022/2001097388.html;   http:
//www.loc.gov/catdir/toc/cam027/2001097388.html. Fore-

word by Lynn Arthur Steen.

Mandelbrot:2003:F

[MF03] Benoît Mandelbrot and Michael Frame. Fractals. In Robert A. Meyers, editor, Encyclopedia of Physical Science and Technology (Third Edition), pages 185–207. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 2003. ISBN 0-12-227410-5.

LCCN ???? URL http://www.sciencedirect.com/science/
article/pii/B0122274105002593.

Mandelbrot:2009:PNT

[MF09] Benoît B. Mandelbrot and Michael Frame. A primer of negative test dimensions and degrees of emptiness for latent sets. Fractals, 17(1):1–14, ???? 2009. CODEN FRACEG. ISSN 0218-348X.

McGregor:1996:FFC

[MFCM96] D. R. McGregor, R. J. Fryer, P. Cockshott, and P. Murray. Faster fractal compression. Dr. Dobb’s Journal of Software Tools, 21(1): 34, 36, 38–40, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Mandelbrot:1984:PPN

[MG84] Benoît B. Mandelbrot and James A. Given. Physical properties of a new fractal model of percolation clusters. Physical Review Letters, 52:1853–??, 1984. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.

aps.org/abstract/PRL/v52/e1853.

Mandelbrot:1998:FSF

[MG98] Benoît B. Mandelbrot and Nigel Goldenfeld. Fractals and scaling in finance: Discontinuity, concentration, risk. Physics Today, 51(8):59–??, August 1998. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/phtoad/v51/i8/p59/s1.

Mandelbrot:1985:FTT

[MGAP85] Benoît B. Mandelbrot, Yuval Gefen, Amnon Aharony, and Jacques Peyrière. Fractals, their transfer matrices and their eigendimensional sequences. Journal of Physics A (Mathematical and General), 18(2):335–354, ???? 1985. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic). URL http://

stacks.iop.org/0305-4470/18/335.

McFarlane:1994:ODR

[MH94] I. McFarlane and S. G. Hoggar. Optimal drivers for the ‘random’ iteration algorithm. The Computer Journal, 37(7): 629–640, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/Volume_37/Issue_07/Vol37_07.body.
html#AbstractMcFarlane.

Mandelbrot:2004:MBM

[MH04a] Benoît B. Mandelbrot and Richard L. Hudson. The (mis)behavior of markets: a fractal view of risk, ruin, and reward. Basic Books, New York, NY, USA, 2004. ISBN 0-465-04355-0, 0-465-04357-7 (paperback). xxiv + 328 pp. LCCN HG4523 .M257

2004.  URL http://www.loc.gov/catdir/enhancements/
fy0830/2004011400-b.html;   http://www.loc.gov/catdir/
enhancements/fy0830/2004011400-d.html. A fractal view of

risk, ruin, and reward.

Mandelbrot:2004:MMF

[MH04b] Benoît B. Mandelbrot and Richard L. Hudson. The misbehaviour of markets: a fractal view of risk, ruin and reward. Profile, London, UK, 2004. ISBN 1-86197-765-4 (hardcover). xxiv + 328 pp. LCCN ???? £18.99.

Mandelbrot:2005:FFM

[MH05a] Benoît B. Mandelbrot and Richard L. Hudson. Fraktale und Finanzen. Märkte zwischen Risiko, Rendite und Ruin. (German) [Fractals and finance. Markets between risk, reward and ruin]. Piper, München, Germany, 2005. ISBN 3-492-04632-0. LCCN ???? Translation by Helmut Reuter of [MH04a].

Mandelbrot:2005:MBM

[MH05b] Benoît B. Mandelbrot and Richard L. Hudson. The (mis)behaviour of markets: a fractal view of risk, ruin, and reward. Profile Books, London, UK, 2005. ISBN 1-86197-790-5. xxiv + 328 pp. LCCN HG4523.

Mandelbrot:2005:AFM

[MH05c] Benoît B. Mandelbrot and Richard. L. Hudson. Une approche fractale des marchés: risquer, perdre et gagner. (French) [A fractal approach to markets: risk, ruin, and reward]. O. Jacob, Paris, France, 2005. ISBN 2-7381-1536-5 (paperback). 361 pp. LCCN ???? Translated by Marcel Filoche from the American edition [MH04a].

Mandelbrot:2008:FFM

[MH08a] Benoît B. Mandelbrot and Richard L. Hudson. Fraktale und Finanzen: Märkte zwischen Risiko, Rendite und Ruin. (German) [Fractals and finance. Markets between risk, reward and ruin], volume 4861 of Serie Piper. Piper, München, Germany, second edition, 2008. ISBN 3-492-24861-6. 445 pp. LCCN ???? EUR 12.95

(DE). URL http://www.gbv.de/dms/zbw/511599846.pdf.

Mandelbrot:2008:MBM

[MH08b] Benoît B. Mandelbrot and Richard L. Hudson. The (Mis)Behaviour of Markets: a Fractal View of Risk, Ruin and Reward. Profile, London, UK, 2008. ISBN 1-84668-262-2, 1-84765-155-0. 353 pp. LCCN HG4523 .M257 2008. URL http:/

/www.gbv.eblib.com/patron/FullRecord.aspx?p=584650;
http://www.loc.gov/catdir/toc/fy0905/2009437381.html.

Mandelbrot:2009:AFM

[MH09] Benoît B. Mandelbrot and Richard. L. Hudson. Une approche fractale des marchés: risquer, perdre et gagner. (French) [A fractal approach to markets: risk, ruin, and reward]. O. Jacob, Paris, France, new edition, 2009. v + 361 pp. LCCN ???? Translated by Marcel Filoche from the American edition [MH04a].

Miller:1957:SEI

[Mil57] George A. Miller. Some effects of intermittent silence. American Journal of Psychology, 70(2):311–314, June 1957. CODEN AJPCAA. ISSN 0002-9556 (print), 1939-8298 (electronic). URL

http://www.jstor.org/stable/1419346.

Miller:1986:DRT

[Mil86] Gavin S. P. Miller. The definition and rendering of terrain maps. Computer Graphics, 20(4):39–48, August 1986. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569

(electronic). URL http://www.acm.org:80/pubs/citations/
proceedings/graph/15922/p39-miller/.

Milne:1988:MFG

[Mil88] Bruce T. Milne. Measuring the fractal geometry of landscapes. Applied Mathematics and Computation, 27(1):67–79, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mirowski:1990:MCE

[Mir90] Philip Mirowski. From Mandelbrot to chaos in economic theory. Southern Economic Journal, 57(2):289–307, October 1990.

CODEN ???? ISSN ???? URL http://www.jstor.org/stable/
1060611.

Mirowski:2001:RBS

[Mir01] Philip E. Mirowski. Review: Fractals and Scaling in Finance: Discontinuity, Concentration, Risk by Benoît B. Mandelbrot. Journal of Economic Literature, 39(2):585–587, June 2001. CO-DEN JECLB3. ISSN 0022-0515 (print), 1547-1101 (electronic).

URL http://www.jstor.org/stable/2698262.

Mitzenmacher:2003:BHG

[Mit03] Michael Mitzenmacher. A brief history of generative models for power law and lognormal distributions. Internet Mathematics, 1(2):226–251, December 2003. CODEN ???? ISSN 1542-7951 (print), 1944-9488 (electronic). URL http://

projecteuclid.org/euclid.im/1089229510;   http://www.
eecs.harvard.edu/~michaelm/CS223/powerlaw.pdf;   http:
//www.tandfonline.com/toc/uinm20/1/2.

Miyata:1990:MGS

[Miy90] Kazunori Miyata. A method of generating stone wall patterns. Computer Graphics, 24(4):387–394, August 1990. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569

(electronic). URL http://www.acm.org:80/pubs/citations/
proceedings/graph/97879/p387-miyata/.

MacDonald:1986:FFD

[MJ86] Mairi MacDonald and Naeem Jan. Fractons and the fractal dimension of proteins. Canadian Journal of Physics = Journal canadien de physique, 64(10):1353–1355, 1986. CO-DEN CJPHAD. ISSN 0008-4204 (print), 1208-6045 (elec-

tronic). URL http://www.nrcresearchpress.com/doi/abs/
10.1139/p86-239.

Mandelbrot:1997:PPM

[MJ97] B. B. Mandelbrot and Stéphane Jaffard. Peano–Pólya motions, when time is intrinsic or binomial (uniform or multifractal). The Mathematical Intelligencer, 19(4):21–26, December 1997. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414

(electronic). URL http://www.springerlink.com/content/
9120830561261547/.

Mandelbrot:2002:AGR

[MKA02] Benoît B. Mandelbrot, Boaz Kol, and Amnon Aharony. Angular gaps in radial diffusion-limited aggregation: Two fractal dimensions and nontransient deviations from linear self-similarity. Physical Review Letters, 88:055501, 2002. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL

http://link.aps.org/abstract/PRL/v88/e055501.

Musgrave:1989:SRE

[MKM89] F. Kenton Musgrave, Craig E. Kolb, and Robert S. Mace. The synthesis and rendering of eroded fractal terrains. Computer Graphics, 23(3):41–50, July 1989. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

URL http://www.acm.org:80/pubs/citations/proceedings/
graph/74333/p41-musgrave/.

Morse:1985:FDV

[MLDW85] D. R. Morse, J. H. Lawton, M. M. Dodson, and M. H. Williamson. Fractal dimension of vegetation and the distribution of arthropod body lengths. Nature, 314(6013):731–733, April 25, 1985.

CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v314/
n6013/pdf/314731a0.pdf.

Mandelbrot:1970:SPM

[MM70] Benoît B. Mandelbrot and Keith McCamy. On the secular pole motion and the Chandler wobble. Geophysical Journal of the Royal Astronomical Society, 21(2):217–232, November 1970. CO-DEN GEOJAN. ISSN 0956-540X (print), 1365-246X (electronic).

Musgrave:1989:NEM

[MM89] F. Kenton Musgrave and Benoît B. Mandelbrot. Natura ex machina. IEEE Computer Graphics and Applications, 9(1):4– 7, January 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Musgrave:1991:AFL

[MM91] F. K. Musgrave and B. B. Mandelbrot. The art of fractal landscapes. IBM Journal of Research and Development, 35(4):535– 536, 539, July 1991. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

McIntyre-Mills:2008:MBM

[MM08] Janet J. McIntyre-Mills. Möbius bands and Mandelbrot sets as metaphors for systemic praxis. Systems Research and Behavioral Science, 25(2):323–329, March 2008. CODEN SRBSF9. ISSN 1092-7026 (print), 1099-1743 (electronic).

McNally:2010:FGN

[MM10] James G. McNally and Davide Mazza. Fractal geometry in the nucleus. The EMBO Journal, 29(1):2–3, January 6, 2010. CODEN EMJODG. ISSN 0261-4189 (print), 1460-2075 (elec-

tronic). URL http://www.nature.com/emboj/journal/v29/
n1/full/emboj2009375a.html.

Malacarne:2001:QED

[MML01] L. C. Malacarne, R. S. Mendes, and E. K. Lenzi. q-exponential distribution in urban agglomeration. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 65(??):017106, December 21, 2001. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.

aps.org/doi/10.1103/PhysRevE.65.017106.

Montemurro:2001:BZM

[Mon01] Marcelo A. Montemurro. Beyond the Zipf–Mandelbrot law in quantitative linguistics. Physica A, 300(3–4):567–578, November 15, 2001. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378437101003557.

Morcrette:1996:EDF

[Mor96] Michelle Morcrette. Sur l’équivalence de descriptions de figures iterées. (French) [On the equivalence of descriptions of iterative figures]. Theoretical Computer Science, 165(2):325– 354, October 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.elsevier.

com/cgi-bin/cas/tree/store/tcs/cas_sub/browse/browse.
cgi?year=1996&volume=165&issue=2&aid=2167.

Mandelbrot:1984:FAM

[MP84] Benoît B. Mandelbrot and Dann E. Passoja, editors. Fractal aspects of materials: metal and catalyst surfaces, powders and aggregates: extended abstracts, volume EA–4 of Materials Research Society extended abstracts. Materials Research Society, Pittsburgh, PA, USA, 1984. LCCN QA447 .F72 1984.

Mandelbrot:1998:NFI

[MPB^{+}98] Benoît B. Mandelbrot, Peter Pfeifer, Ofer Biham, Ofer Malcai, Daniel A. Lidar, and David Avnir. Is nature fractal? Science, 279(5352):783, 785–786, February 6, 1998. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.jstor.org/stable/2894997.

Manaris:2002:PTR

[MPM02] B. Manaris, T. Purewal, and C. McCormick. Progress towards recognizing and classifying beautiful music with computers: MIDI-encoded music and the Zipf-Mandelbrot Law. In IEEE, editor, Proceedings: IEEE SoutheastCon 2002: April 5–7, 2002, Embassy Suites Hotel, Columbia, South Carolina, USA, pages 52–57. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7803-7252-2.

LCCN TK7801 .I56 2002. URL http://ieeexplore.ieee.org/
servlet/opac?punumber=7814.

Mandelbrot:1984:FCF

[MPP84] Benoît B. Mandelbrot, Dann. E. Passoja, and Alvin J. Paullay. Fractal character of fracture surfaces of metals. Nature, 308 (5961):721–722, April 19, 1984. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v308/n5961/pdf/308721a0.pdf.

Mandelbrot:1997:IMI

[MR97] Benoît B. Mandelbrot and Rudolf H. Riedi. Inverse measures, the inversion formula, and discontinuous multifractals. Advances in Applied Mathematics, 18(1):50–58, January 1997. CO-DEN ???? ISSN 0196-8858 (print), 1090-2074 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019688589690500X.

Manaris:2005:ZLM

[MRM^{+}05] Bill Manaris, Juan Romero, Penousal Machado, Dwight Krehbiel, Timothy Hirzel, Walter Pharr, and Robert B. Davis. Zipf’s law, music classification, and aesthetics. Computer Music Journal, 29 (1):55–69, Spring 2005. CODEN CMUJDY. ISSN 0148-9267 (print), 1531-5169 (electronic). URL http://muse.jhu.edu/

journals/cmj/summary/v029/29.1manaris.html;   http://
muse.jhu.edu/journals/computer_music_journal/v029/29.
1manaris.pdf.

Miles:1978:GPB

[MS78] R. E. (Roger Edmund) Miles and J. (Jean) Serra, editors. Geometrical probability and biological structures: Buffon’s 200th anniversary, proceedings of the Buffon Bicentenary Symposium on Geometrical Probability, Image Analysis, Mathematical Stereology, and their Relevance to the Determination of Biological Structures, held in Paris, June 1977, volume 23 of Lecture notes in biomathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. ISBN 3-540-08856-3 (Berlin), 0-387-08856-3 (New York). LCCN ????

Mandelbrot:1994:ACT

[MS94] Benoît B. Mandelbrot and Dietrich Stauffer. Antipodal correlations and the texture (fractal lacunarity) in critical percolation clusters. Journal of Physics A (Mathematical and General), 27 (9):L237–L242, ???? 1994. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic). URL http://stacks.iop.org/

0305-4470/27/L237.

Mandelbrot:2000:CS

[MSLG00] Benoît B. Mandelbrot, Martin Shaw, and Nigel Lesmoir-Gordon. Clouds are not spheres. Gordon Films, Inc., ????, 2000. ISBN ???? LCCN ???? Video recording. Written by Nigel Lesmoir-Gordon, John Tate, Bill Rood. Produced and directed by Nigel Lesmoir-Gordon. Cameraman: Jo Zafar. Editor: Gabriel Lesmoir-Gordon. Music: James Watson and Charles Wuorinen.

Ma:2008:PLS

[MSW08] D. Ma, A. D. Stoica, and X.-L. Wang. Power-law scaling and fractal nature of medium-range order in metallic glasses. Nature Materials, 8(1):30–34, December 7, 2008. CODEN NMAACR. ISSN 1476-1122 (print), 1476-4660 (electronic). URL http://

www.nature.com/nmat/journal/v8/n1/full/nmat2340.html.

Mandelbrot:1967:DSP

[MT67] Benoît Mandelbrot and Howard M. Taylor. On the distribution of stock price differences. Operations Research, 15(6):1057–1062, December 1967. CODEN OPREAI. ISSN 0030-364X (print),

1526-5463 (electronic). URL http://www.jstor.org/stable/
168611.

Mandelbrot:1979:RAL

[MT79] Benoît B. Mandelbrot and Murad S. Taqqu. Robust R/S analysis of long-run serial correlation. Bull. Inst. Internat. Statist., 48(2): 69–99, ???? 1979. CODEN BISQA3. ISSN ???? Proceedings of the 42nd session of the International Statistical Institute, Vol. 2 (Manila, 1979).

Mandelbrot:2010:MVW

[MT10] Benoît B. Mandelbrot and Nassim Nicholas Taleb. Mild vs. wild randomness: focusing on those risks that matter. In Diebold et al. [DDH10], pages 47–58. ISBN 0-691-12883-9 (hardcover). LCCN ????

Mudge:2002:RBM

[Mud02] Michael R. Mudge. Review: The Mandelbrot Set, Theme and Variations by Tan Lei. Mathematical Gazette, 86(505):185, March 2002. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL http://www.jstor.org/stable/3621636.

Murphy:1993:PFH

[Mur93] Nancey Murphy. Philosophical fractals: Or, history as metaphilosophy. Studies in History and Philosophy of Science Part A, 24(3):501–508, August 1993. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic). URL http://www.

sciencedirect.com/science/article/pii/003936819390041H.

Mandelbrot:1968:FBM

[MV68] Benoît B. Mandelbrot and John W. Van Ness. Fractional Brownian motions, fractional noises and applications. SIAM Review, 10 (4):422–437, October 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mandelbrot:1989:DRM

[MV89] Benoît B. Mandelbrot and Tamás Vicsek. Directed recursive models for fractal growth. Journal of Physics A (Mathematical and General), 22(9):L377–L383, ???? 1989. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic). URL http://stacks.

iop.org/0305-4470/22/L377.

Meson:2002:KLG

[MV02] Alejandro M. Mesón and Fernando Vericat. On the Kolmogorovlike generalization of Tsallis entropy, correlation entropies and multifractal analysis. Journal of Mathematical Physics, 43(2): 904–917, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Manaris:2003:EMZ

[MVW^{+}03] Bill Manaris, Dallas Vaughan, Christopher Wagner, Juan Romero, and Robert B. Davis. Evolutionary music and the Zipf–Mandelbrot Law: Developing fitness functions for pleasant music. In Raidl [Rai03], pages 522–534. CODEN LNCSD9. ISBN 3-540-00976-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618 .E899 2003. URL http:

//link.springer-ny.com/link/service/series/0558/tocs/
t2611.htm;   http://www.springerlink.com/content/978-
3-540-00976-4;   http://www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&volume=2611.

Mandelbrot:1968:NJO

[MW68] Benoît Mandelbrot and J. R. Wallis. Noah, Joseph and operational hydrology. Water Resources Research, 4(??):909–918, ????

1968. CODEN WRERAQ. ISSN 0043-1397 (print), 1944-7973 (electronic).

Mandelbrot:1969:SLR

[MW69a] B. B. Mandelbrot and J. R. Wallis. Some long-run properties of geophysical records. Water Resources Research, 5(??):321–340, ???? 1969. CODEN WRERAQ. ISSN 0043-1397 (print), 1944-7973 (electronic).

Mandelbrot:1969:CEF

[MW69b] Benoît Mandelbrot and J. R. Wallis. Computer experiments with fractional Gaussian noises. Water Resources Research, 5(??):228– ??, ???? 1969. CODEN WRERAQ. ISSN 0043-1397 (print), 1944-7973 (electronic).

Mandelbrot:1969:RRR

[MW69c] Benoît Mandelbrot and J. R. Wallis. Robustness of the rescaled range R/S in the measurement of noncyclic long run statistical dependence. Water Resources Research, 5(??):967–988, ???? 1969. CODEN WRERAQ. ISSN 0043-1397 (print), 1944-7973 (electronic).

Narayaninsamy:2007:MCI

[Nar07] T. Narayaninsamy. A method to construct irregular fractal curves. Applied Mathematics and Computation, 192(1):260–273, September 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300307002925.

Nathan:1991:RBC

[Nat91] Kathirgama Nathan. Review: Fractals for the Classroom. Part One: Introduction to Fractals and Chaos by H. O. Peitgen, H. Jürgens, D. Saupe; Fractals for the Classroom. Part Two: Complex Systems and Mandelbrot Set by H. O. Peitgen, H. Jürgens, D. Saupe. College Mathematics Journal, 22(5):455–456, November 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.jstor.org/stable/2686611.

Nguyen:2004:HPF

[NCS04] Hung Phi Nguyen, Bastien Chopard, and Serge Stoll. Hydrodynamic properties of fractal aggregates in 2D using Lattice Boltzmann simulation. Future Generation Computer Systems, 20(6):

981–991, August 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Nittmann:1985:FGV

[NDS85] Johann Nittmann, Gérard Daccord, and H. Eugene Stanley. Fractal growth viscous fingers: quantitative characterization of a fluid instability phenomenon. Nature, 314(6007):141–144, March 14, 1985. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v314/n6007/pdf/314141a0.pdf.

Nakano:2001:TON

[NFT^{+}01] M. Nakano, H. Fujita, M. Takahata, S. Kiribayashi, and K. Yamaguchi. Third-order nonlinear optical properties of dendritic molecular aggregates: Effects of fractal architecture. International Journal of Quantum Chemistry, 84(6):649–659, ???? 2001. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/85008257/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext/85008257/FILE?TPL=ftx_start;
 http://www3.interscience.wiley.com/cgi-bin/fulltext?
ID=85008257&PLACEBO=IE.pdf.

Nettleton:1995:EF

[NG95] David John Nettleton and Roberto Garigliano. Evolving fractals. Computers and Graphics, 19(5):779–782, September– October 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http://www.elsevier.com/

cgi-bin/cas/tree/store/cag/cas_sub/browse/browse.cgi?
year=1995&volume=19&issue=5&aid=9500058.

Nikiel:2003:GVE

[NG03] Slawek Nikiel and Adam Goinski. Generation of volumetric escape time fractals. Computers and Graphics, 27(6):977–982, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Naschitz:2003:ACRa

[NIS^{+}03a] J. E. Naschitz, R. Itzhak, N. Shaviv, I. Khorshidi, S. Sundick, H. Isseroff, M. Fields, R. M. Priselac, D. Yeshurun, E. Sabo, et al. Assessment of cardiovascular reactivity by fractal and recurrence quantification analysis of heart rate and pulse transit time. Journal of Human Hypertension, 17(2):111–118, February 4, 2003. CODEN ???? ISSN ???? URL http://www.nature.

com/jhh/journal/v17/n2/full/1001517a.html.

Naschitz:2003:ACRb

[NIS^{+}03b] J. E. Naschitz, R. Itzhak, N. Shaviv, I. Khorshidi, S. Sundick, H. Isseroff, M. Fields, R. M. Priselac, D. Yeshurun, E. Sabo, et al. Assessment of cardiovascular reactivity by fractal and recurrence quantification analysis of heart rate and pulse transit time. Journal of Human Hypertension, 17(8):585, July 22, 2003.

CODEN ???? ISSN ???? URL http://www.nature.com/jhh/
journal/v17/n8/full/1001615a.html.

Navarro:2009:UFD

[NM09] Juan F. Navarro and M. C. Martínez. Universality of fractal dimension on time-independent Hamiltonian systems. Applied Mathematics and Computation, 214(2):462–467, August 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300309003415.

Norton:1982:GDG

[Nor82] A. Norton. Generation and display of geometric fractals in 3-D. Computer Graphics, 16(3):61–67, July 1982. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Novak:2004:TPF

[Nov04] M. M. (Mirosłav Michał) Novak, editor. Thinking in patterns: fractals and related phenomena in nature. World Scientific Publishing, Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2004. ISBN 981-238-822-2. LCCN QC20.7.F73 T45 2004.

Nehlig:1995:FQA

[NR95] P. W. Nehlig and J.-P. Reveilles. Fractals and quasi-affine transformations. Computer Graphics Forum, 14(2):147–158 (or 147– 157??), June 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Newkome:2006:NFP

[NWM^{+}06] George R. Newkome, Pingshan Wang, Charles N. Moorefield, Tae Joon Cho, Prabhu P. Mohapatra, Sinan Li, Seok-Ho Hwang, Olena Lukoyanova, Luis Echegoyen, Judith A. Palagallo, Violeta Iancu, and Saw-Wai Hla. Nanoassembly of a fractal polymer:

A molecular ‘Sierpinski Hexagonal Gasket’. Science, 312(5781): 1782–1785, June 23, 2006. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/312/5781/1782.full.pdf.

Nakayama:1994:DPF

[NYO94] Tsuneyoshi Nakayama, Kousuke Yakubo, and Raymond L. Orbach. Dynamical properties of fractal networks: Scaling, numerical simulations, and physical realizations. Reviews of Modern Physics, 66(2):381–443, April 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.66.381;   http:
//rmp.aps.org/abstract/RMP/v66/i2/p381_1.

Ortega:2003:GED

[ODA03] A. Ortega, A. A. Dalhoum, and M. Alfonseca. Grammatical evolution to design fractal curves with a given dimension. IBM Journal of Research and Development, 47(4):483–493, ???? 2003. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556

(electronic). URL http://www.research.ibm.com/journal/
rd/474/ortega.pdf.

Ortega:2002:CGP

[OdlCA02] Alfonso Ortega, Marina de la Cruz, and Manuel Alfonseca. Chaos and graphics: Parametric 2-dimensional L systems and recursive fractal images: Mandelbrot set, Julia sets and biomorphs. Computers and Graphics, 26(1):143–149, February 2002. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

URL http://www.elsevier.com/gej-ng/10/13/20/68/27/40/
abstract.html.

Ohnishi:1990:CRH

[Ohn90] Teruaki Ohnishi. Is the cosmic-ray hadron shower a fractal? Canadian Journal of Physics = Journal canadien de physique, 68 (9):906–911, 1990. CODEN CJPHAD. ISSN 0008-4204 (print),

1208-6045 (electronic). URL http://www.nrcresearchpress.
com/doi/abs/10.1139/p90-127.

Oldershaw:2001:SMA

[Old01] Robert L. Oldershaw. M-set as metaphor: The abuse of algebra. Physics Education, 36(5):433, ???? 2001. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic). URL http://

stacks.iop.org/0031-9120/36/i=5/a=605.

Oliver:1991:FRW

[Oli91] Dick Oliver. Fractals in the real world. Dr. Dobb’s Journal of Software Tools, 16(4):68, 71–72, 74–75, 101–107, April 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Oku:2005:DMC

[OMK^{+}05] Naohiko Oku, Kenya Murase, Kazuo Kitagawa, Yasuyuki Kimura, Katsufumi Kajimoto, Ashik Bin Anser, Makiko Tanaka, Hiroki Katoh, Masatsugu Hori, Jun Hatazawa, et al. Diagnosis of mild cognitive impairment using fractal analysis in CBF SPECT. Journal of Cerebral Blood Flow and Metabolism, 25 (1s):S357, July 15, 2005. CODEN JCBMDN. ISSN 0271-678X (print), 1559-7016 (electronic). URL http://www.nature.com/

jcbfm/journal/v25/n1s/full/9591524.0357a.html.

Orbach:1986:DFN

[Orb86] R. Orbach. Dynamics of fractal networks. Science, 231(4740): 814–819, February 21, 1986. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/231/4740/814.full.pdf.

Orlov:1970:GZM

[Orl70] Ju. K. Orlov. A generalization of the Zipf–Mandelbrot law. Sakharth. SSR Mech. Akad. Moambe, 57(??):37–40, ???? 1970. CODEN ???? ISSN ????

Orlov:1976:CBP

[Orl76] Ju. K. Orlov. The connection between the Pareto distribution and the generalized Zipf–Mandelbrot law. Sakharth. SSR Mech. Akad. Moambe, 83(1):57–60, ???? 1976. CODEN ???? ISSN ????

Parsons:1989:IRU

[Par89] Michael S. Parsons. Image representations using Miranda Laws. Computer Graphics Forum, 8(2):99–106, June 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Platzman:1989:SCP

[PB89] Loren K. Platzman and John J. Bartholdi, III. Spacefilling curves and the planar travelling salesman problem. Journal of the Association for Computing Machinery, 36(4):719–737, October 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/76361.html.

Park:2006:UOS

[PC06] Moongyu Park and John H. Cushman. On upscaling operatorstable Lévy motions in fractal porous media. Journal of Computational Physics, 217(1):159–165, September 1, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999106000362.

Pe:2004:XF

[Pe04] Joseph L. Pe. The 3 x +1 fractal. Computers and Graphics, 28(3):431–435, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Peitgen:2010:BBM

[Pei10a] Heinz-Otto Peitgen. Benoît B. Mandelbrot (1924–2010). Science, 330(6006):926, November 12, 2010. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/330/6006/926.full.pdf.

Peitgen:2010:RBB

[Pei10b] Heinz-Otto Peitgen. Retrospective. Benoît B. Mandelbrot (1924– 2010). Science, 330(6006):926, November 12, 2010. CODEN SCI-EAS. ISSN 1095-9203.

Pendry:1988:STD

[Pen88] John Pendry. Symmetry and transport in disordered systems. IBM Journal of Research and Development, 32(1):137–143, January 1988. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Perline:1996:ZLC

[Per96] Richard Perline. Zipf’s law, the Central Limit Theorem, and the random division of the unit interval. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 54(1):220–223, July 1, 1996. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRevE.54.220.

Persson:2010:BM

[Per10] Ulf Persson. Benoît Mandelbrot (1924–2010). Normat: Nordisk matematisk tidskrift, 58(3):139–140, 144, ???? 2010. CODEN ???? ISSN 0801-3500.

Peyriere:1991:MM

[Pey91] Jacques Peyrière. Multifractal measures. In Byrnes et al. [BBBH92], pages 175–186. ISBN 0-7923-1804-8. LCCN ????

Philip:1994:WMM

[PFR94] A. G. Davis Philip, Michael Frame, and Adam Robucci. Warped midgets in the Mandelbrot set. Computers and Graphics, 18(2): 239–248, March–April 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Palmore:1990:CAC

[PH90] J. Palmore and C. Herring. Computer arithmetic, chaos and fractals. Physica D, Nonlinear phenomena, 42(1–3):99–110, June 1990. CODEN PDNPDT. ISSN 0167-2789 (print), 1872-8022 (electronic). Ninth Annual International Conference of the Center for Nonlinear Studies on Self-Organizing, Collective and Cooperative Phenomena in Natural and Artificial Networks.

Pentland:1991:PAF

[PH91] Alexander P. Pentland and Bradley Horowitz. Practical approach to fractal-based image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 1605(??):467– 474, ???? 1991. CODEN PSISDG. ISSN ????

Philip:1992:FLM

[Phi92] Kenelm W. Philip. Field lines in the Mandelbrot set. Computers and Graphics, 16(4):443–447, Winter 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pickover:1986:MCA

[Pic86] C. A. Pickover. A Monte Carlo approach for ϵ placement in fractal-dimension calculations for waveform graphs. Computer Graphics Forum, 5(3):203–209, September 1986. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pickover:1987:GBO

[Pic87] C. A. Pickover. Graphics, bifurcation, order and chaos. Computer Graphics Forum, 6(1):26–33, January 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pickover:1994:APG

[Pic94] Clifford A. Pickover. Automatic parallel generation of Aeolian fractals on the IBM Power Visualization System. Computers and Graphics, 18(3):407–416, May–June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pickover:1995:CFB

[Pic95] Clifford A. Pickover. The crying of fractal Batrachion 1,489. Computers and Graphics, 19(4):611–615, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.elsevier.com/cgi-bin/cas/tree/
store/cag/cas_sub/browse/browse.cgi?year=1995&volume=
19&issue=4&aid=9500039.

Pickstone:2007:BRM

[Pic07] John Pickstone. Book review: Margaret Pelling and Scott Mandelbrote (eds.), The Practice of Reform in Health, Medicine, and Science, 1500–2000: Essays for Charles Webster. Aldershot: Ashgate, 2005. Pp. xv + 376. ISBN 0-7546-3933-9. £60.00 (hardback). British Journal for the History of Science, 40(3):465–466, September 2007. CODEN BJHSAT. ISSN 0007-0874 (print),

1474-001X (electronic). URL http://www.jstor.org/stable/
4500767.

Peitgen:1991:FC

[PJS91] Heinz-Otto Peitgen, Hartmut Jürgens, and Dietmar Saupe. Fractals for the Classroom. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97346-X (vol. 1, New York), 3-540-97346-X (vol. 1, Berlin), 0-387-97722-8 (vol. 2, New York), 3-540-97722-8 (vol. 2, Berlin). xii + 128 (vol.

1) pp. LCCN QA614.86 .P45 1991. US$24.95 (vol. 1), US$39.95 (vol. 2). Issued in 1991–1992, with corrected second printing in 1993–1994.

Peitgen:1992:CFNa

[PJS92a] Heinz-Otto Peitgen, Hartmut Jürgens, and Dietmar Saupe. Chaos and fractals: New frontiers of science. In Chaos and Fractals: New Frontiers of Science [PJS92b], chapter 5, pages 229– 296. ISBN 0-387-97903-4, 3-540-97903-4. LCCN QA614.86 .P43 1992. US$49. With 686 illustrations, 40 in color.

Peitgen:1992:CFNb

[PJS92b] Heinz-Otto Peitgen, Hartmut Jürgens, and Dietmar Saupe. Chaos and Fractals: New Frontiers of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97903-4, 3-540-97903-4. xvi + 984 pp. LCCN QA614.86 .P43 1992. US$49. With 686 illustrations, 40 in color.

Peitgen:1992:EIS

[PJS92c] Heinz-Otto Peitgen, Hartmut Jürgens, and Dietmar Saupe. Encoding images by simple transformations. In Chaos and Fractals: New Frontiers of Science, chapter 5, pages 229–296. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97903-4.

Peitgen:1992:FCPa

[PJS92d] Heinz-Otto Peitgen, Hartmut Jürgens, and Dietmar Saupe. Fractals for the Classroom. Part One: Introduction to Fractals and Chaos. Springer-Verlag, New York, NY, USA, 1992. ISBN 0-387-97041-X. xiv + 450 pp. LCCN ????

Peitgen:1992:FCPb

[PJS92e] Heinz-Otto Peitgen, Hartmut Jürgens, and Dietmar Saupe. Fractals for the Classroom. Part Two: Complex Systems and Mandelbrot Set. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97722-8. xii + 500 pp. LCCN ????

Peitgen:1993:CFN

[PJS93] Heinz-Otto Peitgen, Hartmut Jürgens, and Dietmar Saupe. Chaos and Fractals: New Frontiers of Science (corrected second printing). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-97903-4, 3-540-97903-4. xvi + 984, with 686 illustrations, 40 in color pp. LCCN QA614.86 .P43 1993. US$49.

Peitgen:1999:FCS

[PJS99] Heinz-Otto Peitgen, Hartmut Jürgens, and Dietmar Saupe. Fractals for the classroom: strategic activities. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98420-8 (vol. 3). xiv + 107 pp. LCCN QA614.86 .F724 1999. URL http://www.loc.gov/

catdir/enhancements/fy0815/2003535060-d.html;   http:

//www.loc.gov/catdir/enhancements/fy0815/2003535060-t.
html.

Peitgen:2004:CFN

[PJS04a] Heinz-Otto Peitgen, H. (Hartmut) Jürgens, and Dietmar Saupe. Chaos and fractals: new frontiers of science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2004. ISBN 0-387-20229-3. xiii + 864 pp. LCCN Q172.5.C45 P45 2004. URL http://www.loc.gov/

catdir/enhancements/fy0818/2003063341-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0818/2003063341-t.
html.

Peitgen:2004:CGH

[PJS04b] Heinz-Otto Peitgen, H. (Hartmut) Jürgens, and Dietmar Saupe. The chaos game: How randomness creates deterministic shapes. In Chaos and fractals: new frontiers of science [PJS04a], chapter 7, pages 277–327. ISBN 0-387-20229-

3. LCCN Q172.5.C45 P45 2004. URL http://www.loc.gov/

catdir/enhancements/fy0818/2003063341-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0818/2003063341-t.
html.

Peskin:1994:MED

[PM94] C. S. Peskin and D. M. McQueen. Mechanical equilibrium determines the fractal fiber architecture of the aortic heart valve leaflets. American Journal of Physiology, 266(1):H319–H328, January 1994. CODEN AJPHAP. ISSN 0002-9513 (print), 2163-5773 (electronic).

Pool:1990:FF

[Poo90] Robert Pool. Fractal fracas. Science, 249(4967):363–364, July 27, 1990. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/249/
4967/363.full.pdf.

Popova:2022:FHO

[Pop22] Maria Popova. Fractals, the hidden order beneath chaos, and the story of the refugee who revolutionized the mathematics of reality. Web site, February 22, 2022. URL https://getpocket.com/

explore/item/the-pattern-inside-the-pattern-fractals-
the-hidden-order-beneath-chaos-and-the-story-of-the-refugee.

Pountain:1986:TM

[Pou86] Dick Pountain. Turbocharging Mandelbrot. BYTE Magazine, 11 (9):359–366, September 1986. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Peitgen:1984:HCK

[PR84a] Heinz-Otto Peitgen and Peter H. Richter. Harmonie in Chaos und Kosmos. Bilder aus d. Theorie dynamischer Systeme. (German) [Harmony in chaos and the cosmos. Pictures from the theory of dynamical systems]. Die Sparkasse in Bremen, Bremen, West Germany, 1984. 63 pp. LCCN ????

Peitgen:1984:MKG

[PR84b] Heinz Otto Peitgen and Peter H. Richter, editors. Morphologie komplexer Grenzen: Bilder aus der Theorie dynamischer Systeme. Ausstellung vom 27. Mai–9. Juni 1984 im Max-Planck-Institut für Biophysikalische Chemie, Verwaltungsgebäude am Fassberg, Göttingen u.a.. (German) [The morphology of complex limits: Pictures from the theory of dynamical systems. Exhibition from 27 May to 9 June 1984 at the Max Planck Institute for Biophysical Chemistry, Administration building at Fassberg, Göttingen]. Forschungsgruppe “Komplexe Dynamik”, Universität Bremen, Bremen, West Germany, 1984. 47 pp. LCCN ???? URL

http://www.gbv.de/dms/ohb-opac/01903640X.pdf.

Peitgen:1985:MSM

[PR85] Heinz-Otto Peitgen and Peter H. Richter. The Mandelbrot set in a model for phase transitions. Lecture Notes in Mathematics, 1111:111–134, 1985. CODEN LNMAA2. ISBN 3-540-15195-8 (print), 3-540-39298-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0084587/.

Peitgen:1986:BF

[PR86] Heinz-Otto Peitgen and Peter H. Richter. The Beauty of Fractals. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 0-387-15851-0. xii + 199 pp. LCCN QA447 .P45 1986.

Pastor:2004:CBM

[PRÁM04] G. Pastor, M. Romera, G. Álvarez, and F. Montoya. Chaotic bands in the Mandelbrot set. Computers and Graphics, 28(5):

779–784, October 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Prodinger:1995:SPY

[Pro95] H. Prodinger. Solution of a problem of Yekutieli and Mandelbrot. Lecture Notes in Computer Science, 911(??):461–??, ???? 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prodinger:1997:PYM

[Pro97] Helmut Prodinger. On a problem of Yekutieli and Mandelbrot about the bifurcation ratio of binary trees. Theoretical Computer Science, 181(1):181–194, July 15, 1997. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.com/cgi-bin/cas/tree/
store/tcs/cas_sub/browse/browse.cgi?year=1997&volume=
181&issue=1&aid=2471.

Peitgen:1988:SFI

[PS88a] Heinz-Otto Peitgen and Dietmar Saupe, editors. The Science of Fractal Images. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-96608-0 (New York), 1-4612-8349-3, 3-540-96608-0 (Berlin). xiii + 312 pp. LCCN QA614.86 .S35 1988.

Prusinkiewicz:1988:KCA

[PS88b] Przemyslaw Prusinkiewicz and Glen Sandness. Koch curves as attractors and repellers. IEEE Computer Graphics and Applications, 8(6):26–40, November 1988. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Peitgen:1984:CPJ

[PSvH84] H.-O. Peitgen, D. Saupe, and F. von Haeseler. Cayley’s problem and Julia sets. The Mathematical Intelligencer, 6(??):11–20, ???? 1984. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Pietronero:1986:FPP

[PT86] L. (Luciano) Pietronero and E. (Erio) Tosatti, editors. Fractals in physics: proceedings of the Sixth Trieste International Symposium on Fractals in Physics, ICTP, Trieste, Italy, July 9–12, 1985. North-Holland Publishing Co., Amsterdam, The Netherlands, 1986. ISBN 0-444-86995-6 (U.S.). LCCN QC20.7.G44 I585 1985.

Paramanathan:2008:ACF

[PU08] P. Paramanathan and R. Uthayakumar. An algorithm for computing the fractal dimension of waveforms. Applied Mathematics and Computation, 195(2):598–603, February 1, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300307005899.

Pumar:1996:ZTI

[Pum96] M. A. Pumar. Zooming of terrain imagery using fractalbased interpolation. Computers and Graphics, 20(1):171–176, January–February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http://www.elsevier.

com/cgi-bin/cas/tree/store/cag/cas_sub/browse/browse.
cgi?year=1996&volume=20&issue=1&aid=9500102.

Romera:2008:DCE

[RAA^{+}08] M. Romera, G. Alvarez, D. Arroyo, A. B. Orue, V. Fernandez, and G. Pastor. Drawing and computing external rays in the multiple-spiral medallions of the Mandelbrot set. Computers and Graphics, 32(5):597–610, October 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0097849308000484.

Rachev:2003:HHT

[Rac03] Svetlozar T. Rachev, editor. Handbook of heavy tailed distributions in finance, volume 1 of Handbooks in Finance. Elsevier, Amsterdam, The Netherlands, 2003. ISBN 0-444-50896-1. xxiv + 680 pp. LCCN ????

Radai:1996:CAM

[Rad96] Yisrael Radai. Computer art from the Mandelbrot set. Computers and Graphics, 20(6):925–926, November–December 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.elsevier.com/cgi-bin/cas/tree/
store/cag/cas_sub/browse/browse.cgi?year=1996&volume=
20&issue=6&aid=9600062.

Raidl:2003:AEC

[Rai03] Gunther Raidl, editor. Applications of evolutionary computing: EvoWorkshops 2003: EvoBIO, EvoCOP, EvoIASP, Evo-MUSART, EvoROB, and EvoSTIM, Essex, UK, April 14–16, 2003: Proceedings, volume 2611 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNCSD9. ISBN 3-540-00976-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.618 .E899 2003. URL http:

//link.springer-ny.com/link/service/series/0558/tocs/
t2611.htm;   http://www.springerlink.com/content/978-
3-540-00976-4;   http://www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&volume=2611.

Rankin:1991:RBL

[Ran91] John R. Rankin. Recursive bisection line algorithm. Computers and Graphics, 15(1):1–8, ???? 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rarity:1989:CSF

[Rar89] John Rarity. Colloids stick to fractal rules. Nature, 339(6223): 340–341, June 1, 1989. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v339/n6223/pdf/339340a0.pdf.

Ratliff:1993:MFI

[Rat93] Michael I. Ratliff. Mandelbrot’s functional iteration and continued fractions. Fibonacci Quarterly, 31(3):263–267, August 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/31-3/ratliff.pdf.

Robinson:1979:PBB

[RB79] Peter D. Robinson and Michael F. Barnsley. Pointwise bivariational bounds on solutions of Fredholm integral equations. SIAM Journal on Numerical Analysis, 16(1):135–144, February 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Redmill:1996:GAF

[RBM96] David W. Redmill, David R. Bull, and Ralph R. Martin. Genetic algorithms for fast search in fractal image coding. Proceedings of the SPIE — The International Society for Optical Engineering, 2727(??):1367–1376, ???? 1996. CODEN PSISDG. ISSN ????

Ritter:2006:RTI

[RHD^{+}06] Felix Ritter, Christian Hansen, Volker Dicken, Olaf Konrad, Bernhard Preim, and Heinz-Otto Peitgen. Real-time illustration of vascular structures. IEEE Transactions on Visualization and Computer Graphics, 12(5):877–884, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rice:1994:FC

[Ric94] Sean Rice. Fractal colonies. Science, 266(5185):664–665, October 28, 1994. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/266/
5185/664.2.full.pdf.

Riedi:1995:MFI

[RM95] Rudolf H. Riedi and Benoît B. Mandelbrot. Multifractal formalism for infinite multinomial measures. Advances in Applied Mathematics, 16(2):132–150, ???? 1995. CODEN ???? ISSN 0196-8858 (print), 1090-2074 (electronic).

Riedi:1997:IFC

[RM97] Rudolf H. Riedi and Benoît B. Mandelbrot. Inversion formula for continuous multifractals. Advances in Applied Mathematics, 19 (3):332–354, ???? 1997. CODEN ???? ISSN 0196-8858 (print), 1090-2074 (electronic).

Riedi:1998:EMF

[RM98] Rudolf H. Riedi and Benoît B. Mandelbrot. Exceptions to the multifractal formalism for discontinuous measures. Mathematical Proceedings of the Cambridge Philosophical Society, 123(1):133– 157, ???? 1998. CODEN MPCPCO. ISSN 0305-0041 (print), 1469-8064 (electronic).

Robinson:1985:FFV

[Rob85] Arthur L. Robinson. Fractal fingers in viscous fluids. Science, 228(4703):1077–1080, May 31, 1985. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/228/4703/1077.full.pdf.

Rojas:1991:TEC

[Roj91] Raul Rojas. A tutorial on efficient computer graphic representations of the Mandelbrot set. Computers and Graphics, 15(1): 91–100, ???? 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rood:2010:FLM

[Roo10] Will Rood. Fractal limits: The Mandelbrot set and the selfsimilar tilings of M. C. Escher. In Lesmoir-Gordon [LG10b], chapter 5, pages 74–87. ISBN 1-84996-485-8. LCCN QA614.86

.C65 2010. URL http://www.springerlink.com/content/978-

1-84996-485-2. With contributions by Ian Stewart, Sir Arthur C. Clarke, Benoît B. Mandelbrot, Michael and Louisa Barnsley, Will Rood, Gary Flake, and David Pennock.

Romera:2004:EAD

[RPÁM04] M. Romera, G. Pastor, G. Álvarez, and F. Montoya. External arguments of Douady cauliflowers in the Mandelbrot set. Computers and Graphics, 28(3):437–449, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Romera:2006:EAM

[RPAM06] M. Romera, G. Pastor, G. Alvarez, and F. Montoya. External arguments in the multiple-spiral medallions of the Mandelbrot set. Computers and Graphics, 30(3):460–469, June 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0097849306000690.

Risser:2006:HFN

[RPS^{+}06] Laurent Risser, Franck Plouraboué, Alexandre Steyer, Peter Cloetens, Géraldine Le Duc, Caroline Fonta, et al. From homogeneous to fractal normal and tumorous microvascular networks in the brain. Journal of Cerebral Blood Flow and Metabolism, 27 (2):293–303, May 24, 2006. CODEN JCBMDN. ISSN 0271-678X (print), 1559-7016 (electronic). URL http://www.nature.com/

jcbfm/journal/v27/n2/full/9600332a.html.

Ribeiro:2006:MQA

[RRB06] Vinay J. Ribeiro, Rudolf H. Riedi, and Richard G. Baraniuk. Multiscale queueing analysis. IEEE/ACM Transactions on Networking, 14(5):1005–1018, October 2006. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Roughan:2000:RTE

[RVA00] Matthew Roughan, Darryl Veitch, and Patrice Abry. Realtime estimation of the parameters of long-range dependence. IEEE/ACM Transactions on Networking, 8(4):467–478, August 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-

2566 (electronic). URL http://www.acm.org/pubs/citations/
journals/ton/2000-8-4/p467-roughan/.

Resch:1989:CAC

[RW89] Mark Resch and Deborah Williams, editors. Computer Art in Context: SIGGRAPH ’89 Art Show Catalog. July 31–August 4, 1989, SIGGRAPH ’89, Boston, Massachusetts and June 28–September 5 1989, The Computer Museum, Boston, Massachusetts. Pergamon, New York, NY, USA, 1989. CODEN ???? ISBN 0-08-037921-4. ISSN 0024-094X (print), 1530-9282 (electronic). LCCN N 7433.8 C66 1989. URL http://info.acm.org/

pubs/contents/proceedings/graph/. Also published as 1989

Supplement to Leonardo: Journal of the International Society for the Arts, Sciences and Technology.

Sahimi:1993:FPR

[Sah93] Muhammad Sahimi. Flow phenomena in rocks: from continuum models to fractals, percolation, cellular automata, and simulated annealing. Reviews of Modern Physics, 65(4):1393–1534, October 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.65.1393;   http://rmp.aps.org/
abstract/RMP/v65/i4/p1393_1.

Stauffer:1993:SSC

[SAM93] D. Stauffer, A. Aharony, and B. B. Mandelbrot. Self-similarity and covered neighborhoods of fractals: a random walk test. Physica A, 196(1):1–5, ???? 1993. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

Sander:1985:SIP

[San85] L. M. Sander. Scale-invariant phenomena: Viscous fingers and fractal growth. Nature, 314(6010):405–406, April 4, 1985. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v314/
n6010/pdf/314405a0.pdf.

Sander:1986:FGP

[San86] Leonard M. Sander. Fractal growth processes. Nature, 322(6082): 789–793, August 28, 1986. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v322/n6082/pdf/322789a0.pdf.

Sander:1987:FGa

[San87a] Leonard M. Sander. Fractal growth. Scientific American, 256 (1):82–88, January 1987. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Sander:1987:FGb

[San87b] Leonard M. Sander. Fractal growth. Scientific American, 256 (1):94–100, January 1987. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v256/n1/pdf/scientificamerican0187-
94.pdf.

Saupe:1987:ECJ

[Sau87] Dietmar Saupe. Efficient computation of Julia sets and their fractal dimension. Physica D, Nonlinear phenomena, 28(??):358– 370, ???? 1987. CODEN PDNPDT. ISSN 0167-2789 (print), 1872-8022 (electronic).

Saupe:1996:FSI

[Sau96a] Dietmar Saupe. The futility of square isometries in fractal image compression. In Proceedings ICIP-96 (IEEE International Conference on Image Processing), volume I, pages 161–164. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, September 1996.

Saupe:1996:LDP

[Sau96b] Dietmar Saupe. Lean domain pools for fractal image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2669(??):150–157, January 1996. CODEN PSISDG. ISSN ????

Samuel:2012:IBM

[SBvB^{+}12] Nina Samuel, Matthias Bruhn, Jan von Brevern, Juliet Koss, Margarete Pratschke, and Wlaldimir Velminski, editors. The islands of Benoît Mandelbrot: fractals, chaos, and the materiality of thinking. Bard Graduate Center: Decorative Arts, Design History, Material Culture, New Haven, CT, USA, 2012. ISBN 0-300-18643-6. ???? pp. LCCN Q223 .I85 2012.

Stepoway:1988:PRF

[SC88] Stephen L. Stepoway and Michael Christiansen. Parallel rendering of fractal surfaces. International Journal of Parallel Programming, 17(1):43–58, February 1988. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0885-
7458&volume=17&issue=1&spage=43.

Soehle:2007:VFA

[SCC^{+}07] Martin Soehle, Marek Czosnyka, Doris A. Chatfield, Andreas Hoeft, and Alonso Peña. Variability and fractal analysis of middle cerebral artery blood flow velocity and arterial blood pressure in subarachnoid hemorrhage. Journal of Cerebral Blood Flow and Metabolism, 28(1):64–73, May 2, 2007. CODEN JCBMDN. ISSN 0271-678X (print), 1559-7016 (electronic). URL http://www.

nature.com/jcbfm/journal/v28/n1/full/9600506a.html.

Schroeder:1986:PMS

[Sch86] Peter B. Schroeder. Plotting the Mandelbrot set. BYTE Magazine, 11(13):207–211, December 1986. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schroeder:1987:TST

[Sch87] Manfred R. Schroeder. Is there such a thing as fractal music? Nature, 325(6107):765–766, February 26, 1987. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v325/
n6107/pdf/325765c0.pdf.

Schaefer:1989:PFC

[Sch89] Dale W. Schaefer. Polymers, fractals, and ceramic materials. Science, 243(4894):1023–1027, February 24, 1989. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/243/4894/1023.full.pdf.

Schechter:1990:CHM

[Sch90] Dan Schechter. CUG305 HGA Mandelbrot Explorer and card games. C Users Journal, 8(3):131–??, March 1990. CODEN ???? ISSN 0898-9788.

Schilling:2005:RBC

[Sch05] René L. Schilling. Review: Fractals and Chaos. The Mandelbrot Set and beyond by Benoît B. Mandelbrot. Mathematical Gazette, 89(514):168, March 2005. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL http://www.jstor.org/

stable/3620693.

Schleicher:2011:CDM

[Sch11] Dierk Schleicher. Complex dynamics, the Mandelbrot set, and Newton’s method: or, on useless and useful mathematics. In Schleicher and Lackmann [SL11], pages 207–220. ISBN 3-642-19532-6 (soft cover), 3-642-19533-4 (e-book). LCCN

????  URL http://link.springer.com/chapter/10.1007/
978-3-642-19533-4_14.

Spehar:2003:UAF

[SCNT03] Branka Spehar, Colin W. G. Clifford, Ben R. Newell, and Richard P. Taylor. Universal aesthetic of fractals. Computers and Graphics, 27(5):813–820, October 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Selinker:2010:NTJ

[Sel10] Mike Selinker. Never trend away: Jonathan Coulton on Benoît Mandelbrot. Wired, ??(??):??, October 18, 2010. CODEN WREDEM. ISSN ???? URL http://www.

wired.com/magazine/2010/10/never-trend-away-jonathan-
coulton-on-benoit-mandelbrot/.

Senn:1990:MSB

[Sen90] Peter Senn. The Mandelbrot set for binary numbers. American Journal of Physics, 58:1018–??, 1990. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.

aip.org/link/ajpias/v58/i10/p1018/s1.

Setti:1992:GEH

[Set92] Alberto G. Setti. Gaussian estimates for the heat kernel of the weighted Laplacian and fractal measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(??):1061– 1078, ???? 1992. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Seuront:2011:BBM

[Seu11] Laurent Seuront. Benoît B. Mandelbrot (1924–2010) [obituary]. Journal of Plankton Research, 33(6):983–988, June 2011. CO-DEN JPLRD9. ISSN 0142-7873 (print), 1464-3774 (electronic).

Seuront:2012:OBB

[Seu12] Laurent Seuront. Obituary: Benoît B. Mandelbrot (1924–2010) (vol 33, pg 983). Journal of Plankton Research, 34(2):190, February 2012. CODEN JPLRD9. ISSN 0142-7873 (print), 1464-3774 (electronic).

Sole:2000:SLH

[SG00] Ricard Vicente Solé and Brian Goodwin. Signs of Life: How Complexity Pervades Biology. Basic Books, New York, NY, USA, 2000. ISBN 0-465-01927-7, 0-465-01928-5 (paperback). xi + 322 pp. LCCN QH501 .S65 2000.

Sleeman:1992:ABC

[SH92] B. D. Sleeman and Chen Hua. An analogue of Berry’s conjecture for the phase in fractal obstacle scattering. IMA Journal of Applied Mathematics, 49(3):193–202, ???? 1992. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Saupe:1996:LAF

[SH96] Dietmar Saupe and Hannes Hartenstein. Lossless acceleration of fractal image compression by fast convolution. In Proceedings ICIP-96 (IEEE International Conference on Image Processing), volume I, pages 185–188. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, September 1996.

Shirriff:1993:IFG

[Shi93] Ken W. Shirriff. An investigation of fractals generated by z → z^{−n} + c. Computers and Graphics, 17(5):603–607, September– October 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Shishikura:1998:HDB

[Shi98] Mitsuhiro Shishikura. The Hausdorff dimension of the boundary of the Mandelbrot set and Julia sets. Annals of Mathematics, 147 (2):225–267, March 1998. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL http://www.jstor.org/

stable/121009.

Shlyk:2005:SSE

[Shl05] V. A. Shlyk. He has scratched the surface of everything: on the eightieth birthday of Benoît Mandelbrot. Izv. Chelyabinsk. Nauchn. Tsentra, 3(29):107–124, ???? 2005. CODEN ???? ISSN 1727-7434.

Signes:1996:RCS

[Sig96] Julien Signes. Reducing the codebook size in fractal image compression by geometrical analysis. Proceedings of the SPIE — The International Society for Optical Engineering, 2727(??): 1400–1409, ???? 1996. CODEN PSISDG. ISSN ????

Simon:1955:CSD

[Sim55] Herbert A. Simon. On a class of skew distribution functions. Biometrika, 42(3–4):425–440, December 1955. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333389.

Simon:1960:SFN

[Sim60] Herbert A. Simon. Some further notes on a class of skew distribution functions. Information and Control, 3(1):80–88, March 1960. CODEN IFCNA4. ISSN 0019-9958 (print), 1878-2981

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0019995860903028.

Simon:1961:RDM

[Sim61a] Herbert A. Simon. Reply to Dr. Mandelbrot’s post scriptum. Information and Control, 4(2–3):305–308, September 1961. CO-DEN IFCNA4. ISSN 0019-9958 (print), 1878-2981 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0019995861800260.

Simon:1961:RFN

[Sim61b] Herbert A. Simon. Reply to “Final note” by Benoît Mandelbrot. Information and Control, 4(2–3):217–223, September 1961. CO-DEN IFCNA4. ISSN 0890-5401 (print), 1090-2651 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0019995861800090.

Singh:2007:VU

[Sin07] Gary Singh. Vamping on ultrafractal. IEEE Computer Graphics and Applications, 27(1):4–5, January/February 2007. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic).  URL http://csdl.computer.org/comp/mags/cg/
2007/01/g1004.pdf.

Sirvent:1997:SDS

[Sir97] Víctor F. Sirvent. On some dynamical subsets of the Rauzy fractal. Theoretical Computer Science, 180(1–2):363–370, June 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.com/cgi-bin/cas/tree/
store/tcs/cas_sub/browse/browse.cgi?year=1997&volume=
180&issue=1-2&aid=2533.

Skarbek:1999:PCD

[Ska99] W. Skarbek. Perceptual convergence of discrete clamped fractal operator. Lecture Notes in Computer Science, 1609(??):395–??, ???? 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smithe:1997:ESA

[SL97] David Smithe and Larry Ludeking. An eigenmode solution algorithm based on high-order power iteration with fractally ordered shifts. Computer Physics Communications, 106(1–2):95– 104, October 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465597000817.

Schleicher:2011:IMC

[SL11] Dierk Schleicher and Malte Lackmann, editors. An Invitation to Mathematics: From Competitions to Research. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-19532-6 (soft cover), 3-642-19533-4 (e-book). xiv + 220 pp. LCCN ????

Slater:2007:FFR

[Sla07] P. B. Slater. Fractal fits to Riemann zeros. Canadian Journal of Physics = Journal canadien de physique, 85(4):345–357, 2007. CODEN CJPHAD. ISSN 0008-4204 (print), 1208-6045 (elec-

tronic). URL http://www.nrcresearchpress.com/doi/abs/
10.1139/p07-050.

Sloan:1994:LBR

[Slo94] Alan D. Sloan. Low-bit-rate fractal image coding. Proceedings of the SPIE — The International Society for Optical Engineering, 2239(??):210–213, ???? 1994. CODEN PSISDG. ISSN ????

Scholz:1989:I

[SM89a] C. H. Scholz and B. B. Mandelbrot. Introduction. Pure and Applied Geophysics, 131(1–2):1–4, ???? 1989. CODEN PAGYAV.

ISSN ???? URL http://www.springerlink.com/content/
x26l140400101035/.

Scholz:1989:FG

[SM89b] C. H. (Christopher H.) Scholz and Benoît B. Mandelbrot, editors. Fractals in geophysics, volume 131(1/2) of Pure and applied geophysics. Birkhäuser Verlag, Basel, Switzerland, 1989. ISBN 0-8176-2206-3 (U.S.). 313 pp. LCCN QC801 .G37 v.131, no.1/2; QE33.2.M3 F73 1989. Reprinted from Pure and applied geophysics (PAGEOPH), volume 131 (1989), no. 1.

Smith:1984:PFF

[Smi84] Alvy Ray Smith III. Plants, fractals, and formal languages. Computer Graphics, 18(3):1–10, July 1984. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic). Applications of formal grammars for generating graphic fractals — “graftals” — in the modeling of plants.

Smirnov:1991:NMC

[Smi91] B. M. Smirnov. From Newton to Mandelbrot: concepts in theoretical physics. Soviet Physics. Uspekhi, 34(11):1021, November 1991. CODEN SOPUAP. ISSN 0038-5670. URL http://ufn.

ru/en/articles/1991/11/h/.

Shlesinger:1984:FPS

[SMR84] Michael F. Shlesinger, Benoît B. Mandelbrot, and Robert J. Rubin, editors. Fractals in the physical sciences. Proceedings of the Symposium on Fractals in the Physical Sciences, held at the U.S. National Bureau of Standards, Gaithersburg, Maryland, November 21–23, 1983., volume 36(5/6) of Journal of Statistical Physics. Plenum Press, New York, NY, USA; London, UK, 1984. LCCN ????

Sommerer:1993:PFM

[SO93] John C. Sommerer and Edward Ott. Particles floating on a moving fluid: A dynamically comprehensible physical fractal. Science, 259(5093):335–339, January 15, 1993. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/259/5093/335.full.pdf.

Sparrow:1984:BRB

[Spa84] Colin Sparrow. Book review: The Fractal Geometry of Nature, by B. Mandelbrot. Journal of the Royal Statistical Society. Series A (General), 147(4):616–618, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.org/stable/2981858.

Saupe:1996:EFI

[SR96] Dietmar Saupe and Matthias Ruhl. Evolutionary fractal image compression. In Proceedings ICIP-96 (IEEE International Conference on Image Processing), volume I, pages 129–132. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, September 1996.

Shallit:1989:TMG

[SS89] Jeffrey Shallit and Jorge Stolfi. Two methods for generating fractals. Computers and Graphics, 13(2):185–191, ???? 1989. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Steacy:1991:AFP

[SS91] Sandra J. Steacy and Charles G. Sammis. An automaton for fractal patterns of fragmentation. Nature, 353(6341):250–252, September 19, 1991. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v353/n6341/pdf/353250a0.pdf.

Stark:2000:HSF

[SS00] Colin P. Stark and Marc Stieglitz. Hydrology: The sting in a fractal tail. Nature, 403(6769):493–495, February 3, 2000. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v403/
n6769/full/403493a0.html.

Saka:1993:PSF

[SSK93] G. Saka, F. Schroder, and H.-J. Koppert. Pseudo-satellite film using fractal clouds to enhance animated weather forecasting. Computer Graphics Forum, 12(3):C329–C338, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szeliski:1989:SF

[ST89] Richard Szeliski and Demetri Terzopoulos. From splines to fractals. Computer Graphics, 23(3):51–60, July 1989. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569

(electronic). URL http://www.acm.org:80/pubs/citations/
proceedings/graph/74333/p51-szeliski/.

Staiger:1996:CSW

[Sta96] Ludwig Staiger. Codes, simplifying words, and open set condition. Information Processing Letters, 58(6):297–301, June 24, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stewart:1988:BFD

[Ste88] Ian Stewart. The beat of a fractal drum. Nature, 333(6170):206– 207, May 19, 1988. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v333/n6170/pdf/333206a0.pdf.

Stewart:1998:DJN

[Ste98] Ian Stewart. Dieu joue-t-il aux dés ?: les nouvelles mathématiques du chaos. Flammarion, Paris, France, second edition, 1998. ISBN 2-08-081411-7 (hardcover), 2-08-081302-1 (paperback). vii + 602 pp. LCCN ???? Preface by Benoît Mandelbrot. Translation by Marianne Robert from the English edition. Remarks for the second edition by Marcel Filoche.

Stewart:2000:FGT

[Ste00a] Ian Stewart. A fractal guide to tic-tac-toe. Scientific American, 283(2):86–88, August 2000. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v283/n2/pdf/scientificamerican0800-
86.pdf.

Stewart:2000:MRF

[Ste00b] Ian Stewart. Mathematical recreations: a fractal guide to tic-tac-toe. Scientific American, 283(2):86–??, August 2000. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Stewart:2010:NFG

[Ste10] Ian Stewart. The nature of fractal geometry. In Lesmoir-Gordon [LG10b], chapter 1, pages 2–23. ISBN 1-84996-485-8. LCCN

QA614.86 .C65 2010. URL http://www.springerlink.com/

content/978-1-84996-485-2. With contributions by Ian Stewart, Sir Arthur C. Clarke, Benoît B. Mandelbrot, Michael and Louisa Barnsley, Will Rood, Gary Flake, and David Pennock.

Sato:1999:NCS

[STI99] Yuzuru Sato, Makoto Taiji, and Takashi Ikegami. NPcompleteness of k SAT and multifractals. Computer Physics Communications, 121–122(??):51–53, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0010465599002787.

Stoyan:1979:BRM

[Sto79] D. Stoyan. Book review: Mandelbrot, B. B., Fractals: Form, Chance, and Dimension. San Francisco. W. H. Freeman and Company. 1977. 352 S., 68 Abb., $14.95. Zeitschrift für Angewandte Mathematik und Mechanik, 59(8):402–403, ???? 1979. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Suzuki:1991:GTF

[Suz91] Masuo Suzuki. General theory of fractal path integrals with applications to many-body theories and statistical physics. Journal of Mathematical Physics, 32(2):400–407, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i2/
p400_s1.

Sastre-Vazquez:2009:IGZ

[SVVRGA09] Patricia Sastre-Vazquez, Yolanda Villacampa, Jose A. Reyes, and Fernando Garcia-Alonso. Interpretation of the generalized Zipf– Mandelbrot law parameters. Cybernetics and Systems, 40(4):326– 336, ???? 2009. CODEN CYSYDH. ISSN 0196-9722 (print), 1087-6553 (electronic).

Sakas:1992:FAV

[SW92] G. Sakas and R. Westermann. A functional approach to the visual simulation of gaseous turbulence. Computer Graphics Forum, 11(3):C107–C117, C467–C468, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Swehosky:1991:RBC

[Swe91] Frank J. Swehosky. Review: Transition to Chaos: The Orbit Diagram and the Mandelbrot Set by Robert L. Devaney. The Mathematics Teacher, 84(9):773–774, ???? 1991. CODEN ???? ISSN

0025-5769. URL http://www.jstor.org/stable/27967426.

Taqqu:2013:BMF

[Taq13] Murad S. Taqqu. Benoît Mandelbrot and fractional Brownian motion. Statistical Science, 28(1):131–134, February 2013. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1359468412.

Taubes:1998:FRN

[Tau98] Gary Taubes. Fractals reemerge in the new math of the Internet. Science, 281(5385):1947–1948, September 25, 1998. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL

http://www.sciencemag.org/content/281/5385/1947.full.

Taylor:1985:F

[Tay85] S. James Taylor. Fractals. Science, 229(4720):1381, September 27, 1985. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/229/
4720/1381.1.full.pdf.

Taylor:2011:BMO

[Tay11] Richard Taylor. Benoît Mandelbrot [obituary]. Physics Today, 64(6):63–??, June 2011. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v64/i6/p63/s1.

Taylor:2013:BRF

[Tay13] Richard Taylor. Book review: The Fractalist: Memoir of a Scientific Maverick, Benoît B. Mandelbrot, Pantheon Books, 2012. $30.00 (352 pp.). ISBN 978-0-307-37735-7. Physics Today, 66 (11):52–55, November 2013. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Timbo:2009:CCC

[TdRGD09] C. Timbó, L. A. R. da Rosa, M. Gonçalves, and S. B. Duarte. Computational cancer cells identification by fractal dimension analysis. Computer Physics Communications, 180(6):850–853, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508004190.

Tsonis:1992:NPW

[TE92] A. A. Tsonis and J. B. Elsner. Nonlinear prediction as a way of distinguishing chaos from random fractal sequences. Nature, 358 (6383):217–220, July 16, 1992. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v358/n6383/pdf/358217a0.pdf.

Tempczyk:1996:FGC

[Tem96] Michal Tempczyk. Fractal geometry — the case of a rapid career. International Studies in the Philosophy of Science, 10(1): 53–65, 1996. CODEN ???? ISSN 0269-8595 (print), 1469-

9281 (electronic). URL http://www.tandfonline.com//doi/
abs/10.1080/02698599608573529.

Tenner:2010:BMM

[Ten10] Edward Tenner. Benoît Mandelbrot the maverick, 1924– 2010. The Atlantic, ??(??):??, October 16, 2010. CO-

DEN ???? ISSN ???? URL http://www.theatlantic.
com/technology/archive/2010/10/benoit-mandelbrot-the-
maverick-1924-2010/64684/.

Terdik:2009:LFF

[TG09] György Terdik and Tibor Gyires. Lévy flights and fractal modeling of Internet traffic. IEEE/ACM Transactions on Networking, 17(1):120–129, February 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Thiebaut:1988:FDI

[Thi88] Dominique Thiebaut. From the fractal dimension of the intermiss gaps to the cache-miss ratio. IBM Journal of Research and Development, 32(6):796–803, November 1988. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Thiebaut:1989:FDC

[Thi89] D. Thiebaut. On the fractal dimension of computer programs and its application to the prediction of the cache miss ratio. IEEE Transactions on Computers, 38(7):1012–1026, July 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-

9956 (electronic). URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=30852.

Tiihonen:1997:CBR

[TKR^{+}97] J. Tiihonen, J. Kuikka, P. Räsänen, U. Lepola, H. Koponen, A. Liuska, A. Lehmusvaara, P. Vainio, M. Könönen, K. Bergström, M. Yu, I. Kinnunen, K. Åkerman, and J. Karhu.

Cerebral benzodiazepine receptor binding and distribution in generalized anxiety disorder: a fractal analysis. Molecular Psychiatry, 2(6):463–471, October 23, 1997. CODEN MOPSFQ. ISSN 1359-4184 (print), 1476-5578 (electronic). URL http://www.

nature.com/mp/journal/v2/n6/pdf/4000329a.pdf.

Termonia:1986:FFC

[TM86] Yves Termonia and Paul Meakin. Formation of fractal cracks in a kinetic fracture model. Nature, 320(6061):429–431, April 3, 1986. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v320/n6061/pdf/320429a0.pdf.

Taylor:1999:FAP

[TMJ99] Richard P. Taylor, Adam P. Micolich, and David Jonas. Fractal analysis of Pollock’s drip paintings. Nature, 399(6735):422, June 3, 1999. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v399/n6735/full/399422a0.html.

Taylor:2006:FAR

[TMJ06] R. P. Taylor, A. P. Micolich, and D. Jonas. Fractal analysis: Revisiting Pollock’s drip paintings (reply). Nature, 444(7119): E10–E11, November 29, 2006. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v444/n7119/full/nature05399.html.

Turiel:2006:NME

[TPVG06] Antonio Turiel, Conrad J. Pérez-Vicente, and Jacopo Grazzini. Numerical methods for the estimation of multifractal singularity spectra on sampled data: a comparative study. Journal of Computational Physics, 216(1):362–390, July 20, 2006. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999105005565.

Tang:2010:FIR

[TQ10] Xiutao Tang and Cuilu Qu. Facial image recognition based on fractal image encoding. Bell Labs Technical Journal, 15(1):209– 214, June 2010. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Turcotte:1985:CLF

[TSS85] D. L. Turcotte, R. F. Smalley, and Sara A. Solla. Collapse of loaded fractal trees. Nature, 313(6004):671–672, February 21, 1985. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v313/n6004/pdf/313671a0.pdf.

Tsujii:2008:FMT

[Tsu08] Kaoru Tsujii. Fractal materials and their functional properties. Polymer Journal (Tokyo, Japan), 40(9):785–799, May 28, 2008. CODEN POLJB8. ISSN 0032-3896 (print), 1349-0540

(electronic). URL http://www.nature.com/pj/journal/v40/
n9/abs/pj2008129a.html.

Titkov:1996:SBF

[TTMB96] Bronislav Titkov, Anatoli Tikhotskij, Alexandr Myboroda, and Helmut Buley. Structure-based fractal image coding. Proceedings of the SPIE — The International Society for Optical Engineering, 2952(??):652–663, ???? 1996. CODEN PSISDG. ISSN ????

Tang:2011:ESP

[TW11] Lian Tang and Da Hui Wang. An explanation of the shift parameter ρ in the Zipf–Mandelbrot law. Beijing Shifan Daxue Xuebao, 47(1):97–100, ???? 2011. CODEN BSDKDH. ISSN 0476-0301.

Uhl:1996:PIC

[UH96] A. Uhl and J. Hammerle. Parallel image compression on a workstation cluster using PVM. In Bode et al. [BDLS96], pages 301– ?? CODEN ???? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Uthayakumar:2007:ACF

[UP07] R. Uthayakumar and P. Paramanathan. An algorithm for computing fractal dimension of rectifiable irregular graphs. Applied Mathematics and Computation, 190(1):305–308, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300307000501.

Vallee:2011:BMF

[Val11] Robert Vallée. Benoît Mandelbrot — further notes. Ky-

bernetes, 40(9–10):??, ????  2011.  CODEN KBNTA3.

ISSN 0368-492X (print), 1758-7883 (electronic). URL http:

//www.emeraldinsight.com/journals.htm?issn=0368-492X&
volume=40&issue=9/10&articleid=17003220.

Vanraan:1990:FDC

[Van90] A. F. J. Vanraan. Fractal dimension of co-citations. Nature, 347 (6294):626, October 18, 1990. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v347/n6294/pdf/347626a0.pdf.

VanLoocke:2009:NLI

[Van09] Philip Van Loocke. Non-linear iterated function systems and the creation of fractal patterns over regular polygons. Computers and Graphics, 33(6):698–704, December 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S009784930900082X.

VanLoocke:2010:CBO

[Van10] Philip Van Loocke. Combination of basic origami with fractal iteration. Computers and Graphics, 34(1):66–71, February 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0097849309001149.

Verklan:1998:QNH

[VBM98] M. Terese Verklan, David R. Bickel, and Jon Moon. Quantification of neonatal heart rate variability from 26–35 weeks postconceptional age using the fractal dimension as estimated by dispersional analysis 1172. Pediatric Research, 43(4s):201, April 1998. CODEN ???? ISSN ???? URL http://www.nature.

com/pr/journal/v43/n4s/full/pr19981320a.html.

Vijayaraman:1996:PTS

[VBMS96] T. M. Vijayaraman, Alejandro P. Buchmann, C. Mohan, and Nandlal L. Sarda, editors. Proceedings of the twenty-second international Conference on Very Large Data Bases, September 3–6, 1996, Mumbai (Bombay), India. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1996. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996.

Voss:1975:NMS

[VC75] R. F. Voss and J. Clarke. 1/f noise in music and speech. Nature, 258(5533):317–318, November 27, 1975. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v258/n5533/pdf/258317a0.pdf.

Voss:1978:NMM

[VC78] R. F. Voss and J. Clarke. 1/f noise in music: Music from 1/f noise. Journal of the Acoustical Society of America, 63(??):258– 263, ???? 1978. CODEN JASMAN. ISSN 0001-4966.

Vlcek:1986:FAL

[VC86] J. Vlcek and E. Cheung. Fractal analysis of leaf shapes. Canadian Journal of Forest Research = Journal canadien de la recherche forestière, 16(1):124–127, 1986. CODEN CJFRAR. ISSN

???? URL http://www.nrcresearchpress.com/doi/abs/10.
1139/x86-020.

Vailati:2011:FFD

[VCM^{+}11] Alberto Vailati, Roberto Cerbino, Stefano Mazzoni, Christopher J. Takacs, David S. Cannell, Marzio Giglio, et al. Fractal fronts of diffusion in microgravity open. Nature Communications, 2(4):290, April 19, 2011. CODEN NCAOBW. ISSN

2041-1723 (electronic). URL http://www.nature.com/ncomms/
journal/v2/n4/full/ncomms1290.html.

Viennot:1989:CAR

[VEJA89] Xavier Gerard Viennot, Georges Eyrolles, Nicolas Janey, and Didier Arqués. Combinatorial analysis of ramified patterns and computer imagery of trees. Computer Graphics, 23(3):31–40, July 1989. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print),

1558-4569 (electronic). URL http://www.acm.org:80/pubs/
citations/proceedings/graph/74333/p31-viennot/.

Veitch:2005:MTI

[VHA05] Darryl Veitch, Nicolas Hohn, and Patrice Abry. Multifractality in TCP/IP traffic: the case against. Computer Networks (Amsterdam, Netherlands: 1999), 48(3):293–313, June 21, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

VanLoocke:2006:SBF

[VJ06] Philip Van Loocke and Yannick Joye. Symmetry breaking in fields as a methodology for three-dimensional fractal form generation. Computers and Graphics, 30(5):843–853, October 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0097849306001324.

Vlad:1994:SRG

[Vla94] Marcel Ovidiu Vlad. Stochastic renormalization group approach to random point processes. A fractal generalization of Poisson statistics. Journal of Mathematical Physics, 35(2):796–807, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i2/p796_s1.

Vlad:1995:GFA

[VM95] Marcel Ovidiu Vlad and Michael C. Mackey. Generating functional approach to multichannel parallel relaxation with application to the problem of direct energy transfer in fractal systems with dynamic disorder. Journal of Mathematical Physics, 36(4):1834–1853, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vuori:1998:MES

[VM98] Harri Vuori and Benoît B. Mandelbrot. The Mandelbrot echoes: for symphony orchestra, 1995. Edition Love, Helsinki, Finland, 1998. LCCN M1045.V96 M36 1998. 1 score (65).

Voldman:1983:FNS

[VMH^{+}83] Jean Voldman, Benoît Mandelbrot, Lee W. Hoevel, Joshua Knight, and Philip L. Rosenfeld. Fractal nature of software-cache interaction. IBM Journal of Research and Development, 27(2): 164–170, March 1983. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Vojta:1989:BRB

[Voj89] G. Vojta. Book review: Benoît B. Mandelbrot. Die fraktale Geometrie der Natur. Birkhäuser Verlag, Basel, Boston, 1987, 491 S., zahlreiche Abb. und Tab., SFr98.–, ISBN 3-7643-1771-X. Crystal Research and Technology, 24(2):166, February 1989. CODEN CRTEDF. ISSN ???? URL http://onlinelibrary.

wiley.com/doi/10.1002/crat.2170240207/abstract.

VanDamme:1986:FVF

[VOL^{+}86] Henri Van Damme, François Obrecht, Pierre Levitz, L. Gatineau, and Claude Laroche. Fractal viscous fingering in clay slurries. Nature, 320(6064):731–733, April 24, 1986. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v320/n6064/pdf/320731a0.pdf.

Vormoor:2002:LSF

[Vor02] Oliver Vormoor. Large scale fractal aggregates using the tunable dimension cluster-cluster aggregation. Computer Physics Communications, 144(2):121–129, April 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S001046550200142X.

Voss:1979:FNB

[Vos79] R. F. Voss. 1/f (flicker) noise: a brief review. In ????, editor, Proceedings of the Annual Symposium on Frequency Control (32nd) Held in Atlantic City, New Jersey on 31 May–2 June 1978, pages 40–46. Electronic Industries Association, Washington, DC, USA, 1979. ISBN ???? LCCN TK7872.O7 F72 1978.

Voss:1986:RFC

[Vos86] R. F. Voss. Random fractals: Characterization and measurement. Physica Scripta, T13(??):27–32, ???? 1986. CODEN PHSTBO. ISSN 0031-8949 (print), 1402-4896 (electronic).

Voss:1988:FNC

[Vos88] R. F. Voss. Fractals in nature from characterization to simulation. In Peitgen and Saupe [PS88a], pages 21–70. ISBN 0-387-96608-0 (New York), 1-4612-8349-3, 3-540-96608-0 (Berlin). LCCN QA614.86 .S35 1988.

Vrscay:1986:JSM

[Vrs86] Edward R. Vrscay. Julia sets and Mandelbrot-like sets associated with higher order Schröder rational iteration functions: a computer assisted study. Mathematics of Computation, 46(173): 151–169, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.jstor.org/

stable/2008220.

vanWijk:1991:SNT

[vW91] Jarke J. van Wijk. Spot noise: texture synthesis for data visualization. Computer Graphics, 25(4):309–318, July 1991. CO-DEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569

(electronic). URL http://www.acm.org:80/pubs/citations/
proceedings/graph/122718/p309-van_wijk/.

Wainer:1988:GFL

[Wai88] M. Wainer. Generating fractal-like surfaces on general purpose mesh-connected computers. IEEE Transactions on Computers, 37(7):882–886, July 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=2238.

Walter:1996:OFC

[Wal96] Malcolm Walter. Old fossils could be fractal frauds. Nature, 383 (6599):385–386, October 3, 1996. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v383/n6599/pdf/383385a0.pdf.

West:1999:FDL

[WBE99] Geoffrey B. West, James H. Brown, and Brian J. Enquist. The fourth dimension of life: Fractal geometry and allometric scaling of organisms. Science, 284(5420):1677–1679, June 4, 1999. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/284/
5420/1677.full.pdf.

Wang:2006:RFS

[WC06] Xingyuan Wang and Peijun Chang. Research on fractal structure of generalized M–J sets utilized Lyapunov exponents and periodic scanning techniques. Applied Mathematics and Computation, 175(2):1007–1025, April 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300305006582.

Wang:2003:DRT

[WCH03] Ching-Te Wang, Tung-Shou Chen, and Shao-Hau He. Detecting and restoring the tampered images based on iteration-free fractal compression. The Journal of Systems and Software, 67(2):131– 140, August 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Weibel:2005:MFGa

[Wei05] E. R. Weibel. Mandelbrot’s fractals and the geometry of life: A tribute to Benoît Mandelbrot on his 80^{th} birthday. In Losa et al. [LMNW05], pages 3–16. ISBN 3-7643-7172-2 (hardcover). LCCN ????

Westgaard:1994:MA

[Wes94] Odin Westgaard. A Mandelbrot analog. Performance + Instruction, 33(8):14–16, September 1994. CODEN PEISEL. ISSN 0884-1985 (print), 1932-0116 (electronic). URL http:

//onlinelibrary.wiley.com/doi/10.1002/pfi.4160330805/
abstract.

Werner:2012:MDK

[WGKS12] T. R. Werner, T. Gubiec, R. Kutner, and D. Sornette. Mechanisms of Dragon–Kings modeling of super-extreme events: An application to the hierarchical Weierstrass–Mandelbrot continuoustime random walk. European Physical Journal — Special Topics, 205:27–52, ???? 2012. CODEN EPJSAC. ISSN 1951-6355 (print), 1951-6401 (electronic).

Wheeler:1983:BRB

[Whe83] John A. Wheeler. Book review: The Fractal Geometry of Nature, Benoît B. Mandelbrot. 460 pp. Freeman, San Francisco, 1982. Price: $32.50. American Journal of Physics, 51(3):286–287, March 1983. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://link.aip.org/link/ajpias/
v51/i3/p286/s1.

Williams:1997:BFG

[Wil97] Nigel Williams. Biology: Fractal geometry gets the measure of life’s scales. Science, 276(5309):34, April 4, 1997. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//www.sciencemag.org/content/276/5309/34.full.

Winslett:2005:CFS

[Win05] Marianne Winslett. Christos Faloutsos speaks out: on power laws, fractals, the future of data mining, sabbaticals, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 34(4):85–89, December 2005. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wang:2022:BNF

[WLWL22] Yupin Wang, Xiaodi Li, Da Wang, and Shutang Liu. A brief note on fractal dynamics of fractional Mandelbrot sets. Applied Mathematics and Computation, 432(??):??, November 1, 2022. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300322004271.

Woolley:1994:RDP

[WM94] Stuart J. Woolley and Donald M. Monro. Rate-distortion performance of fractal transforms for image compression. Fractals, 2(3 (or 6??)):395–398, ???? 1994. CODEN FRACEG. ISSN

0218-348X. URL http://dmsun2.bath.ac.uk:8080/people/
papers/montreal.ps.gz.

Wolfram:2016:IMP

[Wol16] Stephen Wolfram. Idea makers: personal perspectives on the lives and ideas of some notable people. Wolfram Media, Inc., Champaign, IL, USA, 2016. ISBN 1-57955-003-7 (hardcover), 1-57955-005-3 (e-book), 1-57955-011-8. 250 (est.) pp. LCCN

Q141 .W678562 2016. URL http://www.wolfram-media.com/
products/idea-makers.html.

Wood:1994:GDC

[Woo94] Christopher M. Wood. General data compression algorithm for space images using fractal techniques. Proceedings of the SPIE — The International Society for Optical Engineering, 2198(??): 1336–1341, ???? 1994. CODEN PSISDG. ISSN ????

Weitz:1988:FAM

[WSM88] David A. Weitz, Leonard M. (Leonard Michael) Sander, and Benoît B. Mandelbrot, editors. Fractal aspects of materials: disordered systems, volume EA–17 of Materials Research Society extended abstracts. Materials Research Society, Pittsburgh, PA, USA, 1988. LCCN QA447 .F72 1988.

Wolf:2010:FOH

[WT10] E. T. Wolf and O. B. Toon. Fractal organic hazes provided an ultraviolet shield for early Earth. Science, 328(5983):1266–1268, June 4, 2010. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.sciencemag.org/content/
328/5983/1266.full.pdf.

Wu:1995:RSR

[Wu95] En-Hua Wu. A radiosity solution for random fractal surfaces. Journal of Visualization and Computer Animation, 6(4):219–230 (or 219–229??), October–December 1995. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Wu:1999:MCC

[Wu99] P.-Y. Wu. Minimum communication cost fractal image compression on PVM. In Dongarra et al. [DLM99], pages 434–441. CO-DEN ???? ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Wang:2008:INS

[WW08] Xing-Yuan Wang and Shu-Guo Wang. An improved no-search fractal image coding method based on a modified gray-level transform. Computers and Graphics, 32(4):445–450, August 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0097849308000320.

Wong:2005:FEF

[WWGF05] Angeline Wong, Leejay Wu, Phillip B. Gibbons, and Christos Faloutsos. Fast estimation of fractal dimension and correlation integral on stream data. Information Processing Letters, 93(2): 91–97, January 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1992:FNE

[WY92] Bennan Wang and Shi Yin. The fractal nature of an ecological model. Computer Graphics Forum, 11(1):61–64, March 1992. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:1996:PIF

[WY96] Zhou Wang and Ying-Lin Yu. Partial iterated function systembased fractal image coding. Proceedings of the SPIE — The International Society for Optical Engineering, 2751(??):42–49, ???? 1996. CODEN PSISDG. ISSN ????

Xingyuan:2009:GMS

[XRZ09] Wang Xingyuan, Jia Ruihong, and Zhang Zhenfeng. The generalized Mandelbrot set perturbed by composing noise of additive and multiplicative. Applied Mathematics and Computation, 210(1):107–118, April 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0096300308009119.

Xingyuan:2007:VGF

[XX07] Wang Xingyuan and Yu Xuejing. Visualizing generalized 3x + 1 function dynamics based on fractal. Applied Mathematics and Computation, 188(1):234–243, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300306013439.

Ye:2002:ACO

[Ye02] Ruisong Ye. Another choice for orbit traps to generate artistic fractal images. Computers and Graphics, 26(4):629–633, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-

7684 (electronic). URL http://www.elsevier.com/gej-ng/10/
13/20/68/55/44/abstract.html.

Young:1993:ICP

[YHHS93] Bong Kim Young, Sub Kim Hoi, Oh Kim Hong, and Yong Shin Sung. Infinite-corner-point fractal image generation by Newton’s method for solving exp(−αζ + Z/ζ − Z) − 1 = 0. Computers and Graphics, 17(6):705–711, November–December 1993. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Yu:2009:SFM

[YJ09] Zhibin Yu and Hai Jin. Simple and fast micro-architecture simulation: a trisection Cantor fractal approach. ACM SIGMETRICS Performance Evaluation Review, 37(2):61–62, September 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Yekutieli:1994:HSO

[YM94] Iddo Yekutieli and Benoît B. Mandelbrot. Horton–Strahler ordering of random binary trees. Journal of Physics A (Mathematical and General), 27(2):285–293, ???? 1994. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic). URL http://stacks.

iop.org/0305-4470/27/285.

Yekutieli:1994:SSB

[YMK94] Iddo Yekutieli, Benoît B. Mandelbrot, and Henry Kaufman. Selfsimilarity of the branching structure in very large DLA clusters and other branching fractals. Journal of Physics A (Mathematical and General), 27(2):275–284, ???? 1994. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic). URL http://stacks.

iop.org/0305-4470/27/275.

Young:2013:ATL

[You13] D. S. Young. Approximate tolerance limits for Zipf–Mandelbrot distributions. Physica A, 392(7):1702–1711, April 1, 2013. CO-DEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378437112010497.

Yue:2010:FFR

[Yue10] Hong Yue. A fractal function related to the John–Nirenberg inequality for Q_{α}(R^{n}). Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5):1182–1200, October 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wang:2007:APG

[yWjCnG07] Xing yuan Wang, Pei jun Chang, and Ni ni Gu. Additive perturbed generalized Mandelbrot–Julia sets. Applied Mathematics and Computation, 189(1):754–765, June 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300306016535.

Zaanen:2010:HTS

[Zaa10] Jan Zaanen. High-temperature superconductivity: The benefit of fractal dirt. Nature, 466(7308):825–827, August 11, 2010. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v466/
n7308/full/466825a.html.

Zakeri:2006:BPM

[Zak06] Saeed Zakeri. On biaccessible points of the Mandelbrot set. Proceedings of the American Mathematical Society, 134(8):2239– 2250, August 2006. CODEN PAMYAR. ISSN 0002-9939 (print),

1088-6826 (electronic). URL http://www.jstor.org/stable/
4098261.

Zhao:2009:SNF

[Zha09]  Hong Zhao.  Some new fractals behaviors in (2 + 1)-

dimensional integrable system. Applied Mathematics and Computation, 215(5):1653–1658, November 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300309006651.

Zahlten:1990:FHE

[ZPSJ90] Cornelia Zahlten, Heinz-Otto Peitgen, Dietmar Saupe, and H. (Hartmut) Jürgens. Fractals: an animated discussion. Films for the humanities and sciences, Princeton, NJ, USA, 1990. ISBN 0-7365-6093-9. LCCN QA614.86 .F76 2003. One DVD (63 min).

Zair:1996:FMU

[ZT96] Chems Eddine Zair and Eric Tosan. Fractal modeling using free form techniques. Computer Graphics Forum, 15(3):C269–C278, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhu:2006:AFT

[ZWZL06] Xuegui Zhu, Yi Wang, Jianming Zan, and Chuanbing Li. Application of fractal theory in generation and refinement of finite element mesh. Applied Mathematics and Computation, 175(2):1039–1045, April 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0096300305006600.

Zhao:1994:ICU

[ZY94] Y. Zhao and B. Yuan. Image compression using fractals and discrete cosine transform. Electronics Letters, 30(6):474–475, March 1994. CODEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

Zhang:1996:HFI

[ZZZ^{+}96] Zhengbing Zhang, Yaoting Zhu, Guangxi Zhu, Hanqiang Cao, and Donghui Xue. Hybrid fractal image coding method. Proceedings of the SPIE — The International Society for Optical Engineering, 2727(??):1360–1366, ???? 1996. CODEN PSISDG. ISSN ????